A Bibliography of Publications of the USENIX Association: 2000–2009

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

06 February 2024 Version 2.14

Title word cross-reference

3 [CB00]. >> [Pat02a]. N [Sal00d]. O(1) [NVZ01]. x = 1 [Cou08a].

.profile [How03].

/dev [Kam02]. /dev/random [Fer06a, Fer06b, Fer06c, Fer06d, Fer07a, Fer07b, Fer07c, Fer07d, Fer07e, Fer07f, Fer08a, Fer08b, Fer08c, Fer08d, Fer08e, Fer08f, Fer09f, Fer09a, Fer09b, Fer09c, Fer09d, Fer09e]. /etc [Sim03].

’00s [Zad02]. ’01 [Ano00g, Ano00f, Ano00d, Ano01e, Ano01j, USE01g, USE01c]. ’02 [Ano02f, USE02c, USE02f]. ’03 [Ano03f, Ano04d, Kam04, Ano03g]. ’04 [Ano04h, Ano04e, Ano04i, Ano04c, Ano04f, Ano05b, Ano05l]. ’05 [Ano05e, Ano05k, Ano05m, Ano05n, Ano06a, Ano06b, Ano06c].

1 [App07a, Bur06b]. 101 [Gal08d, Nic02b, Uhl03]. 102 [Gal08e]. 106461 [LaB00]. 10th [Ano01b]. 11th [Ano02a, USE02b]. 12th [Ano03a]. 13th [Ano01i, Ano04c]. 14th [Ano00a, Ano01c, Ano05a]. 15th [Ano02g, Ano03b, USE01a]. 17th [Ano04d]. 180x86 [Law00]. 18th [Ano05b]. 19th [Ano06b]. 1st [Ano01d, Ano01e].

2 [App07b, Bur06c, Luc00, Mar01b]. 2.0 [Bla01a, Cor00]. 2.4 [vR01a]. 2.4.17 [BFH02]. 2000 [BCGD00, Luc00, USE00a]. 2001 [Ano01g]. 2002 [Jos02, Ano02e, Ano02h]. 2003 [Ano03o, Ano03j, Ano03k, Ano04g, Ano04d]. 2004 [Ano03p, Ano04a, Ano05g]. 2006 [Ano05s, Ano06d]. 216-Processor [MDH00]. 21st [Joh06]. 28-30 [USE02c]. 2k [USE00e]. 2nd [Ano02d, Ano03e, Ano04b, Ano05e, Chu04b].

3 [Bea09, Bur06d, HBG^{+}06]. 3.0 [MS00]. 32-bit [Min00]. 36-User [Sol08]. 390 [Tho00b]. 3D [KD02]. 3rd [Ano00d, Ano01j, Ano04e, Ano04f, USE00b, USE01c].

4 [Bur07, Gal09f]. 4.x [HC00, Siv03]. 4000 [Fly03c]. 4th [Ano00e, Ano03f, Ano06a, USE00c, USE00d].

5 [CL01]. 5.2 [McK04a]. 5th [Ano01f, USE01d].

6 [Pis08, Sal01b]. 64 [vdL02]. 6th [Ano00f, Ano05f, USE01e].

7th [USE00e].

802.1 [Nik02].

9 [CGP^{+}02, Pik00a, Siv03].

a.k.a [Far04e]. Aberrant [Bru00b]. Ability [RI00b]. Abstract [McC04c]. Abstraction [RMSK00]. Abyss [EKNG^{+}05]. Accelerator [RCK02]. Access [BSF02, Bye02b, Bye02a, Fou02, HRF00, Jos09c, KHDC01, Mag02, MAF^{+}02, MoCA^{+}02, SGLG02, SV02b, SZ02, VB00]. Access-Control [BSF02]. Accessory [Won01]. Account [Hug00]. Accuracy [HR01a]. Accurate [BBD^{+}01, GSS^{+}02]. A’Changin [Hum00b]. Achievement [Coc01]. Achieving [McC08]. ACM [Ano01d, Ano05m, Jac04]. ACM/IFIP/USENIX [Jac04]. ACM/USENIX [Ano05m]. ACPI [WS02]. Act [App07a, Eck01, NW04]. Action [Tyl02d]. Actions [MY04, Mul05a, Mul05b, MY06, Veg07, You06c, You07e, You08a, You09b]. Activations [Wil02]. Active [Ark02a, Die02, JBH^{+}00, Orm00a]. Activities [Sto06a]. Activity [MVS01, Orm00b]. Ad [Bur06d, Cha01d, SLR02, SDN^{+}08]. Ad-Hoc [SLR02]. Ad-Hocracy [Bur06d]. Adaptation [ABC^{+}00, dLWZ00, dLWZ01]. Adaptive [CKV^{+}02, FSAK01, HBC^{+}01, JZ01, MW05, MFJPPMK04, NSK04]. Add [WS01]. Added [Dar03a]. Adding [BC06, Wat01]. Address [BM00a, NSZ01, Sch00c, And01]. Addressable [Mud00a]. Addresses [GB01]. Addressing [BCD^{+}00, Che00b, GB01]. Admin [Jos08a, Llo07]. Administration [AHK^{+}02, Ano00a, Ano01c, Ano02g, Ano04d, Ano05b, Ano06b, Bru09a, Bur00b, Bur00c, Bur00d, Bur00e, Bur05, Car00b, Cou08c, ECF^{+}00, Kan05b, MD00, MNH00, See06, Sta00, Tra02, USE00g, USE00b, USE01a]. Administrative [SZ02]. Administrator [Cou08b, Dij00d]. Administrators [Gal09a, Mil05, Pei05]. Advanced [Ger02, GHH06, Dan01]. Adventures [Man02]. AES [Dan01]. Affiliated [Sch01]. Affordable [KHF01]. AFS [FZ01]. After [SS05]. Against [BST00, Hoh01, KS00, KM01, Mue07a, Pro01, Syv01]. Age [CK01]. Agent [DD07, McC00g, RB01]. Agentk [McC00g]. Agents [Fly04e]. AGFL [KV02]. Aggressive [JN00]. Agile [Sch02b]. Aging [Whi00]. AGNI [RBPM00]. Ago [Sal00d, Sal00c, Sal00a, Sal00b, Sal03f, Sal03e, Sal03c, Sal04i, Sal04a, Sal04k, Sal05d, Sal05h, Sal05i, Sal05c, Sal05e]. Aid [BS02]. Air [RMS02]. Airline [OSE^{+}00]. Alain [Kol04a]. Alden [Far00i]. Alert [ABQ02]. Algorithm [CAGS00, KL01, ZPL01]. Algorithms [HPS02, Joh06, SLR02, ZABN01]. Aligned [SGLG02]. Alike [Eck01]. Alistair [Sch02b].

Allocation [CML00, MW05, Sac08]. Allocator [BA01]. ALMI [PSVW01]. alone [LB00]. Alpine [ESW01]. Already [Sal04b]. Alternative [LR04, SSB00]. Alternatives [Cha07c]. Amazon [Gar07, Wal08]. AMD [vdL02, Zad02]. Among [PF04]. AMPing [Gal09b]. Analysis [AC08, Bal03, BS01, CWVL01, Dij00d, Gal08d, Gal08e, Gol06, Hun00, KM01, LM00, Lev00, Mee00, MDH00, MNH00, RI00b, SWT01, SDHD07, Syv01, ZEJ02]. Analytics [Gal09a]. Analyzer [Woo01b]. Analyzing [ABQ02, DLD00, Fly03c, PLT00]. Animals [BE01]. Animated [MZ00]. Anniversaries [Sal03a]. Anniversary [Sal01a, Sal05g]. Announcement [Ano00e, Ano00g, Ano00a, Ano00b, Ano00c, Ano00f, Ano00d]. Annual [Ano00e, Ano00b, Ano00c, Ano00h, Ano00p, Ano01i, Ano01k, Ano02c, Ano03b, Ano03k, Ano04i, Ano05d, Ano05t, USE00c, USE00a, USE01d, USE01f, USE01b, USE02d, USE02e, USE06]. Anonymizing [MS02]. Anonymous [BC00]. Answers [Bea01a]. Antfarm [JADAD06]. Anti [App04b, Has06a, Sam02]. Anti-Circumvention [Sam02]. Anti-Spam [App04b, Has06a]. Anton [Far04a]. Antonio [USE01e]. Anyway [Bel02]. Apache [Pie03]. Appia [WOSW02]. Applet [Ros00]. Appliance [Ban00, Cha07b, Sha00, Van09b]. Application [CV02, Ing00, Kri00b, LGW04, Mar01b, OESV04, PSVW01, PN05, PW00, Sch02a, Sch00b, SNC00, Tie01, YMC02, ZS01, APH00, BE09e, BE09f]. Application-level [SNC00]. Application-Specific [YMC02, ZS01, CV02]. Applications [AR00, ABD^{+}08, ABC^{+}00, Ano05f, Ano05j, Ano05n, Bou01, BR00a, Dam03, FM00, Gal09c, Kal09, KD02, LB00, LR04, Mah04, PA04, PMN^{+}04, PS01, RAMC^{+}04, Sto02, VKK^{+}01, WRA00, ZM00]. Applying [Fly00e]. Approach [Apt01, BJP00, CD01a, Fin00a, MV04, NSZ01, PRS06, QD02]. Appropriate [Dar03b]. Approved [Jos02]. Approximate [BC00]. Apps [Ash02]. April [USE01g]. Apropos [Dar00a, Dar00b, Dar00c, Dar00d, Dar00e, Dar00f, Dar01a, Dar01b, Dar01c, Dar01d, Dar01e, Dar01f, Dar01g, Dar02a, Dar02b, Dar02c, Dar02d, Dar04a, Dar04c]. Aqueduct [LAW02]. Arbitrary [BA01]. Archipelago [JFWS00]. Architect [Ano04b]. Architecting [Woo01b]. Architectural [JR02]. Architecture [And00, AC08, CW04, EKNG^{+}05, GMAM01, GC01, Hun00, Pac01b, RI00a, SMF^{+}01, Tra01, YMC02, Yen00, KW00, WT00]. Architectures [BLF^{+}00, LR04]. Archival [MB02a, QD02, SGMV08]. Area [KCN02, Mee00, PGCR01, SVF01, Sch02c, Tie01, WOSW02]. Arena [Fer09f]. Argos [BTS08]. Aroma [Sur01]. Array [McC01b, YGC^{+}00, ZYKW02]. Arrays [ASV^{+}02, McC01g]. Arrived [Gra02]. Art [Rus00b, Slu05]. Article [Kol05d]. Artifacts [Too09]. Artist [Ham00]. Arts [Ham00]. Ask [Har02b]. ASR [Cha07a]. Assembly [Bro06]. Assessment [Iam01]. Assisted [sLC02]. Association [Ano03o, Ano05u, Sal05g, The04, You04b, You06b, You07d, You09d]. Assurance [SV02b]. Asterisk [Cha05a, Cha05b, Cha07b, Cha07a, Cha07c, Cha07d, Sol08]. Asynchronous [PMF^{+}02, SGM^{+}00]. Atlanta [USE00c, Ano00e]. Atoms [Mor02]. atque [You05b]. Attached [MLFR02]. Attachments [BE08b]. Attack [GP01b, KS00, Loz05b]. Attacks [BST00, BGM00, BU02, Has05d, Kan05a, Lem02, Oli01, SWT01]. Attributes [Sto08a]. AUC [EKNG^{+}05]. AUC-Abyss [EKNG^{+}05]. Audit [CMP00, CMP00]. Audit/Logging [CMP00]. Auditing [Lan06]. August [USE00b, USE00i, USE00d, USE01h, USE02f, USE02b, Jos02]. Auralizer [GC00, Gil01]. Austin [USE00e, Jos00a, Jos00c, Jos01a, Jos02, See01]. Authenticated [BU02]. Authentication [BH00a, DP00, FSSF01, Way02, Coc01]. Authorisation [RJ01]. Authorization [HK00b, Nik02, Ras06, ZEJ02]. Auto [Ban00]. Auto-Diagnosis [Ban00]. Autoconfiguring [Kno00]. Automated [All02b, MB00, McC05b, RV00, SF00, Tra02]. Automatic [CBB^{+}01, KYH^{+}01, Kri00b, San01, SMES01, WOSW02]. Automating [AFE00, Fre06, Hem00, She00]. Automation [MD00, Tur04d]. Autonomic [Bur06a, FF07]. Autotools [Zad02]. Availability [Ash02, BP00, CDGN01, Hob02, LR01, LS00, Pat02a, Zwi01]. Available [Dic09, JFWS00, ZSB04]. Ave [You05b]. Average [Cou08b]. Avoid [Nic03a]. Avoiding [Ale04a]. Awards [Ano05t, Ano04a]. Aware [CW04, FADAD02, HS00b, MV04, ASDZ00, CV02]. Awareness [CDRV00, MZ00, RAMC^{+}04, ZM00]. Awash [Lea04]. AX [Fly03c]. AX-4000 [Fly03c]. AxTcl [Fly03c].

Babble [Bur06c]. Babel [Bur06c]. Back [BE06c, BE08c, Gee05, Gra04]. Backbone [Ald05]. Backdoors [ZP00a]. Background [McK02]. Backup [Hos03, VSV09]. Backus [Aik07]. Bad [Bea09]. Balancing [App07a]. Balkan [Pie00]. Bandwidth [ABC^{+}00, GP01b, LB01, LSG^{+}00, dLWZ00]. Base [Jos01b]. Based [Ark02b, BGM00, BBLo04, Bon01, Chu03, EG01, GHH06, GSAA04, JGKvS04, JN00, JFWS00, JKN^{+}01, KP01, Lam00, MSO00, OESV04, OP01, PG01, PMF^{+}02, Ran02, Sch02a, Sha02a, Spi00, WJL04, ASDZ00, FSLM02, GSGN00, Hem00, Hob05, SYC^{+}01, WG00, dLWZ01]. Basics [Van09a]. Basis [KPv00]. Bastion [Nor00]. Batch [BHK^{+}00]. Battery [CKV^{+}02]. Battery-Operated [CKV^{+}02]. Bayesian [HK03]. Bcfg2 [DBH06]. Be [BE09a, Bro06, Mud00b, Tyl01, Tyl02b, Tyl02c]. Bean [WCD^{+}01]. Beaten [Cha01b]. Becoming [HLU^{+}00]. Been [ECF^{+}00, Nic01a]. Beginning [Mar01a]. Behavior [AR00, Bru00b, Ken02c, Lev00, PLT00, Reg02, VKK^{+}01, YD00]. Behavioral [Ken02b]. Behaviour [HS01]. Behavioural [WS01]. Being [Cou09, Lea04, Sch06]. Belong [Sim03]. Benchmark [BRS01, ZS01]. Benchmarking [Wal08, YNN01]. Benchmarks [BP00]. Benefits [MD05]. Beowulf [Bec01, Bro00, Bro01b, MCB^{+}01, MDH00, Ste01]. Berkeley [USE01g]. Better [Fiu05, Sit02, WRPK02]. Between [CWL^{+}01, MCD^{+}02, MB01, Pri01]. Beyond [Lan08, Luc00, SNOK01]. BGP [Jos06a]. Big [Hel09]. Biglook [GS02]. Billing [HDM00]. Binary [YST02]. Bind [PN05]. BIND8 [Siv01]. Binding [Mas01]. Biological [GB04]. Biometric [Way02]. BIOS [MHW00]. Bit [ASS02, Cha01b, Min00]. Bit-Rate [ASS02]. Bits [HSW09]. BitTorrent [PIA^{+}07]. BitTyrant [PIA^{+}07]. BL [Lan08]. Blacklisting [ZPU08]. Blaine [Fen00]. Blame [JHZK08]. Blanket [Gal09c]. BLAS [Tho00a]. BLASTH [Tho00a]. Blocking [BGM00, Has06b]. Blog [Cha02d]. Blogs [App07a, App07b]. Blood [Tyl01, Tyl02b]. Blotter [Ran00a, Ran00b, Ran00c, Ran00d, Ran00e, Ran00f]. Blunders [Bur01g]. Board [Ano03o, Ano04r, MY04, Mul04b, Mul05a, Mul05b, MY06, Veg07, YM05, You06c, You07a, You07e, You08a, You09a, You09b, Ano05r, BY00a, BY00b, You03]. Bombs [JM03]. Book [Ano01a, Ano04b, Chu02a, Chu02b, Chu04a, Chu04b, Far04a, Far04b, Far04c, Far04d, FH05, Far05a, Fly00e, Fri05a, Fri05b, FKS08, Har02a, Has03a, Has03b, Joh02, Kar02, Nat04, Pur02, Sal00e, Sal00f, Sal00g, Sal00h, Sal00i, Sal00j, Sal00k, Sal01c, Sal01d, Sal01e, Sal01f, Sal01g, Sal01h, S^{+}01, Sal01i, Sal02b, Sal02a, Sal02c, Sal02d, Sal02e, Sal02f, Sal03b, Sal03d, Sal03f, Sal03e, Sal03c, Sal03a, Sal04j, Sal04c, Sal04d, Sal04e, Sal04f, Sal04b, Sal04h, Sal04g, Sal04i, Sal05b, Sal05a, Sch02b, Sch02c, Sch02d, ZT05, Zwi05, Z^{+}06a, ZSF06, Z^{+}06b, Z^{+}06c, Z^{+}06d, Z^{+}06e, Z^{+}07a, Z^{+}07b, Z^{+}07c, Z^{+}07d, Z^{+}07e, Z^{+}07f, Z^{+}08a, Z^{+}08b, Z^{+}08c, Z^{+}08d, Z^{+}08e, Z^{+}09a, Z^{+}09b, Z^{+}09c, Z^{+}09d, Z^{+}09e, Z^{+}09f]. Booklets [Far05h]. Bookshelf [Bla00b, Bla00c]. Bookworm [Fri05a, Fri05b, Sal00e, Sal00f, Sal00g, Sal00h, Sal00i, Sal00j, Sal00k, Sal01c, Sal01d, Sal01e, Sal01f, Sal01g, Sal01h, S^{+}01, Sal01i, Sal02b, Sal02a, Sal02c, Sal02d, Sal02e, Sal02f, Sal03b, Sal03d, Sal03f, Sal03e, Sal03c, Sal03a, Sal04j, Sal04c, Sal04d, Sal04e, Sal04f, Sal04h, Sal05b]. Boost [WH02]. Boot [AFE00]. Bootcamp [Ano04b]. Booting [Min00]. Boots [Min00]. Border [Han03]. Boston [USE01f, USE01b, USE06]. Both [JW01a]. Botnets [DAK08]. Bots [Hol05]. Bottle [Jos09b]. Bottleneck [LB01]. Boundaries [Ken02b]. Box [Bal03, Ber02, BBADAD02]. Boxes [Rus03]. Branch [DVWW01]. Breaches [App04a]. Breaking [Par02a, PN05]. Brews [Jos07c]. Brian [Har02a, Sel08]. Bricks [Gra02]. Bridging [DADAD02]. Brief [McK07]. Briefcases [Mue07b]. Bringing [Adl00]. Brings [Kol05e]. Broadband [Sch02a, Nat04]. Broken [Tur03c]. Broker [WPSO01]. Browse [ABQ02]. Browser [BE09a, GTK08, IB01, Mey09, MM00, Uhl03]. Browsers [DEH^{+}09, YS02]. BSD [Gla02, McK07, iiH01]. BSDCon [Ano02e, Ano03g, USE02a]. Bubble [SS05]. Buffer [Ale04a, Ale05a, BGM00, BBADAD02, JN00, KCK^{+}00, LE01, sLC02, WH02, ZPL01]. Buffer-Cache [BBADAD02]. Build [Gol00, Guy01, Kri00b, LM00, MVS00]. Building [AB04, GTK08, IB01, Kar02, Kno00, Kno08, Nor00, PIA^{+}07, Uri06, WRA00, Woo01a]. Built [Van09b]. Bundling [HP01]. Burden [MD00]. Burgess [Far07g]. Burnham [Fen00]. Business [HD02, Sch02e]. Businesses [Bro01c, DVWW01]. Busy [LSG^{+}00]. BUTS [JW01e]. Buy [Van01]. Buzzword [LM00]. Buzzword-Compliant [LM00]. Bypass [LHAP06]. Byte [WS01]. Bytecode [GH01, SW01]. Byzantine [CL00]. Byzantine-Fault-Tolerant [CL00].

C [Ano04b, BPL07, DLGD01, Gsc01, JMG^{+}02, Kri00b, Mas01, Sto07a, dD00]. C# [McC03a, McC03b, McC03c, McC03d, McC03e, McC04a, McC04b, McC04c, McC04d, McC04e, McC05a]. C99 [McC02a]. C9X [McC01a, McC01b, McC01e, McC01f, McC02b, McC02c, McC02d]. CA [USE00a, USE00j, USE01d, USE01g, USE01a, USE02a, USE02b]. Cabbage [Bur06b]. Cache [BBADAD02, CK01, JN00, Lam00, Lem02, Sea00, WW02, YHS^{+}02]. Cache-Based [Lam00]. Caches [Tom00, ZPL01]. Caching [FADAD02, YC01]. Cailin [SvM05]. Calendaring [Sif00]. California [USE00h, USE00f, USE01g, USE01c, USE02c, USE02d, USE02e, USE02f, App04a, NW04]. Caliper [Hun00]. Call [Com00, GJM02, SF00, Ano00c]. Called [Tyl01, Tyl02b]. Campus [AC08]. Can [BE01, BS01, Bro06, Bru09b, Fra01, GW07, Ran05, RR00, Smi01, Swa03b, App04b]. CAN-SPAM [App04b]. Canada [Jac04]. CANS [FSAK01]. Capabilities [BDTW01, Mal05]. Capability [RJ01]. Capacity [YGC^{+}00]. Capsule [Mas02]. Capture [Mee00, Sur01]. Capturing [Fly03c, HcFG01]. Car [BE06a]. Card [Los02]. Cards [Fou02, Kal09]. Cars [Tyl02c]. Case [BU02, BV00, BP00, DLD00, Ezo01, Jon02, Zad02, vdL02]. Castle [Har02a]. Catch [Dit09b]. Categorization [LV02]. Cathedral [USE02a]. Causes [Mur01]. CDs [GS06]. CE [Sta00]. Censorship [FBH^{+}02, WRC00]. Censorship-Resistant [WRC00]. Center [Drz00, Lif01]. CenterTrack [Sto00a]. Central [SS00]. Centralized [SPC08]. Centre [Krz01]. Centric [Hug00]. Century [Joh06, Pat02a]. Certificates [BDTW01, Ren02]. Certification [Dij00a, Hop01]. Certs [But04]. Cerven [Sal03d]. Cfengine [RV00]. cflow [Gol06]. CFS [How04, Ito01]. CGI [BE09e, BE09f, FS00, Tur05b]. Chair [Ano06e]. Challenge [CWL^{+}01]. Challenged [DDB08]. Challenges [Tar02]. Change [How06a, Siv01]. Change-Rooted [Siv01]. Changes [Mur01]. Changing [Hug00, TSW08, ZABN01]. Channel [BU02]. Chapter [Jos06a]. Characteristics [SGLG02]. Characterizing [ABQ02, YD00]. Characters [McC02e]. Charging [Nik02]. Charm [HcC01]. Chat [Cou00]. Checking [ACG01, ECCH00, JJR02]. Checkpointing [DL01]. Chimera [KHF01]. Chips [vD09]. Choosing [Cha06a, Mos03b]. Chosen [KS00]. Chris [Chu04b]. Christmas [Mar01a]. Chunks [JW00a]. Chuvakin [Far04a]. CIO [Dar04c]. Ciphertext [KS00]. Circles [AHK^{+}02]. Circumventing [FBH^{+}02]. Circumvention [Sam02]. CISCO [RFL00, NNW00]. Citrus [iiH01]. City [Kri00a]. Clark [Sch02c]. Class [BLF^{+}00, HP01, Tur04b]. Classes [AR00, McC03e, McC04c]. Classfiles [FC01]. Classloaders [FC01]. Cleaning [Tur03a]. Client [BLF^{+}00, DS01, Fly03d, FSSF01, LH02, PIA^{+}07, RE01, YNN01, YNST02]. Client-Perceived [RE01]. Client-Server [BLF^{+}00]. Clients [ABQ02, MCD^{+}02, McQ00]. Climate [Sch02e]. Clock [GLF^{+}00]. Cloning [Ama00]. Closer [BGS^{+}08]. Cloud [Bur09, VSV09]. Cluster [APGG00, AK01, Ano01d, Ano01f, ASDZ00, CB00, ECF^{+}00, HD02, JM03, Krz01, MCB^{+}01, MDH00, OF00, PBB^{+}00, SYC^{+}01, Ste01, US01]. Cluster-based [ASDZ00, SYC^{+}01]. Clustered [RCK02]. Clustering [Lum01]. Clusters [BHK^{+}00, Bok01, CLRT00, Chi00, HR01b, Hum02, Lif01, Min01, Sch01, SH02]. CNN.com [LeF02]. CoBlitz [PP06]. Cockburn [Sch02b]. Code [DEH^{+}09, Gra00b, Gra00c, Gra00d, Gra01a, HLU^{+}00, Ker06, KM01, PBP02, WS01, Wys04, Chu04a]. Cohort [LP02]. Collaborative [BLF^{+}00, LB00, PMN^{+}04, SMM04]. CollabWiseTk [LB00]. Collect [AHL02]. Collection [BR00a, CKV^{+}02, DKCJ02, FDSZ01, SVF01]. Collections [Kuh00, McC04b]. Collective [MoCA^{+}02]. Collector [Pri01, ZS01]. Collide [EVS^{+}08, Fin05]. Colorado [USE00i]. Column [BE08e]. Com [Lie00]. Combating [Woo02]. Combining [NNW00, Ras04]. Come [Bel02, Ioa02]. Comes [Ano00s, JW01c]. Coming [Dan01, Get02a, Ros01]. Command [DD07, Ing00, MM00, See06]. Command-line [Ing00]. Commentary [Kam04]. Comments [Gee03a]. Commercial [App04b, JBH^{+}00, Pis08]. Commercial-Grade [JBH^{+}00]. Committee [Ano03o, Ano03p, Ano04m, Ano05s, Ano06d, JG08, Jos02, Nem00a, You07b]. Commodity [Gar07]. Common [Nic03a, Ham01]. Communication [HPS02, SMF^{+}01, Whi01]. Communications [Eck01, Sal01a]. Communicative [Cha02d]. Company [JW00e]. Comparison [Cha06i, OF00, RLA00]. Comparisons [BHHV00]. Compatibility [McC02a, SM02, YST02]. compatible [Woo01a]. Competitive [Iam01]. Compilation [CO02, YST02]. Compiler [Ale05a, BM00b, ECCH00, Mid01, PVC01, Sha02b, SYN02]. Compiler-Inserted [BM00b]. Compiler-level [Ale05a]. Compilers [Mar01b]. Complex [KPv00]. Compliance [KA01]. Compliant [LM00]. Comply [Jos08b]. Component [FSLM02, KYH^{+}01, RFS^{+}00, dLWZ01]. Component-based [FSLM02, dLWZ01]. Components [Gra00a, OESV04, WCD^{+}01]. Composable [FSAK01]. Composite [KT01]. Composition [JN04, RFS^{+}00, SNC00, TKM04, WCD^{+}01]. Comprehensive [MNH00]. Compressing [Pie03]. Compression [VK02]. Compromised [Ale05b, SGS^{+}00]. Computer [Ano01i, Apt01, Bag04, Cho05, Chu04b, Dit00, GB04, HS01, Ken02a, MHM^{+}00, Mit02, SF07]. Computers [MMBE00, SGZ^{+}02, Tho00a, Tyl01, Tyl02b, Ver07]. Computing [APGG00, AK01, Ano01d, Ano01h, Ano05f, Bur06a, Cha06i, FF07, FS00, Gar07, Guy01, HBC^{+}01, Hum02, Kol04g, Kol05f, KW00, Lum01, Nic00d, RAMC^{+}04, Rub01a, SH02, Ste01, Thr01b, Wal08, YNST02, dLWZ01]. Concealed [PSW^{+}01]. Concepts [Die02]. Concurrent [Bro08, DKCJ02, Dus08]. Conference [Ano00e, Ano00a, Ano00b, Ano00c, Ano00f, Ano00h, Ano01i, Ano01g, Ano01d, Ano01e, Ano01h, Ano01c, Ano01j, Ano01k, Ano01b, Ano01f, Ano02e, Ano02f, Ano02d, Ano02g, Ano02b, Ano02h, Ano02c, Ano02a, Ano03b, Ano03g, Ano03c, Ano03h, Ano03i, Ano03j, Ano03d, Ano03f, Ano03e, Ano03k, Ano03a, Ano04g, Ano04h, Ano04d, Ano04e, Ano04i, Ano04c, Ano04f, Ano05g, Ano05h, Ano05f, Ano05i, Ano05b, Ano05c, Ano05j, Ano05l, Ano05e, Ano05k, Ano05d, Ano05a, Ano05m, Ano05n, Ano06a, Ano06b, Ano06c, Jac04, Sal01b, Sin05a, USE00c, USE00g, USE00b, USE00e, USE00a, USE01d, USE01f, USE01a, USE01b, USE01e, USE02c, USE02d, USE02e, USE06, Ano00h, Ano01k, Ano02g, Ano02c, Ano03b, Ano03d, Ano05d, Ano06b]. Confessions [Tyl03a]. Configuration [AS00, And06, BE06b, Bur06b, Bur06c, Bur06d, Bur07, Cou08a, Fin00a, GR07, Kan06a, KYH^{+}01, KNRS00, MB01, San01, SV02b, Sha02a, Siv03]. Configurations [SAB^{+}01a]. Configuring [ZYKW02]. Confine [AR00]. Conflicts [Dit09a]. Congestion [SK02]. Connect [NSZ01, MSO00]. Connected [BS00, MMBE00, OF00, SMES01]. Connection [Ayu06, SAB01b]. Connections [BCR07]. Conquer [Ran05]. Conquest [WRPK02]. Consider [Lem00]. Considerations [Rub01b]. Consistency [YV00]. Consolidation [RMSK00]. Consortium [Hoh01]. Constrained [BCH^{+}00, CIK00, MRL^{+}02, Sha02b]. Constraint [BNB01]. Constraint-Enabled [BNB01]. Construction [GBHC00, Krz01, Sha00]. Consulting [Cha01a, Cha02a, Pur06]. Contained [Kno00]. Containers [Gal05]. containing [Jos02]. Containment [PBP02]. Content [ABC^{+}00, ASDZ00, Coo04a, EG01, GC01, GSAA04, OP01]. Content-aware [ASDZ00]. Content-Based [EG01, GSAA04, OP01]. Contest [Kol04h]. Continuous [YV00]. Contract [Cha00b]. Contractors [Cha01c]. Contracts [MMGL01]. Control [BSF02, Cha08a, CM01b, DD07, How05, Jos09c, Kan05b, KHDC01, SK02, See06, SV02b, Sur01, ZD02]. Convergence [BS04]. Conversation [GK04]. Cookies [MSV09]. Cooperative [AHT^{+}02]. Coordinated [For03, TKM04]. Coordination [OMP02]. COOTS [Ano00f, Sal01b]. Copley [USE01f, USE01b]. Coprocessor [Gut00, Ito00, LS01a]. Cops [Per01]. CORBA [BV00, FS00, MMGL01, Rao01a, Rao01b]. Core [Gle02, Joh02, ML06]. Cornell [Lif01]. Corporate [Nic01b]. Correct [Kle09]. Correctness [MB02c]. Correlating [Rom00]. Correspondence [Ano04n]. Corruption [BGS^{+}08]. Cost [Bro01c, CJL^{+}02, Sch02a, Uri06]. Cost-Effective [Uri06, CJL^{+}02]. Costs [MIK02, RCH00]. Could [BE01]. Countering [KFBK00, Oli01]. Counterpoint [DK02]. Counting [Kuh00]. Country [USE01a]. Coupons [GMAM01]. Course [Bru01]. Court [Tyl02d]. Covert [Bal03]. COW [KA01]. Cozens [Sal03d]. CPAN [Tur03b]. CPCMS [Sha02a]. CPU [CAGS00]. CPUs [BA01]. CQUAL [ZEJ02]. Cracked [Nic01a]. Crackproof [Sal03d]. Crap [Lim09]. Crash [Bru01]. Craystack [Bur00a, Bur01b, Bur01c, Bur01d, Bur01e, Bur01f, Bur01a]. Create [Kri00a]. Creating [Joh02, Wal07]. Creation [Cha01f, Eng00]. Creator [Coc01]. Credential [KHDC01]. Crimeware [Dit09b]. Crisis [LeF02, SvM05]. Cross [JR02, SZ02, Tay01, APH00]. Cross-Administrative [SZ02]. Cross-Architectural [JR02]. cross-platform [APH00]. Cross-toolkit [Tay01]. Crucial [And06]. crypt [Per03]. Crypto [Bur01g, Gut02, Ros01]. Cryptographic [CCSP01, Gut00, Ito01, MRL^{+}02, Sha02a]. Cryptography [Bla02, Dre00, Mur01]. CSP [SMF^{+}01]. Cubes [Bur06a]. Cumulus [VSV09]. Curves [Kle00]. Customer [Mil05, Rus00b]. Customers [Pen04]. Cyberconflict [Dit09a]. Cybercrime [App05]. Cybersecurity [Ano04g]. Cyberspace [Ken02b, Gee03a]. Cyclone [JMG^{+}02]. Cyrus [Far04a]. CZAG [Ren02].

D [CB00]. D4 [See01]. DAFS [MAF^{+}02, Mag02]. Dammit [Ioa02]. Danger [Cha01f]. Dangers [Jos09d]. DARPA [Coc01]. Data [APGG00, BGS^{+}08, BP01, Cou02, Drz00, FKV00, GP01b, GBHC00, Gut01, HBC^{+}01, Hum02, Ken03b, LR01, LAW02, MFJPPMK04, Nic02a, PMN^{+}04, QD02, SSCG09, Sel00b, SGS^{+}00, Ver07, VK02, YC01, kc01, SGM^{+}00]. Data-Directed [Sel00b]. Data-Intensive [APGG00]. Data-Structure [GP01b]. Data/Clusters/Resilient [Hum02]. Database [App04a, Fin00a, Kal09, LGJ00, Ley01, MVS00, NNW00, Tur04b]. DataMonster [Uri06]. Date [Tur05a]. Dave [Pur02, Sel08]. David [Sch02d]. Days [Hum05]. DBI [Tur04b]. DC [USE01h]. DDoS [Has05d, Kan05a]. Dead [Hoh01, MI07]. Dealing [BH00a, Loz05b, Per01]. Deals [Nic03a]. Deanonymizing [MS02]. Death [Gal08f]. Debugger [RB01, Wal00]. Debugging [Bea01c, BE08d, CD01b, MK07]. December [USE00g, USE01a]. Decentralized [KC00]. Declarations [McC01a]. Decompilers [Die00]. Decoupled [YMC02]. Defacement [Kol05d]. Defeating [Ale05a, SMJ00]. Defective [Dav01]. Defending [Pro01]. Defense [Ald05, BST00, Die02, Man05]. Defensive [Tur05b]. Defining [Ken02c, Rot01]. Degrading [Gol00]. Déjà [DP00]. Delays [SF00]. Delegating [Lim06]. Deletion [BP01]. Demystified [CWD02, TSW08]. Demystifying [Ark05]. Denial [DLD00, MVS01, Oli01]. Denial-of-Service [MVS01, Oli01]. Denver [USE00i]. Dependability [Pat02b]. Dependable [WRA00]. Depends [Ken02c]. Deploying [Gut02]. Deployme [OM00]. Deployment [Mur01, OM00]. Depot [RS01]. Design [Ano04h, Ano05l, Ano05e, BTS^{+}00, CCSP01, DLGD01, Dus08, DLS^{+}01, EKNG^{+}05, Har02c, KL01, Mag02, Mar01b, MCB^{+}01, Mew01, PG01, Pac01a, PGCR01, Rob00, SA02, Sur01, USE00h, WPSO01, WOSW02, YV00]. Designators [McC01b]. Designing [Drz00, LGJ00, MPC^{+}08, Sel01, Soo08b, Syv01, ZM00]. Desktop [BS00, Fou00, KHF01, MMBE00]. Detecting [GJM02, KFBK00, LE01, Mud00a, STFW01, ZP00a, ZP00b]. Detection [Bru00b, CD01a, Fly01a, GP01b, GHH06, HPK01, sLC02, LV02, NNW00, wCfJRW05]. Detects [SEZ^{+}01]. Deterministic [SW01]. dev [Fer06a, Fer06b, Fer06c, Fer06d, Fer07a, Fer07b, Fer07c, Fer07d, Fer07e, Fer07f, Fer08a, Fer08b, Fer08c, Fer08d, Fer08e, Fer08f, Fer09f, Fer09a, Fer09b, Fer09c, Fer09d, Fer09e, Kam02]. Developers [Ano02b, Ano05c, Far01h]. Developing [DDB08, HC00, OESV04, Sto02]. Development [BV00, CB00, ESW01, Mah04, APH00, Sch02b]. Developments [Has04c]. Deviating [Cha07c]. Deviation [Cou08b]. Device [SG01]. Devices [BCH^{+}00, GSGN00, Gut01, Kam02, MRL^{+}02, Mud00a, Mud00b, PMN^{+}04, Sha02b, Sta00, SVL01, War01]. Devil [Bla01b, MRC^{+}00]. DHCP [Has07a]. Diagnosis [Ban00]. Diagnostic [BE01]. Dialect [JMG^{+}02]. Diary [Sto06c, Sto08b]. Did [Far02f]. Diego [USE00h, USE00f, USE00a, USE01a]. Dies [Coc01]. Difference [GLV^{+}09]. Differencing [VK02]. Differentiation [VTFM01]. Digital [Ken01, Kuh00, Oki07, SD05]. Dima [Far00i]. Direct [SSB00, Mag02, MAF^{+}02]. Directed [Mar04, Sel00b]. Directions [Chi00, Yar08]. Directors [Ano03o, Ano04r, Ano05r, MY04, Mul04b, Mul05a, Mul05b, MY06, Veg07, You03, YM05, You06c, You07a, You07e, You08a, You09a, You09b]. Directory [Phi01]. Disaster [Ama00, CJL^{+}02, PMF^{+}02, SvM05]. Disclosure [Ano04g, App04a]. Disclosures [Mar00]. Disconnected [SGZ^{+}02]. Discovery [Ark05]. Discuss [SvM05]. Discussion [Ano04s, Dic09, Slu07]. Disk [ASS02, ASV^{+}02, Cha06i, LSG^{+}00, LSG02, Ruw00, Sah08, SGLG02, SH02, Siv06, SGMV08, WRPK02, YGC^{+}00, ZYKW02]. Disk/Persistent [WRPK02]. Disk/Persistent-RAM [WRPK02]. Disks [JHZK08, McK06]. Dispatch [CM01a, DLS^{+}01, Won01, ZD02]. Display [YNST02]. Dissemination [NDI04]. Distance [Far06g, Kne01]. Distant [Moo02]. Distributed [ABD^{+}08, Ano01f, Ano06c, CB00, DDB08, DLD00, FS00, FC01, FKM00, GBHC00, KYH^{+}01, LBN01, MW05, RBPM00, RI00a, SVF01, SLR02, SGZ^{+}02, Tie01, TS04, And01]. Distributing [Gra01c]. Distribution [ASDZ00, Hem00, KC00, OP01, RV00]. Distributions [BC01]. DITools [SNC00]. Divide [Ran05]. DMCA [Sam02]. DNS [Hob05, Man02, MCD^{+}02, Nem00b]. DNS-based [Hob05]. DNSBL [Lan08]. Do [Bel02, Dit09b, GW07, HS00a, JW01a, Lim01, Swa02b]. DocBook [Mas00]. Doctors [Tyl01, Tyl02b]. Document [LA04]. Documentation [Luc02]. Documents [MB02a, Wip00, dLWZ00]. Does [Cou08a]. Doesn’t [Lie00, JW00e]. Dog [Mos03a]. Doing [Sch06]. Domain [HK00a, SZ02]. Domains [Org07a, Org07b, Org08a, Org08b]. Done [Kle09]. Don’t [Ioa02, JHZK08, JW00c, Sim03, WE06]. Don’ts [FSSF01]. Dot [Lie00, Pik00a]. Dot-Com [Lie00]. Dot-Dot [Pik00a]. Doubled [Tyl02c]. Douglas [Ano04a]. Down [BE06e, BE06f]. Downloading [TK01]. Downs [Chu03]. Draw [BE07b]. Drive [SGLG02]. Driven [CML00, HS01, HcC01]. Driver [SG01, TK01]. Drivers [HC00]. Drives [Cha06i, LSG^{+}00]. Driving [Bro08]. Dual [AFE00, Tho00a]. Dumps [SD05]. Dynamic [ACG01, Ame01, APGG00, Ber00c, BR00a, CO02, GLF^{+}00, JN00, JCKS04, KYH^{+}01, sLC02, MB01, Moo00, Moo01, SNC00, Won01, ZD02]. Dynamically [RI04].

E-Mail [KS00]. E-Voting [Rub01c]. Eager [ZYKW02]. Eager-Writing [ZYKW02]. Early [Too09]. Easy [Cou09, Thr01a]. Eat [Mos03a]. Eazel [Adl00]. EC2 [Gar07, Wal08]. Echo [Cha06b]. Economy [Cym06]. Ed [Chu04b, Ano04b, Chu04a]. eden [Bro01b]. Edge [BCD^{+}00, BJH^{+}00, Che00b, Dee00, LR04, Nes00, Orm00b, Sch00c, Tom00, USE00j, Vic00]. Edge-Server [LR04]. Editing [MM01]. Editor [Ano05q, Ano05o, Ano05p, Sto00b, Ano01m, Ano00k, Ano00l, Ano00m, Ano00n, Ano01l, Ano03l, Ano04j, Ano04l, Ano04k, Dol06, Far04e, Pre06]. Editors [Ano02i, Ano01n, Ano01o, Ano01p]. Educational [Bro01a]. eEMU [Vol00]. EESR [Ano05n]. Effect [CK01]. Effective [Bro01c, Hal00b, Hal00c, Hal00a, Jos09a, Uri06, CJL^{+}02]. Effects [NW04]. Efficient [ASS01, GH01, HP01, MCD^{+}02, TS04, SGMV08]. Effort [SS02]. Election [Ano00i, Ano04r, Ano05r, You03, You07a, You09a]. Elections [Rub01c]. Electricity [vD09]. Electronic [GMAM01, JJR02, Nic00a, Rub01b, Eck01]. Elements [Sea00]. Ellis [Coc01]. Email [App04b, Has06b, JM03, Jos09b, Rus02, SEZ^{+}01]. Embedded [Bea01c, BR00a, CO02, Far00h, Gra01a, Has05b, Hei05, HPS02, LB02, Mar01a, McG00, Mil02, SMK02]. Embedding [Bea01a, PSW^{+}01, Sal03d]. Empirical [FM00, San01, SYN02]. Employed [Fer09f]. Employees [Cha01c, Rus00a, Rus04a, Rus04b]. Emulation [GSS^{+}02]. Enabled [BNB01, Smi01]. Enabling [MB02a]. Encodings [HEB01]. Encrypt [Dav01]. Encrypting [Pro00]. Encryption [BU02, Dan01, KS00, Siv06]. End [Ano05n, BM00a, CDGN01, FNKR01, FCTV02, HPK01, HK00b, Per03]. End-to-End [Ano05n, BM00a, CDGN01, FCTV02, HPK01, HK00b]. Enemy [Ran05]. Energy [CV02, SGMV08, VKK^{+}01]. Energy-aware [CV02]. Energy-efficient [SGMV08]. Enforce [Bah00]. Enforcement [HK00a, Has06c, Per01]. Engine [GLV^{+}09]. Engineering [HEB01, Llo07, Luc00, VK02]. Enhance [LP02]. Enhancements [BGM00, CMP00]. Enhancing [Ken03a, McC00g, SZ02]. Entanglement [MB02b]. Enterprise [Cha06i, Iam01, KC00, MPC^{+}08, Ott00, She00, Sor06, Wal06, LR04]. Entries [Rom00]. Envelope [Gee05]. Environment [AK01, III01, Fou00, Gal09b, Guy01, JADAD06, KC00, Las01, LM00, NSK04, Pri01, Siv01, War01]. Environments [CKV^{+}02, CO02, HBC^{+}01, LS00, Sac08, Ano05m]. Epilogue [Bur01a]. Equal [Spi06]. EROS [SA02]. Error [Bea01c, Tur05c]. Errors [CMB^{+}01]. Essential [BE09c]. Establish [Ano06e]. Estimating [Apt01]. etc [Sim03]. EtE [FCTV02]. Ethernet [Fly02a, OF00]. Ethical [Sin05a]. Ethics [Par00]. Eugene [Chu02a]. EuroBSDCon [Ano05g]. European [Ano03c]. Evaluating [Coo04a, RKS02, ZS01]. Evaluation [JGKvS04, MCD^{+}02, OK01, YV00]. Evangelizing [Coc01]. Evasion [HPK01]. Event [CM01a, Lem01, Vol00, YD00]. Event-Dispatch [CM01a]. Events [Moo02, OP01, San01]. Every [JHZK08]. Evidence [Ken03a]. Evident [WRC00]. Evolution [BJP00, Bro08, Fiu05, Rot01, SA02]. Evolving [Ken02b, PF04]. EVT [Wal07]. Examiner [Bar00]. Examining [McC03b]. Example [OSE^{+}00, Skv03]. Exception [She00, dD00]. Exceptions [SDN^{+}08]. Excessive [Ezo01]. Exchange [Dit01]. Exclusive [WW02]. Executable [MP01b]. Executables [SEZ^{+}01]. Executing [PBP02]. Execution [Ano05m, GH01, HcC01, KBA02, SW01]. Exedra [ASS01]. expect [Fly03e]. Expectant [Cou00]. Experience [BLF^{+}00, Chi00, III01]. Experiences [ABC^{+}02, Lam00, SS02, USE00f].

Experiment [BLF^{+}00]. Experimental [JGKvS04]. Exploiting [BBADAD02, wRF07, MoCA^{+}02]. Exploits [KCK^{+}00]. Exploration [SMM04]. Exploring [JN04, LCC00]. Exports [NNW00]. Exposed [Ver07]. EXT2 [Phi01, TT02]. Ext2/Ext3 [TT02]. Ext3 [MCD07, TT02]. Ext4 [MCD07]. Extended [FKV00]. Extending [BE00, BA01, CL01, Ger02, JCKS04, MMBE00, Sal03d]. Extensibility [SMRF01]. Extensible [BNB01, DA02, Rig02, WRA00]. Extension [SNC00, WT00, GP01a, Pac01a]. Extensions [Bea01c, ECCH00, Gle02, HC00, TT02]. Extents [SGLG02]. External [Sel01]. Extracting [LSG^{+}00, Sha04]. Extreme [Hén06].

Fabric [WOSW02]. Facilitated [FS01]. Facility [Gra01a, Lem01, McK04a, Thr01b]. Facing [Dit09a, LeF02]. Factories [Tur03e]. Failover [SAB01b]. Failsafe [Phi00]. Failure [HBG^{+}07, JHZK08, LCC00, PBB^{+}00]. Failure-Resilient [HBG^{+}07]. Fair [CAGS00]. Fairy [Dar05]. FAQ [LaB00]. Far [Dit09b]. Farms [Thr01b]. FAST [Ano02f, Ano04e, Ano06a, USE02c, BDTW01, Cor00, Dic01, FKM00, McK07, Orm00a, ZABN01]. Faster [KT01]. Fate [Sin05b]. Fault [CL00]. Faults [Ezo01]. FCIP [Sch02c]. Fear [Abl08]. Feather [Duf00]. Features [McC01e, McC02c, McC02d, Wat01]. featuring [And01]. February [USE00e, USE01e, USE02a]. Federal [App04b]. Feng [Bag04]. Fermi [HR01b]. Fever [Sto07c]. Few [Bea01a, Mal05, NH00]. FFT [CB00]. Fi [MPC^{+}08]. Field [Ban00, Dit01]. Fighting [Chu04a, Has04c]. File [Ano02f, Ano03e, Ano04e, Ano06a, BP01, Bes00, BRS01, CLRT00, CCSP01, CBWKH01, DM07, DDB08, FKM00, FKM02, Hel09, Ito01, JFWS00, KCN02, Kno00, KL01, Mag02, MAF^{+}02, MCD07, Maz01, McC05b, McK06, McK07, MoCA^{+}02, Mor00, PP06, PMF^{+}02, Pik00a, PBB^{+}00, RJ01, Rom00, RLA00, SMRF01, SH02, SGM^{+}00, Spi03, SHS01, USE02c, Uri06, WRPK02, WH02, YMC02, ZN00, ZABN01]. File-I [BRS01]. File-I/O [BRS01]. File-System-Based [PMF^{+}02]. Files [BE06b, Fin00a, Gra01b, Loz05b, Sah08]. Filesystem [BFH02, BC06, DM02, Phi00, TT02, VSV09]. Filter [Har02c, SEZ^{+}01, SEZ^{+}01, vR01b]. Filtering [HK03, LLB02, Wal06, vR01b, Cha03a]. Finals [Kol04f]. Financial [Ano03o, Ano05u, You04b, You06b, You07d, You09d]. Find [Swa03b]. Finding [Ale05b, HS00a, Loz05a]. Fine [Rus02, Slu05, SAB01b]. Fine-Grained [SAB01b]. Fingerprinting [Ark02a, Ras04, SMJ00]. Fingerprints [Dam03]. Firewall [AT05, Bon01, Lim01, MK07, Pre05, Woo01b, Woo02]. Firewalling [GB01]. Firewalls [Dar05, Pis08, Sin03, Sin00b, Zwi01]. Firewalls/VPN [Zwi01]. Firmware [LSG02]. FIRST [Ano01i, Ano03b, Ano04h, Ano05m, JW00b, USE00f]. fish [Hum00a]. Fisherman [USE00j]. FiST [ZN00]. Five [Sal03e]. Fixing [Dzo09, Sch02e, Tur03c]. FLAME [Ano04a]. Flash [Sit03b]. Flat [DM00]. Flexibility [Sul00, SS00]. Flexible [BSF02, CM01b, JCKS04, LLB02, LS01b, McC01b, PM00, PBP02, SNC00]. Flight [Ver07]. Flood [Lem02, Mar04, Oli01]. Flood-Routing [Mar04]. Floods [Sto00a]. Flow [Gol06, Plo00, RFL00]. Flow-tools [RFL00]. Flows [Jos07b]. FlowScan [Plo00]. Fly [CD01a]. Focus [Pat02a]. FOKSTRAUT [BH00a]. Font [Pac01b, TK01]. Food [Mos03a]. Forensics [Ken02a, NNW00, Pei05, Pow01, SD05, Chu04b]. Foreseer [CW04]. Formal [Mit02, MP01b]. Format [CBB^{+}01, STFW01, VK02]. FormatGuard [CBB^{+}01]. Formats [CV02]. Formatting [FKV00, Tur05a]. Forwarding [OW00, YC01]. Fourteenth [USE00g]. Fourth [USE00h]. Fragmentation [BG01]. Framework [FSLM02, GH01, Mar04, Mar01b, PG01, RKS02, San01, SS00, vBBB^{+}02, APH00]. Francisco [USE00j, USE01c, USE02a, USE02f, USE02b]. Free [BG01, Coc01, DH00, HPS02, LSG^{+}00]. Freeblock [LSG02]. FreeBSD [Bal02, Coc01, DM07, DM02, Hop02, Leh01, Los02, Luc02, Luc05, Mag02, Mal05, McK04a, Mur02, Nik02, Sch02a, Siv03, Siv06, WS02, Wat01]. FreeBSD-4.x [Siv03]. FreeBSD-Based [Sch02a]. Freedom [Cha01f, Fel02, FF07]. FREENIX [USE01f, USE02d]. Frequencies [RCH00]. Friendly [OMP02]. Frill [Bal00]. Frill-Necked [Bal00]. Front [Slu07]. fsck [McK02]. FTP [Coh02, NH00]. FUD [Sal04g]. Full [BC06]. Full-Text [BC06]. Fun [Dar03c, Ege04, Hop02, Loz05a, Pur06]. Function [APGG00]. Functions [McC02b, Won01]. Fund [Ano06e]. Funded [Ber00a]. Funds [Coc01]. Future [Bro08, BS04, Cha03b, Chi00, Gal08a, Gee07, Get02a, Gol00, Kri00a, Kry02, Sel08, Yar08]. Fwknop [Ras06, Ras04]. fwsnort [Ras07].

Gadgets [Smi01]. Gaining [WG00]. Game [MV04]. Games [wRF07]. Ganymed [PA04]. Gap [DADAD02]. Garbage [BR00a, CKV^{+}02, DKCJ02, FDSZ01, Pri01, SVF01, ZS01]. Gary [Kar02]. Gates [Har02a]. Gateway [Han03]. Gather [Kne01]. Gauze [Die00]. GCC [MS00]. GCTP [Sif00]. GDBTk [Ing00]. Gecko [HDM00]. Geek [Sto06c, Sto08b]. Geeks [Bag04, Kne01]. Gemini [Wal00]. Genealogy [HSW09]. General [BSF02, BB02, Get02b, Par02b, USE02e, WCS^{+}02].

General/Random [Get02b]. Generalised [Moo00]. Generalize [Won01]. Generated [MRL^{+}02]. Generating [Fin00a, Fly02a]. Generation [GMAM01, Jen04, Kan06a, MCD07]. Generational [DKCJ02, Pri01]. Generators [Swa01]. Generic [DLGD01, Lem01, LCC00, MB00]. Genome [Krz01]. Geoff [Kol00b]. Geographic [SPK02]. Geometry [Bab01]. Georgia [USE00c]. Get [Bur00a, Bur01b, Bur01c, Bur01d, Bur01e, Bur01f, Bur01a, Cha02c, Coc01, JW01b, Jos06b, WD01]. Getting [Gee03b, Pik00a, Sal05f, Slu05]. Gigabit [OF00]. Give [BE06c]. Glass [Bal03]. Glass-Box [Bal03]. Glimpse [Gra01b]. Global [MHM^{+}00]. Glossary [Fer09f]. Glues [LM00]. GNU [Coc01, Gol06, Zad02]. Go [Ano04a, BE08d, Get02a]. Goals [MMS00]. Goat [JW01b]. Goes [Lon09b]. Going [Rus03]. Gold [Kol05e]. Gone [Dit09b, Ros01]. Good [App07b, Bea09, Cha01d, Slu07, Swa03b]. Gossip [JGKvS04]. Gossip-Based [JGKvS04]. Got [Lon09a]. Governance [McK04b]. GPFS [SH02]. Gracie [Cha06g]. Grade [JBH^{+}00]. Graham [Has04b]. Grained [SAB01b]. Grammar [Bur06c, KV02]. Graphic [Ham00]. Graphical [lWS00]. Graphics [KD02]. Gray [BBADAD02]. Gray-Box [BBADAD02]. Green [Cou09]. Greener [Lon09b]. Grid [Gar07, Ano01d]. Grooming [JW01c]. Group [Gwi00, Jos01b, Jos02, Moe02, Sif00, Tyl03b, Jos00a, Jos00c, Jos00d, Jos01b, Jos01a]. Grouping [Ame01]. Groups [Dre00, Wyn00]. Growing [Wyn00]. Grunts [Cha02d]. Gscope [GW02]. GTK [Bla01a]. GUADEC [Ano01g]. Guarantee [RR00]. Guarantees [LAW02]. Guesting [SDN^{+}08]. Guide [Gal09e, Pac01b, Sch02c, Tie01]. Guiding [Coo04b]. GVSAGE [Tyl02a].

Habits [Jos09a]. Hack [Ale05d]. Hackers [Mas00, Chu02b]. Hacking [Dar03c, Gro09, Sch02e, Soo08a]. Haiku [Kol04h]. Haikus [Rea04]. Halprin [Kol00b]. Handbook [Pur02]. Handheld [Guy01, MMBE00, Sta00]. Handler [DD07]. Handler/Agent [DD07]. Handling [Ano01i, Tur05c, dD00]. Happy [Rus00a, Rus04a, Rus04b]. Hard [Sim03]. Hardening [Mao09, Mue07a]. Hardware [AE01, FS01, HWF07, Has05b, MRC^{+}00]. Hash [Lev00, VB00]. HBench [ZS01]. Head [Joh02, LSG^{+}00]. Headers [Kri00b]. Health [Sch01]. Heartbeat [Rob00]. Heimdal [WD01]. Hello [Cha06g, McC03b]. Help [Bru09b]. Hénon [Kol04a]. Here [JW01c, ML06, SDHD07]. Heterogeneity [CL01]. Heterogeneous [FADAD02, Hug00, KC00, NSZ01, PMN^{+}04, Uri06]. Hey [Lim09]. Hi [BE08d]. Hi-ho [BE08d]. Hiding [Bru04]. High [Ash02, FNKR01, HR01a, Hob02, HcC01, JKN^{+}01, KCK^{+}00, LHAP06, Lum01, McC08, Min01, MPC^{+}08, NSH^{+}02, RB01, SV02b, Uri06, Wal08, YC01, Zwi01]. High-Assurance [SV02b]. High-Availability [Ash02]. High-Level [RB01]. High-Performance [HcC01, JKN^{+}01, KCK^{+}00, Lum01, Min01, MPC^{+}08, Wal08]. High-Speed [Uri06, YC01]. Higher [LSG^{+}00]. Highly [BH00b, JFWS00, Jos09a, LY01, ZPU08, ZSB04]. Highly-Available [ZSB04]. Hijacking [Jos06a]. Hill [USE02a]. Hippodrome [AHK^{+}02]. Historian [Her06]. Historical [Ros01]. History [MB02b, McK07, Sal03a, Sal04h, Sal04g, Sal04i]. Hitz [Sel08]. HMO [Tyl03b]. Hoarding [Ame01]. Hoc [SLR02, SDN^{+}08]. Hocracy [Bur06d]. Hogs [BM00b]. Hold [Jos08c]. Holiday [USE00j]. Home [Bec01, Bok01, Bro01b, Kol05e, SSCG09, Sin00b, Van09b, WRA00]. Home-Built [Van09b]. Homebrew [Ale04b]. Homeless [Jos06a]. Honeypot [GHH06, Spi02]. Honeypot-Based [GHH06]. Honeypots [Chu02b]. Hook [ZEJ02]. Hooks [Moo00]. Host [Bal03, Chu03, GB01, MB01, Nor00]. Host-Based [Chu03]. Hosted [SVL01]. Hostile [Nic00b]. Hosting [LY01]. HOSUS [Spi02]. Hot [Ano05h, Hos03, Pri01]. Hot-Swapping [Pri01]. Hotel [USE01f, USE01a, USE01b, USE02a]. HotOS [Ano05h, Ano03h]. HotOS-IX [Ano03h]. HotSpot^{T M} [PVC01, RB01]. Hourglass [Reg02]. HP [Hun00, KW00, MMS00, MVW00]. HP-UX [MMS00, MVW00]. HPC [Lum03]. HTTP [Kis00, Mal06, Lan08]. Hum [BE06d]. Human [Dre00, Rig02]. Hybrid [WRPK02]. Hype [Way02]. Hyper [Gal09f]. Hyper-V [Gal09f]. Hypermedia [KT01].

I/O [BL00, BRS01, Bur07, HcC01, LHAP06, McC02c, MoCA^{+}02, Ruw00, SVL01]. I/O-Driven [HcC01]. IA32 [ABC^{+}02]. IA64 [dD00]. ICAP [Tom00]. ICMP [Ark02a]. ICSI [Sch02c]. Identifiable [Dam03]. Identified [Fuh06]. Identifying [Fou02]. Identity [Fuh06, GK04, Oki07, TSW08]. IDL [MRC^{+}00]. IDS [Ras07]. IEC [LaB00]. IEEE [Ano01d, Ano05f]. IEEE/ACM [Ano01d]. IETF [Bel02]. If [BE01, Tyl01, Tyl02b, Tyl02c]. IFCP [Sch02c]. IFIP [Jac04]. II [Has03b, Rus04a]. III [Rus04b]. Ill [App07b]. Illusions [Koc02]. Images [DP00]. IMAP [Pen04]. Immersive [Coc01]. Implementation [Ano04h, Ano05l, Ano05e, CCSP01, Cor00, DLS^{+}01, Gle02, JR02, KL01, Leh01, Mag02, Mar01b, Mew01, Mur02, PG01, Pac01a, PGCR01, RI00a, Sko00, Sur01, USE00h, WS02, Wil02]. Implementations [JGKvS04, Jos09a, LGW04]. Implementing [Dic01, Gut02, Las01, NS01]. Implications [Sam02, San01]. Importance [Ale05c]. Important [HLU^{+}00]. Impractical [BE07a]. Improve [CMP00]. Improved [Fin00a, GB01]. Improvements [Lev00]. Improving [Ale04b, Cho00, HPS02, LS00, Leh01, LS01a, MSV09]. In-lining [SYN02]. Incident [Ano01i, Chu02a, Chu04b, For03, Jon01a, RCH00]. Incremental [MVW00]. Independent [BBLo04]. Independently [PF04]. Index [Phi01]. Indexing [ZABN01]. Industrial [USE00f]. Industry [Cam00, Nic03b, OSE^{+}00]. Ineffective [CD01b]. Inevitability [CFH^{+}05]. Inferring [MVS01, Reg02]. Influence [Sto08d]. Influencing [Bla01b]. InfoBeacons [Coo04b]. Informant [Ost00]. Informant/Killer [Ost00]. Informatics [Kol09a, Pie00]. Information [BS02, Coo04b, DADAD02, Fer09f, HRF00, KC00, MZ00, Mun02, NDI04, OSE^{+}00, Sch01, SMM04, Sin02, ZM00]. Informative [BH00b]. Infranet [FBH^{+}02]. Infrastructure [And09, DA02, ESW01, FSAK01, HS00b, Jon02, LGW04, PSVW01, Sel01, Tra01]. Infrastructures [Kno08]. Initiative [Coc01, WG00]. Inline [Tur04c]. Inn [USE00j]. Innovation [Sam02]. Inodes [BE07e]. Ins [Car02]. Insecurities [Soo08a, Soo08b]. Insecurity [Bla02, Cam00]. Inserted [BM00b]. Insider [Mud03]. Install [Lim09]. Installation [Ano04d, Ano05b, Ano06b, GR07, MB00, Sol08, USE00b]. Installations [AFE00]. Installed [RS01]. Instance [BCGD00]. Instead [SGMV08]. Institute [Col00]. Institutions [Bro01a]. Instruction [HEB01]. Instrumentation [BH00b, Drz00]. Insurance [Van01]. Integrated [AK01, LGJ00]. Integrating [Ing00, JBH^{+}00, LS01b, MM00]. Integration [CB00, Fin05, Ham01, Ito00, Lei00, MSO00, Tur04c]. Integrity [SV02b]. Intel [RI00b]. Intelligence [Nes00, USE00j]. Intelligent [Orm00b]. Intelligently [BM00b]. Intensive [APGG00]. Intentions [Cha01e]. Inter [MoCA^{+}02, Whi01]. Inter-File [MoCA^{+}02]. Inter-Process [Whi01]. Interaction [Gib00, lWS00]. Interactive [KD02, MM01]. Interface [BR00b, FBMZ04, Moo00, MMBE00, RI00a, Rig02, Ros00, Whi01]. Interfaces [McC04e, McC00g]. Intermediaries [YC01]. Internal [EKNG^{+}05]. International [Ano01d, Ano05i, Ano05j, Jac04, Kol04f, Kol09a]. Internet [Ano00d, Ano01j, Ano03f, Ano05k, Ker03, USE01c, ABC^{+}00, And05, BCK06, BCD^{+}00, Che00b, CBB00, Che02, DVWW01, FCTV02, GBHC00, GC01, Has07b, HRF00, JFWS00, MVS01, NSZ01, PH01, RJ01, Rub01b, SMF^{+}01, SPW02, SPK02, kc01]. Interoperability [Hil00, PF04]. Interposed [ACV00]. Interpreter [SMK02]. Interprocess [HPS02]. Interview [Far00i, wRF07, Far07g, Fen00, Has04b, Has05c, Kol00b, Kol00c, Sel08]. Interviewing [Cla00]. Interviews [Cha00a]. Intranet [Che02]. Introducing [CDRV00, Gil01, Ker03, Llo07, Mor00]. Introduction [Han03, Hay06a, Ley01, McC03c, Org07a, Org07b, Org08a, Org08b, Pat02b, RMS02, Tur04a]. Introductions [Bla00a]. Intrusion [Fly01a, GHH06, HPK01, LV02, NNW00]. Intrusions [KFBK00]. Intrusive [Wal00]. Investigating [BTS08]. Investment [Swa03b]. Invited [Jos09f, Tar02]. Involved [Jos06b]. iOverlay [LGW04]. IP [Ark02b, Cha05b, Cha06a, Fly03c, KNRS00, Pis08, Sen07, SMJ00, Sto00a, vR01b]. Ip-filter [vR01b]. IPSans [Sch02c]. IPSEC [SAB^{+}01a, Ioa02, ZS00]. IPSECvalidate [SAB^{+}01a]. iptables [Ras07]. IPV6 [GB01, AB04, And05, BCK06, Cha08b, Dee00, KK08, Siv03]. ISA [Los02]. Island [JFWS00]. Island-Based [JFWS00]. Isn’t [Swa03a]. ISO [LaB00]. ISO/IEC [LaB00]. ISO9660 [Kno00]. Isolation [BHL00, PN05, Sul00, SS00]. ISP [Has03c].

ISPAdmin [Has01a, Has01b, Has01c, Has01d, Has01e, Has01f, Has02a, Has02b, Has02c, Has02d, HK03, Has03c, Has03a, Has03b, Has04a, Has04b, Has04c, Has05a, Has05b, Has05c, Has05d, Has06a, Has06b, Has06c, Has06d, Has06e, Has07a]. Issue [Far00a, Far01a, Far02a, Far03a, See01, TSW08, Ano00j]. Issues [Bea01a, BH00a, Hil00, Kim01, Lei00, Sin05a, Yar08]. ITA [FBMZ04]. Itch [BE09e, BE09f]. iVoyeur [Jos07b, Jos07c, Jos07a, Jos08a, Jos08b, Jos08c, Jos08d, Jos08e, Jos08f, Jos09d, Jos09a, Jos09b, Jos09c, Jos09e, Jos09f]. IX [Ano03h].

J [Ano04b]. J2EE [JCKS04]. Jack [Sho01, Tyl02d]. Jack-of-All-Trades [Sho01]. Jail [McK04a]. Jails [Hop02]. Jalapeño [ACG01, NS01]. January [USE01e, USE02c]. Japan [SS02]. JaRec [Chr01]. JaRTS [Gle02]. Java [Ano00g, Ano01e, Ano02d, USE01g, USE02f, Gos00, BHL00, BTS^{+}00, Bou01, Chr01, CD01a, DLS^{+}01, FC01, GH01, Gle02, JR02, Mah04, McC00a, McC00b, McC00c, McC00d, McC00e, McC00f, McC01c, McC01d, Mid01, MP01b, NS01, PVC01, Ran02, Rao00a, Rao00b, Rao00c, Rao00d, Rao00e, Rao00f, Rao01a, Rao01b, Ros00, SMK02, SW01, SYN02, Sur01, TS04, VKK^{+}01, WS01, ZD02]. JavaBeans [LR04, WCD^{+}01]. Java^{T M} [SMES01, Dic01, RB01, vD00]. Jay [You09c]. Jenness [Sal03d]. JFS [Bes00]. JGC [ZS01]. Jikes [ABC^{+}02]. Jim [Coc01, Joh02]. JIT [Mar01b]. JNI [NS01]. Jobs [Dar04c, JW01d]. John [Aik07, Ano06e, Kar02]. Journaled [Bes00]. Journaling [SGM^{+}00]. Joy [Par02a]. Jr. [Sch02d]. Judgment [JW02]. JumpStart [Ama00]. June [USE00a, USE01f, USE01b, USE02d, USE02e, USE06]. Just [JW00d, Sha02b, SYN02]. Just-in-Time [Sha02b, SYN02]. JVM [Ano00g, Ano01e, USE01g, USE02f, And01, CMB^{+}01, DA02, ZS01]. JX [GFWK02].

Kaffemik [And01]. KaffeOS [BHL00]. Kasetsart [US01]. Kava [WS01]. KDE [Gra00a, Whi01]. Keeping [Rus00a, Rus04a, Rus04b, Slu06, Tur03d]. Kerberized [KHDC01]. Kerberos [Coc01, Hil00, Ito00, WD01]. Kernel [Ano02b, Ano05c, Bal02, BGM00, BH00b, CBWKH01, CMP00, Far01h, Fiu05, Jon01b, JKN^{+}01, Kam02, Kle09, Lev00, Mag02, Min00, Moo00, Sac08, VTFM01, WCS^{+}02, YD00, YHS^{+}02, Dik00]. Kernel-Level [YD00]. Kernelized [Per02]. Kernels [BHHV00, FSLM02, Mil02]. Kevin [Chu04b, Far00i]. Key [BDTW01]. Keynote [Ber02]. Keys [Lea04, MRL^{+}02]. Keystrokes [SWT01]. Keyword [CW04]. Kick [JW01e]. Kiddies [Mue07b]. Killer [Ost00]. Kiss [Bur06b]. KLAT2 [DM00]. Knit [RFS^{+}00]. Knocking [Ras04]. Know [Lim01, Lim09, Har02b]. Know-It-All [Har02b]. Knowing [Can00]. Knowledge [BBADAD02]. Known [Mal05, SF07]. KParts [Gra00a]. Kqueue [Lem01]. KSOS [Per02].

L [Sch02d]. Lab [AB04, KV02, HR01b]. Laboratory [Pre07]. Lack [Bur06d]. LAM [OF00]. LAM/MPI [OF00]. Lance [Chu02b]. Language [Bea01c, Dus08, Dzo09, GS02, Gos00, Ham01, MP01a, NZ00, Vol00, WCD^{+}01, ZN00, WCD^{+}01]. Languages [Bro08, Won01]. Laptop [Gar02]. Laptop/PDA [Gar02]. Large [AS00, Ano01h, Ano04d, Ano05b, Ano06b, Ano06c, BH00a, BL00, HR01b, HDM00, Hum02, Kuh00, LSB^{+}09, PP06, SH02, USE00b]. Large-File [PP06]. Large-Scale [Ano01h, BH00a, LSB^{+}09]. Larger [CWSC01]. Larry [Chu04a]. Latency [HR01a, VG00]. Law [App04b, Dan02, Per01]. Laws [App05, Nic00c, SF07]. Layer [SM02, ZS00].

LCFG [AS00]. LDAP [Car01, Qui07]. Leak [Ren02]. Learn [BE01]. Learned [Gut02, Qui07]. Learning [AHL02, Wys03]. Legacy [DEH^{+}09, LM00]. Legal [Ken02c, Mue07b, Sin05a]. Length [McC01g]. Lepreau [You09c]. Lesser [Mal05, SF07]. Lesser-Known [Mal05, SF07]. Lessons [CWL^{+}01, Gut02, Qui07, Sha04]. Let [BE07b]. Letter [Ano04j, Ano04k, Ano05o, Ano05p, Dol06, Pre06, The04]. Letters [Ano00k, Ano00l, Ano00m, Ano00n, Ano01l, Ano01m, Ano01n, Ano01o, Ano01p, Ano02i, Ano03l, Ano04l, Ano05q]. Level [CL01, DL01, ESW01, Jos09c, Maz01, OMP02, PSVW01, RB01, SMRF01, SA02, YD00, ZPL01, Ale05a, CIK00, SNC00]. Levels [ASV^{+}02]. Leverage [Kol05c]. Leveraging [DEH^{+}09]. Lexical [Pik00a]. Liberty [Dav02]. Library [BR00b, GS02, LB02, McC02b, Pac01b, Tho00a, Woo01a, Gsc01]. Licensing [Nic02b]. Life [Sin03, Str02]. Lifetime [Coc01]. Lightweight [Fou00, LGW04, Ran02, SSB00]. like [Mar01a]. Likely [LE01]. Limit [JW00f]. Limits [LCC00]. Line [LBN01, Ing00]. Lines [CWL^{+}01]. lining [SYN02]. Link [LB01]. Linking [FC01, MVW00]. Links [Sim03]. Linux [Ano00e, Ano02b, Ano05c, Min00, USE00c, USE01d, AT05, Adl00, AFE00, AB04, ABC^{+}02, AS00, BP01, Bea01a, BGM00, BHK^{+}00, BR00b, BH00b, BHHV00, BRS01, BFH02, BC06, CLRT00, CM01a, Chu03, III01, Dik00, Dik01, EBC01, Far01h, FNKR01, Gar02, GS06, HK00a, Ham00, Hel09, HR01a, HD02, Iam01, JZ01, Kno00, KHF01, Lev00, LH02, LS01b, Lum01, Mar01a, McG00, McQ00, MHW00, Min01, Moe02, MH01, Moo00, Moo01, Mor00, Mor01, OK01, OF00, PG01, PGCR01, PSW^{+}01, PBB^{+}00, Rob00, Ruw00, Sac08, SK02, Sha00, SG01, Sko00, Sor06, TM01, Tho00b, Thr01b, TT02, WG00, WCS^{+}02, vR01a].

Linux-Based [PG01, WG00]. Linux-HA [Rob00]. Linux/IA32 [ABC^{+}02]. LinuxThreads [DL01]. Lions [Ano06e]. LISA [Ano00a, Ano01c, Ano02g, Ano03i, Ano04d, Ano05b, Ano06b, USE00g, USE01a, Kam04]. List [Cur00, Dic09]. Lite [Pow01, Wal00]. Litigation [Mue07a]. Little [BE08a, Eck01, Swa03a]. Live [BE08f, Fra01, GS06]. Lived [Jos06a]. Lives [Ver07]. Living [Qui07]. Liz [Tyl02d]. Lizard [Bal00]. Load [KPv00]. Loadable [And00, Jon01b]. Loading [HP01]. Loaning [Bla01b]. Loathing [Abl08]. LOBOS [Min00]. Local [Dij00e, MCD^{+}02, Tyl02a, Wyn00]. Locality [CW04]. Locality-Aware [CW04]. Locally [RS01]. Location [HS00b, RAMC^{+}04]. Location-Aware [HS00b]. Locked [Har02a]. Locking [Bal02]. Lockmeter [BH00b]. Locks [BH00b, Dic01]. Log [All02b, Bes00, Gla02, Rom00, WH02]. Log-Structured [WH02]. Logging [Bes00, BE00, CMP00, Gra01a, YD00]. Logical [Org07a, Org07b, Org08a, Org08b]. login [FL06, Sal04k, Sal05c]. Logs [JHP^{+}09, RFL00]. LOMAC [Fra01]. Loneliness [Kne01]. Long [Far06g, Kne01, Hum00a, Ros01]. Long-Distance [Far06g, Kne01]. Longitudinal [KA01]. Look [BGS^{+}08, Mar01a]. Looks [Gee07]. Looping [KCK^{+}00]. Lot [Mar01a]. Louisiana [USE00g]. Low [HR01a, KCK^{+}00, Sch02a, Sha02b]. Low-Cost [Sch02a]. Low-Overhead [KCK^{+}00]. Low-Power [Sha02b]. LRU [MM03]. LumenVox [Cha07a]. Lumeta [Woo01b].

m [USE01g]. M. [Ano04a]. MA [USE06]. MAC [Fra01, Ger02, SV02a]. Machine [Ano00g, Ano01e, Ano02d, Ano04f, AE01, JADAD06, Law00, Pro00, RI00b, SVL01, USE01g, USE02f, AHL02, DLS^{+}01, JR02, MP01b, Ran02, RB01, SMK02, SMES01, Sur01, vD00]. Machines [Ale05b, Bur00a, Bur01b, Bur01c, Bur01d, Bur01e, Bur01f, Bur01a, LHAP06]. Made [BE07e]. Maelstrom [CD01b]. Magazines [BA01]. Magic [Rus00b]. MAGNeT [HcFG01]. Magnetic [Cou02, Kry02]. Mail [Has00, KS00, SEZ^{+}01]. Mailing [Cur00]. Maintain [Lim09]. Maintainability [Pat02a]. Maintaining [Joh02]. Maintenance [Jos02, MB00]. MAJIC [LBN01]. Make [KK08]. Makeover [Hén06]. Making [Bou01, JJR02, McC04b, WW02]. Malicious [Chu04a, KM01, SEZ^{+}01, SEZ^{+}01]. Malloc [Ezo01]. Mallocs [BG01]. Malveau [Ano04b]. Malware [Ale05b, Chu04a, DD07, Dit09b]. Manage [BM00b]. Manageable [LY01]. Management [AHT^{+}02, AG01, All02a, ASS01, ABC^{+}00, And06, AHL02, BHL00, BR00a, Bro01a, Bur06b, Bur06c, Bur06d, Bur07, BBADAD02, CV02, Che00a, Cou08a, DG09, DBH06, Fuh06, GK04, GSGN00, Hal00d, Hug00, JN00, JW00b, JCKS04, Kan06a, KC00, KCK^{+}00, KYH^{+}01, LGJ00, MHM^{+}00, OM00, Ott00, Par02b, RJ01, RMS02, RS01, SSCG09, SPC08, San01, SMES01, SV02b, Sha02a, She00, SYC^{+}01, SGHS01, SvM05, SS00, Thr01a, VG00, Vol00, vR01a]. Management/General [Par02b]. Manager [BNB01, Cur00, MMS00, Rot01]. Managers [Bab01, SC09, TM01]. Managing [ABD^{+}08, BC01, Bru01, Has03c, MB01, OESV04]. Mandia [Chu04b]. Manipulated [GJM02]. Manual [AHT^{+}02]. Manuals [Dzo09]. Many [BA01, Mid01]. MAPbox [AR00]. Mapping [Bye02b, Bye02a, CBB00, Che02, MK07]. March [USE00j, USE01c]. Mark [III01, Far07g]. Mark&Compact [Pri01]. Mark&Sweep [Pri01]. Market [All02c, All02d]. Marketing [Jos09d, SS05]. Markets [LBN01]. Markup [WCD^{+}01]. Marriott [USE01f, USE01b]. Mars [Per01, SMM04]. Masquerade [SMMM08]. Massachusetts [USE01f, USE01b]. Masses [But04, Gal09a]. Master [Sho01]. Matching [Ras07, SGLG02]. Math [McC02b]. Matsumoto [Sch02d]. Matter [MI07]. Maturity [Cou00]. Maui [BHK^{+}00]. Maximizing [ASS02, Bro00]. May [Jos02, USE06, Jos02]. Maybe [How06b]. Maze [Eck01]. McCarty [Joh02]. McGraw [Kar02]. McIlroy [Ano04a]. Me [BE06e, BE06f, BE06c, BE07b]. Mean [Cou08a]. Measurement [CWVL01, kc01]. Measurements [Ano05i]. Measuring [LB01, Lam00, LR01, RE01, Siv06, YD00, YNN01]. Mechanism [Bea01c, OMP02, PBP02]. Mechanisms [CM01a, McC08, VTFM01, WPSO01, YNST02]. Media [ASS01, CWVL01, Cha08a]. Meeting [Ano00p, Ano03o, BY00a, BY00b, MMS00, Ano03c]. Meetings [Ano04r, Mul04b, MY06, Veg07, YM05, You06c, You07e, You08a]. Meets [BE07f, MB02c, Spi00, Way02]. MEF [SEZ^{+}01]. Meghdoot [GSAA04]. Members [McC01b, McC03d]. Membership [Ano03p, Mul04a, MD05]. Memoir [Ale05d]. Memoriam [Aik07]. Memory [AHL02, BR00a, BM00b, CMB^{+}01, FDSZ01, JKN^{+}01, Mid01, Mil02, MHM^{+}00, PW00, RI00a, SVF01, SMES01, Sha02b, SDS09, Spi06, SD05, VKK^{+}01, vR01a]. Memory-Based [JKN^{+}01]. Memory-Constrained [Sha02b]. MEMS [GSGN00]. MEMS-based [GSGN00]. Merging [RS01]. Merry [BE08d]. Merry-o [BE08d]. Message [Gee00b, Has00, Jos09b]. Messages [App04b]. Messenger [How05]. Meta [SGM^{+}00, WPSO01]. Meta-data [SGM^{+}00]. Meta-Programming [WPSO01]. Metadata [LSB^{+}09]. Metering [Has00]. Method [BDTW01, SYN02]. Methodologies [DBH06]. Methodology [CCG^{+}00]. Methods [Mit02, Mud00a]. Michele [Joh02]. Microkernels [Hei05]. Microprocessor [Ran02]. Microprocessors [ML06]. Microsoft [Gal09f, Sit03a, dLWZ00]. Middleground [WG00]. Middleware [Jac04, LGW04, MFJPPMK04, MV04, PMN^{+}04, RBPM00, RAMC^{+}04, RI04, SMM04, WPSO01, WG00, ZSB04, Jac04]. MiddleWhere [RAMC^{+}04]. Might [Mud00b]. Migration [BCR07, Bru09b, LAW02, PN05, SAB01b]. Milk [OS06]. Mills [Far00i]. MIME [Dav01]. Minders [Bur09]. Minimal [HS00b]. MINIX [HBG^{+}06]. Minor [Ezo01]. Minute [Lon09a]. Miracle [JW00b]. Mirror [NH00]. Mirroring [PMF^{+}02]. Miscellaneous [Ano00o]. Misses [YHS^{+}02]. Missing [Loz05b]. Mission [SMM04]. Mitigating [Has05d]. Mix [JJR02]. Mobile [ABQ02, Ame01, Ano05f, Ano05j, Bou01, Fly04e, SGZ^{+}02, War01, dLWZ01]. Mobility [MB01]. MobiSys [Ano03j, Ano05j]. mod gzip [Pie03]. Mode [AB04, Min00, Tur05b, Dik00, Dik01]. Model [BS00, BBLo04, Coo04a, Gra01c, Hay06b, MP01b, Swa02a, YV00]. Modeling [Ano05i, KHF01, Tur04b]. Models [Bur06b, Bur06c, Bur06d, Bur07, GB04, MCB^{+}01, Mid01]. Modem [SW00]. Modern [Ale04c]. Modifications [Ale04b]. Modular [ASS01, DA02, HBG^{+}06]. Module [Jos08e, Jos08f]. Modules [Jon01b, Tur03b, Tur03a, Tur03c, WCS^{+}02]. Molecular [KHF01]. Mona [SMRF01]. Monitor [Fly03e, RI00b, Sel02, She00, SVL01]. Monitoring [Ark05, Ash02, Bru00b, CWSC01, FCTV02, GC00, Jos09c, Min01, PM00, Sel00a, Skv03, Vol00]. Monitors [Dic01, Gla02]. Monoculture [Gee05].

Monterey

[USE01g, USE02c, USE02d, USE02e]. MOSS [Dav01]. Most [US01]. Mostly [BBD^{+}01]. motd [Dar04b, Kol00d, Kol00e, Kol00f, Kol00g, Kol00h, Kol00i, Kol00j, Kol01a, Kol01b, Kol01c, Kol01d, Kol01e, Kol01f, Kol01g, Kol02a, Kol02b, Kol02c, Kol02d, Kol02e, Kol02f, Kol03a, Kol03b, Kol03c, Kol03d, Kol03e, Kol04b, Kol04c, Kol04d, Kol04e, Kol05c, Kol05a, Kol05b]. Motion [YNN01]. Mouth [Ken03b]. Move [KK08]. Moves [Dan02]. Moving [FS00]. Mowbray [Ano04b]. Mozilla [APH00]. MPI [OF00]. MPICH [OF00]. Mr. [Har02b]. Mudge [Kol00c]. Multi [Chr01, DLS^{+}01, ML06, MoCA^{+}02, MHM^{+}00, RBPM00, RV00, VB00, lWS00, ZS00, ZPL01]. Multi-Collective [MoCA^{+}02]. Multi-Computer [MHM^{+}00]. Multi-Core [ML06]. Multi-Dispatch [DLS^{+}01]. Multi-Layer [ZS00]. Multi-Platform [RV00]. Multi-Queue [ZPL01]. Multi-Threaded [VB00, Chr01, RBPM00]. Multi-User [lWS00]. Multicast [PSVW01]. Multicasting [JGJ^{+}00]. Multicomputer [PW00]. Multicore [Bru09b]. Multics [Sha04]. Multilingual [iiH01]. Multimedia [CV02, Dic08]. Multimillion [DAK08]. Multimillion-node [DAK08]. Multiple [DVWW01, FC01, GB01, McC08, MM01]. Multiprocessor [CM01b]. Multiprocessors [CAGS00, FDSZ01]. Multithreaded [Bal02]. Multithreading [BPL07, Eng00]. MULTOPS [GP01b]. Murky [TSW08]. Music [Bur06a]. Musings [Far00h, Far00b, Far00c, Far00d, Far00e, Far00f, Far00g, Far01b, Far01c, Far01d, Far01e, Far01f, Far01g, Far02f, Far02b, Far02c, Far02d, Far02e, Far03b, Far03c, Far03d, Far03e, Far03f, Far04e, Far04f, Far04g, Far04h, Far04i, Far04j, Far05b, Far05c, Far05d, Far05e, Far05f, Far05g, Far06a, Far06b, Far06c, Far06d, Far06e, Far06f, Far07a, Far07b, Far07c, Far07d, Far07e, Far07f, Far08a, Far08b, Far08c, Far08d, Far08e, Far08f, Far09a, Far09b, Far09c, Far09d, Far09e, Far09f]. MVICH [OF00]. My [BE06c, Cha03a, Far02f, JW01b, WW02]. Myriad [CJL^{+}02]. Mystical [Jos07b]. Myths [Bur06b, Bur06c, Bur06d, Bur07, kc01]. MythTV [Bro06].

Nagios [Jos09a]. Name [DA03, Dij00f, Lie00]. Names [Coc01, Pik00a, RJ01, Sha02a]. NAS [Lev05, Sel08]. NASA [SMM04]. National [Nic00c, Gee03a]. Nationwide [App04b]. Native [Mor01]. Nautilus [Adl00]. ND [MNH00]. Neal [Far04b]. Nearby [Jos07a]. Nearly [Sal03c]. NEB [Jos08e, Jos08f]. Necessity [Slu07]. Necked [Bal00]. Need [Coc01, Hei05]. Needles [Bur00a, Bur01b, Bur01c, Bur01d, Bur01e, Bur01f, Bur01a]. Needs [Ran05]. Neighborhood [DM00]. Neptune [SYC^{+}01]. Nessus [DH00, DG03]. Net [Nic00c]. NetApp [Sel08]. NetBSD [Mew01, Rig02, Wil02, vdL02]. Netfilter [Ayu06]. NetFlow [NNW00, RFL00]. Nets [JJR02]. Nettimer [LB01]. Network [All02a, ACV00, AC08, Ark05, ASDZ00, Bah00, Bru00b, BS02, BJH^{+}00, CV02, Dam03, Dee00, DH00, Die02, DM00, ESW01, FSAK01, FKM02, GC00, Gil01, HPK01, Has07b, HcFG01, HP01, JGJ^{+}00, Jon02, Jos09c, Ken02c, KC00, Mee00, MCZ06, MLFR02, MIK02, MNH00, Moo02, Mor00, Nes00, NNW00, OF00, PM00, Plo00, Ran00a, Ran00b, Ran00c, Ran00d, Ran00e, Ran00f, RS01, Sch02e, Sel00a, SYC^{+}01, Sto00a, Tom00, USE00j, Van01, Vic00, WOSW02, YC01, vBBB^{+}02, Coc01, CD01b]. Network-Attached [MLFR02]. Network-Sharable [RS01]. Networked [Ano04h, Ano05e, Bar00]. Networking [JBH^{+}00, Orm00a, Sal03a, WRA00].

Networks [Ark02b, Che02, DDB08, GB04, GSAA04, KNRS00, MW05, Mar04, MB01, MPC^{+}08, PGCR01, SLR02, Tie01, Sch02c]. Newest [Lie00]. News [Ano00o, Ano00q, Ano00r, Ano00u, Ano00v, Ano00t, Ano01q, Ano01r, Ano01s, Ano01t, Ano01u, Ano01v, Ano01w, Ano01x, Ano01y, Ano01z, Ano01-27, Ano01-28, Ano02j, Ano02k, Ano02l, Ano02m, Ano02n, Ano02p, Ano02q, Ano02r, Ano02s, Ano02t, Ano02u, Ano03m, Ano03n, Ano03o, Ano03p, Coc01, Kol05d, Kol06, Mul04a, Sim00, Sim01]. Next [And05, Kan06a, MCD07, Sit03a]. Next-Generation [Kan06a]. NFS [LH02, SZ02]. Nickle [MP01a]. Ningaui [HD02]. Ninja [vBBB^{+}02]. Ninth [USE00i]. NIX [Sal03c]. Nmap [BE07f]. No [Ber02, Bur07, Can00, Sch04, Sin01, Dar00g]. node [DAK08]. Nominating [Ano03o, Ano03p, Ano05s, Ano06d, JG08, You07b, Ano04m, Nem00a]. Nominations [Com00]. Non [Cou02, Has06b, Sit03b, Wal00]. Non-Email [Has06b]. Non-Intrusive [Wal00]. Non-Magnetic [Cou02]. Non-Proprietary [Sit03b]. Nonblocking [HH01]. None [Sho01]. NOOSE [Bar00]. Normalization [HPK01]. NoScript [Mao09]. Note [Mar00, Sch00b]. Notes [Ano04r, Ano04o, Ano04p, Ano04q, Ano05t, Ano05r, Ano05u, Ano05s, Ano06e, Ano06d, Cha06g, Dol06, Far05h, Kol05f, Kol05e, MD05, Mul05b, MY06, Pal05, Par05, Sal05f, Sal05g, Sal05d, Sal05h, Sal05i, Sal05c, Sal05e, YM05, You05a, You05b]. Notice [Ano00p]. Notification [Lem01]. Novel [CW04, SMF^{+}01, WH02]. November [Ano04g, USE01d]. NSDI [Ano04h, Ano05e]. NT [USE00b, AFE00, BJP00, Car00b, FM00, Luc00, Mor00, Nor00, Ott00, Sta00, Wal00]. NT/2000 [USE00b]. NT5 [Lei00]. NTP [Kno08, Van09b]. Nugache [SDHD07]. Number [Swa01]. Numeric [McC01e].

Nutshell [Sch02d].

o [BE08d, BL00, BRS01, Bur07, LHAP06, McC02c, MoCA^{+}02, Ruw00, SVL01]. O-Driven [HcC01]. Oakland [USE01d]. OASIS [Fre06]. Object [Ano00f, Bar00, Gra01c, HBC^{+}01, LBN01, NZ00, Tur03e, USE01e, WPSO01, Won01]. Object-Oriented [Ano00f, Bar00, NZ00, USE01e, Won01]. Objectbase [BJP00]. Obscurity [Mal05]. Observing [Moo02]. Obtaining [NSH^{+}02]. Occupational [Dij00d]. October [Ano04d, Jac04, See01, USE00c, USE00h, USE00f]. Ode [Hal00d]. Odyssey [Luc00]. Off [Cha01b]. Offensive [Man05]. Office [dLWZ00]. Offices [DVWW01]. Offline [Jen04]. Often [How06a]. Oh [Sch04]. Olympiad [Kol09a, Pie00, Kol04g, Kol05f]. On-Line [LBN01]. On-the-Fly [CD01a]. One [MM03]. Online [wRF07, LAW02, NW04, Zwi05]. Only [FKM00, Zwi05]. Ontology [OESV04, WJL04]. Ontology-Based [OESV04, WJL04]. Opaque [Jos07c]. Open [All02c, All02d, Bro01a, III01, Far06g, Gut00, Hoh01, Jos00d, Jos01b, Kim01, Nes00, SS02, Sif00, Sor06, Str02, The04, WG00, Woo01a]. Open-Source [Bro01a, Gut00, Sif00, WG00, Woo01a]. OpenBSD [AT05, Har02c, OS06]. OpenFlock [Sif00]. OpenJIT [Mar01b]. OpenSolaris [Hay06b]. Operated [CKV^{+}02]. Operating [Ale04c, Ano05h, Ano05l, Ano06e, BTS^{+}00, Ban00, DM07, FSLM02, GFWK02, Gra00d, GSGN00, HBG^{+}07, KM01, LS01b, MB00, Mil02, Moo01, Per02, USE00h, Wat01, Wil02, lWS00, iiH01]. Operational [OSE^{+}00]. Operations [KNRS00, LS01a, SW01]. Operators [App04c, McC04a]. OPPA [App04c]. Opportunities [Kim01, Tar02, dLWZ00]. Oprtions [Cha03a]. Optimisations [DM02]. Optimizations [JN04]. Optimizing [MCZ06, OKN02, SMK02]. Options [HDM08]. Oracle [Hos03, Lev05]. Ordered [San01]. Organizational [Rot01]. Organized [Spi03]. Organizer [SMES01]. Oriented [Ano00f, Bar00, NZ00, USE01e, Won01]. Origin [BS02]. Orleans [USE00g]. ORs [JW01a]. OS-Based [Bon01]. OS/2 [Luc00]. OSDI [Ano05l, USE00h]. OSes [MI07]. OSU [RFL00]. Other [Cha02d, Moo01, Pik00b]. Our [You04a, You05a, You06a, You07c, You08b]. Out-of-Process [RB01]. Output [Pie03]. Outs [Car02]. Outsourcing [DVWW01, NSH^{+}02, Nic03a, Nic03b]. Outwit [Spi00]. Overcast [JGJ^{+}00]. Overflow [Ale05a, BGM00, LE01, sLC02]. Overflows [Ale04a]. Overhauling [Zad02]. Overhead [KCK^{+}00, OKN02]. Overlay [JGJ^{+}00, LGW04, Sto00a]. Overload [PLT00]. Overloaded [McC04a, VTFM01]. Own [Mos03a, SPW02].

P.O. [BE08f]. p0f [BE07f]. P2P

[DD07, GSAA04, SDHD07]. PA [Cor00]. PA-RISC [Cor00]. Package [OM00, RFL00, SPC08, SC09]. Packet [Har02c, Jos09c, LLB02, Mee00, Ras06, vR01b]. Packet-Level [Jos09c]. Packets [BC00, Fly02a, Fly03c]. Page [Dic08, JZ01, vR01a]. Pages [Jen04]. Palm [Cap01, KM01, SMES01]. Pandora [PM00]. Panel [BCD^{+}00, BL00, BJH^{+}00]. Panopticon [YN01]. Parable [All02d]. Paradise [USE00f]. Parallel [CLRT00, FDSZ01]. Parallelism [CM01b, McC08, SDS09]. Parallels [Dus08]. Parameterized [AR00]. Parked [Tyl02c]. Parse [Jos08d]. Parsimonious [CBKH^{+}00]. Part [BE06e, BE06f, BE09e, BE09f, Bur00c, Bur00d, Bur01b, Bur01c, Bur01d, Bur01e, Bur01f, Org07a, Org07b, Org08b, Sto08b, Tyl01, Tyl02b, App07a, App07b, Bur06b, Bur06c, Bur06d, Bur07, Rus04a, Rus04b]. Partial [JJR02, KT01]. Partially [San01]. Parts [Krz01]. Passive [Ark05, FCTV02, Ras04]. Password [Ale04c, BE08e, How06a, Hug00, KP01, Sin05a]. Password-Based [KP01]. Password-Protected [BE08e]. Passwords [Per03, Sin01]. Patch [Bur06b, Gee03c, RV00, Rus03]. Patent [Col00]. Path [Orm00a, SSB00]. Paths [YS02]. PATS [Gal08a, Gal08b, Gal08c, Gal09b]. Patterns [DLGD01, Dus08, MoCA^{+}02, SGLG02, Tur05c, Fly00e]. Paul [Has04b]. Pavol [Sal03d]. Pawlowsky [Sel08]. Payload [YC01]. PC [CM01b]. PCI [Los02]. PCT [BB02]. PDA [Gar02]. PDE [MCB^{+}01]. PDFs [BE07c]. PDM [KP01]. Peck [BE07c]. Peep [GC00, Gil01]. Peer [CW04, Coo04a, PRS06, JGKvS04]. Peer-to-Peer [CW04, Coo04a, PRS06]. Peikari [Far04a]. Pelendur [CFK00]. PEM [Dav01]. Penalty [CK01]. Penetration [GS06]. PenguinOMeter [BRS01]. Pentium [RI00b]. People [JW01d, Mor02, Mun02]. Per-User [Jos09c]. Perceived [RE01]. Performance [AT05, All02a, BH00a, Bro00, BFH02, CK01, CML00, FCTV02, Gra00b, HLU^{+}00, Har02c, HcC01, Hun00, JR02, JN04, JKN^{+}01, KCK^{+}00, KCN02, LP02, LH02, LHAP06, LAW02, Lum01, MAF^{+}02, MMS00, McC00a, McC00b, McC00c, McC00d, McC00e, McC00f, McC01c, McC01d, McC08, Mey09, Min01, Mor01, MPC^{+}08, NSH^{+}02, OK01, OF00, Pat02a, RI00a, Ruw00, Siv06, Wal08, WPSO01, WRPK02, WH02, YNN01, YNST02, YGC^{+}00, ZS01]. Performance-Driven [CML00]. Performing [SW00]. Performs [Bes00]. Pergamum [SGMV08]. Perils [Woo02]. Perl [BE06b, BE06d, BE06a, BE06e, BE06f, BE06c, BE07d, BE07e, BE07a, BE07c, BE07b, BE07f, BE08f, BE08c, BE08a, BE08d, BE08b, BE08e, BE09a, BE09d, BE09b, BE09e, BE09f, BE09c, Hal00b, Hal00c, Hal00a, HAE00, She00, Sit02, Tur02a, Tur02b, Tur02c, Tur03b, Tur03a, Tur03c, Tur03d, Tur03e, Tur04b, Tur04c, Tur04a, Tur04d, Tur05a, Tur05b, Tur05c, Tur05d, Sal03d]. Permission [Jos08d]. Persistent [Bou01, Gsc01, SMES01, SG01]. Persistent-RAM [WRPK02]. Personal [Ale05d]. PersonalRAID [SGZ^{+}02]. Perspective [BC01, McK06, SSCG09, VKK^{+}01, WM00]. Perspectives [Sch00a]. Pervasive [HBC^{+}01]. Pete [Gal08a, Gal08b, Gal08c, Gal08d, Gal08e, Gal08f, Gal09a, Gal09e, Gal09b, Gal09d, Gal09f, Gal09c]. Peter [BE07c]. pf [Har02c]. PGP [BCH^{+}00, Dav01]. Phones [Cha06a]. Physical [BM00b]. Physics [HR01b]. Picked [BE07c]. Picking [Krz01]. Picture [BE07b, LA04]. PIKT [Ost00]. Pilots [Cap01]. Piper [BE07c]. Piranha [CMP00]. PIRUN [US01]. Pitfalls [Che00a, LGJ00]. Pithy [Ann00]. Pixels [Jos08c]. PKCS#7 [Dav01]. PKI [Che00a]. Place [BE08a, USE01f, USE01b]. Placement [APGG00, ZEJ02]. Plaintext [Sin01]. Plan [Swa02b, CGP^{+}02, Pik00a]. Planet [Nat04]. Planned [TT02]. Planners [Kri00a]. Planning [Swa02a]. Plateau [Ste01]. Platform [BBLo04, EBC01, Orm00b, PM00, RV00, APH00]. Platforms [JBH^{+}00]. Play [WD01]. Please [Cha03b, Fiu05, Per03]. Plenary [See01]. Plex86 [Law00]. Plumbing [Pik00b]. Plush [ABD^{+}08]. Pluto [Per01]. PMQS [FNKR01]. Point [Bye02b, Bye02a, DK02, USE00f, Cap01]. Point-to-Point [Cap01]. Point/Counterpoint [DK02]. Points [Fou02]. Police [Ran00a, Ran00b, Ran00c, Ran00d, Ran00e, Ran00f]. Policies [GLF^{+}00, LS01b]. Policy [Bah00, Bla01b, Has06c, MK07, Nic01b]. Politeness [Nic00d]. Politics [Cho05].

Poller [Sel00b]. Polymourphously [BE09d]. Pools [SW00]. Popular [CV02]. PORCMOLSULB [Owe02]. Port [OW00, Ras04, Woo02, Dik00]. Portability [JR02, vdL02]. Portable [Eng00, Gle02, SMK02, Sto02, TS04, BHK^{+}00]. Portal [SMM04]. Porting [ABC^{+}02, DM07, vdL02]. Position [The04]. Positioning [SLR02]. posium [USE01g]. POSIX [Jos02, Bla00b, Bla00c, Gwi00, Jos00b, Jos06b, MB02c]. Possibilities [LGJ00]. Possible [Gro09, Mur01]. PostScript [TK01]. Potential [Cou02]. Power [Adl00, Ama00, MV04, Sha02b, Van09a]. Powerful [US01]. PPPoE [Sko00]. Practical [BE06b, BE06d, BE06a, BE06e, BE06f, BE06c, BE07d, BE07e, BE07a, BE07c, BE07b, BE07f, BE08f, BE08c, BE08a, BE08d, BE08b, BE08e, BE09a, BE09d, BE09b, BE09e, BE09f, BE09c, Dre00, PRS06, Skv03, Tur02a, Tur02b, Tur02c, Tur03b, Tur03a, Tur03c, Tur03d, Tur03e, Tur04b, Tur04c, Tur04a, Tur04d, Tur05a, Tur05b, Tur05c, Tur05d, Vol00, War01]. Pragmatic [HH01]. Pragmatics [MP01a]. Prakash [Has05c]. Precise [MCD^{+}02]. Precision [OKN02]. Precompiled [Kri00b]. Predictable [AG01]. Predictive [KL01, ZPU08]. Prefetch [KT01]. Prefetching [JN00, KL01, YMC02]. Preprocessor [McC02d]. Presence [FC01, MB02c]. Present [Sel08]. Preservation [MB02b]. President [Dij00b, Dij00c, Gee00b, Gee00a]. Presuppose [JW00d]. Prevailed [Hoh01]. Priceless [Cym06]. Primer [App05]. Primitive [SW01]. Principles [Cou02, MP01a]. Printer [TK01]. printf [CBB^{+}01]. Printing [WM00]. Privacy [Cho05, Eck01, Gol00, Nic00a, Nic02a, NW04, RR00, Ren02]. Privacy-Degrading [Gol00]. Private [Cur00, MB01, Tyl03b]. Privileges [Gra01c]. PRNG [Mur02]. PRO [KA01]. PRO-COW [KA01]. Proactive [CL00]. Probabilistic [Apt01, Kuh00]. Probable [Gro09]. Probes [Moo00]. Problem [Gee03b, JHP^{+}09, Owe02, Rus02, Ost00]. Problems [Ale04a, Ban00, Cou02, Los02, Nic03a]. Procedures [CD01b, For03, Jos02]. Proceedings [USE00c, USE00g, USE00h, USE00f, USE00b, USE00i, USE00j, USE00e, USE00d, USE01d, USE01f, USE01g, USE01a, USE01h, USE01b, USE01e, USE01c, USE02a, USE02c, USE02d, USE02e, USE02f, USE02b, USE06, Jac04]. Process [Ber00c, BS02, RB01, SV02a, TSW08, Whi01]. Processes [BHL00, GB01, JADAD06]. Processing [All02b, BGSS00, HcC01, ZYKW02]. Processor [CML00, MDH00]. Processors [OKN02]. Product [Gra01a]. Production [Las01, Lum03]. Profession [BE01, Bur05, Wyn00]. Profiling [BB02]. Profit [Dar03c, Ege04, Hop02, Loz05a, Pur06]. Program [Dit01, KBA02, McC03b, Ano00b]. Programmable [RI04]. Programmatic [Jen04]. Programmer [ECCH00]. Programmer-Written [ECCH00]. Programmers [Coc01]. Programming [And09, BE06d, CDRV00, DLGD01, GS02, Hal00b, Hal00c, Hal00a, HBG^{+}06, Kal09, McC01e, MRC^{+}00, PW00, Sea00, SYC^{+}01, Spi00, Tur05b, WPSO01]. Programs [Chr01, CD01a, DL01, LM00, Tur05d]. Project [All02c, All02d, BLF^{+}00, Jos00b, Luc02, Sto08c, Adl00, Coc01, McQ00, iiH01]. Projects [Ber00a]. Promiscuous [Mud00a]. Promises [Far07g]. Prompt [Cou07]. Proof [FC01, Kle09]. Propagation [BCK06, GB04, ZSB04]. Properties [McC03d, SPK02]. Proportional [CAGS00, NVZ01]. Proportional-Share [CAGS00]. Proposal [Slu05]. Proprietary [Sit03b]. Prosise [Chu04b]. Protect [DS01, JZ01, SC09]. Protected [BE08e].

Protecting [III01, HSW09, SGS^{+}00]. Protection [Ale04c, Ale05a, CBB^{+}01, CBWKH01, FS01, Mil02, SGM^{+}00, SHS01]. Protocol [And05, DADAD02, ESW01, HPK01, Her06, KP01, KA01, Mas01, OMP02, Soo08a, ZS00, Cap01, Cha08a, Han03, Ker03]. Protocols [Joh02, KS00, Lev05, MIK02, Sch02c, Tay01]. ProtoWrap [Wol02]. Prove [Hei07]. Provenance [HSW09]. Provider [Has03a, Has03b, Has07b]. Providing [BS02, Has03c, HRF00, LLB02]. Provisioning [BJH^{+}00, Vic00]. Proxies [SGHS01]. Proximity [MCD^{+}02]. Proxy [Ros00]. Pruning [Spi03]. psad [Ras07]. Pseudonymity [RR00]. Psst [Van01]. PSTL [Gsc01]. Public [BDTW01, MB01, Nik02, Rub01c, SV02b]. Public-Access [SV02b]. Publications [Jos00d]. Publish [EG01, GSAA04, OP01, PRS06, WJL04, ZSB04]. Publish-Subscribe [PRS06]. Publish/Subscribe [EG01, GSAA04, WJL04, ZSB04]. Publishing [WRC00]. Publius [WRC00]. Puppet [Kan06a]. Puppeteer [dLWZ01]. Purpose [Lif01]. Put [Ken03b]. Putting [Wys04]. Puzzles [DS01]. PVFS [CLRT00]. Python [Bea09, Bru09a, HAE00, How06b, Sit02].

QA [Wys03]. qmail [Pur02]. QoS [BJH^{+}00, CDRV00, Orm00a, Vic00]. Qt [TK01]. Qt/X11 [TK01]. QTcl [CDRV00]. Qualifiers [STFW01]. Quant [Gee07]. Queries [Coo04b]. Quest [Guy01]. Questions [Bea01a]. Queue [ZPL01]. Quick [Ley01]. Quicksilver [Far04b]. Quiet [Slu07].

R2 [Gal09f]. Race [CD01a, CBWKH01]. RaceGuard [CBWKH01]. RAID [ASV^{+}02, BP00, CL01]. RAM [WRPK02]. Random [FM00, Get02b, Swa01, Fer06a, Fer06b, Fer06c, Fer06d, Fer07a, Fer07b, Fer07c, Fer07d, Fer07e, Fer07f, Fer08a, Fer08b, Fer08c, Fer08d, Fer08e, Fer08f, Fer09f, Fer09a, Fer09b, Fer09c, Fer09d, Fer09e]. Randomised [San01]. Randomized [JJR02]. RandR [GP01a]. Raphael [Ano04b]. Rapid [BV00, HR01a]. Rate [ASS02]. rc.d [Mew01]. Reaction [HR01a]. Read [FKM00]. Read-Only [FKM00]. Reading [CWL^{+}01]. Ready [Dar00g, Pis08]. Real [Ano06c, AE01, Cha03b, Ger02, Gle02, Gos00, HH01, HPS02, LR01, Mor01, vR01b]. Real-Time [Ger02, Gle02, HH01, HPS02, Mor01]. Real-World [LR01]. Realist [Fer09f]. Reality [Coc01]. Really [Dit00, RR00]. Realtime [Gwi00]. Reasonable [Sch06]. Reasons [Sel02]. Rebutting [The04]. Recalcitrant [Ing00]. received [Jos02]. Receiver [BGSS00]. Recipe [Moe02]. Recipients [Coc01]. Reclaim [Sah08]. Reconfigurable [RI04]. Record [Chr01]. Record/Replay [Chr01]. Recorder [Ver07]. Recording [Kry02]. Recovery [Ama00, Bea01c, CL00, LCC00, PMF^{+}02, PBB^{+}00]. Reducing [Ano05k, YHS^{+}02]. Reference [Gra01c]. References [KCK^{+}00]. Refinement [DKCJ02]. Reflection [EG01, McC04d, WS01]. Reflections [Cha01a, Cha02a, WE04]. Regional [Sin02]. Regions [DDB08, MM01]. Reinvented [Lum03]. Rejection [Hob05]. Related [Ale04a, GB01]. Relational [NNW00]. Relative [MIK02]. Relaxed [Dic01]. Relaxed-Locks [Dic01]. Releases [BM00b]. Reliability [Apt01, Lam00]. Reliable [JGJ^{+}00, LY01]. Relieving [MD00]. Remanence [Gut01]. Remembered [DKCJ02]. Remote [Ark02a, DVWW01, GJM02, Rub01b, Skv03, Sta00, lWS00, YNST02]. Rendering [LB00, Pac01a]. Reordering [WH02]. Repeatable [Bru00a]. Replacement [JZ01, SW00, ZPL01, vR01a]. Replacing [Jos09b]. Replay [Chr01]. Replicated [ASS02, SVF01, YV00]. Replication [MFJPPMK04, PA04, SYC^{+}01]. Report [Ano00h, Ano01i, Ano01g, Ano01d, Ano01e, Ano01h, Ano01c, Ano01j, Ano01k, Ano01b, Ano01f, Ano02e, Ano02f, Ano02d, Ano02g, Ano02b, Ano02h, Ano02c, Ano02a, Ano03b, Ano03g, Ano03c, Ano03h, Ano03i, Ano03j, Ano03d, Ano03f, Ano03e, Ano03k, Ano03a, Ano03o, Ano04g, Ano04h, Ano04d, Ano04e, Ano04i, Ano04c, Ano04f, Ano04m, Ano05g, Ano05h, Ano05f, Ano05i, Ano05b, Ano05c, Ano05j, Ano05l, Ano05e, Ano05k, Ano05d, Ano05a, Ano05m, Ano05n, Ano05u, Ano06a, Ano06b, Ano06c, Ano06d, Bla00c, Col00, Far01h, JG08, Jos00c, Jos00d, Jos01a, Jos02, Jos03, Kol00a, Kol08, Nem00a, Sal01b, Sim01, You04b, You06b, You07d, You09d]. Reporting [Plo00, Slu07]. Reports [Ano02o]. Request [ACV00, ASDZ00, Hem00, WPSO01]. Request-based [Hem00]. Require [Cou07]. Required [Bro06]. Requirements [Mil05]. Requires [App04a]. Research [Ano00g, Ano01e, Ano02d, Ano04g, Ano04f, Bur00b, Bur00c, Bur00d, Dit01, GH01, Sam02, Tar02, USE01g, USE02f, Tar02]. Researchers [EVS^{+}08, Far00i, Wys03]. Resilient [HBG^{+}07, Hum02]. Resistant [WRC00]. Resisting [Lem02]. Resize [GP01a]. Resolution [Sch00c, See01]. Resort [USE00f, USE01a]. Resource [ASS01, BHL00, Bru09a, Cha08a, CIK00, DG09, JCKS04, KYH^{+}01, MW05, MRL^{+}02, NSK04, Sac08, SS00, Sur01]. Resource-Constrained [CIK00, MRL^{+}02]. Resources [AG01, BA01, JW00c]. Responding [Dit09a]. Response [Ano05n, Chu04b, For03, Jon01a, RE01, SF00, Chu02a]. Responsibilities [Dar03b, Vix02]. Restoration [Too09].

Restructuring [McK04b]. Results [Ano00i, Ano04r, Bro06, San01, Way02]. Resumé [Rus00c]. Rethinking [Kam02, Mey09]. Retire [Swa02b]. Retirement [Swa02a]. Retires [Kol04a]. Reverse [HEB01]. Reverse-Engineering [HEB01]. Reversing [YN01]. Review [Ano01a, Ano04b, Chu02a, Chu02b, Chu04a, Chu04b, Far04a, Far04b, Fly00e, Har02a, Joh02, Kar02, Mal05, Nat04, Pur02, Sal00e, Sal02a, Sal04c, Sal04d, Sal04e, Sal04f, Sal04b, Sch02b, Sch02c, Sch02d]. Reviews [Ano01a, Far04c, Far04d, FH05, Far05a, Fri05a, Fri05b, FKS08, Has03a, Has03b, Sal00f, Sal00g, Sal00h, Sal00i, Sal00j, Sal00k, Sal01c, Sal01d, Sal01e, Sal01f, Sal01g, Sal01h, S^{+}01, Sal01i, Sal02b, Sal02c, Sal02d, Sal02e, Sal02f, Sal03b, Sal03d, Sal03f, Sal03e, Sal03c, Sal03a, Sal04j, Sal04h, Sal04g, Sal04i, Sal05b, Sal05a, ZT05, Zwi05, Z^{+}06a, ZSF06, Z^{+}06b, Z^{+}06c, Z^{+}06d, Z^{+}06e, Z^{+}07a, Z^{+}07b, Z^{+}07c, Z^{+}07d, Z^{+}07e, Z^{+}07f, Z^{+}08a, Z^{+}08b, Z^{+}08c, Z^{+}08d, Z^{+}08e, Z^{+}09a, Z^{+}09b, Z^{+}09c, Z^{+}09d, Z^{+}09e, Z^{+}09f]. revised [Jos02]. Revising [TSW08]. Revision [Jos00b, Sto07c]. Revisited [Cha03a, Jos08f]. Revisiting [FADAD02]. Reviving [Hoh01]. Revocation [BDTW01, Hop01]. Révolution [Cha02c]. Rewriting [WS01]. Reynolds [Sch02d]. RFCs [Bel02]. RHIC [Thr01b]. Rich [LLB02]. Right [Gee03b, Pik00a]. Rights [Dre00]. Rik [You05b]. RISC [Cor00]. Risk [Hal00d]. Risky [Swa03a]. Road [Cou07]. Roadmap [HBG^{+}07]. Rob [You05b]. Robin [NVZ01]. Robots [Coc01]. Robust [JJR02, MSV09, SLR02, WRC00]. Robustness [FM00]. Roland [SvM05]. Role [Ker06, Rot01]. Roles [Vix02]. Root [Cou07, Jos08a]. Rooted [Siv01]. Rootkits [BS04]. Rotate [GP01a]. Round [NVZ01]. Round-Robin [NVZ01]. Roundup [Has06a]. Router [CM01b, Sch02a]. Routing [Abl08, ACV00, GC01, JBH^{+}00, Mar04, Sor06, SPK02]. Rovers [SMM04]. RPM [RS01]. RSIP [BM00a]. RSO [Sch04]. RT [Gwi00]. Ruby [Sch02d, Kan06b]. Rule [SDN^{+}08]. Rules [ECCH00]. Run [BST00, BS01]. Run-Time [BST00, BS01]. Running [AHK^{+}02, McK02]. Runtime [Ham01]. Russell [Chu02a]. RVM [ABC^{+}02].

S [Dav01]. S/MIME [Dav01]. S3 [Gar07]. SableVM [GH01]. Safe [JMG^{+}02]. Safety [KCN02]. SafeWeb [MS02]. SAGE [Ano00q, Ano00r, Ano00t, Ano01q, Ano01r, Ano01s, Ano01t, Ano02j, Ano02k, Ano02l, Ano02m, Ano02n, Ano03m, Ber00b, Cha06h, Cha06f, Dij00b, Dij00c, Kol05d, Lon07, LC07, Lon08, McK04b, Pal05, Par05, Sch00a]. SAGE-News [Kol05d]. SAGU [Bro01a]. Salamander [Guy01]. Salary [Sch00a]. Salesman [Tyl03a]. Salesmen [Jos09f]. Sam [Cha02b]. SAMBA [Car02, BH00a, Car00a, Lei00]. Sampling [JGKvS04]. San [USE00h, USE00f, USE00j, USE00a, USE01a, USE01e, USE01c, USE02a, USE02f, USE02b, Lev05, Sel08]. SAN/NAS [Sel08]. Sandboxing [CIK00, PS01]. SANE [Ano02h]. Sanitize [For01]. Sanity [Rot01]. Savvy [Ald05]. Say [Can00, Cha06g]. Sayings [Ann00]. SC-CFS [Ito01]. SCADA [Yar08]. Scalability [BHHV00, BFH02, CM01a, PBB^{+}00]. Scalable [ACV00, Ano01f, ASDZ00, FNKR01, GBHC00, JFWS00, Kno08, KW00, Lem01, LY01, MH01, PP06, PA04, SMF^{+}01, SYC^{+}01]. Scale [AS00, Ano01h, BL00, Ken01, BH00a, LSB^{+}09]. Scalpel [Die00]. Scanner [DH00]. Scanning [BBD^{+}01, PH01]. ScanSSH [PH01]. Scheduler [NVZ01, Wil02, YHS^{+}02, BHK^{+}00]. Scheduling [CAGS00, FNKR01, GLF^{+}00, HS00b, LP02, LSG02, MH01, Reg02, San01, ZYKW02]. Schema [BJP00]. Scheme [JN00, KCK^{+}00, RS01, GS02]. Schemes [BU02]. Schultz [Chu02a]. Science [Rus00b, SF07]. Scientific [AK01, MCB^{+}01, Wal08]. Scientist [Kri00a]. Scientists [Coc01]. SCO [Kol05d, The04]. Scratch [BE09e, BE09f]. Script [Cou00, Mue07b]. Scripted [Wip00]. Scriptics [MSO00]. Scripting [Bea01c, NZ00, RBPM00, War01]. SCSI [PGCR01]. Scwm [BNB01]. SDMI [CWL^{+}01]. SE [Sal03c]. Search [BC06, LSB^{+}09]. Searches [CW04]. Searching [Coo04a, Gra01b]. Seattle [USE00b, USE00d]. Sebek [Bal03, FBMZ04]. sEc [SMK02]. Second [Ano06c, ZPL01]. Secret [Lie00, Ver07]. Secure [BP01, BCR07, Bon01, EBC01, FKM00, FKM02, GMAM01, Gee03a, GTK08, HSW09, Hei05, Hei07, HRF00, IB01, Ito00, Kar02, KBA02, MB02b, McC05b, OP01, Per02, RI00b, Sal02g, vD00]. Secured [Ito01]. Securing [Ale05c, Coh02, Nem00b, SGS^{+}00]. Security [Ale04b, Ano01i, Ano01b, Ano02a, Ano03a, Ano04c, Ano04s, Ano05a, App04a, Ark02b, Bah00, BS00, BJP00, Bar00, BDTW01, BM00a, Bro01c, Bru00a, Bru01, Car00b, Cha06e, Che02, Cho00, Cho05, Chu03, Coc01, CBKH^{+}00, CGP^{+}02, DH00, Dit00, Far04a, Fer09f, For00, For01, Gal08b, Gal09c, GW07, GB04, Jon02, Kim01, KM01, Koc02, LS01b, Mal05, Mar00, MSV09, MLFR02, MIK02, Mit02, Moo02, Mud00b, OS06, Pre07, RKS02, Rub01b, Rub01c, Sam02, Sch02e, USE00i, USE01h, USE02b, WCS^{+}02, Wys03]. Seek [Coc01]. Selecting [ASV^{+}02]. Selection [Fre06]. Self [Kno00, SGS^{+}00]. Self-Contained [Kno00]. Self-Securing [SGS^{+}00]. Selling [Bla01b]. Semantic [SSCG09, Coc01]. Semantics [HPK01]. Semiconductor [Gut01]. Sendmail [All02d]. Sense [Ano05n]. Sense-and-Response [Ano05n]. Sensitive [GW02, PG01]. Sensor [MW05, Mar04, SLR02]. Separating [RJ01]. Sequence [Krz01, MDH00]. Sequencing [KNRS00]. Sequential [KCK^{+}00]. Serialization [McC05a]. Series [Bru00b]. Server [ASS01, BV00, BLF^{+}00, CBKH^{+}00, EKNG^{+}05, Fly03d, Fre06, JCKS04, Kno08, LB02, LP02, LR04, Mag02, McQ00, MSO00, OESV04, PVC01, PLT00, RMSK00, VB00, Wel00, OK01]. Servers [ASDZ00, FNKR01, Gal09d, JKN^{+}01, MCD^{+}02, PH01, She00, SS00, VTFM01]. Service [BH00a, CDGN01, CD01b, DLD00, FCTV02, GBHC00, Has03a, Has03b, Has07b, HBC^{+}01, JN04, LY01, MS02, MVS01, NSZ01, Oli01, PP06, Pen04, Rot01, VTFM01, JGKvS04]. Serviceability [RB01]. Services [ABQ02, Ano05n, Cha00b, DVWW01, FSAK01, Gwi00, Has03c, Has07a, JFWS00, LBN01, Mor00, NSK04, Ott00, PF04, RMSK00, RI04, Sch01, SMF^{+}01, SYC^{+}01, SMM04, TKM04, Tur04a, YV00, vBBB^{+}02, Ano05j]. Session [Jos06a, Nes00]. Set [DKCJ02, Dar00g]. setf [Cou08a]. Sets [App04b, Spi03]. Setting [Siv01]. Setuid [TSW08, CWD02]. Setv [Sha00]. Several [KS00]. Sfio [FKV00]. SFS [FKM02]. Shaft [DLD00]. Shape [BS01]. Shaping [Has06d]. Sharable [RS01]. Share [CAGS00, NVZ01]. Shared [FDSZ01, Joh02, PW00, RI00a, SH02, Uri06, VB00]. Shared-Disk [SH02]. Sharing [BHL00, BM00a, NSK04]. SHARP [BE00]. Shea [Har02a]. Shell [MM00, Tur03b]. Shepherding [KBA02]. Sheriff [Gal08b]. Shibboleth [Cur00]. Shine [Kol09b]. Shocking [Bur06d]. Shoot [How05]. Short [Far05h, Jos06a, LS01a]. Short-Lived [Jos06a]. Should [Bah00, Cou07, Get02a, How06a, How06b, Jos08e, Jos08f, Kan06b]. Show [Tyl02a]. Showcase [Ano00e, USE00c, USE01d]. Shrinking [Sah08]. Shui [Bag04]. Shumway [Chu02a]. Sick [Bur00a, Bur01b, Bur01c, Bur01d, Bur01e, Bur01f, Bur01a].

Side [BU02]. Side-Channel [BU02]. Sided [Fin05]. Sign [Dav01]. Signal [Eng00]. Signaled [BGSS00]. Signals [Van09c]. Signature [Kuh00, Ras07]. Signatures [wCfJRW05, Soo08b]. Signed [MB02a]. Silicon [Bea01b]. Sill [Pur02]. Simple [BS00, BB02, Cha01c, DG09, Gol06, JN00, LB02, Mil02, PLT00, Tur03d]. Simulating [SMMM08]. Simulation [HS01]. Simultaneous [MM01]. Single [BCGD00, Ras06, SA02, Won01, And01]. Single-Dispatch [Won01]. Single-Level [SA02]. Sinkholes [Sor06]. Site [App04c, Dic08]. Sites [CWSC01]. Size [ZABN01]. Size-Changing [ZABN01]. Skills [BE01]. Skoudis [Chu04a]. Slab [BA01]. Slime [Bea01b]. Slow [YNN01]. Slow-Motion [YNN01]. Small [Bro01c, DVWW01, Moo02, PMN^{+}04]. Smart [Far00i]. Smart-Space [Far00i]. Smartbox [Yen00]. Smartcard [Ito01]. Smashing [BST00]. SMP [BHHV00, Ger02, Leh01, Tho00a]. SMPng [Luc05]. SMS [Jos09b]. SMTP [Lan08]. SnapMirror [PMF^{+}02]. Snapshot [SG01]. Snapshotting [Ama00]. Sniffing [Sin05a]. SNMP [Sel00b, Skv03]. Snort [Las01]. Sockets [Fly03d, SSB00]. SockMi [BCR07]. Soft [SGM^{+}00]. Softnet [SNOK01]. Software [Ano04b, AE01, BC01, BP00, Bro08, CCG^{+}00, Cha01f, Coc01, III01, FSLM02, GW02, Gol06, Gut02, Hem00, Joh02, KFBK00, Luc00, Mas02, OESV04, RI00a, RFS^{+}00, RV00, RS01, Sch02b, The04, USE00f, Wys04, Sal03d, Col00, Nic02b, Ano01f, Kar02]. Software-Engineering [Luc00]. Solaris [Ama00, BR00b, Cho00, DM07, Gal05, Gal08d, Gal08e, Gal08f, HDM08, RS01, Siv01, Siv03, Woo01a]. Solaris-compatible [Woo01a]. Solution [KHDC01]. Solutions [Chu03, Sit03b]. Solved [ECF^{+}00]. Some [Bro06, Die02, How05, McC01e, MI07, Sim01, SF07, Spi06, Van01]. Someplace [Jos07a]. Soul [Bla01b]. Sound [GC00]. Source [All02c, All02d, Bro01a, BC00, III01, Far06g, Gra00b, Gra00c, Gra00d, Gra01a, Gut00, Hoh01, Kim01, MS00, SS02, Sif00, Sor06, Str02, The04, WG00, Woo01a]. Sources [Coo04b]. South [Ano06e]. Space [Eng00, Far00i, JKN^{+}01, Sah08, And01]. Spaces [NSZ01]. Spam [Ano03d, App04b, App07a, App07b, HK03, Has04c, Has06a, Has06b, Hob05, How05, Jos06a, Vix02, Wal06, App04b, Cha02b]. Spam-Fighting [Has04c]. Spam-Filtering [HK03]. Spammers [Ege04]. Spare [SPW02]. Spawning [BE07d]. Speaking [Kis00]. Special [Lif01, USE00j]. Specific [ECCH00, YMC02, ZS01, CV02]. Specifications [Mil05]. Spectacular [Bro06]. Speculative [Ros01]. Speech [MRL^{+}02]. Speech-Generated [MRL^{+}02]. Speed [Uri06, YC01]. Speeding [Kri00b, YHS^{+}02]. Spin [BH00b]. Spitzner [Chu02b]. Spot [Fly00a, Fly00b, Fly00c, Fly00d, Fly01b, Fly01c, Fly01d, Fly01e, Fly01f, Fly02b, Fly02c, Fly02d, Fly02e, Fly03c, Fly03d, Fly03e, Fly03a, Fly03b, Fly04e, Fly04a, Fly04d, Fly04b, Fly04c, Fly05]. Spotless [SMES01]. Spyglass [LSB^{+}09]. Spying [Hol05]. Spyware [BS04]. SQL [She00]. SRUTI [Ano05k]. SSH [OW00, PH01, SWT01]. SSL [Car00a, RCK02]. SSLACC [RCK02]. SSO [Sch04]. SSO/RSO/USO/Oh [Sch04]. SSWG [Gwi00]. SSWG-RT [Gwi00]. Stack [AHT^{+}02, AE01, BGS^{+}08, BBD^{+}01, BST00, Eng00, FS01, Ran02, SMJ00]. Stack-Based [Ran02]. Stack-Smashing [BST00]. Stackable [ZN00, ZABN01]. StackGhost [FS01]. Stacks [DADAD02]. Stand [Cha03b, LB00]. Stand-alone [LB00]. Standard [Cou08b, Dzo09, Gsc01, Jos01a, Jos02, Jos03, Sim01, Sto06c, Dan01]. Standardization [Sim00, Sim01]. Standardize [Sto07b]. Standards [Ano02o, App04b, Bla00c, Jos00c, Jos00d, Sto05, Sto06a, Sto06b, Sto06c, Sto07c, Sto08b, Sto08d, Sto08c, Whi01]. Startup [SV02a, Str02]. State [Gal08c, Ham00, MS00, Sur01]. Stateful [Har02c, vR01b]. Statement [Gee03b]. Static [McC03d, YST02, ZEJ02, vD09]. Statically [LE01]. Statistical [BB02, Pro01]. Statistics [NNW00]. Status [Jos00a, Jos01a, Jos02]. Statutes [Eck01]. Steganalysis [Pro01]. Stephenson [Far04b]. Stepping [Ken01, ZP00b]. Steps [Ano05k, NH00]. Steward [CFK00]. Still [Ioa02, MI07, SC09]. Stones [ZP00b]. Stop [Ber02]. Storage [ACV00, AHK^{+}02, Ano02f, Ano03e, Ano04e, Ano06a, BGS^{+}08, BCGD00, Cou02, DADAD02, FADAD02, Gal09a, Gra02, GSGN00, GSS^{+}02, Hel09, JHZK08, JHP^{+}09, Lam00, LSB^{+}09, MVS00, MB02a, MLFR02, Mor02, NSH^{+}02, PG01, PGCR01, QD02, RKS02, Sch02c, SGHS01, SGZ^{+}02, SGMV08, SGS^{+}00, USE02c, VG00, WOSW02, WW02]. Storage-Aware [FADAD02]. Store [SA02]. Stores [PMN^{+}04]. Stories [Mar01a]. Stork [SPC08]. STORM [DG09, EVS^{+}08, SDHD07]. Strategic [PIA^{+}07]. Strategies [Ash02, BCK06]. Strategy [Gee03a, HcC01, Mas02]. Streaming [CV02, CWVL01]. Streams [ASS02, GJM02]. Stress [ZD02]. Stress-testing [ZD02]. Strike [Gra04]. Strikes [SvM05]. String [CBB^{+}01, STFW01]. Striping [ASS02]. Strong [KP01, KPv00, MLFR02]. Structural [EG01]. Structure [GP01b, MAF^{+}02]. Structured [WH02]. Structures [DD07, GBHC00, ZD02]. Stub [HLU^{+}00]. Stub-Code [HLU^{+}00]. Study [BV00, BP00, DP00, Ezo01, FM00, KA01, MIK02, RCH00, SYN02, Zad02, vdL02]. Studying [CCG^{+}00]. Stuff [BE08a]. STUG [Ano04a, Sal04a]. Style [Sea00]. SubDomain [CBKH^{+}00]. Subscribe [EG01, GSAA04, OP01, PRS06, WJL04, ZSB04]. Subscription [ZSB04]. Subsetting [TK01]. Subsystem [JHZK08, Ruw00]. Success [Mar01a]. Successful [Moe02]. Sued [Nic01a]. Suing [Ege04]. Suite [ZS01]. Summary [Ald05, Ano03o, Ano04r, BY00a, BY00b, MY04, Mul04b, Mul05a, Mul05b, MY06, Veg07, YM05, You08a, You09b]. Summer [Far02f]. Summit [Far01h, Ano02b, Ano05c]. Sun [Gal08a, Gal08b, Gal08c, Gal09b, Ano00s, Gal08a, Gal08d, Gal08e, Gal08f, Gal09a, Gal09e, Gal09d, Gal09f, Gal09c]. Super [Gra01c]. Supercomputer [Coc01, US01]. Supercomputing [Ano01f, Chi00]. Supermon [Min01]. Supernetworking [Sch00c]. Superuser [Lan06]. Support [ABC^{+}00, GC01, LS01b, LY01, MD00, RI00b, Rus00b, SNC00, SYC^{+}01, Tyl03b, lWS00, WCS^{+}02, ZABN01, iiH01]. Supporting [MZ00, Pre07, YST02]. Surfing [KT01, Kle00]. Surplus [CAGS00]. Surveillance [FBH^{+}02, Spi02]. Survey [Sch00a]. Survivability [Die01]. Surviving [Kan05a]. SUS [Gra01c]. Susceptibility [CMB^{+}01]. Sushi [Rig02]. Swaddling [Gal09c]. Swapping [Pri01]. SWILL [LB02]. Sword [Fin05]. Sy [USE01g]. SyD [PMN^{+}04]. Symlinks [Sim03]. Symmetric [BU02, CAGS00]. Symposium [Ano00g, Ano00d, Ano01d, Ano01e, Ano01j, Ano03f, Ano04h, Ano04c, Ano04f, Ano05l, Ano05e, USE00h, USE00i, USE00d, USE01h, USE01c, USE02f, USE02b, Ano01b, Ano02d, Ano02a, Ano03a, Ano05a]. SYN [Lem02, Oli01]. Synchronization [Ger02, HH01]. SysAdmin [Cha01d, Ann00, Cha03b, CFK00, Dar04c, Par00, Par02b, Tyl03a, Kne01, Tyl02d, Wyn00]. Sysadmins [BE01, Cla00, Per01]. System [Abl08, Ale04b, AG01, All02a, Ama00, ABC^{+}00, Ano05b, Ano06b, Apt01, Ayu06, Ban00, BSF02, Bes00, BE00, Bro01a, Bru09a, Bur00b, Bur00c, Bur00d, Bur00e, Bur05, CW04, CLRT00, Car00b, CL00, CCSP01, Cou08c, Cou08b, DM07, DDB08, DBH06, Dij00d, Dit00, Drz00, Dus08, ECCH00, FSLM02, Fly03e, For01, FKM00, FKM02, Gal08d, Gal08e, GFWK02, GSGN00, Gwi00, Hei07, Hel09, HBG^{+}06, HBG^{+}07, HDM00, Ito01, JFWS00, JHP^{+}09, Kan05b, KCN02, KM01, Kno00, KFBK00, Lam00, LBN01, Llo07, LS01b, Los02, LY01, MAF^{+}02, MVS00, MB00, MCD07, McK06, Mew01, MD00, Mil02, Mil05, MHM^{+}00, Mue07b, OM00, PG01, PMF^{+}02, Pei05, Per02, PBB^{+}00, RJ01, RKS02, Rob00, RS01, RLA00, SMRF01, SH02, See06, Sel00a, Sel02, SMF^{+}01, SV02b, Sha02a, SF00, SHS01, Tra02, USE00b, Uri06]. System [Vol00, WRC00, WRPK02, WH02, WJL04, Wat01, Wil02, lWS00, YD00, Coc01, Get02a, Mag02, McK07, Spi02, Tho00b]. System-Call [SF00]. System-Specific [ECCH00]. System/390 [Tho00b]. SystemImager [Fin00b]. Systems [Ale04c, Ano00a, Ano00f, Ano00d, Ano01c, Ano01j, Ano02g, Ano03f, Ano04h, Ano04d, Ano05f, Ano05j, Ano05l, Ano05e, Ano05n, Ark05, BTS^{+}00, BP01, Bec01, BL00, BP00, Cha06b, Far00h, FADAD02, Gra00d, Hei05, HH01, HPS02, KPv00, KYH^{+}01, LSB^{+}09, Ley01, Maz01, Moo01, Mue07a, Mun02, OSE^{+}00, OP01, RFS^{+}00, RMS02, SC09, See06, SGM^{+}00, SGS^{+}00, Tyl02c, USE00g, USE00h, USE00f, USE00d, USE01a, USE01e, USE01c, VG00, ZN00, ZABN01, iiH01, Ano06e, Sch01, Ano05h, Ano06c]. SystemStarter [SV02a].

T [Gal09d]. Table [Lev00, VB00]. Tableau [BE09b]. Taint [Tur05b]. Taking [All02c, All02d, Lan08]. Tale [Lev05]. Tales [Dar05]. Talk [BE01, Ros01, Tar02]. Talking [Bur04]. Taming [BM00b, Nic00e]. Tamper [WRC00]. Tamper-Evident [WRC00]. Tape [SGMV08]. Tapping [SDS09]. Target [PGCR01]. Targeting [CO02, McC08]. Task [AHT^{+}02]. Tcl [Ano03c, USE00e, Bab01, BV00, BLF^{+}00, CDRV00, Fly04e, KD02, KPv00, LM00, MSO00, Ros00, WT00, Wel00, ZM00]. Tcl/2k [USE00e]. Tcl/Tk [Ano03c, USE00e, BLF^{+}00, KPv00, ZM00]. Tclsh [Fly00a, Fly00b, Fly00c, Fly00d, Fly01b, Fly01c, Fly01d, Fly01e, Fly01f, Fly02b, Fly02c, Fly02d, Fly02e, Fly03c, Fly03d, Fly03e, Fly03a, Fly03b, Fly04e, Fly04a, Fly04d, Fly04b, Fly04c, Fly05]. TCP [MSV09, OMP02, SK02, SMJ00, Tie01, vR01b]. TCP-Friendly [OMP02]. TCP/IP [SMJ00]. tcpdump [For01]. Teaching [Cho05, Gra00d, Ker06]. Team [Kol05e]. Teams [Kol09b]. Technical [Ano00b, Ano00h, Ano01k, Ano02c, Ano03k, Ano04i, Ano05d, Cha00a, USE00a, USE01f, USE01b, USE02d, USE02e, USE06]. Technique [Ama00, SMK02]. Techniques [BTS^{+}00, BE09c, Coo04a, Die02, Gro09, HK03, How05, LV02]. Technologies [Ano00f, Ano00d, Ano01j, Ano02f, Ano03f, Ano04e, Gol00, Kry02, USE01e, USE01c, USE02c, Way02, Ano03e, Ano06a]. Technology [Ano00g, Ano01e, Ano02d, Ano04f, Dav02, Fin05, Hoh01, Ken03a, Ker03, Kle00, NNW00, Tyl02a, USE01g, USE02f]. Teenagers [Ale05d]. Telemedicine [Sch02a]. Telephony [Ark02b, Sen07]. Telephony-Based [Ark02b]. Telescopes [Moo02]. Tellme [OM00]. Temperature [PG01]. Temperature-Sensitive [PG01]. Template [Gsc01]. Temporary [CBWKH01]. Tempting [Sin05b]. Ten [Lem00, Sel02]. Tenth [Ano05h, USE01h]. Terminal [McQ00]. Terms [Fer09f]. Test [AT05, Cou07, Way02]. Testbed [PMN^{+}04]. Testing [FM00, For00, GS06, Ker06, KPv00, ZD02]. Texas [USE00e, USE01e]. Text [BC06, LV02, MM01, Sch02d]. Thailand [US01]. Thanks [You04a, You05a, You06a, You07c, You08b, Hum00a]. Their [BC00, MCD^{+}02]. Them [Nic03a, Rus00a, Rus04a, Rus04b, Smi01, WD01]. Theoretic [MV04]. Theoretical [Bur00e, HR01b]. Theory [Lif01]. There [Ber02, Dit00, Bur07]. Thermodynamics [Cou08c]. These [BE07e, Hum05, Rus03]. Thin [YNN01, YNST02, McQ00]. Thin-Client [YNN01, YNST02]. Things [Gal08a, Gal08b, Gal08c, Gal08d, Gal08e, Gal08f, Gal09a, Gal09e, Gal09b, Gal09d, Gal09f, Gal09c, Lem00, Par02a]. Think [FSLM02]. Thinking [Ber02]. Third [Ano05j, Ste01]. Thirtieth [Sal05g]. Thomas [Ano04b]. Those [Bel02, Dit09a, JW01e]. Thoughts [Far04e]. Thousand [SMMM08]. Thrashing [JZ01]. Thread [Eng00, MP01b]. Threaded [VB00, Chr01, RBPM00]. Threads [MB02c, TS04, Woo01a]. Threat [AC08, Dit00, FF07, Mud03, Mue07b]. Threats [Ark02b, KM01]. Three [Ano01a]. Thresh [Sel00b]. Threshold [Sel00b]. Throughput [ASS02, LS01a]. Thyself [Jos08a]. Tie [BE06e, BE06f]. TierStore [DDB08]. Ties [PN05]. Time [BST00, Bru00b, Ger02, Gle02, GW02, HS00a, HH01, HPS02, Kno06, KK08, Mal06, Mas02, Mor01, NVZ01, RE01, Ros01, Sha02b, Slu06, SPW02, SYN02, Tur05a, BS01]. Time-Sensitive [GW02]. Timeline [BE08c, MB02b]. Times [Hum00b, Kri00b]. Timing [GSS^{+}02, HR01a, SWT01]. Timing-accurate [GSS^{+}02]. Tinker [Fel02]. Tk [Ano03c, USE00e, Bab01, BLF^{+}00, Ing00, KPv00, Ros00, ZM00]. TkGecko [Bal00]. TLS [DS01]. Together [LM00, WD01]. Toil [Bur06c]. Tolerance [CJL^{+}02]. Tolerant [CL00]. Tom [Sch02c]. Tomorrow [Kri00a]. Too [Lim09]. Tool [Ark02a, Cho00, GW02, Kim01, LB01, MNH00, Plo00, SAB^{+}01a, Spi00, Vol00, Ost00]. Tool-Based [Spi00]. Toolbox [LLB02]. Toolkit [LB00, Maz01, McC00g, Ros00, WRA00, Tay01]. Toolman [Sin00a]. Tools [BE06b, BE06d, BE06a, BE06e, BE06f, BE06c, BE07d, BE07e, BE07a, BE07c, BE07b, BE07f, BE08f, BE08c, BE08a, BE08d, BE08b, BE08e, BE09a, BE09d, BE09b, BE09e, BE09f, BE09c, Bru09b, CWSC01, DG09, DLD00, Dij00e, Fiu05, Ham00, Hun00, Sor06, RFL00]. Top [Gro09, Jos09e]. Topics [Ano05h, Far05h]. Topological [CD01a]. TORBA [MMGL01]. Toronto [Jac04]. Town [USE01a]. Trace [Moo01]. Traceback [BS02]. Tracing [BC00, Gra01a]. Track [Cha01b, PSW^{+}01, SGLG02, Slu06, USE01f, USE02d, USE02e]. Track-Aligned [SGLG02]. Tracker [NH00]. Tracking [Ayu06, Chu02b, Fou02, HDM00, JADAD06, Sto00a]. TRaDe [CD01a]. Trademarks [III01]. Trades [Sho01, Tyl02d]. Trading [MMGL01, YGC^{+}00]. Traffic [Ano05i, Ano05k, HPK01, Has06d, HcFG01, Plo00, RMS02, Syv01]. Transaction [HcC01, ZYKW02]. Transactional [PA04, SDS09]. Transfer [McC05b, PP06]. Transformation [BBLo04]. Transient [GB01]. translated [Sch02d]. Translation [KHDC01, SS02]. Translucent [Pac00]. Transparency [LCC00]. Transparent [BST00, CCSP01, NDI04]. Transport [OMP02]. Transport-Level [OMP02]. Travails [How04]. Tree [Ber00c, Bla01a]. Trees [Bru04]. Trends [Nic03b]. TRIAD [Che00b]. Tribute [You09c]. Trick [Eng00]. Tricks [Cha05a]. Trinux [For00]. TRIPLE [BBLo04, Cor00]. Triple-DES [Cor00]. Trojans [Loz05a, SDHD07]. Troubleshooting [JHP^{+}09]. True [iiH01]. Trust [Che00a, Smi01, Wys04]. Trust-Management [Che00a]. Trusted [MVS00, Wat01, YS02]. TrustedBSD [Wat01]. Tuning [All02a, Gra00b, Tie01]. Turf [JW01f]. Turned [Tyl03a]. Tutorial [Mas00]. Tux2 [Phi00]. TV [Gib00]. Twenty [Sal03f, Sal03e, Sal04i, Sal04k].

Twenty-Five [Sal03e]. Twist

[Die01, Swa01]. Twisty [Eck01]. Two

[Com00, Fin05, Lev05]. Two-Sided [Fin05]. Type [ACG01, HK00a, sLC02, STFW01]. Type-Assisted [sLC02]. Types [McC01f, McC03a, Spi06].

U [Bur07, Sal03c]. Ubiquitous [RAMC^{+}04]. Ugly [Bea09]. UML [Fly00e]. Unauthorized [Fou02]. Uncle [Cha02b]. Underground [Cym06]. Understanding [Has05d]. Undue [Sto08d]. Unicode [LaB00]. Unified [Bro01a, KCK^{+}00]. Unifying [SHS01]. Universal [Moo01]. University [Ano06e, US01]. UNIX [Ann00, BE00, CCSP01, Chu03, Dzo09, Ger02, Gla02, Gra00b, Gra00c, Gra00d, Gra01a, Gra01c, Kam02, Loz05b, Sal00a, Sal00b, Sal03e, SEZ^{+}01, Spi00, Too09]. UNIX/Linux [Chu03]. Unleashing [Ama00]. Unstructured [JGKvS04]. Untaint [Tur05b]. Untapped [Bru09a]. Untraceable [JM03]. Untrusted [AR00, CCG^{+}00, MVS00, PBP02]. Unwanted [Ano05k, Mal06]. Update [Ber00b, Cha06h, Cha06f, Dit01, Jos00a, Jos00c, Jos01b, Jos01a, Jos02, Kol07a, Kol07b, Lon07, LC07, Lon08, Pal05, Par05, Sto05, Sto06b, Sto06c, Sto07c, Sto08b, Sto08d, Sto08c]. Updates [Jos03, SGM^{+}00]. Updating [Los02]. Ups [Chu03]. Urban [Kri00a]. URN [NH00]. USA [USE00a, USE01g, Kol04g, Kol05f, USE00c, USE00g, USE00h, USE00f, USE00b, USE00e, USE00d, USE01f, USE01h, USE01b, USE01e, USE01c, USE02c, USE02d, USE02e, USE06]. USACO [Ano04n, Kol00a, Kol04f, Kol04g, Kol05f, Kol05e, Kol06, Kol07a, Kol07b, Kol08, Kol09b]. Usage [Lan06, NNW00]. Use [Dij00e, Dre00, How06b, Kan06b, McC04b, RV00]. Useful [Guy01]. Usenet [Coc01]. USENIX [Ano00b, Ano00c, Ano00f, Ano00d, Ano00h, Ano01j, Ano01k, Ano01b, Ano02c, Ano02a, Ano03f, Ano03e, Ano03k, Ano03a, Ano03o, Ano03p, Ano04e, Ano04i, Ano04c, Ano05d, Ano05a, Ano05m, Ano05s, Ano06a, Ano06d, Jac04, USE00i, USE00e, USE00d, USE00a, USE01f, USE01h, USE01b, USE01e, USE01c, USE02d, USE02e, USE02b, USE06, Ano00s, Ano00u, Ano00v, Ano00t, Ano01u, Ano01v, Ano01w, Ano01x, Ano01y, Ano01z, Ano01-27, Ano01-28, Ano02p, Ano02q, Ano02r, Ano02s, Ano02t, Ano02u, Ano03n, Ano03o, Ano03p, Ano04r, Ano04o, Ano04p, Ano04q, Ano05t, Ano05r, Ano05u, Ano05s, Ano06e, Ano06d, Cha06g, Coc01, Dic08, Dic09, Dol06, Far05h, JG08, Kol04h, Kol05f, Kol05e, Lon09b, MY04, Mul04b, Mul05a, MD05, Mul05b, MY06, Pal05, Par05, Sal00c]. USENIX [Sal03f, Sal04h, Sal04g, Sal04i, Sal05f, Sal05g, Sal05d, Sal05h, Sal05i, Sal05c, Sal05e, Sto06a, Sto08c, The04, Veg07, You04b, YM05, You05a, You05b, You06b, You06c, You07a, You07d, You07e, You07b, You08a, You09a, You09b, You09d]. User [AB04, AK01, Ano03c, CIK00, DP00, DL01, Dik01, ESW01, Eng00, Gra01c, HS01, Hug00, Jos09c, JKN^{+}01, LGJ00, Maz01, Moe02, RJ01, Ros00, SMRF01, Sol08, lWS00, Dik00]. User-Centric [Hug00]. User-Driven [HS01]. User-Level [DL01, ESW01, Maz01, SMRF01, CIK00]. User-Mode [AB04, Dik01, Dik00]. User-Space [Eng00, JKN^{+}01]. Usernames [Mos03b]. Users [Adl00, MS02, Pac01b]. Using [AR00, AB04, Ark02a, BCR07, BM00b, DS01, DP00, ECCH00, Fly03c, FM00, For00, For01, FKV00, FKM02, GB01, GS06, Gol06, Gra00a, HRF00, Hop02, Kan05b, KFBK00, LP02, LM00, LV02, McC01f, McC03d, McC04c, McC04d, MZ00, MoCA^{+}02, Nik02, Pie03, Rao00a, Rao00b, Rao00c, Rao00d, Rao00e, Rao00f, Rao01a, Rao01b, SPC08, SAB01b, SF00, Sor06, Sta00, SD05, Tur03e, WS01, Won01, YD00, YNN01, ZEJ02]. USITS [Ano00d, Ano01j, Ano03f, USE01c].

USO [Sch04]. Utilities [CMP00, Pik00b]. Utilization [LSG^{+}00]. UX [MMS00, MVW00].

V [Gal09f]. V5 [Ito00]. VA [Fin00b].

Vacation [Far02f]. Vale [You05b]. Validate [SAB^{+}01a]. Value [Dar03a]. Values [Ken02c, McC01f]. Variable [ASS02, McC01g]. Ved [Has05c]. VEE [Ano05m]. Vendor [BC01, Cha00b]. Vendors [Kim01]. Venti [QD02]. Verification [FC01, GMAM01]. Verifying [MB02c]. VeriSign [Ren02]. VERITAS [LS00]. Version [Kan05b, Pis08]. Versioned [BE09d]. Versioning [BR00b]. Versus [LaB00, SGM^{+}00]. Very [HDM00]. via [CD01b, KBA02, DVWW01]. Victims [SMMM08]. Video [Kal09]. Videos [Dic09]. Viega [Kar02]. View [Jos07a]. Views [Coc01]. Vikings [Jos06a]. Vipul [Has05c]. Virtual [AB04, And09, Ano00g, Ano01e, Ano02d, Ano04f, Ano05m, AE01, DLS^{+}01, JR02, JADAD06, Law00, LHAP06, MP01b, NVZ01, OK01, Pre05, Pro00, Ran02, RI00a, RMSK00, RI00b, RB01, SMK02, SMES01, Sch00c, SVL01, Sur01, USE01g, USE02f, vD00]. Virtual-Time [NVZ01]. Virtualization [Gal09e, GW07, HWF07, HDM08, MCZ06]. Virtualized [Sac08]. Virtualizing [Cha07d, SVL01]. Virus [GB04]. Visibility [KCN02]. Vision [Joh02]. Visits [Tyl02a]. Visualization [GW02, Mun02, Plo00]. Visualizing [CBB00]. Vive [Cha02c]. VM [Ano01e, Ano04f]. Vmem [BA01]. VMM [LHAP06]. VMM-Bypass [LHAP06]. VMware [Gal09f, SVL01]. Voice [Cha05b, Ker03]. VoIP [Cha06b, Cha06e, Cha06c, Cha06d, Cha08b, Has07b]. Volume [TM01]. Volunteers [You04a, You05a, You06a, You07c, You08b]. Voting [JJR02, Rub01b, Rub01c]. VPN [Sch02a, Zwi01]. VRRP [Zwi01]. vs [AT05, Bea01b, Gal09f, JW01d, Lev05, Tyl02d]. vSphere [Gal09f]. Vu [DP00]. Vulnerabilities [CBB^{+}01, CBWKH01, LE01, STFW01, Uhl03, Yar08]. Vulnerability [Wys03]. Vulnerable [SC09].

W [Aik07]. Wait [HPS02]. Wait-Free [HPS02]. Wales [Ano06e]. Walkin [BE07e]. Walls [Bur04]. WAN [CDGN01]. Wanna [Van01]. Want [Slu05]. Wanted [Cha01d]. Warnings [Jos09b]. Warrior [Far04a]. Wars [Jos06a]. Was [Rus02]. Washington [USE00b, USE00d, USE01h, Dav02]. Watch [Cha06e, Cha06c, Cha06d, Cha03a, Sin02]. Watching [Ber00c]. Water [JW01a]. Waypoint [NSZ01]. Weaknesses [KM01]. Weapon [Lie00]. Weapons [PSW^{+}01]. Web [App04c, Ash02, BSF02, Coc01, Dic08, DEH^{+}09, EKNG^{+}05, FBH^{+}02, FBMZ04, FSSF01, Gal09b, GTK08, Gro09, IB01, Jen04, JN04, JKN^{+}01, KHDC01, KA01, LB02, Lim06, LY01, MCD^{+}02, Mao09, MM00, NSK04, Nic01b, Pie03, PA04, PF04, PLT00, SGHS01, SMM04, Smi01, TKM04, Tur04a, Tur04d, Uhl03, VTFM01, WRC00, Wel00]. Web-Enabled [Smi01]. Webapp [BE09e, BE09f]. We’d [Tyl01, Tyl02b]. Were [BE07e, Tyl02c]. West [Nic00e]. Western [Slu07]. Wharf [USE00j]. Whazzup [And00]. Where [Bel02, BE06a, Get02a, Ken03b, Kne01, Ros01, Swa03b]. Which [Sit02]. White [WE06]. Whither [Sto07a]. Who [Jos09f, Mal06, Ran05]. Wi [MPC^{+}08]. Wide [KCN02, McC02e, Mee00, SVF01, Tie01]. Wide-Area [KCN02]. Widely [Fer09f]. Widespread [NW04]. Widget [Bla01a, GS02]. Widgets [MZ00]. WIESS [USE00f]. Wiki [Cha02d]. Wild [Nic00e]. Will [Cha03b, BE08d]. Window [BNB01, Get02a]. WindowBox [BS00]. Windows [Pac00, USE00b, USE00d, Ale05b, BJP00, BCGD00, Chi00, FM00, Luc00, Mor00, MMBE00, Nor00, Ott00, SEZ^{+}01, SSB00, Spi00, Sta00, Ver07, Wal00, WD01].

Wine [OS06, Whi00]. Wireless [Ano04s, Ano05i, AC08, BCH^{+}00, Bye02b, Bye02a, Far06g, Has06e, Mah04, Mar04, SLR02]. WISPER [Far06g]. Within [Bru04]. Without [AHT^{+}02, Bla01b, Bur07, Sin03]. Withstanding [DAK08]. WitMeMo’05 [Ano05i]. Witty [WE04]. WLANs [Nik02]. WMCSA [Ano05f]. WOLF [WH02]. Women [Rub01a]. Woobie [BE06c]. Work [BS01, Gee03c, KV02, WE06]. Working [Gwi00, Jos01b, McC03e, McC04e, McC05a]. Workload [CWVL01, MMS00, ZYKW02]. Workloads [RLA00]. Workplace [Nic00a, Nic00b]. Workshop [Ano01f, Ano05h, Ano05f, Ano05i, Ano05k, Ano05n, Ano06c, USE00f, USE00j, Wal07, Ano01h]. Workstation [SVL01]. Workstations [Ale05c, CB00, KHF01]. World [Gos00, LR01, LeF02, McC03b, Spi00]. WORLDS [Ano06c, Fin05]. Worm [BCK06]. Worms [WE06]. Would [Tyl02c]. Wrappers [KFBK00]. Write [Jos08e, Jos08f, LH02, Tur05d, WH02]. Writers [Mas00]. Writing [Bab01, Cha01d, FL06, wCfJRW05, Rus00c, Slu05, ZYKW02]. Written [ECCH00]. Wrong [Sal05f]. WSDLite [SSB00]. WWW [RE01].

X [Ano05h, Cou08a, Ber00c, Ger02, GP01a, Get02a, Hoh01, Mas01, MB02c, Pac00, Pac01a, SV02a]. X.509 [Ren02]. X11 [TK01]. x86 [OKN02, vdL02]. x86-64 [vdL02]. XCB [Mas01, SM02]. XCL [SM02]. XEMBED [Tay01]. Xen [CFH^{+}05, GR07, HWF07, MCZ06]. Xfce [Fou00]. XFree86 [And00, HC00, Hoh01]. XFree86-4.x [HC00]. XFS [Hel09]. Xft [Pac01b]. XIV [USE00g]. Xlib [SM02]. XML [Dav01, MSO00, Soo08b]. XOTcl [NZ00]. xps [Ber00c]. XSETTINGS [Tay01]. XSETTINGS/XEMBED [Tay01]. XVI [Ano03i].

Ya [Lim09]. Yarrow [Mur02]. YASSP [Cho00]. Years [Sal00d, Sal00c, Sal00a, Sal00b, Sal03f, Sal03e, Sal03c, Sal04b, Sal04i, Sal04a, Sal04k, Sal05d, Sal05h, Sal05i, Sal05c, Sal05e]. Yep [JW00g]. You’re [Nic01a]. Yours [Tyl02c, WW02]. Yourself [Ran05]. You’ve [Nic01a]. Yukihiro [Sch02d].

Z [MB02c]. Zeitser [Chu04a]. ZFS [DM07, Gal08c, Hay06a]. Zope [Cha02d, HAE00, Pel01].

References

Al-Buraiky:2004:BVI

[AB04] Salah M. S. Al-Buraiky.

Building a virtual IPv6 lab using user-mode Linux. ;login: the USENIX Association newsletter, 29(1):??, February 2004. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2004-02/pdfs/al-buraiky.
pdf.

Andersen:2000:SSB

[ABC^{+}00] David Andersen, Deepak Bansal, Dorothy Curtis, Srinivasan Seshan, and Hari Balakrishnan. System support for bandwidth management and content adaptation in Internet applications. In USENIX [USE00h], page ?? ISBN 1-880446-16-2. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/osdi2000/
andersonbansal.html.

Alpern:2002:EPJ

[ABC^{+}02] Bowen Alpern, Maria Butrico, Anthony Cocchi, Julian Dolby, Stephen Fink, David Grove, and Ton Ngo. Experiences porting the Jikes RVM to Linux/IA32. In USENIX [USE02f], page ?? ISBN 1-931971-01-3. LCCN QA76.73 .J38 J42 2002.

URL http://www.usenix.
org/publications/library/
proceedings/javavm02/alpern.
html.

Albrecht:2008:MDA

[ABD^{+}08] Jeannie Albrecht, Ryan Braud, Ren Dao, Nikolay Topilski, Christopher Tuttle, Alex C. Snoeren, and Amin Vahdat. Managing distributed applications with Plush. ;login: the USENIX Association newsletter, 33(1): ??, February 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
february-2008-volume-33-
number-1/managing-distributed-
applications-plush.

Abley:2008:FLR

[Abl08] Joe Abley. Fear and loathing in the routing system. ;login: the USENIX Association newsletter, 33(1):??, February 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
february-2008-volume-33-

number-1/fear-and-loathing-
routing-system.

Adya:2002:CAA

[ABQ02] Atul Adya, Paramvir Bahl, and Lili Qiu. Characterizing and analyzing alert and browse services of mobile clients. In USENIX [USE02e], page ?? ISBN 1-880446-00-6. LCCN QA76.8.U65

U84 2002. URL http://www.
usenix.org/publications/
library/proceedings/usenix02/
adyabahl.html.

Ardi:2008:ATA

[AC08] Calvin Ardi and Daniel Chen.

Architecture and threat analysis of a campus wireless network. ;login: the USENIX Association newsletter, 33(6): ??, December 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
december-2008-volume-33-
number-6/architecture-and-
threat-analysis-campus-
wireless.

Alpern:2001:DTC

[ACG01] Bowen Alpern, Anthony Cocchi, and David Grove. Dynamic type checking in Jalapeño. In USENIX [USE01g], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38

J42 2001. URL http://www.
usenix.org/publications/
library/proceedings/jvm01/
alpern.html.

Anderson:2000:IRR

[ACV00] Darrell Anderson, Jeff Chase, and Amin Vahdat. Interposed request routing for scalable network storage. In USENIX [USE00h], page ?? ISBN 1-880446-16-2. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/osdi2000/
                                        andersonchase.html.

Adler:2000:ENP

[Adl00] Darin Adler. Eazel and The Nautilus Project: Bringing the power of Linux to more users, 2000. URL http://

                                        www.usenix.org/publications/
                                        library/proceedings/als2000/

. Unpublished invited talk at Annual Linux Showcase 2000.

Aoki:2001:SVM

[AE01] Takashi Aoki and Takeshi Eto. On the software virtual machine for the real hardware stack machine. In USENIX [USE01g], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38

                                        J42 2001. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/jvm01/
                                        aoki.html.

Agrawala:2000:ADB

[AFE00] Rajeev Agrawala, Rob Fulmer, and Shaun Erickson. Automating dual boot (Linux and NT) installations. In USENIX [USE00g], page ?? ISBN 1-880446-13-8. LCCN

                                        ????  URL http://www.

usenix.org/publications/
library/proceedings/lisa2000/
agrawala.html.

Alicherry:2001:PMS

[AG01] Mansoor Alicherry and K. Gopinath. Predictable management of system resources. In USENIX [USE01f], page ?? ISBN 1-880446-10-3. LCCN QA76.8.U65 U84 2001. URL

http://www.usenix.org/publications/
library/proceedings/usenix01/
freenix01/alicherry.html.

Anderson:2002:HRC

[AHK^{+}02] Eric Anderson, Michael Hobbs, Kimberly Keeton, Susan Spence, Mustafa Uysal, and Alistair Veitch. Hippodrome: Running circles around storage administration. In USENIX [USE02c], page ?? ISBN 1-880446-03-

0. LCCN QA76.9.F5 C65

2002. URL http://www.
usenix.org/publications/
library/proceedings/fast02/
andersonHIP.html.

Andreasson:2002:CCM

[AHL02] Eva Andreasson, Frank Hoffmann, and Olof Lindholm. To collect or not to collect? Machine learning for memory management. In USENIX [USE02f], page ?? ISBN 1-931971-01-3. LCCN QA76.73 .J38

J42 2002. URL http://www.
usenix.org/publications/
library/proceedings/javavm02/
andreasson.html.

Adya:2002:CTM

[AHT^{+}02] Atul Adya, Jon Howell, Marvin Theimer, Bill Bolosky, and John Douceur. Cooperative task management without manual stack management. In USENIX [USE02e], page ?? ISBN 1-880446-00-6. LCCN QA76.8.U65

                                        U84 2002. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/usenix02/
                                        adyahowell.html.

Aiken:2007:MJW

[Aik07]  Alex Aiken.  In memo-

riam: John W. Backus, 1924–2007. ;login: the USENIX Association newsletter, 32(3):68–69, June 2007. CODEN LOGNEM. ISSN 1044-6397. URL

                                        https://www.usenix.org/
                                        publications/login/june-
                                        2007-volume-32-number-3/
                                        memoriam-john-w-backus-
                                        1924%E2%80%932007.

Andersson:2001:IUE

[AK01] Niclas Andersson and Peter Kjellström. An integrated user environment for scientific cluster computing. In USENIX [USE01d], page ?? ISBN ???? LCCN ???? URL http:

                                        //www.linuxshowcase.org/
                                        andersson.html.

Alder:2005:SSB

[Ald05] Raven Alder. A summary of savvy backbone defense.

;login: the USENIX Association newsletter, 30(6):??, December 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-12/openpdfs/alder.
pdf.

Alexander:2004:ABO

[Ale04a] Steven Alexander. Avoiding buffer overflows and related problems. ;login: the USENIX Association newsletter, 29(1):??, February 2004. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2004-02/pdfs/alexander.
pdf.

Alexander:2004:ISH

[Ale04b] Steven Alexander. Improving security with homebrew system modifications. ;login: the USENIX Association newsletter, 29(6):??, December 2004. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2004-12/pdfs/homebrew.pdf.

Alexander:2004:PPM

[Ale04c] Steven Alexander. Password protection for modern operating systems. ;login: the USE-NIX Association newsletter, 29(3):??, June 2004. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/

                                        2004-06/pdfs/alexander.
                                        pdf.

Alexander:2005:DCL

[Ale05a] Steven Alexander. Defeating compiler-level buffer overflow protection. ;login: the USE-NIX Association newsletter, 30(3):??, June 2005. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2005-06/pdfs/alexander0506.
                                        pdf.

Alexander:2005:FMC

[Ale05b] Steven Alexander. Finding malware on compromised Windows machines. ;login: the USENIX Association newsletter, 30(2): ??, April 2005. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2005-04/pdfs/malware0504.
                                        pdf.

Alexander:2005:ISW

[Ale05c] Steven Alexander. The importance of securing workstations. ;login: the USENIX Association newsletter, 30(1): ??, February 2005. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2005-02/pdfs/workstations.
                                        pdf.

Alexander:2005:WTH

[Ale05d]  Steven Alexander.  Why

teenagers hack: a personal memoir. ;login: the USENIX Association newsletter, 30(1): ??, February 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-02/pdfs/teenagers.
pdf.

Allen:2002:NMS

[All02a]  Jeff R. Allen.  Network

management, system performance tuning. In USE-NIX [USE02e], page 355. ISBN 1-880446-00-6. LCCN QA76.8.U65 U84 2002. URL

http://www.usenix.org/publications/
library/proceedings/usenix02/
tech/techonefile.html. Un-

published guru session talk, 2002 USENIX Annual Technical Conference, June 10-15, 2002, Monterey Conference Center, Monterey, CA.

Allison:2002:ALP

[All02b] Jared Allison. Automated log processing. ;login: the USENIX Association newsletter, 27(6):??, December 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-12/pdfs/allison.pdf.

Allman:2002:TOSa

[All02c] Eric Allman. Taking an open source project to market. In USENIX [USE02e], page ?? ISBN 1-880446-00-6. LCCN QA76.8.U65

U84 2002. URL http://www.

                                        usenix.org/publications/
                                        library/proceedings/usenix02/
                                        tech/techonefile.html. Un-

published invited talk, 2002 USENIX Annual Technical Conference, June 10-15, 2002, Monterey Conference Center, Monterey, CA.

Allman:2002:TOSb

[All02d] Eric Allman. Taking an open source project to market: a parable of sendmail. In USE-NIX [USE02a], pages viii +

151. ISBN 1-880446-02-2. LCCN QA76.76.O63 B736

                                        2002. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/bsdcon02/

tech.html. Unpublished invited talk, BSDCON2002: Growing the BSD Community, February 11–14, 2002, Cathedral Hill Hotel, San Francisco, CA.

Amatangelo:2000:UPJ

[Ama00] Lee “Leonardo” Amatangelo. Unleashing the power of JumpStart: a new technique for disaster recovery, cloning, or snapshotting a Solaris system. In USENIX [USE00g], page ?? ISBN 1-880446-13-8. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/lisa2000/
                                        amatangelo.html.

Amer:2001:MHD

[Ame01] Ahmed Amer. Mobile hoarding and dynamic grouping.

;login: the USENIX Association newsletter, 26(2): ??, April 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-04/pdfs/amer.pdf.

Anderson:2000:XLA

[And00]  Stuart Anderson.  The

XFree86 loadable architecture & whazzup with XFree86?, 2000. URL http:/

/www.usenix.org/publications/
library/proceedings/als2000/

. Unpublished talk at Annual Linux Showcase 2000.

Andersson:2001:KDJ

[And01]  Johan Andersson.  Kaf-

femik — a distributed JVM featuring a single address space. In USENIX [USE01g], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38

J42 2001. URL http://www.
usenix.org/publications/
library/proceedings/jvm01/
JVM_wips/S09.pdf.

Andress:2005:INI

[And05] Jason G. Andress. IPv6:

The next Internet protocol. ;login: the USENIX Association newsletter, 30(2): ??, April 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-04/pdfs/andress0504.
pdf.

Anderson:2006:WCM

[And06] Paul Anderson. Why configuration management is crucial. ;login: the USENIX Association newsletter, 31(1): ??, February 2006. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2006-02/openpdfs/
                                        anderson.pdf.

Anderson:2009:PVI

[And09] Paul Anderson. Programming the virtual infrastructure. ;login: the USENIX Association newsletter, 34(1): ??, February 2009. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        february-2009-volume-34-
                                        number-1/programming-virtual-
                                        infrastructure.

Annis:2000:PSU

[Ann00] William S. Annis. Pithy sayings for the UNIX sysadmin. ;login: the USENIX Association newsletter, 25(5): ??, August 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-8/features/pithy.html.

Anonymous:2000:ASA

[Ano00a] Anonymous. Announcement: 14th Systems Administration Conference (LISA 2000). ;login: the USENIX Association newsletter, 25(1):??, February 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/events/lisa00.

Anonymous:2000:AUAa

[Ano00b] Anonymous. Announcement:

2000 USENIX Annual Technical Conference Program. ;login: the USENIX Association newsletter, 25(2): ??, April 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/events/usenix2000.

Anonymous:2000:AUAb

[Ano00c] Anonymous. Announcement:

2001 USENIX Annual Conference Call. ;login: the USE-NIX Association newsletter, 25(4):??, July 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/events/usenix01.

Anonymous:2000:AUS

[Ano00d] Anonymous. Announcement:

3rd USENIX Symposium on Internet Technologies and Systems (USITS ’01). ;login: the USENIX Association newsletter, 25(3):??, June 2000. CODEN LOGNEM. ISSN 1044-6397. URL http:

//www.usenix.org/events/
usits01.

Anonymous:2000:AAL

[Ano00e] Anonymous. Announcement:

4th Annual Linux Showcase and Conference, Atlanta. ;login: the USENIX Association newsletter, 25(2):??, April 2000. CODEN LOGNEM. ISSN 1044-6397. URL http:

                                        //www.linuxshowcase.org.

Anonymous:2000:AUC

[Ano00f] Anonymous. Announcement: 6th USENIX Conference on Object-Oriented Technologies and Systems (COOTS ’01). ;login: the USENIX Association newsletter, 25(2): ??, April 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/events/coots01.

Anonymous:2000:AJV

[Ano00g] Anonymous. Announcement: Java Virtual Machine Research and Technology Symposium (JVM ’01). ;login: the USENIX Association newsletter, 25(3):??, June 2000. CODEN LOGNEM. ISSN 1044-6397. URL http:

                                        //www.usenix.org/events/
                                        jvm01.

Anonymous:2000:CRU

[Ano00h] Anonymous. Conference report: USENIX 2000 Annual Technical Conference. ;login: the USENIX Association newsletter, 25(6): ??, October 2000. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/usenix2000/
                                        loginconfrpts.pdf.

Anonymous:2000:ER

[Ano00i] Anonymous. Election results. ;login: the USE-NIX Association newsletter, 25(3):??, June 2000. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2000-6/usenixnews.
html#election.

Anonymous:2000:I

[Ano00j] Anonymous. In this issue. . ..

;login: the USENIX Association newsletter, 25(1):??, February 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-2/inthisissue.html.

Anonymous:2000:LEa

[Ano00k] Anonymous. Letters to the Editor. ;login: the USENIX Association newsletter, 25(1): ??, February 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-2/letters.html.

Anonymous:2000:LEb

[Ano00l] Anonymous. Letters to the Editor. ;login: the USE-NIX Association newsletter, 25(2):??, April 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-4/letters.html.

Anonymous:2000:LEc

[Ano00m] Anonymous. Letters to the Editor. ;login: the USENIX Association newsletter, 25(6): ??, October 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-10/letterstotheeditor.
pdf.

Anonymous:2000:LEd

[Ano00n] Anonymous. Letters to the Editor. ;login: the USENIX Association newsletter, 25(8): ??, December 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-12/letters.pdf.

Anonymous:2000:MN

[Ano00o] Anonymous. Miscellaneous news. ;login: the USENIX Association newsletter, 25 (4):??, July 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-7/sagenews.html#news.

Anonymous:2000:NAM

[Ano00p] Anonymous. Notice of annual meeting. ;login: the USE-NIX Association newsletter, 25(3):??, June 2000. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2000-6/usenixnews.
html#annual.

Anonymous:2000:SNa

[Ano00q] Anonymous. SAGE news.

;login: the USENIX Association newsletter, 25(6): ??, October 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-10/sagenews.pdf.

Anonymous:2000:SNb

[Ano00r] Anonymous. SAGE news.

;login: the USENIX Association newsletter, 25(8):??, December 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-12/sagenews.pdf.

Anonymous:2000:SCT

[Ano00s]  Anonymous.  Sun comes

through for USENIX. ;login: the USENIX Association newsletter, 25(5):??, August 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-8/usenixnews.html#
sun.

Anonymous:2000:USN

[Ano00t]  Anonymous.  USENIX

and SAGE news. ;login: the USENIX Association newsletter, 25(7):??, November 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-11/news.pdf.

Anonymous:2000:UNa

[Ano00u] Anonymous. USENIX news. ;login: the USENIX Association newsletter, 25(6): ??, October 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-10/usenixnews.pdf.

Anonymous:2000:UNb

[Ano00v] Anonymous. USENIX news. ;login: the USENIX Association newsletter, 25(8):??, December 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-12/unews.pdf.

Anonymous:2001:BRT

[Ano01a] Anonymous. Book review: Three book reviews. ;login: the USENIX Association newsletter, 26(4): ??, July 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-07/bookreviews.pdf.

Anonymous:2001:CRUc

[Ano01b] Anonymous. Conference report: 10th USENIX Security Symposium. ;login: the USENIX Association newsletter, 26(7):??, November 2001. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
library/proceedings/sec01/
confrpts.pdf.

Anonymous:2001:CRS

[Ano01c] Anonymous. Conference report: 14th Systems Administration Conference (LISA 2000). ;login: the USE-NIX Association newsletter, 26(2):??, April 2001. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
library/proceedings/lisa2000/
confrpts.pdf.

Anonymous:2001:CRI

[Ano01d] Anonymous. Conference report: 1st IEEE/ACM International Symposium on Cluster Computing and the Grid. ;login: the USENIX Association newsletter, 26(5): ??, August 2001. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:2001:CRJ

[Ano01e] Anonymous. Conference report: 1st Java VM Virtual Machine Research and Technology Symposium (JVM ’01). ;login: the USENIX Association newsletter, 26(5): ??, August 2001. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:2001:CRW

[Ano01f] Anonymous. Conference report: 5th Workshop on Distributed Supercomputing Scalable Cluster Software. ;login: the USENIX Association newsletter, 26(5):??, August 2001. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:2001:CRG

[Ano01g] Anonymous. Conference report: GUADEC 2001. ;login: the USENIX Association newsletter, 26(5):??, August 2001. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:2001:CRL

[Ano01h] Anonymous. Conference report: Large-Scale Computing Workshop. ;login: the USE-NIX Association newsletter, 26(5):??, August 2001. CO-DEN LOGNEM. ISSN 1044-6397.

Anonymous:2001:CRA

[Ano01i] Anonymous. Conference report: The 13th Annual Computer Security Incident Handling Conference (FIRST). ;login: the USENIX Association newsletter, 26(6): ??, October 2001. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2001-10/openpdfs/
                                        FIRSTconfrpts.pdf.

Anonymous:2001:CRUa

[Ano01j] Anonymous. Conference report: The 3rd USENIX Symposium on Internet Technologies and Systems (USITS ’01). ;login: the USENIX Association newsletter, 26(5): ??, August 2001. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:2001:CRUb

[Ano01k] Anonymous. Conference report: USENIX 2001 Annual Technical Conference. ;login: the USENIX Association newsletter, 26(6): ??, October 2001. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
library/proceedings/usenix01/
confrpts.pdf.

Anonymous:2001:LEa

[Ano01l] Anonymous. Letters to the Editor. ;login: the USENIX Association newsletter, 26(1): ??, February 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-02/letters.pdf.

Anonymous:2001:LEb

[Ano01m] Anonymous. Letters to the editor. ;login: the USE-NIX Association newsletter, 26(3):??, June 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-06/letters.pdf.

Anonymous:2001:LEc

[Ano01n] Anonymous. Letters to the Editors. ;login: the USE-NIX Association newsletter, 26(5):??, August 2001. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2001-08/openpdfs/
letters.pdf.

Anonymous:2001:LEd

[Ano01o] Anonymous. Letters to the Editors. ;login: the USE-NIX Association newsletter, 26(6):??, October 2001. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2001-10/openpdfs/
                                        letters.pdf.

Anonymous:2001:LEe

[Ano01p] Anonymous. Letters to the Editors. ;login: the USENIX Association newsletter, 26(8): ??, December 2001. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2001-12/openpdfs/
                                        letters.pdf.

Anonymous:2001:SNa

[Ano01q] Anonymous. SAGE news. ;login: the USENIX Association newsletter, 26(1):??, February 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-02/sagenews.pdf.

Anonymous:2001:SNb

[Ano01r] Anonymous. SAGE news. ;login: the USENIX Association newsletter, 26(2): ??, April 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-04/sagenews.pdf.

Anonymous:2001:SNc

[Ano01s] Anonymous. SAGE news.

;login: the USENIX Association newsletter, 26(3): ??, June 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-06/sagenews.pdf.

Anonymous:2001:SNd

[Ano01t] Anonymous. SAGE news.

;login: the USENIX Association newsletter, 26(5): ??, August 2001. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2001-08/openpdfs/
sagenews.pdf.

Anonymous:2001:UNa

[Ano01u] Anonymous. USENIX news.

;login: the USENIX Association newsletter, 26(1):??, February 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-02/usenixnews.pdf.

Anonymous:2001:UNb

[Ano01v] Anonymous. USENIX news.

;login: the USENIX Association newsletter, 26(2): ??, April 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-04/usenixnews.pdf.

Anonymous:2001:UNc

[Ano01w] Anonymous. USENIX news. ;login: the USENIX Association newsletter, 26(3): ??, June 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-06/usenixnews.pdf.

Anonymous:2001:UNd

[Ano01x] Anonymous. USENIX news. ;login: the USENIX Association newsletter, 26(4): ??, July 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-07/usenixnews.pdf.

Anonymous:2001:UNe

[Ano01y] Anonymous. USENIX news. ;login: the USENIX Association newsletter, 26(5): ??, August 2001. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2001-08/openpdfs/
usenixnews.pdf.

Anonymous:2001:UNf

[Ano01z] Anonymous. USENIX news. ;login: the USENIX Association newsletter, 26(6): ??, October 2001. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2001-10/openpdfs/
usenixnews.pdf.

Anonymous:2001:UNg

[Ano01-27] Anonymous. USENIX news.

;login: the USENIX Association newsletter, 26(7): ??, November 2001. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2001-11/openpdfs/
usenixnews.pdf.

Anonymous:2001:UNh

[Ano01-28] Anonymous. USENIX news.

;login: the USENIX Association newsletter, 26(8): ??, December 2001. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2001-12/openpdfs/
usenixnews.pdf.

Anonymous:2002:CRUb

[Ano02a] Anonymous. Conference report: 11th USENIX Security Symposium. ;login: the USENIX Association newsletter, 27(6):??, December 2002. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
library/proceedings/sec02/
confrpts.pdf.

Anonymous:2002:CRLb

[Ano02b] Anonymous. Conference report: 2002 Linux Kernel Developers Summit. ;login: the USENIX Association newsletter, 27(5):??, October 2002. CODEN LOGNEM.

ISSN 1044-6397. URL http:

//www.usenix.org/events/
kernel02/2002Summit.pdf.

Anonymous:2002:CRUa

[Ano02c] Anonymous. Conference report: 2002 USENIX Annual Technical Conference. ;login: the USENIX Association newsletter, 27(5): ??, October 2002. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
library/proceedings/usenix02/
USENIX2002.pdf.

Anonymous:2002:CRJ

[Ano02d] Anonymous. Conference report: 2nd Java Virtual Machine Research and Technology Symposium. ;login: the USENIX Association newsletter, 27(5):??, October 2002. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
library/proceedings/javavm02/
JavaVM2002.pdf.

Anonymous:2002:CRB

[Ano02e] Anonymous. Conference report: BSDCon 2002. ;login: the USENIX Association newsletter, 27(3): ??, June 2002. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
library/proceedings/bsdcon02/
confrpts.pdf.

Anonymous:2002:CRC

[Ano02f] Anonymous. Conference report: Conference on File and Storage Technologies (FAST ’02). ;login: the USENIX Association newsletter, 27 (2):??, April 2002. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
library/proceedings/fast02/
fast.pdf.

Anonymous:2002:CRLa

[Ano02g] Anonymous. Conference report: LISA 2001 — The 15th Systems Administration Conference. ;login: the USE-NIX Association newsletter, 27(1):??, February 2002. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
library/proceedings/lisa2001/
lisa2001confrpts.pdf.

Anonymous:2002:CRS

[Ano02h] Anonymous. Conference report: SANE 2002. ;login: the USENIX Association newsletter, 27(4):??, August 2002. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2002-08/openpdfs/
SANEconf.pdf.

Anonymous:2002:LE

[Ano02i] Anonymous. Letters to the editors. ;login: the USE-NIX Association newsletter, 27(1):??, February 2002. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2002-02/openpdfs/
                                        letters.pdf.

Anonymous:2002:SNa

[Ano02j] Anonymous. SAGE news. ;login: the USENIX Association newsletter, 27(1): ??, February 2002. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2002-02/openpdfs/
                                        sagenews.pdf.

Anonymous:2002:SNb

[Ano02k] Anonymous. SAGE news. ;login: the USENIX Association newsletter, 27(2): ??, April 2002. CODEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2002-04/openpdfs/
                                        sagenews.pdf.

Anonymous:2002:SNc

[Ano02l] Anonymous. SAGE news. ;login: the USENIX Association newsletter, 27(3): ??, June 2002. CODEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2002-06/openpdfs/
                                        sagenews.pdf.

Anonymous:2002:SNd

[Ano02m] Anonymous. SAGE news. ;login: the USENIX Association newsletter, 27(4): ??, August 2002. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2002-08/openpdfs/
sagenews.pdf.

Anonymous:2002:SNe

[Ano02n] Anonymous. SAGE news.

;login: the USENIX Association newsletter, 27(5): ??, October 2002. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2002-10/openpdfs/
sagenews.pdf.

Anonymous:2002:SR

[Ano02o] Anonymous. Standards reports. ;login: the USE-NIX Association newsletter, 27(2):??, April 2002. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
standards/52.standards.
pdf.

Anonymous:2002:UNa

[Ano02p] Anonymous. USENIX news.

;login: the USENIX Association newsletter, 27(1): ??, February 2002. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2002-02/openpdfs/
usenixnews.pdf.

Anonymous:2002:UNb

[Ano02q] Anonymous. USENIX news. ;login: the USENIX Association newsletter, 27(2): ??, April 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-04/openpdfs/usenix.
pdf.

Anonymous:2002:UNc

[Ano02r] Anonymous. USENIX news. ;login: the USENIX Association newsletter, 27(3): ??, June 2002. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2002-06/openpdfs/
usenixnews.pdf.

Anonymous:2002:UNd

[Ano02s] Anonymous. USENIX news. ;login: the USENIX Association newsletter, 27(4): ??, August 2002. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2002-08/openpdfs/
usenixnews.pdf.

Anonymous:2002:UNe

[Ano02t] Anonymous. USENIX news. ;login: the USENIX Association newsletter, 27(5): ??, October 2002. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2002-10/openpdfs/
usenixnews.pdf.

Anonymous:2002:UNf

[Ano02u] Anonymous. USENIX news.

;login: the USENIX Association newsletter, 27(6):??, December 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-08/openpdfs/usenix.
pdf.

Anonymous:2003:CRUd

[Ano03a] Anonymous. Conference report: 12th USENIX Security Symposium. ;login: the USENIX Association newsletter, 28(6):??, December 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-12/openpdfs/sec03.
pdf.

Anonymous:2003:CRA

[Ano03b] Anonymous. Conference report: 15th Annual FIRST Conference. ;login: the USE-NIX Association newsletter, 28(5):??, October 2003. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2003-10/openpdfs/
first.pdf.

Anonymous:2003:CRE

[Ano03c]  Anonymous.  Conference

report: 2003 European Tcl/Tk User Meeting. ;login: the USENIX Association newsletter, 28(4):??, August 2003. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2003-08/openpdfs/
tcl_Tkuser.pdf.

Anonymous:2003:CRS

[Ano03d] Anonymous. Conference report: 2003 MIT Spam Conference. ;login: the USE-NIX Association newsletter, 28(3):??, June 2003. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2003-06/openpdfs/
spam.pdf.

Anonymous:2003:CRUb

[Ano03e] Anonymous. Conference report: 2nd USENIX Conference on File and Storage Technologies. ;login: the USENIX Association newsletter, 28(3): ??, June 2003. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2003-06/openpdfs/
FASTconf.pdf.

Anonymous:2003:CRUa

[Ano03f] Anonymous. Conference report: 4th USENIX Symposium on Internet Technologies and Systems (USITS ’03). ;login: the USENIX Association newsletter, 28 (3):??, June 2003. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.

usenix.org/publications/
login/2003-06/openpdfs/
USITSconf.pdf.

Anonymous:2003:CRB

[Ano03g] Anonymous. Conference report: BSDCon ’03. ;login: the USENIX Association newsletter, 28(6):??, December 2003. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2003-12/openpdfs/
bsdcon03.pdf.

Anonymous:2003:CRH

[Ano03h]  Anonymous.  Conference

report: HotOS-IX. ;login: the USENIX Association newsletter, 28(4):??, August 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-08/openpdfs/HotOS-
IX.pdf.

Anonymous:2003:CRL

[Ano03i] Anonymous. Conference report: LISA XVI. ;login: the USENIX Association newsletter, 28(1):??, February 2003. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2003-02/openpdfs/
confrepts.pdf.

Anonymous:2003:CRM

[Ano03j] Anonymous. Conference report: MobiSys 2003. ;login: the USENIX Association newsletter, 28(4):??, August 2003. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2003-08/openpdfs/
mobysis03.pdf.

Anonymous:2003:CRUc

[Ano03k] Anonymous. Conference report: USENIX Annual Technical Conference 2003. ;login: the USENIX Association newsletter, 28(5):??, October 2003. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2003-10/openpdfs/
usenix03.pdf.

Anonymous:2003:LE

[Ano03l] Anonymous. Letters to the Editor. ;login: the USE-NIX Association newsletter, 28(1):??, February 2003. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2003-02/openpdfs/
letters.pdf.

Anonymous:2003:SN

[Ano03m] Anonymous. SAGE news. ;login: the USENIX Association newsletter, 28(1): ??, February 2003. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2003-02/openpdfs/
sagenews.pdf.

Anonymous:2003:UN

[Ano03n] Anonymous. USENIX news.

;login: the USENIX Association newsletter, 28(1): ??, February 2003. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2003-02/openpdfs/
usenixnews.pdf.

Anonymous:2003:UNUa

[Ano03o] Anonymous. USENIX news:

2003 USENIX Nominating Committee; summary of USENIX Board of Directors meeting; USENIX Association’s financial report for 2002. ;login: the USE-NIX Association newsletter, 28(4):??, August 2003. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2003-08/openpdfs/
usenixnews.pdf.

Anonymous:2003:UNUb

[Ano03p] Anonymous. USENIX news:

2004 USENIX Nominating Committee; membership news. ;login: the USENIX Association newsletter, 28 (5):??, October 2003. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2003-10/openpdfs/
usenixnews.pdf.

Anonymous:2004:SFA

[Ano04a] Anonymous. 2004 STUG and FLAME Awards go to M.

Douglas McIlroy. ;login: the USENIX Association newsletter, 29(4):??, August 2004. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:2004:BRR

[Ano04b] Anonymous. Book review: Raphael C. Malveau and Thomas J. Mowbray, Software Architect Bootcamp, 2nd ed. ;login: the USE-NIX Association newsletter, 29(3):??, June 2004. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2004-06/openpdfs/
reviews.pdf.

Anonymous:2004:CRUc

[Ano04c] Anonymous. Conference report: 13th USENIX Security Symposium (Security ’04). ;login: the USENIX Association newsletter, 29(6): ??, December 2004. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2004-12/openpdfs/
sec04reports.pdf.

Anonymous:2004:CRL

[Ano04d] Anonymous. Conference report: 17th Large Installation Systems Administration Conference (LISA ’03), October 26–31, 2003. ;login: the USE-NIX Association newsletter, 29(1):??, February 2004. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.

usenix.org/publications/
login/2004-02/openpdfs/
lisa03.pdf.

Anonymous:2004:CRUa

[Ano04e] Anonymous. Conference report: 3rd USENIX Conference on File and Storage Technologies (FAST ’04). ;login: the USENIX Association newsletter, 29(3): ??, June 2004. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2004-06/openpdfs/
fast04reports.pdf.

Anonymous:2004:CRV

[Ano04f]  Anonymous.  Conference

report: 3rd Virtual Machine Research and Technology Symposium (VM ’04). ;login: the USENIX Association newsletter, 29(4): ??, August 2004. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2004-08/openpdfs/
vm04reports.pdf.

Anonymous:2004:CRC

[Ano04g] Anonymous. Conference report: Cybersecurity, Research, and Disclosure Conference, November 22, 2003. ;login: the USENIX Association newsletter, 29(1): ??, February 2004. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/

                                        login/2004-02/openpdfs/
                                        stanford.pdf.

Anonymous:2004:CRF

[Ano04h] Anonymous. Conference report: First Symposium on Networked Systems Design and Implementation (NSDI ’04). ;login: the USENIX Association newsletter, 29 (3):??, June 2004. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2004-06/openpdfs/
                                        nsdi04reports.pdf.

Anonymous:2004:CRUb

[Ano04i] Anonymous. Conference report: USENIX Annual Technical Conference (USENIX ’04). ;login: the USENIX Association newsletter, 29 (5):??, October 2004. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2004-10/openpdfs/
                                        atc04reports.pdf.

Anonymous:2004:LEa

[Ano04j] Anonymous. Letter to the Editor. ;login: the USE-NIX Association newsletter, 29(4):??, August 2004. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2004-08/openpdfs/
                                        mail.pdf.

Anonymous:2004:LEc

[Ano04k] Anonymous. Letter to the Editor. ;login: the USENIX Association newsletter, 29(6): ??, December 2004. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2004-12/openpdfs/letter.
pdf.

Anonymous:2004:LEb

[Ano04l] Anonymous. Letters to the Editor. ;login: the USE-NIX Association newsletter, 29(5):??, October 2004. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2004-10/openpdfs/
mail.pdf.

Anonymous:2004:RNC

[Ano04m] Anonymous. Report from the Nominating Committee. ;login: the USENIX Association newsletter, 29(1):??, February 2004. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2004-02/openpdfs/report.
pdf.

Anonymous:2004:UC

[Ano04n] Anonymous. USACO correspondence. ;login: the USE-NIX Association newsletter, 29(4):??, August 2004. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/

                                        login/2004-08/openpdfs/
                                        usacomail.pdf.

Anonymous:2004:UNa

[Ano04o] Anonymous. USENIX notes. ;login: the USENIX Association newsletter, 29(4): ??, August 2004. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2004-08/openpdfs/unotes.
                                        pdf.

Anonymous:2004:UNb

[Ano04p] Anonymous. USENIX notes. ;login: the USENIX Association newsletter, 29(5): ??, October 2004. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2004-10/openpdfs/unotes.
                                        pdf.

Anonymous:2004:UNc

[Ano04q] Anonymous. USENIX notes. ;login: the USENIX Association newsletter, 29(6):??, December 2004. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2004-12/openpdfs/unotes.
                                        pdf.

Anonymous:2004:UNE

[Ano04r] Anonymous. USENIX notes: Election results; summary of the USENIX Board of Directors meetings. ;login: the USENIX Association newsletter, 29(3):

??, June 2004. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2004-06/openpdfs/USENIX.
pdf.

Anonymous:2004:WSD

[Ano04s] Anonymous. Wireless security: a discussion. ;login: the USENIX Association newsletter, 29(6):??, December 2004. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2004-12/openpdfs/
wireless.pdf.

Anonymous:2005:CRUb

[Ano05a] Anonymous. Conference report: 14th USENIX Security Symposium. ;login: the USENIX Association newsletter, 30(6):??, December 2005. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2005-12/openpdfs/
sec05summaries.pdf.

Anonymous:2005:CRLa

[Ano05b] Anonymous. Conference report: 18th Large Installation System Administration Conference (LISA ’04). ;login: the USENIX Association newsletter, 30(1):??, February 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/

                                        2005-02/openpdfs/lisa04.
                                        pdf.

Anonymous:2005:CRLb

[Ano05c] Anonymous. Conference report: 2005 Linux Kernel Developers Summit. ;login: the USENIX Association newsletter, 30(5):??, October 2005. CODEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2005-10/openpdfs/
                                        kernel05.pdf.

Anonymous:2005:CRUa

[Ano05d] Anonymous. Conference report: 2005 USENIX Annual Technical Conference. ;login: the USENIX Association newsletter, 30(4): ??, August 2005. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2005-08/openpdfs/atc05.
                                        pdf.

Anonymous:2005:CRSb

[Ano05e] Anonymous. Conference report: 2nd Symposium on Networked Systems Design and Implementation (NSDI ’05). ;login: the USENIX Association newsletter, 30(4): ??, August 2005. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2005-08/openpdfs/nsdi05.
                                        pdf.

Anonymous:2005:CRIa

[Ano05f] Anonymous. Conference report: 6th IEEE Workshop on Mobile Computing Systems and Applications (WM-CSA 2004). ;login: the USE-NIX Association newsletter, 30(4):??, August 2005. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2005-08/openpdfs/
wmcsa2004.pdf.

Anonymous:2005:CRE

[Ano05g] Anonymous. Conference report: EuroBSDCon 2004. ;login: the USENIX Association newsletter, 30(1): ??, February 2005. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2005-02/openpdfs/
eurobsdcon.pdf.

Anonymous:2005:CRH

[Ano05h]  Anonymous.  Conference

report: HotOS X: Tenth Workshop on Hot Topics in Operating Systems. ;login: the USENIX Association newsletter, 30(5):??, October 2005. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2005-10/openpdfs/
hotos05.pdf.

Anonymous:2005:CRIb

[Ano05i] Anonymous. Conference report: International Workshop on Wireless Traffic Measurements and Modeling (WitMeMo’05). ;login: the USENIX Association newsletter, 30(5):??, October 2005. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2005-10/openpdfs/
witmemo05.pdf.

Anonymous:2005:CRM

[Ano05j] Anonymous. Conference report: MobiSys 2005: The Third International Conference on Mobile Systems, Applications, and Services. ;login: the USENIX Association newsletter, 30(5): ??, October 2005. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2005-10/openpdfs/
mobisys05.pdf.

Anonymous:2005:CRSc

[Ano05k] Anonymous. Conference report: Steps to Reducing Unwanted Traffic on the Internet Workshop (SRUTI ’05). ;login: the USENIX Association newsletter, 30(5): ??, October 2005. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2005-10/openpdfs/
sruti05.pdf.

Anonymous:2005:CRSa

[Ano05l] Anonymous. Conference report: Symposium on Operating Systems Design and Implementation (OSDI ’04). ;login: the USENIX Association newsletter, 30(2): ??, April 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-04/openpdfs/osdi04.
pdf.

Anonymous:2005:CRV

[Ano05m] Anonymous. Conference report: VEE ’05: First ACM/ USENIX Conference on Virtual Execution Environments. ;login: the USENIX Association newsletter, 30(5): ??, October 2005. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2005-10/openpdfs/
vee05.pdf.

Anonymous:2005:CRW

[Ano05n] Anonymous. Conference report: Workshop on Endto-End, Sense-and-Response Systems, Applications, and Services (EESR ’05). ;login: the USENIX Association newsletter, 30(5):??, October 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-10/openpdfs/eesr05.
pdf.

Anonymous:2005:LEb

[Ano05o] Anonymous. Letter to the editor. ;login: the USE-NIX Association newsletter, 30(2):??, April 2005. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2005-04/openpdfs/
                                        letter0504.pdf.

Anonymous:2005:LEc

[Ano05p] Anonymous. Letter to the editor. ;login: the USE-NIX Association newsletter, 30(3):??, June 2005. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2005-06/openpdfs/
                                        letter0506.pdf.

Anonymous:2005:LEa

[Ano05q] Anonymous. Letters to the editor. ;login: the USE-NIX Association newsletter, 30(1):??, February 2005. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2005-02/openpdfs/
                                        letters.pdf.

Anonymous:2005:UNE

[Ano05r] Anonymous. USENIX notes: 2006 election for Board of Directors. ;login: the USENIX Association newsletter, 30(6): ??, December 2005. CODEN LOGNEM. ISSN 1044-6397. Anonymous:2005:UNUb

[Ano05s] Anonymous. USENIX notes: 2006 USENIX Nominating Committee. ;login: the USE-NIX Association newsletter, 30(5):??, October 2005. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2005-10/openpdfs/
nomcomm.pdf.

Anonymous:2005:UNA

[Ano05t]  Anonymous.  USENIX

notes: Annual awards. ;login: the USENIX Association newsletter, 30(3): ??, June 2005. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2005-06/openpdfs/
awards0506.pdf.

Anonymous:2005:UNUa

[Ano05u] Anonymous. USENIX notes:

The USENIX Association financial report for 2004. ;login: the USENIX Association newsletter, 30(5): ??, October 2005. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2005-10/openpdfs/
financials05.pdfa.

Anonymous:2006:CRF

[Ano06a] Anonymous. Conference report: FAST ’05: 4th USE-NIX Conference on File and Storage Technologies. ;login: the USENIX Association newsletter, 31(2):103– 115, April 2006. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/

                                        login/2006-04/openpdfs/
                                        fast05summaries.pdf.

Anonymous:2006:CRL

[Ano06b] Anonymous. Conference report: LISA ’05: The 19th Large Installation System Administration Conference. ;login: the USENIX Association newsletter, 31(2):81– 98, April 2006. CODEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2006-04/openpdfs/
                                        lisa05summaries.pdf.

Anonymous:2006:CRW

[Ano06c] Anonymous. Conference report: WORLDS ’05: Second Workshop on Real, Large Distributed Systems. ;login: the USENIX Association newsletter, 31(2):98– 102, April 2006. CODEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2006-04/openpdfs/
                                        worlds05summaries.pdf.

Anonymous:2006:UNU

[Ano06d] Anonymous. USENIX notes: 2006 USENIX Nominating Committee report. ;login: the USENIX Association newsletter, 31(1):??, February 2006. CODEN LOGNEM. ISSN 1044-6397.

Anonymous:2006:UNF

[Ano06e] Anonymous. USENIX notes: Fund to establish the John Lions Chair in Operating Systems at the University of New South Wales. ;login: the USENIX Association newsletter, 31(2): 80, April 2006. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2006-04/openpdfs/lions.
pdf.

Amiri:2000:DFP

[APGG00] Khalil Amiri, David Petrou, Gregory R. Ganger, and Garth A. Gibson. Dynamic function placement for dataintensive cluster computing. In USENIX [USE00a], pages 307–322. ISBN 1-880446-22-7. LCCN ???? URL http:

//www.usenix.org/events/
usenix2000/general/amiri.
html.

Ascher:2000:MCP

[APH00] David Ascher, Eric Promislow, and Dick Hardt. Mozilla as a cross-platform application development framework. In USENIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN ????

URL http://www.usenix.
org/publications/library/
proceedings/als2000/ascher.
html.

Appelman:2004:CRD

[App04a]  Dan Appelman.  Califor-

nia requires disclosure of database security breaches. ;login: the USENIX Association newsletter, 29(3):

??, June 2004. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2004-06/pdfs/appelman.pdf.

Appelman:2004:CSN

[App04b] Dan Appelman. CAN-SPAM: New federal anti-spam law sets nationwide standards for commercial email messages. ;login: the USENIX Association newsletter, 29(2): ??, April 2004. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2004-04/pdfs/Appelman.pdf.

Appelman:2004:OWS

[App04c] Dan Appelman. OPPA and Web site operators. ;login: the USENIX Association newsletter, 29(4):??, August 2004. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2004-08/pdfs/appelman.pdf.

Appelman:2005:PCL

[App05] Daniel L. Appelman. Primer on cybercrime laws. ;login: the USENIX Association newsletter, 30(4):??, August 2005. CODEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2005-08/openpdfs/
                                        appelman.pdf.

Appelman:2007:SBPa

[App07a] Daniel L. Appelman. Spam and blogs. Part 1: Spam: a balancing act. ;login: the USENIX Association newsletter, 32(2): ??, April 2007. CODEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/april-2007-volume-
32-number-2/spam-and-blogs-
part-1-spam-balancing-act.

Appelman:2007:SBPb

[App07b] Daniel L. Appelman. Spam and blogs. Part 2: Blogs, for good or ill. ;login: the USENIX Association newsletter, 32(3):??, June 2007. CODEN LOGNEM. ISSN 1044-6397. URL

https://www.usenix.org/
publications/login/june-
2007-volume-32-number-3/
spam-and-blogs-part-2-blogs-
good-or-ill.

Apthorpe:2001:PAE

[Apt01] Robert Apthorpe. A probabilistic approach to estimating computer system reliability. In USENIX [USE01a], page ?? ISBN 1-880446-05-7. LCCN ????

URL http://db.usenix.
org/publications/library/
proceedings/lisa2001/tech/
apthorpe.html.

Acharya:2000:MUP

[AR00] Anurag Acharya and Mandar Raje. MAPbox: Using parameterized behavior classes to confine untrusted applications. In USENIX [USE00i], page ?? ISBN 1-880446-18-9. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/sec2000/
acharya.html.

Arkin:2002:RAF

[Ark02a] Ofir Arkin. A remote active OS fingerprinting tool using ICMP. ;login: the USE-NIX Association newsletter, 27(2):??, April 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-04/pdfs/arkin.pdf.

Arkin:2002:STI

[Ark02b] Ofir Arkin. Security threats to IP telephony-based networks. ;login: the USENIX Association newsletter, 27(6): ??, December 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-12/pdfs/arkin.pdf.

Arkin:2005:DPN

[Ark05]  Ofir Arkin.  Demystify-

ing passive network discovery and monitoring systems. ;login: the USENIX Association newsletter, 30(3): ??, June 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/

2005-06/pdfs/arkin0506.
pdf.

Anderson:2000:LSL

[AS00] Paul Anderson and Alastair Scobie. Large scale Linux configuration with LCFG. In USENIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/als2000/
anderson.html.

Aron:2000:SCA

[ASDZ00] Mohit Aron, Darren Sanders, Peter Druschel, and Willy Zwaenepoel. Scalable contentaware request distribution in cluster-based network servers. In USENIX [USE00a], pages 323–336. ISBN 1-880446-22-7. LCCN ????

URL http://www.usenix.
org/events/usenix2000/general/
aron.html.

Ashizawa:2002:MSH

[Ash02]  Neil Ashizawa.  Moni-

toring strategies for highavailability Web apps. ;login: the USENIX Association newsletter, 27(4):??, August 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-08/pdfs/sellens.pdf.

Anastasiadis:2001:MER

[ASS01] Stergios V. Anastasiadis, Kenneth C. Sevcik, and Michael Stumm. Modular and efficient resource management in the Exedra media server. In USENIX [USE01c], page ?? ISBN 1-880446-14-6. LCCN TK5105.875.I57

U96 2001. URL http://www.
usenix.org/publications/
library/proceedings/usits01/
anastasiadis.html.

Anastasiadis:2002:MTR

[ASS02] Stergios V. Anastasiadis, Kenneth C. Sevcik, and Michael Stumm. Maximizing throughput in replicated disk striping of variable bit-rate streams. In USENIX [USE02e], page ?? ISBN 1-880446-00-6. LCCN QA76.8.U65 U84 2002. URL

http://www.usenix.org/publications/
library/proceedings/usenix02/
anastaiadis.html.

Anderson:2002:SRL

[ASV^{+}02] Eric Anderson, Ram Swaminathan, Alistair Veitch, Guillermo A. Alvarez, and John Wilkes. Selecting RAID levels for disk arrays. In USENIX [USE02c], page ?? ISBN 1-880446-03-

0. LCCN QA76.9.F5 C65

2002. URL http://www.
usenix.org/publications/
library/proceedings/fast02/
andersonRAID.html.

Adamo:2005:LVO

[AT05] Massimiliano Adamo and Mauro Tablò. Linux vs. OpenBSD: a firewall performance test. ;login: the USENIX Association newsletter, 30(6):??, December 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-12/pdfs/adamo.pdf.

Ayuso:2006:NCT

[Ayu06] Pablo Neira Ayuso. Netfilter’s connection tracking system. ;login: the USE-NIX Association newsletter, 31(3):??, June 2006. CO-DEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/june-2006-volume-
31-number-3/netfilters-
connection-tracking-system.

Bonwick:2001:MVE

[BA01] Jeff Bonwick and Jonathan Adams. Magazines and Vmem: Extending the slab allocator to many CPUs and arbitrary resources. In USENIX [USE01b], page ?? ISBN 1-880446-09-X. LCCN QA76.8.U65 U84 2001. URL

http://www.usenix.org/publications/
library/proceedings/usenix01/
bonwick.html.

Babkin:2001:WTG

[Bab01] Sergei Babkin. Writing the Tk geometry managers in Tcl. ;login: the USENIX Association newsletter, 26(6): ??, October 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-10/pdfs/babkin.pdf.

Bagwell:2004:FSC

[Bag04]  Christine Bagwell.  Feng

shui for computer geeks. ;login: the USENIX Association newsletter, 29(2): ??, April 2004. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2004-04/pdfs/Bagwell.pdf.

Bahadur:2000:WSY

[Bah00] Gary Bahadur. Why should you enforce a network security policy? ;login: the USENIX Association newsletter, 25(8):??, December 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-12/pdfs/bahadur.pdf.

Ball:2000:TFN

[Bal00] Steve Ball. TkGecko: a frill-necked lizard. In USE-NIX [USE00e], page ?? ISBN 1-880446-24-3. LCCN

                                        ????  URL http://db.
                                        usenix.org/publications/
                                        library/proceedings/tcl2k/
                                        ball.html.

Baldwin:2002:LMF

[Bal02] John H. Baldwin. Locking in the multithreaded FreeBSD kernel. In USE-NIX [USE02a], pages 27– 35. ISBN 1-880446-02-2. LCCN QA76.76.O63 B736

                                        2002. URL http://www.
                                        usenix.org/publications/

library/proceedings/bsdcon02/
baldwin.html.

Balas:2003:SCG

[Bal03] Edward Balas. Sebek: Covert glass-box host analysis. ;login: the USENIX Association newsletter, 28(6):??, December 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-12/pdfs/balas.pdf.

Banga:2000:ADF

[Ban00]  Gaurav Banga.  Auto-

diagnosis of field problems in an appliance operating system. In USE-NIX [USE00a], pages 293–

306. ISBN 1-880446-22-7. LCCN ???? URL http:

//www.usenix.org/events/
usenix2000/general/banga.
html.

Barnett:2000:NNO

[Bar00] Bruce Barnett. NOOSE — Networked Object-Oriented Security Examiner. In USENIX [USE00g], page ?? ISBN 1-880446-13-8. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/lisa2000/
barnett.html.

Blake:2002:SGS

[BB02] Charles L. Blake and Steven Bauer. Simple and general statistical profiling with PCT. In USENIX [USE02e], page ?? ISBN 1-880446-00-6. LCCN QA76.8.U65

U84 2002. URL http://www.
usenix.org/publications/
library/proceedings/usenix02/
blake.html.

Burnett:2002:EGB

[BBADAD02] Nathan C. Burnett, John Bent, Andrea C. Arpaci-Dusseau, and Remzi H. Arpaci-Dusseau. Exploiting gray-box knowledge of buffer-cache management. In USENIX [USE02e], page ?? ISBN 1-880446-00-6. LCCN QA76.8.U65 U84 2002. URL

http://www.usenix.org/publications/
library/proceedings/usenix02/
burnett.html.

Barabash:2001:MAS

[BBD^{+}01] Katherine Barabash, Niv Buchbinder, Tamar Domani, Elliot K. Kolodner, Yoav Ossia, Shlomit S. Pinter, Janice Shepherd, Ron Sivan, and Victor Umansky. Mostly accurate stack scanning. In USENIX [USE01g], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38

J42 2001. URL http://www.
usenix.org/publications/
library/proceedings/jvm01/
barabash.html.

Billig:2004:PIM

[BBLo04] Andreas Billig, Susanne Busse, Andreas Leicher, and other. Platform independent model transformation based on TRIPLE. In Jacobsen [Jac04], pages 493– ?? CODEN LNCSD9. ISBN 3-540-23428-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.M54 I33 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.springerlink.
com/openurl.asp?genre=issue&
issn=0302-9743&volume=3231;
http://www.springerlink.
com/openurl.asp?genre=volume&
id=doi:10.1007/b101561.

Burch:2000:TAP

[BC00] Hal Burch and Bill Cheswick.

Tracing anonymous packets to their approximate source. In USENIX [USE00g], page ?? ISBN 1-880446-13-8. LCCN ????

URL http://www.usenix.
org/publications/library/
proceedings/lisa2000/burch.
html.

Bottomley:2001:MDS

[BC01] James Bottomley and Paul Clements. Managing distributions, from the software vendor’s perspective. In USE-NIX [USE01d], page ?? ISBN ???? LCCN ???? URL http:

//www.linuxshowcase.org/
bottomley.html.

Buttcher:2006:AFT

[BC06] Stefan Büttcher and Charles L. A. Clarke. Adding fulltext filesystem search to Linux. ;login: the USE-NIX Association newsletter, 31(3):??, June 2006. CO-DEN LOGNEM. ISSN 1044-

6397. URL https://www.

usenix.org/publications/
login/june-2006-volume-
31-number-3/adding-full-
text-filesystem-search-
linux.

Borella:2000:PIA

[BCD^{+}00] Michael S. Borella, Dave Cheriton, Steve Deering, Gabriel Montenegro, and Christoph Schuba. Panel: Internet addressing at the edge, 2000. URL http://db.

usenix.org/publications/
library/proceedings/ine2000/
technical.html. Unpub-

lished panel talk at the Special Workshop on Intelligence at the Network Edge: March 20, 2000, Holiday Inn — Fisherman’s Wharf, San Francisco, CA.

Bolosky:2000:SIS

[BCGD00] William J. Bolosky, Scott Corbin, David Goebel, and John R. Douceur. Single instance storage in Windows 2000. In USE-NIX [USE00d], page ?? ISBN 1-880446-20-0. LCCN

????  URL http://db.
usenix.org/publications/
library/proceedings/usenix-
win2000/bolosky.html.

Brown:2000:PCW

[BCH^{+}00] Michael Brown, Donny Cheung, Darrel Hankerson, Julio Lopez Hernandez, Michael Kirkup, and Alfred Menezes. PGP in constrained wireless devices. In USE-NIX [USE00i], page ??

ISBN 1-880446-18-9. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/sec2000/
brown.html.

Bellovin:2006:WPS

[BCK06] Steven M. Bellovin, Bill Cheswick, and Angelos D. Keromytis. Worm propagation strategies in an IPv6 Internet. ;login: the USENIX Association newsletter, 31(1): ??, February 2006. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2006-02/pdfs/bellovin.pdf.

Bernaschi:2007:MSC

[BCR07] Massimo Bernaschi, Francesco Casadei, and Samuele Ruco. Migration of secure connections using SockMi. ;login: the USENIX Association newsletter, 32(4):??, August 2007. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2007-volume-32-number-
4/migration-secure-connections-
using-sockmi.

Boneh:2001:MFR

[BDTW01] Dan Boneh, Xuhua Ding, Gene Tsudik, and Chi Ming Wong. A method for fast revocation of public key certificates and security capabilities. In USENIX [USE01h], page ?? ISBN 1-880446-18-9, 1-880446-07-

3. LCCN QA76.9.A25 U565

                                        2001. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/sec01/
                                        boneh.html.

Bing:2000:EUS

[BE00] Matthew Bing and Carl Erickson. Extending UNIX system logging with SHARP. In USENIX [USE00g], page ?? ISBN 1-880446-13-8. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/lisa2000/
                                        bing.html.

Blank-Edelman:2001:ICT

[BE01] David N. Blank-Edelman. If I could talk to the animals — what sysadmins can learn about diagnostic skills from another profession, 2001. URL http://db.

                                        usenix.org/publications/
                                        library/proceedings/lisa2001/

tech/. Unpublished invited talk, LISA 2001: 15th Systems Administration Conference, December 2–7, 2001, Town and Country Resort Hotel, San Diego, CA.

Blank-Edelman:2006:PPTc

[BE06a] David Blank-Edelman. Practical Perl tools: Car 10.0.0.54, where are you? ;login: the USENIX Association newsletter, 31(3):??, June 2006. CODEN LOGNEM. ISSN 1044-6397. URL

                                        https://www.usenix.org/
                                        publications/login/june-
                                        2006-volume-31-number-3/

practical-perl-tools-car-
100054-where-are-you.

Blank-Edelman:2006:PPTa

[BE06b] David Blank-Edelman. Practical Perl tools: Configuration files. ;login: the USENIX Association newsletter, 31(1): ??, February 2006. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2006-02/pdfs/blank-edelman.
pdf.

Blank-Edelman:2006:PPTf

[BE06c] David Blank-Edelman. Practical Perl tools: Give me my woobie back. ;login: the USENIX Association newsletter, 31(6):??, December 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
december-2006-volume-31-
number-6/practical-perl-
tools-give-me-my-woobie-
back.

Blank-Edelman:2006:PPTb

[BE06d] David Blank-Edelman. Practical Perl tools: Programming, ho hum. ;login: the USENIX Association newsletter, 31(2):60– 65, April 2006. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2006-04/pdfs/blank-edelman.
pdf.

Blank-Edelman:2006:PPTd

[BE06e] David Blank-Edelman. Practical Perl tools: Tie me up, tie me down (part 1). ;login: the USENIX Association newsletter, 31(4): ??, August 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2006-volume-31-number-
4/practical-perl-tools-
tie-me-tie-me-down-part-
1.

Blank-Edelman:2006:PPTe

[BE06f] David Blank-Edelman. Practical Perl tools: Tie me up, tie me down (part 2). ;login: the USENIX Association newsletter, 31(5): ??, October 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
october-2006-volume-31-
number-5/practical-perl-
tools-tie-me-tie-me-down-
part-2.

Blank-Edelman:2007:PPTc

[BE07a] David Blank-Edelman. Practical Perl tools: Impractical Perl tools. ;login: the USENIX Association newsletter, 32(3):??, June 2007. CODEN LOGNEM. ISSN 1044-6397. URL

https://www.usenix.org/
publications/login/june-
2007-volume-32-number-3/
practical-perl-tools-impractical-
perl-tools.

Blank-Edelman:2007:PPTe

[BE07b] David Blank-Edelman. Practical Perl tools: Let me draw you a picture. ;login: the USENIX Association newsletter, 32(5):??, October 2007. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
october-2007-volume-32-
number-5/practical-perl-
tools-let-me-draw-you-picture.

Blank-Edelman:2007:PPTd

[BE07c] David Blank-Edelman. Practical Perl tools: Peter Piper picked a peck of PDFs. ;login: the USENIX Association newsletter, 32(4):37– 41, August 2007. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2007-volume-32-number-
4/practical-perl-tools-
peter-piper-picked-peck-
pdfs.

Blank-Edelman:2007:PPTa

[BE07d] David Blank-Edelman. Practical Perl tools: Spawning. ;login: the USENIX Association newsletter, 32(1):??, February 2007. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
february-2007-volume-32-
number-1/practical-perl-
tools-spawning.

Blank-Edelman:2007:PPTb

[BE07e] David Blank-Edelman. Practical Perl tools: These inodes were made for walkin’. ;login: the USENIX Association newsletter, 32(2): ??, April 2007. CODEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/april-2007-volume-
32-number-2/practical-perl-
tools-these-inodes-were-
made-walkin.

Blank-Edelman:2007:PPTf

[BE07f] David N. Blank-Edelman. Practical Perl tools: Perl meets Nmap and p0f. ;login: the USENIX Association newsletter, 32(6):??, December 2007. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
december-2007-volume-32-
number-6/practical-perl-
tools-perl-meets-nmap-and-
p0f.

Blank-Edelman:2008:PPTc

[BE08a] David N. Blank-Edelman. Practical Perl tools: a little place for your stuff. ;login: the USENIX Association newsletter, 33(3):??, June 2008. CODEN LOGNEM. ISSN 1044-6397. URL

https://www.usenix.org/
publications/login/june-
2008-volume-33-number-3/
practical-perl-tools-little-
place-your-stuff.

Blank-Edelman:2008:PPTe

[BE08b] David N. Blank-Edelman.

Practical Perl tools: Attachments. ;login: the USENIX Association newsletter, 33(5): ??, October 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
october-2008-volume-33-
number-5/practical-perl-
tools-attachments.

Blank-Edelman:2008:PPTb

[BE08c] David N. Blank-Edelman.

Practical Perl tools: Back in timeline. ;login: the USE-NIX Association newsletter, 33(2):??, April 2008. CO-DEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/april-2008-volume-
33-number-2/practical-perl-
tools-back-timeline.

Blank-Edelman:2008:PPTd

[BE08d] David N. Blank-Edelman.

Practical Perl tools: Hi-ho the merry-o, debugging we Will go. ;login: the USENIX Association newsletter, 33(4): ??, August 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2008-volume-33-number-
4/practical-perl-tools-
hi-ho-merry-o-debugging-
we.

Blank-Edelman:2008:PPTf

[BE08e] David N. Blank-Edelman.

Practical Perl tools: This column is password-protected. ;login: the USENIX Association newsletter, 33(6):??, December 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
december-2008-volume-33-
number-6/practical-perl-
tools-column-password-protected.

Blank-Edelman:2008:PPTa

[BE08f] David N. Blank-Edelman. Practical Perl tools: Why I live at the P.O. ;login: the USENIX Association newsletter, 33(1):??, February 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
february-2008-volume-33-
number-1/practical-perl-
tools-why-i-live-po.

Blank-Edelman:2009:PPTa

[BE09a] David N. Blank-Edelman. Practical Perl tools: Be the browser. ;login: the USENIX Association newsletter, 34(1): ??, February 2009. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
february-2009-volume-34-
number-1/practical-perl-
tools-be-browser.

Blank-Edelman:2009:PPTc

[BE09b] David N. Blank-Edelman. Practical Perl tools: En tableau. ;login: the USE-NIX Association newsletter, 34(3):??, June 2009. CO-DEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/june-2009-volume-
34-number-3/practical-perl-
tools-en-tableau.

Blank-Edelman:2009:PPTf

[BE09c] David N. Blank-Edelman.

Practical Perl tools: Essential techniques. ;login: the USENIX Association newsletter, 34(6):??, December 2009. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
december-2009-volume-34-
number-6/practical-perl-
tools-essential-techniques.

Blank-Edelman:2009:PPTb

[BE09d] David N. Blank-Edelman.

Practical Perl tools: Polymourphously versioned. ;login: the USENIX Association newsletter, 34(2): ??, April 2009. CODEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/april-2009-volume-
34-number-2/practical-perl-
tools-polymourphously-versioned.

Blank-Edelman:2009:PPTd

[BE09e] David N. Blank-Edelman.

Practical Perl tools: Scratch the Webapp itch with CGI::Application, part 1. ;login: the USENIX Association newsletter, 34(4): ??, August 2009. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        august-2009-volume-34-number-
                                        4/practical-perl-tools-
                                        scratch-webapp-itch.

Blank-Edelman:2009:PPTe

[BE09f] David N. Blank-Edelman. Practical Perl tools: Scratch the Webapp itch with CGI::Application, part 2. ;login: the USENIX Association newsletter, 34(5): ??, October 2009. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        october-2009-volume-34-
                                        number-5/practical-perl-
                                        tools-scratch-webapp-itch.

Beal:2001:ELI

[Bea01a]  David Beal.  Embedding

Linux: Issues, questions, and a few answers. In USENIX [USE01d], page ?? ISBN ???? LCCN ???? URL http:

                                        //www.linuxshowcase.org/
                                        beal.html.

Bear:2001:SVS

[Bea01b] Greg Bear. Slime vs. silicon. In USENIX [USE01a], page ?? ISBN 1-880446-05-7. LCCN ????

                                        URL http://db.usenix.
                                        org/publications/library/
                                        proceedings/lisa2001/tech/

. Unpublished keynote ad-

dress.

Beazley:2001:EER

[Bea01c] David M. Beazley. An embedded error recovery and debugging mechanism for scripting language extensions. In USENIX [USE01b], page ?? ISBN 1-880446-09-X. LCCN QA76.8.U65

U84 2001. URL http://www.
usenix.org/publications/
library/proceedings/usenix01/
beazley.html.

Beazley:2009:PGB

[Bea09] David Beazley. Python 3:

The good, the bad, and the ugly. ;login: the USE-NIX Association newsletter, 34(2):??, April 2009. CO-DEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/april-2009-volume-
34-number-2/python-3-good-
bad-and-ugly.

Becker:2001:HBS

[Bec01] Don Becker. Home Beowulf systems, 2001. URL http:

//www.linuxshowcase.org/

tech.html. Unpublished invited talk, 5th Annual Linux Showcase and Conference, November 5–10, Oakland, CA.

Bellovin:2002:IWD

[Bel02] Steve Bellovin. The IETF, or, where do all those RFCs come from, anyway?,

2002. URL http://www.
usenix.org/publications/
library/proceedings/usenix02/
tech/bellovin/ietf-usenix.

htm. Unpublished invited talk, 2002 USENIX Annual Technical Conference, June 10-15, 2002, Monterey Conference Center, Monterey, CA.

Berkowitz:2000:NPF

[Ber00a] Gale Berkowitz. New projects funded. ;login: the USENIX Association newsletter, 25(1): ??, February 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-2/projects.html.

Berkowitz:2000:SU

[Ber00b] Gale Berkowitz. SAGE update. ;login: the USENIX Association newsletter, 25(1): ??, February 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-2/update.html.

Bernstein:2000:XDP

[Ber00c] Rocky Bernstein. xps — dynamic process tree watching under X. In USE-NIX [USE00g], page ?? ISBN 1-880446-13-8. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/lisa2000/
                                        bernstein.html.

Berry:2002:KST

[Ber02]  Robert Berry.  Keynote:

Stop thinking outside the box. there is no box, 2002.

                                        URL http://www.usenix.
                                        org/publications/library/
                                        proceedings/javavm02/keynote/

index.htm. Invited talk, not published in proceedings.

Best:2000:JLH

[Bes00] Steve Best. JFS log: How the journaled file system performs logging. In USE-NIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/als2000/
best.html.

Bryant:2002:FPS

[BFH02] Ray Bryant, Ruth Forester, and John Hawkes. Filesystem performance and scalability in Linux 2.4.17. In USENIX [USE02d], page ?? ISBN 1-880446-01-4. LCCN QA76.8.U65 P765 2002.

URL http://www.usenix.
org/publications/library/
proceedings/usenix02/tech/
freenix/bryant.html.

Bohra:2001:MFF

[BG01] Aniruddha Bohra and Eran Gabber. Are mallocs free of fragmentation? In USENIX [USE01f], page ?? ISBN 1-880446-10-3. LCCN QA76.8.U65 U84 2001. URL

http://www.usenix.org/publications/
library/proceedings/usenix01/
freenix01/bohra.html.

Bernaschi:2000:ELK

[BGM00] Massimo Bernaschi, Emanuele Gabrielli, and Luigi V. Mancini. Enhancements to the Linux kernel for blocking buffer overflow based attacks. In USE-NIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/als2000/
bernaschi.html.

Bairavasundaram:2008:DCS

[BGS^{+}08] Lakshmi Bairavasundaram, Garth Goodson, Bianca Schroeder, Andrea Arpaci-Dusseau, and Remzi Arpaci-Dusseau. Data corruption in the storage stack: a closer look. ;login: the USENIX Association newsletter, 33(3):??, June 2008. CODEN LOGNEM. ISSN 1044-6397. URL

https://www.usenix.org/
publications/login/june-
2008-volume-33-number-3/
data-corruption-storage-
stack-closer-look.

Brustoloni:2000:SRP

[BGSS00] José Brustoloni, Eran Gabber, Abraham Silberschatz, and Amit Singh. Signaled receiver processing. In USE-NIX [USE00a], pages 211–

223. ISBN 1-880446-22-7. LCCN ???? URL http:

//www.usenix.org/events/
usenix2000/general/brustoloni.
html.

Beck:2000:FSD

[BH00a] Robert Beck and Steve Holstead. FOKSTRAUT and Samba — dealing with authentication and performance issues on a largescale Samba service. In USENIX [USE00g], page ?? ISBN 1-880446-13-8. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/lisa2000/
beck.html.

Bryant:2000:LHI

[BH00b] Ray Bryant and John Hawkes.

Lockmeter: Highly informative instrumentation for spin locks in the Linux kernel. In USENIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/als2000/
bryantlock.html.

Bryant:2000:SSC

[BHHV00] Ray Bryant, Bill Hartner, Qi He, and Ganesh Venkitachalam. SMP scalability comparisons of Linux kernels 2.2.14 and 2.3.99. In USENIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/als2000/
bryantscale.html.

Bode:2000:PBS

[BHK^{+}00] Brett Bode, David M. Halstead, Ricky Kendall, Zhou Lei, and David Jackson. The Portable Batch Scheduler and the Maui Scheduler on Linux clusters. In USENIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/als2000/
                                        bode.html.

Back:2000:PKI

[BHL00] Godmar Back, Wilson H. Hsieh, and Jay Lepreau. Processes in KaffeOS: Isolation, resource management, and sharing in Java. In USENIX [USE00h], page ?? ISBN 1-880446-16-2. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/osdi2000/
                                        back.html.

Burge:2000:PQP

[BJH^{+}00] Michel Burge, Rob Jaeger, Hilarie, Jonathan Smith, and John Vicente. Panel: QoS provisioning at the network edge, 2000. URL http://db.

                                        usenix.org/publications/
                                        library/proceedings/ine2000/
                                        technical.html. Unpub-

lished panel talk at the Special Workshop on Intelligence at the Network Edge: March 20, 2000, Holiday Inn — Fisherman’s Wharf, San Francisco, CA.

Barker:2000:OSE

[BJP00] K. Barker, Raj Jayaplan, and R. Peters. An objectbase schema evolution approach to Windows NT security. In USENIX [USE00d], page ?? ISBN 1-880446-20-0. LCCN

                                        ????  URL http://db.
                                        usenix.org/publications/

library/proceedings/usenix-
win2000/barker.html.

Braam:2000:LSS

[BL00] Peter Braam and Walt Ligon. I/O for large scale systems panel, 2000.

URL http://www.usenix.
org/publications/library/
proceedings/als2000/. Un-

published panel talk at Annual Linux Showcase 2000.

Blackwood:2000:I

[Bla00a] David Blackwood. Introductions. ;login: the USENIX Association newsletter, 25(1): ??, February 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
standards/32.introductions.
html.

Blackwood:2000:PB

[Bla00b]  David Blackwood.  The

POSIX bookshelf. ;login: the USENIX Association newsletter, 25(3): ??, June 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
standards/39.POSIX.html.

Blackwood:2000:SRP

[Bla00c] David Blackwood. Standards report: The POSIX bookshelf. ;login: the USENIX Association newsletter, 25(2): ??, April 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.

                                        org/publications/login/
                                        standards/34.posix.html.

Blandford:2001:NTW

[Bla01a]  Jonathan Blandford.  A

new tree widget for GTK+ 2.0. In USENIX [USE01d], page ?? ISBN ???? LCCN ???? URL http:

                                        //www.linuxshowcase.org/
                                        blandford.html.

Blaze:2001:LYS

[Bla01b] Matt Blaze. Loaning your soul to the devil: Influencing policy without selling out, 2001. Unpublished invited talk, Tenth USENIX Security Symposium, August 13– 17, 2001, Washington, DC, USA.

Blaze:2002:CI

[Bla02] Matt Blaze. Cryptography and insecurity. In USE-NIX [USE02a], pages viii +

151. ISBN 1-880446-02-2. LCCN QA76.76.O63 B736

                                        2002. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/bsdcon02/

tech.html. Unpublished invited talk, BSDCON2002: Growing the BSD Community, February 11–14, 2002, Cathedral Hill Hotel, San Francisco, CA.

Brglez:2000:CCS

[BLF^{+}00] Franc Brglez, Hemang Lavana, Zhi Fu, Debabrata Ghosh, Lorie Moffitt, Steve Nelson, J. Marshall Smith, and Jun Zhou. Collaborative client-server architectures in Tcl/Tk: a class project experiment and experience. In USENIX [USE00e], page ?? ISBN 1-880446-24-3. LCCN

????  URL http://db.
usenix.org/publications/
library/proceedings/tcl2k/
brglez.html.

Borella:2000:RAS

[BM00a] Michael Borella and Gabriel E.

Montenegro. RSIP: Address sharing with end-to-end security. In USENIX [USE00j], page ?? ISBN ???? LCCN

????  URL http://db.
usenix.org/publications/
library/proceedings/ine2000/
borella.html.

Brown:2000:TMH

[BM00b] Angela Demke Brown and Todd C. Mowry. Taming the memory hogs: Using compiler-inserted releases to manage physical memory intelligently. In USE-NIX [USE00h], page ?? ISBN 1-880446-16-2. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/osdi2000/
demkebrown.html.

Badros:2001:SEC

[BNB01] Greg J. Badros, Jeffrey Nichols, and Alan Borning. Scwm: An extensible constraint-enabled window manager. In USE-NIX [USE01f], page ?? ISBN 1-880446-10-3. LCCN QA76.8.U65 U84 2001. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/usenix01/
                                        freenix01/badros.html.

Boklund:2001:HC

[Bok01] Andreas Boklund. Home clusters. ;login: the USENIX Association newsletter, 26(5): ??, August 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-08/pdfs/boklund.pdf.

Bonilla:2001:SBF

[Bon01] Oscar Bonilla. A secure OS-based firewall. ;login: the USENIX Association newsletter, 26(7):??, November 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-11/pdfs/bonilla.pdf.

Bouchenak:2001:MJA

[Bou01]  Sara Bouchenak.  Mak-

ing Java applications mobile or persistent. In USE-NIX [USE01e], page ?? ISBN 1-880446-12-X. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/coots01/
                                        bouchenak.html.

Brown:2000:TAB

[BP00] Aaron Brown and David A. Patterson. Towards availability benchmarks: a case study of software RAID systems. In USENIX [USE00a], pages 263–276. ISBN 1-880446-22-7. LCCN ???? URL http:

//www.usenix.org/events/
usenix2000/general/brown.
html.

Bauer:2001:SDD

[BP01] Steven Bauer and Nissanka B. Priyantha. Secure data deletion for Linux file systems. In USENIX [USE01h], page ?? ISBN 1-880446-18-9, 1-880446-07-3. LCCN QA76.9.A25 U565

2001. URL http://www.
usenix.org/publications/
library/proceedings/sec01/
bauer.html.

Boehm:2007:MCC

[BPL07] Hans Boehm, Bill Pugh, and Doug Lea. Multithreading in C and C++. ;login: the USENIX Association newsletter, 32(1):??, February 2007. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
february-2007-volume-32-
number-1/multithreading-
c-and-c.

Brega:2000:DMM

[BR00a] Roberto Brega and Gabrio Rivera. Dynamic memory management with garbage collection for embedded applications. In USE-NIX [USE00f], page ?? ISBN 1-880446-15-4. LCCN QA76.76.S95 W67 2000.

URL http://www.usenix.
org/publications/library/

                                        proceedings/osdi2000/wiess2000/
                                        brega.html.

Brown:2000:LIV

[BR00b] David J. Brown and Karl Runge. Library interface versioning in Solaris and Linux. In USE-NIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/als2000/
                                        browndavid.html.

Brown:2000:MBP

[Bro00] Robert G. Brown. Maximizing Beowulf performance. In USENIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/als2000/
                                        brownrobert.html.

Brod:2001:SOS

[Bro01a] Cesar Brod. SAGU: Opensource unified management system for educational institutions. In USENIX [USE01d], page ???? ISBN ???? LCCN ???? URL http:

                                        //www.linuxshowcase.org/

brod.html. Unpublished invited talk, 5th Annual Linux Showcase and Conference, November 5–10, Oakland, CA.

Brown:2001:EHB

[Bro01b] Robert G. Brown. eden: a home Beowulf. ;login: the USENIX Association newsletter, 26(5):??, August 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-08/pdfs/brown.pdf.

Brown:2001:CES

[Bro01c] Sean R. Brown. Cost effective security for small businesses. In USENIX [USE01f], page ?? ISBN 1-880446-10-3. LCCN QA76.8.U65

U84 2001. URL http://www.
usenix.org/publications/
library/proceedings/usenix01/
freenix01/brown.html.

Brown:2006:MSA

[Bro06] Dave Brown. MythTV: Some assembly required … but the results can be spectacular. ;login: the USENIX Association newsletter, 31(1): ??, February 2006. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2006-02/pdfs/brown.pdf.

Brownsword:2008:DES

[Bro08] Andrew Brownsword. Driving the evolution of software languages to a concurrent future. ;login: the USENIX Association newsletter, 33(3):??, June 2008. CODEN LOGNEM. ISSN 1044-6397. URL

https://www.usenix.org/
publications/login/june-
2008-volume-33-number-3/
driving-evolution-software-
languages-concurrent.

Bryant:2001:PNF

[BRS01] Ray Bryant, Dave Raddatz, and Roger Sunshine. PenguinOMeter: a new file-I/O benchmark for Linux. In USENIX [USE01d], page ?? ISBN ???? LCCN ???? URL http:/

                                        /www.linuxshowcase.org/
                                        bryant.html.

Brumley:2000:RS

[Bru00a] David Brumley. Repeatable security. ;login: the USENIX Association newsletter, 25(7): ??, November 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-11/pdfs/repeat.pdf.

Brutlag:2000:ABD

[Bru00b] Jake D. Brutlag. Aberrant behavior detection in time series for network monitoring. In USE-NIX [USE00g], page ?? ISBN 1-880446-13-8. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/lisa2000/
                                        brutlag.html.

Brumley:2001:CCM

[Bru01] David Brumley. A crash course in managing security. ;login: the USENIX Association newsletter, 26(7):76– 82, November 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-11/pdfs/brumley.pdf.

Brunette:2004:HWT

[Bru04] Glenn M. Brunette, Jr. Hiding within the trees. ;login: the USENIX Association newsletter, 29(1):??, February 2004. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2004-02/pdfs/brunette.pdf.

Brune:2009:PUR

[Bru09a] Corey Brune. Python: An untapped resource in system administration. ;login: the USENIX Association newsletter, 34(1):??, February 2009. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
february-2009-volume-34-
number-1/python-untapped-
resource-system-administration.

Brutch:2009:MMT

[Bru09b] Tasneem G. Brutch. Migration to multicore: Tools that can help. ;login: the USENIX Association newsletter, 34(5): ??, October 2009. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
october-2009-volume-34-
number-5/migration-multicore-
tools-can-help.

Balfanz:2000:WSS

[BS00] Dirk Balfanz and Daniel R.

Simon. WindowBox: a simple security model for the connected desktop. In USENIX [USE00d], page ?? ISBN 1-880446-20-0. LCCN

????  URL http://db.
usenix.org/publications/
library/proceedings/usenix-
win2000/balfanz.html.

Bogda:2001:CSA

[BS01] Jeff Bogda and Ambuj Singh. Can a shape analysis work at run-time? In USE-NIX [USE01g], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38 J42 2001. URL

http://www.usenix.org/publications/
library/proceedings/jvm01/
bogda.html.

Buchholz:2002:PPO

[BS02] Florian Buchholz and Clay Shields. Providing process origin information to aid in network traceback. In USENIX [USE02e], page ?? ISBN 1-880446-00-6. LCCN QA76.8.U65 U84 2002. URL

http://www.usenix.org/publications/
library/proceedings/usenix02/
buchholz.html.

Butler:2004:SRF

[BS04] James Butler and Sherri Sparks. Spyware and rootkits: The future convergence. ;login: the USENIX Association newsletter, 29(6):??, December 2004. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2004-12/pdfs/spyware.pdf.

Bauer:2002:GFA

[BSF02] Lujo Bauer, Michael A.

Schneider, and Edward W. Felten. A general and flexible access-control system for the Web. In USE-NIX [USE02b], pages 93–108. ISBN 1-931971-00-5. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/sec02/
bauer.html.

Baratloo:2000:TRT

[BST00] Arash Baratloo, Navjot Singh, and Timothy Tsai. Transparent run-time defense against stack-smashing attacks. In USENIX [USE00a], pages 251–262. ISBN 1-880446-22-7. LCCN ????

URL http://www.usenix.
org/events/usenix2000/general/
baratloo.html.

Back:2000:TDJ

[BTS^{+}00] Godmar Back, Patrick Tullmann, Leigh Stoller, Wilson C. Hsieh, and Jay Lepreau. Techniques for the design of Java operating systems. In USE-NIX [USE00a], pages 197–

210. ISBN 1-880446-22-7. LCCN ???? URL http:

//www.usenix.org/events/
usenix2000/general/back.
html.

Berg:2008:IA

[BTS08] Jeffrey Berg, Evan Teran, and Sam Stover. Investigating argos. ;login: the USENIX Association newsletter, 33(6): ??, December 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
december-2008-volume-33-
number-6/investigating-
argos.

Black:2002:SCA

[BU02] John Black and Hector Urtubia. Side-channel attacks on symmetric encryption schemes: The case for authenticated encryption. In USENIX [USE02b], pages 327–338. ISBN 1-931971-00-5. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/sec02/
black.html.

Burgess:2000:NCW

[Bur00a] Mark Burgess. Needles in the Craystack: When machines get sick. ;login: the USENIX Association newsletter, 25(8): ??, December 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-12/pdfs/burgess.pdf.

Burgess:2000:SARa

[Bur00b] Mark Burgess. System administration research. ;login: the USENIX Association newsletter, 25(3): ??, June 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/

2000-6/features/research.
html.

Burgess:2000:SARb

[Bur00c] Mark Burgess. System administration research, part

2. ;login: the USENIX Association newsletter, 25(4): ??, July 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-7/features/system.
html.

Burgess:2000:SARc

[Bur00d] Mark Burgess. System administration research, part

3. ;login: the USENIX Association newsletter, 25(5): ??, August 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-8/features/sysadminresearch.
html.

Burgess:2000:TSA

[Bur00e] Mark Burgess. Theoretical system administration. In USENIX [USE00g], page ?? ISBN 1-880446-13-8. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/lisa2000/
burgess.html.

Burgess:2001:NCWf

[Bur01a] Mark Burgess. Needles in the Craystack: When machines get sick, epilogue. ;login: the USENIX Association newsletter, 26(8):??, December 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-12/pdfs/burgess.pdf.

Burgess:2001:NCWa

[Bur01b] Mark Burgess. Needles in the Craystack: When machines get sick, part 2. ;login: the USENIX Association newsletter, 26(1):??, February 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-02/pdfs/burgess.pdf.

Burgess:2001:NCWb

[Bur01c]  Mark Burgess.  Needles

in the Craystack: When machines get sick, part 3. ;login: the USENIX Association newsletter, 26(2): ??, April 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-04/pdfs/burgess.pdf.

Burgess:2001:NCWc

[Bur01d]  Mark Burgess.  Needles

in the Craystack: When machines get sick, part 4. ;login: the USENIX Association newsletter, 26(3): ??, June 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-06/pdfs/burgess.pdf.

Burgess:2001:NCWd

[Bur01e]  Mark Burgess.  Needles

in the Craystack: When machines get sick, part 5. ;login: the USENIX Association newsletter, 26(4): ??, July 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-07/pdfs/burgess.pdf.

Burgess:2001:NCWe

[Bur01f] Mark Burgess. Needles in the Craystack: When machines get sick, part 7. ;login: the USENIX Association newsletter, 26(6):??, October 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-10/pdfs/burgess.pdf.

Burnett:2001:CB

[Bur01g] Steve Burnett. Crypto blunders, 2001. URL http://db.

usenix.org/publications/
library/proceedings/lisa2001/

tech/. Unpublished invited talk, LISA 2001: 15th Systems Administration Conference, December 2–7, 2001, Town and Country Resort Hotel, San Diego, CA.

Burgess:2004:TW

[Bur04] Mark Burgess. Talking to the walls. ;login: the USE-NIX Association newsletter, 29(3):??, June 2004. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2004-06/pdfs/burgess.pdf.

Burgess:2005:PSA

[Bur05] Mark Burgess. The profession of system administration. ;login: the USE-NIX Association newsletter, 30(1):??, February 2005. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2005-02/openpdfs/
burgess.pdf.

Burgess:2006:ACM

[Bur06a]  Mark Burgess.  Auto-

nomic computing — the music of the cubes. ;login: the USENIX Association newsletter, 31(3):??, June 2006. CODEN LOGNEM. ISSN 1044-6397. URL

https://www.usenix.org/
publications/login/june-
2006-volume-31-number-3/
autonomic-computing%C2%
97-music-cubes.

Burgess:2006:CMMa

[Bur06b]  Mark Burgess.  Configu-

ration management: Models and myths. Part 1: Cabbage patch kiss. ;login: the USENIX Association newsletter, 31(4):??, August 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2006-volume-31-number-
4/configuration-management-
models-and-myths-part-1.

Burgess:2006:CMMb

[Bur06c] Mark Burgess. Configuration management: Models and myths. Part 2: Babel, babble, toil, and grammar. ;login: the USENIX Association newsletter, 31(5): ??, October 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
october-2006-volume-31-
number-5/configuration-
management-models-and-myths-
part-2.

Burgess:2006:CMMc

[Bur06d] Mark Burgess. Configuration management: Models and myths. Part 3: a shocking lack of ad-hocracy. ;login: the USENIX Association newsletter, 31(6):??, December 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
december-2006-volume-31-
number-6/configuration-
management-models-and-myths-
part-3.

Burgess:2007:CMM

[Bur07] Mark Burgess. Configuration management: Models and myths. Part 4: There’s no I/O without U. ;login: the USENIX Association newsletter, 32(1):??, February 2007. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/

                                        february-2007-volume-32-
                                        number-1/configuration-
                                        management-models-and-myths-
                                        part-4.

Burgess:2009:CM

[Bur09] Mark Burgess. The Cloud minders. ;login: the USE-NIX Association newsletter, 34(2):??, April 2009. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL https://www.
                                        usenix.org/publications/
                                        login/april-2009-volume-
                                        34-number-2/cloud-minders.

Butler:2004:CM

[But04] Adam Butler. Certs for the masses. ;login: the USE-NIX Association newsletter, 29(3):??, June 2004. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2004-06/pdfs/butler.pdf.

Brazile:2000:RCS

[BV00] Jason Brazile and Andrej Vckovski. Rapid CORBA server development in Tcl: a case study. In USE-NIX [USE00e], page ?? ISBN 1-880446-24-3. LCCN

                                        ????  URL http://db.
                                        usenix.org/publications/
                                        library/proceedings/tcl2k/
                                        brazile.html.

Berkowitz:2000:BMSa

[BY00a] Gale Berkowitz and Ellie Young. Board meeting summary. ;login: the USENIX Association newsletter, 25(1):

??, February 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-2/meeting.html.

Berkowitz:2000:BMSb

[BY00b] Gale Berkowitz and Ellie Young. Board meeting summary. ;login: the USENIX Association newsletter, 25(2): ??, April 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-4/boardmeeting.html.

Byers:2002:WAPb

[Bye02a] Simon Byers. Wireless access point mapping, 2002. ISBN 1-880446-02-2. URL http://

www.usenix.org/publications/
library/proceedings/bsdcon02/

byers.pdf. Unpublished invited talk, BSDCON2002: Growing the BSD Community, February 11–14, 2002, Cathedral Hill Hotel, San Francisco, CA.

Byers:2002:WAPa

[Bye02b]  Simon D. Byers.  Wire-

less access point mapping,

2002. URL http://www.
usenix.org/publications/
library/proceedings/sec02/

tech.html. Unpublished.

Chandra:2000:SFS

[CAGS00] Abhishek Chandra, Micah Adler, Pawan Goyal, and Prashant Shenoy. Surplus fair scheduling: a proportionalshare CPU scheduling algorithm for symmetric multiprocessors. In USE-NIX [USE00h], page ?? ISBN 1-880446-16-2. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/osdi2000/
chandra.html.

Campbell:2000:II

[Cam00]  Duncan Campbell.  The

insecurity industry, 2000.

URL http://www.usenix.
org/publications/library/
proceedings/sec2000/invitedtalks.

html. Unpublished invited talk at Ninth USENIX Security Symposium.

Cantos:2000:KWS

[Can00]  Allan Cantos.  Knowing

when to say no. In USE-NIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/als2000/
cantos.html.

Caple:2001:PPP

[Cap01] James Caple. Palm Pilots and the Point-to-Point Protocol. ;login: the USENIX Association newsletter, 26 (4):??, July 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-07/pdfs/caple.pdf.

Carrere:2000:SS

[Car00a] Daniel Carrere. Samba and SSL, 2000. URL http://

www.usenix.org/publications/
library/proceedings/als2000/

. Unpublished talk at Annual Linux Showcase 2000.

Carvey:2000:SSA

[Car00b] Harlan Carvey. System security administration for NT. In USENIX [USE00b], page ?? ISBN 1-880446-19-7. LCCN

????  URL http://db.
usenix.org/publications/
library/proceedings/lisa-
nt2000/carvey.html.

Carter:2001:L

[Car01]  Gerald Carter.  LDAP,

2001.  URL http://db.
usenix.org/publications/
library/proceedings/lisa2001/

tech/. Unpublished invited talk, LISA 2001: 15th Systems Administration Conference, December 2–7, 2001, Town and Country Resort Hotel, San Diego, CA.

Carter:2002:SIO

[Car02]  Gerald Carter.  SAMBA

— ins and outs. In USENIX [USE02e], page

355. ISBN 1-880446-00-

6. LCCN QA76.8.U65 U84

2002. URL http://www.
usenix.org/publications/
library/proceedings/usenix02/
tech/techonefile.html. Un-

published guru session talk, 2002 USENIX Annual Technical Conference, June 10-15, 2002, Monterey Conference Center, Monterey, CA.

Cramer:2000:DID

[CB00] Christopher E. Cramer and John A. Board. The development and integration of a distributed 3D FFT for a cluster of workstations. In USENIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/als2000/
cramer.html.

Cheswick:2000:MVI

[CBB00] Bill Cheswick, Hal Burch, and Steve Branigan. Mapping and visualizing the Internet. In USENIX [USE00a], pages 1–12. ISBN 1-880446-22-7. LCCN ???? URL http:

//www.usenix.org/events/
usenix2000/general/cheswick.
html.

Cowan:2001:FAP

[CBB^{+}01] Crispin Cowan, Matt Barringer, Steve Beattie, Greg Kroah-Hartman, Mike Frantzen, and Jamie Lokier. Format-Guard: Automatic protection from printf format string vulnerabilities. In USENIX [USE01h], page ?? ISBN 1-880446-18-9, 1-880446-07-3. LCCN QA76.9.A25 U565

2001. URL http://www.
usenix.org/publications/
library/proceedings/sec01/
cowanbarringer.html.

Cowan:2000:SPS

[CBKH^{+}00] Crispin Cowan, Steve Beattie, Greg Kroah-Hartman, Calton Pu, Perry Wagle, and Virgil Gligor. SubDomain: Parsimonious server security. In USE-NIX [USE00g], page ?? ISBN 1-880446-13-8. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/lisa2000/
cowan.html.

Cowan:2001:RKP

[CBWKH01] Crispin Cowan, Steve Beattie, Chris Wright, and Greg Kroah-Hartman. Race-Guard: Kernel protection from temporary file race vulnerabilities. In USENIX [USE01h], page ?? ISBN 1-880446-18-9, 1-880446-07-

3. LCCN QA76.9.A25 U565

2001. URL http://www.
usenix.org/publications/
library/proceedings/sec01/
cowanbeattie.html.

Castro:2000:MSU

[CCG^{+}00] J. C. Hernández Castro, J. M. Sierra Cámara, A. Ribagorda Garnacho, B. Ramos álvarez, and A. Muñoz Cuenca. A methodology for studying untrusted software. ;login: the USENIX Association newsletter, 25(3): ??, June 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/

                                        2000-6/features/method.
                                        html.

Cattaneo:2001:DIT

[CCSP01] Giuseppe Cattaneo, Luigi Catuogno, Aniello Del Sorbo, and Pino Persiano. The design and implementation of a transparent cryptographic file system for UNIX. In USENIX [USE01f], page ?? ISBN 1-880446-10-3. LCCN QA76.8.U65

                                        U84 2001. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/usenix01/
                                        freenix01/cattaneo.html.

Christiaens:2001:TTA

[CD01a] Mark Christiaens and Koen De Bosschere. TRaDe: a topological approach to onthe-fly race detection in Java programs. In USE-NIX [USE01g], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38 J42 2001. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/jvm01/
                                        christiaens.html.

Couch:2001:MNS

[CD01b] Alva Couch and Noah Daniels. The Maelstrom: Network service debugging via “ineffective procedures”. In USENIX [USE01a], page ?? ISBN 1-880446-05-7. LCCN

                                        ????  URL http://db.
                                        usenix.org/publications/
                                        library/proceedings/lisa2001/
                                        tech/couch.html.

Chandra:2001:EEW

[CDGN01] Bharat Chandra, Mike Dahlin, Lei Gao, and Amol Nayate. End-to-end WAN service availability. In USE-NIX [USE01c], page ?? ISBN 1-880446-14-6. LCCN TK5105.875.I57 U96 2001.

URL http://www.usenix.
org/publications/library/
proceedings/usits01/chandra.
html.

Canonico:2000:IQA

[CDRV00] Roberto Canonico, Maurizio D’Arienzo, Simon P. Romanao, and Giorgio Ventre. Introducing QoS awareness in Tcl programming: QTcl. In USENIX [USE00e], page ?? ISBN 1-880446-24-3. LCCN

????  URL http://db.
usenix.org/publications/
library/proceedings/tcl2k/
canonico.html.

Crowcroft:2005:IX

[CFH^{+}05] Jon Crowcroft, Keir Fraser, Steven Hand, Ian Pratt, and Andrew Warfield. The inevitability of Xen. ;login: the USENIX Association newsletter, 30(4):??, August 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-08/pdfs/crowcroft.
pdf.

Curtin:2000:PSS

[CFK00] Matt Curtin, Sandy Farrar, and Tami King. Pelendur:

Steward of the sysadmin. In USENIX [USE00g], page ?? ISBN 1-880446-13-8. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/lisa2000/
curtin.html.

Cox:2002:SP9

[CGP^{+}02] Russ Cox, Eric Grosse, Rob Pike, Dave Presotto, and Sean Quinlan. Security in Plan 9. In USE-NIX [USE02b], pages 3–16. ISBN 1-931971-00-5. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/sec02/
cox.html.

Chalup:2000:TI

[Cha00a] Strata Rose Chalup. Technical interviews. ;login: the USENIX Association newsletter, 25(8):??, December 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-12/pdfs/chalup.pdf.

Chalup:2000:VCS

[Cha00b] Strata Rose Chalup. Vendor contract services. ;login: the USENIX Association newsletter, 25(6):??, October 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-10/pdfs/vendorcontractservices.
pdf.

Chalup:2001:CR

[Cha01a] Strata R. Chalup. Consulting reflections. ;login: the USENIX Association newsletter, 26(8):??, December 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-12/pdfs/chalup.pdf.

Chalup:2001:BBT

[Cha01b] Strata Rose Chalup. A bit off the beaten track. ;login: the USENIX Association newsletter, 26(6):??, October 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-10/pdfs/chalup.pdf.

Chalup:2001:CES

[Cha01c] Strata Rose Chalup. Contractors or employees? It’s not that simple! ;login: the USE-NIX Association newsletter, 26(4):??, July 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-07/pdfs/chalup.pdf.

Chalup:2001:WGS

[Cha01d] Strata Rose Chalup. Writing a good “SysAdmin Wanted” ad. ;login: the USENIX Association newsletter, 26(3): ??, June 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-06/pdfs/chalup.pdf.

Chapman:2001:WYI

[Cha01e]  Brent Chapman.  What

are your intentions? ;login: the USENIX Association newsletter, 26(4): ??, July 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-07/pdfs/chapman.pdf.

Chassell:2001:SFC

[Cha01f] Robert Chassell. Software freedom: Creation and danger, 2001. URL http://www.

                                        linuxshowcase.org/tech.

html. Unpublished invited talk, 5th Annual Linux Showcase and Conference, November 5–10, Oakland, CA.

Chalup:2002:CR

[Cha02a] Strata R. Chalup. Consulting reflections. ;login: the USE-NIX Association newsletter, 27(1):??, February 2002. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2002-02/openpdfs/
                                        chalup.pdf.

Chalup:2002:USU

[Cha02b]  Strata R. Chalup.  Un-

cle Sam as Uncle Spam. ;login: the USENIX Association newsletter, 27(3): ??, June 2002. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2002-06/pdfs/chalup.pdf.

Chalup:2002:VRN

[Cha02c] Strata R. Chalup. Vive la Révolution! Now get over it! ;login: the USENIX Association newsletter, 27(5): ??, October 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-10/pdfs/chalup.pdf.

Chalup:2002:WBZ

[Cha02d] Strata R. Chalup. Wiki, Blog, Zope, and other communicative grunts. ;login: the USE-NIX Association newsletter, 27(2):??, April 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-04/pdfs/chalup.pdf.

Chalup:2003:MWF

[Cha03a] Strata R. Chalup. Not on my watch! Filtering oprtions revisited. ;login: the USE-NIX Association newsletter, 28(2):??, April 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-04/pdfs/chalup.pdf.

Chalup:2003:WRS

[Cha03b] Strata R. Chalup. Will the real “sysadmin of the future” please stand up? ;login: the USENIX Association newsletter, 28(1):??, February 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.

                                        org/publications/login/
                                        2003-02/pdfs/chalup.pdf.

Chak:2005:MAT

[Cha05a] Heison Chak. More Asterisk tricks. ;login: the USE-NIX Association newsletter, 30(3):??, June 2005. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2005-06/pdfs/chak0506.pdf.

Chak:2005:VIA

[Cha05b] Heison Chak. Voice over IP with Asterisk. ;login: the USENIX Association newsletter, 30(1):??, February 2005. CODEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2005-02/openpdfs/
                                        asterisk.pdf.

Chak:2006:CIP

[Cha06a] Heison Chak. Choosing IP phones. ;login: the USENIX Association newsletter, 31(4): ??, August 2006. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        august-2006-volume-31-number-
                                        4/choosing-ip-phones.

Chak:2006:EVS

[Cha06b] Heison Chak. Echo in VoIP systems. ;login: the USENIX Association newsletter, 31(5): ??, October 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
october-2006-volume-31-
number-5/echo-voip-systems.

Chak:2006:VWa

[Cha06c] Heison Chak. VoIP watch.

;login: the USENIX Association newsletter, 31(2):66– 69, April 2006. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2006-04/pdfs/chak.pdf.

Chak:2006:VWb

[Cha06d] Heison Chak. VoIP watch.

;login: the USENIX Association newsletter, 31(3): ??, June 2006. CODEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/june-2006-volume-
31-number-3/voip-watch.

Chak:2006:VWS

[Cha06e] Heison Chak. VoIP watch:

Security. ;login: the USENIX Association newsletter, 31(6): ??, December 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
december-2006-volume-31-
number-6/voip-watch-security.

Chalup:2006:SUb

[Cha06f] Strata Chalup. SAGE update. ;login: the USENIX Association newsletter, 31(5): ??, October 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
october-2006-volume-31-
number-5/sage-update.

Chalup:2006:UNS

[Cha06g] Strata R. Chalup. USE-NIX notes: “Say Hello, Gracie.” “Hello, Gracie.”. ;login: the USENIX Association newsletter, 31(1):??, February 2006. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2006-02/openpdfs/
sage.pdf.

Chalup:2006:SUa

[Cha06h] Strata Rose Chalup. SAGE update. ;login: the USE-NIX Association newsletter, 31(3):??, June 2006. CO-DEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/june-2006-volume-
31-number-3/sage-update.

Chan:2006:CDD

[Cha06i] Kurt Chan. A comparison of disk drives for enterprise computing. ;login: the USE-NIX Association newsletter, 31(3):??, June 2006. CO-DEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/june-2006-volume-
31-number-3/comparison-
disk-drives-enterprise-
computing.

Chak:2007:ALA

[Cha07a]  Heison Chak.  Asterisk

and LumenVox ASR. ;login: the USENIX Association newsletter, 32(6):??, December 2007. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
december-2007-volume-32-
number-6/asterisk-and-lumenvox-
asr.

Chak:2007:AA

[Cha07b] Heison Chak. Asterisk appliance. ;login: the USE-NIX Association newsletter, 32(2):??, April 2007. CO-DEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/april-2007-volume-
32-number-2/asterisk-appliance.

Chak:2007:DAA

[Cha07c] Heison Chak. Deviating alternatives to Asterisk. ;login: the USENIX Association newsletter, 32(4):??, August 2007. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2007-volume-32-number-
4/deviating-alternatives-
asterisk.

Chak:2007:VA

[Cha07d] Heison Chak. Virtualizing Asterisk. ;login: the USENIX Association newsletter, 32(1): ??, February 2007. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        february-2007-volume-32-
                                        number-1/virtualizing-asterisk.

Chak:2008:MRC

[Cha08a] Heison Chak. Media Resource Control Protocol. ;login: the USENIX Association newsletter, 33(4):??, August 2008. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        august-2008-volume-33-number-
                                        4/media-resource-control-
                                        protocol.

Chak:2008:VI

[Cha08b] Heison Chak. VoIP and IPv6. ;login: the USE-NIX Association newsletter, 33(2):??, April 2008. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL https://www.
                                        usenix.org/publications/
                                        login/april-2008-volume-
                                        33-number-2/voip-and-ipv6.

Chen:2000:TMP

[Che00a] Mark Chen. Trust-management pitfalls of PKI, 2000. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/sec2000/
                                        invitedtalks/chen.pdf;
                                        http://www.usenix.org/publications/
                                        library/proceedings/sec2000/
                                        invitedtalks/chen_html.

Unpublished invited talk at Ninth USENIX Security Symposium.

Cheriton:2000:TIA

[Che00b] Dave Cheriton. TRIAD and Internet addressing at the edge. In USENIX [USE00j], page ?? ISBN ???? LCCN

????  URL http://db.
usenix.org/publications/
library/proceedings/ine2000/
cheriton_slides.html;
http://db.usenix.org/publications/
library/proceedings/ine2000/
cheriton_slides.ppt.

Cheswick:2002:ISI

[Che02]  Bill Cheswick.  Internet

security, intranet security, mapping networks. In USENIX [USE02e], page

355. ISBN 1-880446-00-

6. LCCN QA76.8.U65 U84

2002. URL http://www.
usenix.org/publications/
library/proceedings/usenix02/
tech/techonefile.html. Un-

published guru session talk, 2002 USENIX Annual Technical Conference, June 10-15, 2002, Monterey Conference Center, Monterey, CA.

Chien:2000:SWC

[Chi00] Andrew A. Chien. Supercomputing on Windows clusters: Experience and future directions. In USENIX [USE00d], page ?? ISBN 1-880446-20-

0. LCCN ???? URL http:

//db.usenix.org/events/
usenix-win2000/invitedtalks/
chien.ppt.

Chouanard:2000:YTI

[Cho00] Jean Chouanard. YASSP! A tool for improving Solaris security. In USE-NIX [USE00g], page ?? ISBN 1-880446-13-8. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/lisa2000/
                                        chouanard.html.

Chow:2005:TCS

[Cho05] Ming Chow. Teaching computer security, privacy, and politics. ;login: the USENIX Association newsletter, 30(6): ??, December 2005. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2005-12/pdfs/chow.
                                        pdf.

Christiaens:2001:JRR

[Chr01] Mark Christiaens. JaRec: Record/replay for multithreaded Java programs. In USENIX [USE01g], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38 J42 2001. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/jvm01/
                                        JVM_wips/S07.pdf.

Chuvakin:2002:BRE

[Chu02a] Anton Chuvakin. Book review: Eugene Schultz and Russell Shumway: Incident Response. ;login: the USE-NIX Association newsletter, 27(5):??, October 2002. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.

usenix.org/publications/
login/2002-10/openpdfs/
books.pdf.

Chuvakin:2002:BRL

[Chu02b] Anton Chuvakin. Book review: Lance Spitzner: Honeypots: Tracking Hackers. ;login: the USENIX Association newsletter, 27(5): ??, October 2002. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2002-10/openpdfs/
books.pdf.

Chuvakin:2003:UDU

[Chu03] Anton Chuvakin. Ups and downs of UNIX/Linux hostbased security solutions. ;login: the USENIX Association newsletter, 28(2): ??, April 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-04/pdfs/chuvakin.pdf.

Chuvakin:2004:BRE

[Chu04a] Anton Chuvakin. Book review: Ed Skoudis (with Larry Zeitser), Malware: Fighting Malicious Code. ;login: the USENIX Association newsletter, 29(2):??, April 2004. CODEN LOGNEM. ISSN 1044-6397.

Chuvakin:2004:BRK

[Chu04b] Anton Chuvakin. Book review: Kevin Mandia and Chris Prosise, Incident Response: Computer Forensics, 2nd ed. ;login: the USE-NIX Association newsletter, 29(1):??, February 2004. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2004-02/openpdfs/
history.pdf.

Chang:2000:ULR

[CIK00] Fangzhe Chang, Ayal Itzkovitz, and Vijay Karamcheti. Userlevel resource-constrained sandboxing. In USE-NIX [USE00d], page ?? ISBN 1-880446-20-0. LCCN

????  URL http://db.
usenix.org/publications/
library/proceedings/usenix-
win2000/chang.html.

Chang:2002:MCE

[CJL^{+}02] Fay Chang, Minwen Ji, Shun-Tak Leung, John Mac-Cormick, Sharon Perl, and Li Zhang. Myriad: Costeffective disaster tolerance. In USENIX [USE02c], page ?? ISBN 1-880446-03-0. LCCN QA76.9.F5 C65 2002. URL

http://www.usenix.org/publications/
library/proceedings/fast02/
chang.html.

Cohen:2001:APE

[CK01] Edith Cohen and Haim Kaplan. The age penalty and its effect on cache performance. In USENIX [USE01c], page ?? ISBN 1-880446-14-6. LCCN TK5105.875.I57

U96 2001. URL http://www.
usenix.org/publications/

library/proceedings/usits01/
cohen.html.

Chen:2002:AGC

[CKV^{+}02] G. Chen, M. Kandemir, N. Vijaykrishnan, M. J. Irwin, and M. Wolczko. Adaptive garbage collection for battery-operated environments. In USENIX [USE02f], page ?? ISBN 1-931971-01-3. LCCN QA76.73 .J38

J42 2002. URL http://www.
usenix.org/publications/
library/proceedings/javavm02/
chen_g.html.

Castro:2000:PRB

[CL00] Miguel Castro and Barbara Liskov. Proactive recovery in a Byzantine-fault-tolerant system. In USE-NIX [USE00h], page ?? ISBN 1-880446-16-2. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/osdi2000/
castro.html.

Cortes:2001:EHR

[CL01] T. Cortes and J. Laborta.

Extending heterogeneity to RAID Level 5. In USE-NIX [USE01b], page ?? ISBN 1-880446-09-X. LCCN QA76.8.U65 U84 2001. URL

http://www.usenix.org/publications/
library/proceedings/usenix01/
cortes.html.

Clark:2000:IS

[Cla00]  Dave Clark.  Interview-

ing for sysadmins. ;login: the USENIX Association newsletter, 25(1):??, February 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-2/features/interviewing.
html.

Carns:2000:PPF

[CLRT00] Philip H. Carns, Walter B. Ligon III, Robert B. Ross, and Rajeev Thakur. PVFS: a parallel file system for Linux clusters. In USE-NIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/als2000/
carns.html.

Chandra:2001:SLE

[CM01a] Abhishek Chandra and David Mosberger. Scalability of Linux event-dispatch mechanisms. In USENIX [USE01b], page ?? ISBN 1-880446-09-X. LCCN QA76.8.U65

U84 2001. URL http://www.
usenix.org/publications/
library/proceedings/usenix01/
chandra.html.

Chen:2001:FCP

[CM01b] Benjie Chen and Robert Morris. Flexible control of parallelism in a multiprocessor PC router. In USENIX [USE01b], page ?? ISBN 1-880446-09-X. LCCN QA76.8.U65

U84 2001. URL http://www.
usenix.org/publications/
library/proceedings/usenix01/
chen.html.

Chen:2001:JSM

[CMB^{+}01] Deqing Chen, Alan Messer, Philippe Bernadat, Guangrui Fu, Zoran Dimitrijevic, David Jeun Fung Lie, Durga Mannaru, Alma Riska, and Dejan Milojicic. JVM susceptibility to memory errors. In USENIX [USE01g], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38

J42 2001. URL http://www.
usenix.org/publications/
library/proceedings/jvm01/
chen.html.

Corbaln:2000:PDP

[CML00] Julita Corbalán, Xavier Martorell, and Jesús Labarta. Performance-driven processor allocation. In USE-NIX [USE00h], page ?? ISBN 1-880446-16-2. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/osdi2000/
corbalan.html.

Cutello:2000:PAK

[CMP00] Vincenzo Cutello, Emilio Mastriani, and Francesco Pappalardo. Piranha Audit: Kernel enhancements and utilities to improve audit/logging. In USE-NIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/als2000/
cutello.html.

Chen:2002:TDC

[CO02] Michael Chen and Kunle Olukotun. Targeting dynamic compilation for embedded environments. In USENIX [USE02f], page ?? ISBN 1-931971-01-3. LCCN QA76.73 .J38 J42 2002.

                                        URL http://www.usenix.
                                        org/publications/library/
                                        proceedings/javavm02/chen_
                                        m.html.

Cochran:2001:NVS

[Coc01] Shannon Cochran. News and views: Scientists seek immersive reality; USENIX names lifetime achievement recipients [the GNU Project and the Kerberos network authentication system]; robots need programmers; evangelizing the Semantic Web; get your supercomputer software free; Usenet creator Jim Ellis dies; DARPA funds FreeBSD security initiative. Dr. Dobbs Journal, 26(9):18, September 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:

                                        //www.ddj.com/.

Cohen:2002:SF

[Coh02] Gary Cohen. Securing FTP. ;login: the USENIX Association newsletter, 27(4): ??, August 2002. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2002-08/pdfs/cohen.pdf.

Cole:2000:SPI

[Col00] Roland J. Cole. Software Patent Institute report. ;login: the USENIX Association newsletter, 25(2): ??, April 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-4/patent.html.

Committee:2000:CNT

[Com00] The USENIX Awards Committee. Call for nominations for two awards. ;login: the USENIX Association newsletter, 25(2): ??, April 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-4/nominations.html.

Cooper:2004:CME

[Coo04a] Brian F. Cooper. A content model for evaluating peer-to-peer searching techniques. In Jacobsen [Jac04], pages 18–?? CODEN LNCSD9. ISBN 3-540-23428-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.M54 I33 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.springerlink.
com/openurl.asp?genre=issue&
issn=0302-9743&volume=3231;
 http://www.springerlink.
com/openurl.asp?genre=volume&
id=doi:10.1007/b101561.

Cooper:2004:GQI

[Coo04b] Brian F. Cooper. Guiding queries to information sources with InfoBeacons. In Jacobsen [Jac04], pages 59– ?? CODEN LNCSD9. ISBN 3-540-23428-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.M54 I33 2004. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.springerlink.
                                        com/openurl.asp?genre=issue&
                                        issn=0302-9743&volume=3231;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=volume&
                                        id=doi:10.1007/b101561.

Corella:2000:FIT

[Cor00] Francisco Corella. A fast implementation of DES and Triple-DES on PA-RISC 2.0. In USENIX [USE00f], page ?? ISBN 1-880446-15-

4. LCCN QA76.76.S95 W67

                                        2000. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/osdi2000/
                                        wiess2000/corella.html.

Couch:2000:ECA

[Cou00] Alva L. Couch. An expectant chat about script maturity. In USENIX [USE00g], page ?? ISBN 1-880446-13-8. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/lisa2000/
                                        couch.html.

Coufal:2002:NMD

[Cou02]  Hans J. Coufal.  Non-

magnetic data storage: Principles, potential, and problems. In USENIX [USE02c], page ?? ISBN 1-880446-03-

0. LCCN QA76.9.F5 C65

2002. URL http://www.
usenix.org/publications/
library/proceedings/fast02/
coufal.pdf.

Couch:2007:SRP

[Cou07] Alva Couch. Should the root prompt require a road test? ;login: the USENIX Association newsletter, 32(4): ??, August 2007. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2007-volume-32-number-
4/should-root-prompt-require-
road-test.

Couch:2008:SXW

[Cou08a] Alva Couch. From x = 1 to (setf x 1): What does configuration management mean? ;login: the USENIX Association newsletter, 33(1): ??, February 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
february-2008-volume-33-
number-1/x1-setf-x-1-what-
does-configuration-management.

Couch:2008:SDA

[Cou08b] Alva Couch. ‘standard deviations’ of the average system administrator. ;login: the USENIX Association newsletter, 33(4):??, August 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2008-volume-33-number-
4/standard-deviations-average-
system-administrator.

Couch:2008:SAT

[Cou08c] Alva L. Couch. System administration Thermodynamics. ;login: the USENIX Association newsletter, 33(5): ??, October 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
october-2008-volume-33-
number-5/system-administration-
thermodynamics.

Couch:2009:IEB

[Cou09]  Alva L. Couch.  Is it

easy being green? ;login: the USENIX Association newsletter, 34(3):??, June 2009. CODEN LOGNEM. ISSN 1044-6397. URL

https://www.usenix.org/
publications/login/june-
2009-volume-34-number-3/
it-easy-being-green.

Curtin:2000:SPM

[Cur00] Matt Curtin. Shibboleth: Private mailing list manager. In USENIX [USE00i], page ?? ISBN 1-880446-18-9. LCCN

????  URL http://www.
usenix.org/publications/

library/proceedings/sec2000/
curtin.html.

Chandra:2002:ASN

[CV02] Surendar Chandra and Amin Vahdat. Application-specific network management for energy-aware streaming of popular multimedia formats. In USENIX [USE02e], page ?? ISBN 1-880446-00-6. LCCN QA76.8.U65

U84 2002. URL http://www.
usenix.org/publications/
library/proceedings/usenix02/
chandra.html.

Cai:2004:FNL

[CW04] Hailong Cai and Jun Wang.

Foreseer: a novel, localityaware peer-to-peer system architecture for keyword searches. In Jacobsen [Jac04], pages 38–?? CODEN LNCSD9. ISBN 3-540-23428-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.M54 I33 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.springerlink.
com/openurl.asp?genre=issue&
issn=0302-9743&volume=3231;
http://www.springerlink.
com/openurl.asp?genre=volume&
id=doi:10.1007/b101561.

Chen:2002:SD

[CWD02] Hao Chen, David Wagner, and Drew Dean. Setuid demystified. In USE-NIX [USE02b], pages 171–

190. ISBN 1-931971-00-5. LCCN ???? URL

http://www.cs.berkeley.
edu/~daw/papers/setuid-
usenix02.pdf;   http://
www.usenix.org/publications/
library/proceedings/sec02/
chen.html.

Craver:2001:RBL

[CWL^{+}01] Scott A. Craver, Min Wu, Bede Liu, Adam Stubblefield, Ben Swartzlander, Dan S. Wallach, Drew Dean, and Edward W. Felten. Reading between the lines: Lessons from the SDMI challenge. In USE-NIX [USE01h], page ?? ISBN 1-880446-18-9, 1-880446-07-3. LCCN QA76.9.A25 U565

2001. URL http://www.
usenix.org/publications/
library/proceedings/sec01/
craver.html.

Chan:2001:MTL

[CWSC01] Stephen Chan, Cary Whitney, Iwona Sakreja, and Shane Canon. Monitoring tools for larger sites. ;login: the USENIX Association newsletter, 26(5): ??, August 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-08/pdfs/chan.pdf.

Chesire:2001:MAS

[CWVL01] Maureen Chesire, Alec Wolman, Geoffrey M. Voelker, and Henry M. Levy. Measurement and analysis of a streaming media workload. In USENIX [USE01c], page ?? ISBN 1-880446-14-6. LCCN TK5105.875.I57

U96 2001. URL http://www.
usenix.org/publications/
library/proceedings/usits01/
chesire.html.

Cymru:2006:UEP

[Cym06] Team Cymru. The underground economy: Priceless. ;login: the USENIX Association newsletter, 31(6):??, December 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
december-2006-volume-31-
number-6/underground-economy-
priceless.

Doyle:2002:MEJ

[DA02] Patrick Doyle and Tarek S.

Abdelrahman. A modular and extensible JVM infrastructure. In USE-NIX [USE02f], page ?? ISBN 1-931971-01-3. LCCN QA76.73 .J38 J42 2002.

URL http://www.usenix.
org/publications/library/
proceedings/javavm02/doyle.
html.

Darmohray:2003:WN

[DA03] Tina Darmohray and Dan Appelman. What’s in a name? ;login: the USENIX Association newsletter, 28(4): ??, August 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-08/pdfs/darmohray.
pdf.

Denehy:2002:BIG

[DADAD02] Timothy E. Denehy, Andrea C. Arpaci-Dusseau, and Remzi H. Arpaci-Dusseau. Bridging the information gap in storage protocol stacks. In USENIX [USE02e], page ?? ISBN 1-880446-00-6. LCCN QA76.8.U65

                                        U84 2002. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/usenix02/
                                        denehy.html.

Dixon:2008:WMN

[DAK08] Colin Dixon, Thomas Anderson, and Arvind Krishnamurthy. Withstanding multimillion-node botnets. ;login: the USENIX Association newsletter, 33(4): ??, August 2008. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        august-2008-volume-33-number-
                                        4/withstanding-multimillion-
                                        node-botnets.

Damron:2003:IFN

[Dam03] Jason Damron. Identifiable fingerprints in network applications. ;login: the USENIX Association newsletter, 28(6): ??, December 2003. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2003-12/pdfs/damron.pdf.

Danielyan:2001:AAE

[Dan01]  Edgar Danielyan.  AES:

Advanced Encryption Standard is coming. ;login: the USENIX Association newsletter, 26(1):??, February 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-02/pdfs/danielyan.
pdf.

Danielyan:2002:LM

[Dan02] Edgar Danielyan. The law moves in. ;login: the USENIX Association newsletter, 27(1): ??, February 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-02/pdfs/danielyan.
pdf.

Darmohray:2000:Aa

[Dar00a]  Tina Darmohray.  Apro-

pos. ;login: the USENIX Association newsletter, 25(2): ??, April 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-4/apropos.html.

Darmohray:2000:Ab

[Dar00b]  Tina Darmohray.  Apro-

pos. ;login: the USENIX Association newsletter, 25(3): ??, June 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-6/apropos.html.

Darmohray:2000:Ac

[Dar00c]  Tina Darmohray.  Apro-

pos. ;login: the USENIX Association newsletter, 25 (4):??, July 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-7/apropos.html.

Darmohray:2000:Ad

[Dar00d]  Tina Darmohray.  Apro-

pos. ;login: the USENIX Association newsletter, 25(5): ??, August 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-8/apropos.html.

Darmohray:2000:Ae

[Dar00e]  Tina Darmohray.  Apro-

pos. ;login: the USENIX Association newsletter, 25(6): ??, October 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-10/apropos.pdf.

Darmohray:2000:Af

[Dar00f] Tina Darmohray. Apropos. ;login: the USENIX Association newsletter, 25(8):??, December 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-12/apropos.pdf.

Darmohray:2000:RSN

[Dar00g] Tina Darmohray. Ready! Set. No? ;login: the USE-NIX Association newsletter, 25(1):??, February 2000. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2000-2/ready.html.

Darmohray:2001:Aa

[Dar01a] Tina Darmohray. Apropos.

;login: the USENIX Association newsletter, 26(1):??, February 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-02/apropos.pdf.

Darmohray:2001:Ab

[Dar01b]  Tina Darmohray.  Apro-

pos. ;login: the USENIX Association newsletter, 26(2): ??, April 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-04/apropos.pdf.

Darmohray:2001:Ac

[Dar01c]  Tina Darmohray.  Apro-

pos. ;login: the USENIX Association newsletter, 26(3): ??, June 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-06/apropos.pdf.

Darmohray:2001:Ad

[Dar01d]  Tina Darmohray.  Apro-

pos. ;login: the USENIX Association newsletter, 26 (4):??, July 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-07/apropos.pdf.

Darmohray:2001:Ae

[Dar01e]  Tina Darmohray.  Apro-

pos. ;login: the USENIX Association newsletter, 26 (5):??, August 2001. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2001-08/openpdfs/
                                        apropos.pdf.

Darmohray:2001:Af

[Dar01f]  Tina Darmohray.  Apro-

pos. ;login: the USENIX Association newsletter, 26 (6):??, October 2001. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2001-10/openpdfs/
                                        apropos.pdf.

Darmohray:2001:Ag

[Dar01g]  Tina Darmohray.  Apro-

pos. ;login: the USENIX Association newsletter, 26(8): ??, December 2001. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2001-12/openpdfs/
                                        apropos.pdf.

Darmohray:2002:Aa

[Dar02a]  Tina Darmohray.  Apro-

pos. ;login: the USENIX Association newsletter, 27(1): ??, February 2002. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2002-02/openpdfs/
apropos.pdf.

Darmohray:2002:Ab

[Dar02b]  Tina Darmohray.  Apro-

pos. ;login: the USENIX Association newsletter, 27 (3):??, June 2002. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2002-06/openpdfs/
apropos.pdf.

Darmohray:2002:Ac

[Dar02c]  Tina Darmohray.  Apro-

pos. ;login: the USENIX Association newsletter, 27 (4):??, August 2002. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2002-08/openpdfs/
apropos.pdf.

Darmohray:2002:Ad

[Dar02d]  Tina Darmohray.  Apro-

pos. ;login: the USENIX Association newsletter, 27 (5):??, October 2002. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2002-10/openpdfs/
apropos.pdf.

Darmohra:2003:VA

[Dar03a]  Tina Darmohra.  Value

added. ;login: the USENIX Association newsletter, 28(5): ??, October 2003. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2003-10/pdfs/darmohra.pdf.

Darmohray:2003:AR

[Dar03b] Tina Darmohray. Appropriate responsibilities. ;login: the USENIX Association newsletter, 28(1):??, February 2003. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2003-02/pdfs/darmohray.
                                        pdf.

Darmohray:2003:HFP

[Dar03c] Tina Darmohray. Hacking for fun and profit. ;login: the USENIX Association newsletter, 28(2): ??, April 2003. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2003-04/pdfs/darmohray.
                                        pdf.

Darmohray:2004:A

[Dar04a]  Tina Darmohray.  Apro-

pos. ;login: the USENIX Association newsletter, 29 (2):??, April 2004. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2004-04/openpdfs/
                                        Apropos.pdf.

Darmohray:2004:M

[Dar04b]  Tina Darmohray.  motd.

;login: the USENIX Association newsletter, 29(1): ??, February 2004. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2004-02/openpdfs/
motd.pdf.

Darmohray:2004:ASJ

[Dar04c] Tina M. Darmohray. Apropos: Sysadmin jobs: The CIO. ;login: the USENIX Association newsletter, 29 (4):??, August 2004. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2004-08/openpdfs/
apropos.pdf.

Darmohray:2005:FFT

[Dar05]  Tina Darmohray.  Fire-

walls and fairy tales. ;login: the USENIX Association newsletter, 30(1):??, February 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-02/pdfs/firewalls.
pdf.

Davis:2001:DSA

[Dav01] Don Davis. Defective Sign & Encrypt in S/MIME, PKCS#7, MOSS, PEM, PGP, and XML. In USE-NIX [USE01b], page ?? ISBN 1-880446-09-X. LCCN QA76.8.U65 U84 2001. URL

http://www.usenix.org/publications/
library/proceedings/usenix01/
davis.html.

Davidson:2002:TLW

[Dav02] Alan Davidson. Technology, liberty, and Washington. In USENIX [USE02e], page ?? ISBN 1-880446-00-6. LCCN QA76.8.U65

                                        U84 2002. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/usenix02/
                                        tech/techonefile.html. Un-

published invited talk, 2002 USENIX Annual Technical Conference, June 10-15, 2002, Monterey Conference Center, Monterey, CA.

Desai:2006:SMM

[DBH06] Narayan Desai, Rick Bradshaw, and Joey Hagedorn. System management methodologies with Bcfg2. ;login: the USENIX Association newsletter, 31(1):??, February 2006. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2006-02/pdfs/desai.pdf.

deDinechin:2000:CEH

[dD00] Christophe de Dinechin. C++ exception handling for IA64. In USENIX [USE00f], page ?? ISBN 1-880446-15-

4. LCCN QA76.76.S95 W67

                                        2000. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/osdi2000/
                                        wiess2000/dinechin.html.

Dittrich:2007:CCS

[DD07] David Dittrich and Sven Dietrich. Command and control structures in malware: From handler/agent to P2P. ;login: the USENIX Association newsletter, 32(6):??, December 2007. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
december-2007-volume-32-
number-6/command-and-control-
structures-malware.

Demmer:2008:TDF

[DDB08] Michael Demmer, Bowei Du, and Eric Brewer. Tier-Store: a distributed file system for challenged networks in developing regions. ;login: the USENIX Association newsletter, 33(3):??, June 2008. CODEN LOGNEM. ISSN 1044-6397. URL

https://www.usenix.org/
publications/login/june-
2008-volume-33-number-3/
tierstore-distributed-file-
system-challenged.

Deering:2000:INE

[Dee00] Steve Deering. IPv6 for the network edge. In USENIX [USE00j], page ?? ISBN ???? LCCN

????  URL http://db.
usenix.org/publications/
library/proceedings/ine2000/
deering_slides.html;   http:
//db.usenix.org/publications/
library/proceedings/ine2000/
deering_slides.ppt.

Douceur:2009:LLC

[DEH^{+}09] John Douceur, Jeremy Elson, Jon Howell, Jacob R. Lorch, and Rik Farrow. Leveraging legacy code for Web browsers. ;login: the USE-NIX Association newsletter, 34(2):??, April 2009. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL https://www.
                                        usenix.org/publications/
                                        login/april-2009-volume-
                                        34-number-2/leveraging-
                                        legacy-code-web-browsers.

Deraison:2003:N

[DG03] Renaud Deraison and Ron Gula. Nessus. ;login: the USENIX Association newsletter, 28(6):??, December 2003. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2003-12/pdfs/deraison.pdf.

Dehus:2009:SST

[DG09] Mark Dehus and Dirk Grunwald. STORM: Simple tools for resource management. ;login: the USENIX Association newsletter, 34(1):??, February 2009. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        february-2009-volume-34-
                                        number-1/storm-simple-tools-
                                        resource-management.

Deraison:2000:NFN

[DH00] Renaud Deraison and Jordan Hrycaj. Nessus: The free network security scanner. ;login: the USENIX Association newsletter, 25(7):

??, November 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-11/pdfs/nessus.pdf.

Dice:2001:IFJ

[Dic01]  David Dice.  Implement-

ing fast Java^{T M} monitors with relaxed-locks. In USE-NIX [USE01g], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38 J42 2001. URL

http://www.usenix.org/publications/
library/proceedings/jvm01/
dice.html.

Dickison:2008:NUW

[Dic08] Anne Dickison. New on the USENIX Web site: The multimedia page. ;login: the USENIX Association newsletter, 33(1):??, February 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
february-2008-volume-33-
number-1/new-usenix-web-
site-multimedia-page.

Dickison:2009:UDL

[Dic09] Anne Dickison. USENIX discussion list and new videos available! ;login: the USENIX Association newsletter, 34(1):??, February 2009. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
february-2009-volume-34-
number-1/usenix-discussion-
list-and-new-videos-available.

Dietrich:2000:SGD

[Die00]  Sven Dietrich.  Scalpel,

gauze, and decompilers. ;login: the USENIX Association newsletter, 25(7):??, November 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-11/pdfs/dietrich.pdf.

Dietrich:2001:ST

[Die01] Sven Dietrich. Survivability with a twist. ;login: the USENIX Association newsletter, 26(7):??, November 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-11/pdfs/dietrich.pdf.

Dietrich:2002:AND

[Die02] Sven Dietrich. Active network defense: Some concepts and techniques. ;login: the USENIX Association newsletter, 27(6):??, December 2002. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2002-12/pdfs/dietrich.pdf.

Dijker:2000:C

[Dij00a] Barbara Dijker. Certification 2000. ;login: the USE-NIX Association newsletter, 25(2):??, April 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-4/certification.html.

Dijker:2000:SPa

[Dij00b]  Barbara Dijker.  From

the SAGE President. ;login: the USENIX Association newsletter, 25(1):??, February 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-2/president.html.

Dijker:2000:SPb

[Dij00c]  Barbara Dijker.  From

the SAGE President. ;login: the USENIX Association newsletter, 25(2): ??, April 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-4/president.html.

Dijker:2000:SAO

[Dij00d]  Barbara Dijker.  System

administrator occupational analysis. ;login: the USE-NIX Association newsletter, 25(2):??, April 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-4/occupational.html.

Dijker:2000:UYL

[Dij00e] Barbara Dijker. Use your local tools. ;login: the USENIX Association newsletter, 25(6): ??, October 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-10/pdfs/useyourlocaltools.
pdf.

Dijker:2000:WN

[Dij00f] Barbara Dijker. What’s in a name? ;login: the USE-NIX Association newsletter, 25(4):??, July 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-7/sagenews.html#what.

Dike:2000:UMP

[Dik00] Jeff Dike. A user-mode port of the Linux kernel. In USENIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/als2000/
                                        dike.html.

Dike:2001:UML

[Dik01]  Jeff Dike.  User-mode

Linux. In USENIX [USE01d], page ?? ISBN ???? LCCN

                                        ????  URL http://www.
                                        linuxshowcase.org/dike.
                                        html.

Dittrich:2000:CSS

[Dit00] Dave Dittrich. Computer system security: Is there really a threat?, 2000.

                                        URL http://www.usenix.
                                        org/publications/library/
                                        proceedings/sec2000/invitedtalks/
                                        dittrich_html.  Unpub-

lished invited talk at Ninth USENIX Security Symposium.

Dittmer:2001:REP

[Dit01]  Wilfred Dittmer.  Re-

search exchange program update from the field. ;login: the USENIX Association newsletter, 26(3): ??, June 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-06/pdfs/dittmer.pdf.

Dittrich:2009:CFR

[Dit09a] David Dittrich. The conflicts facing those responding to cyberconflict. ;login: the USENIX Association newsletter, 34(6):??, December 2009. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
december-2009-volume-34-
number-6/conflicts-facing-
those-responding-cyberconflict.

Dittrich:2009:MCH

[Dit09b] David Dittrich. Malware to crimeware: How far have they gone, and how do we catch up? ;login: the USENIX Association newsletter, 34(4): ??, August 2009. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2009-volume-34-number-
4/malware-crimeware-how-
far-have-they-gone-and-
how.

Darmohray:2002:PC

[DK02] Tina Darmohray and Rob Kolstad. Point/counterpoint. ;login: the USENIX Association newsletter, 27(2):

??, April 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-04/openpdfs/point.
pdf.

Detlefs:2002:CRS

[DKCJ02] David Detlefs, Ross Knippel, William D. Clinger, and Matthias Jacob. Concurrent remembered set refinement in generational garbage collection. In USENIX [USE02f], page ?? ISBN 1-931971-01-3. LCCN QA76.73 .J38

J42 2002. URL http://www.
usenix.org/publications/
library/proceedings/javavm02/
detlefs.html.

Dieter:2001:ULC

[DL01] William R. Dieter and James E. Lumpp, Jr. Userlevel checkpointing for LinuxThreads programs. In USENIX [USE01f], page ?? ISBN 1-880446-10-3. LCCN QA76.8.U65 U84 2001. URL

http://www.usenix.org/publications/
library/proceedings/usenix01/
freenix01/dieter.html.

Dietrich:2000:ADD

[DLD00] Sven Dietrich, Neil Long, and David Dittrich. Analyzing distributed denial of service tools: The Shaft case. In USENIX [USE00g], page ?? ISBN 1-880446-13-8. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/lisa2000/
dietrich.html.

Duret-Lutz:2001:DPG

[DLGD01] Alexandre Duret-Lutz, Thierry Géraud, and Akim Demaille. Design patterns for generic programming in C++. In USENIX [USE01e], page ?? ISBN 1-880446-12-X. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/coots01/
duret.html.

Dutchyn:2001:MDJ

[DLS^{+}01] Christopher Dutchyn, Paul Lu, Duane Szafron, Steven Bromling, and Wade Holst. Multi-dispatch in the Java Virtual Machine: Design and implementation. In USENIX [USE01e], page ?? ISBN 1-880446-12-X. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/coots01/
dutchyn.html.

deLara:2000:OBA

[dLWZ00] Eyal de Lara, Dan S. Wallach, and Willy Zwaenepoel. Opportunities for bandwidth adaptation in Microsoft Office documents. In USE-NIX [USE00d], page ?? ISBN 1-880446-20-0. LCCN

????  URL http://db.
usenix.org/publications/
library/proceedings/usenix-
win2000/delara.html.

deLara:2001:PCB

[dLWZ01] Eyal de Lara, Dan S. Wallach, and Willy Zwaenepoel.

Puppeteer: Componentbased adaptation for mobile computing. In USE-NIX [USE01c], page ?? ISBN 1-880446-14-6. LCCN TK5105.875.I57 U96 2001.

URL http://www.usenix.
org/publications/library/
proceedings/usits01/delara.
html.

Dietz:2000:KFN

[DM00] H. G. Dietz and T. I. Mattox. KLAT2’s flat neighborhood network. In USE-NIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/als2000/
dietz.html.

Dowse:2002:RFO

[DM02] Ian Dowse and David Malone. Recent filesystem optimisations on FreeBSD. In USENIX [USE02d], page ?? ISBN 1-880446-01-4. LCCN QA76.8.U65 P765 2002.

URL http://www.usenix.
org/publications/library/
proceedings/usenix02/tech/
freenix/dowse.html.

Dawidek:2007:PSZ

[DM07] Pawel Jakub Dawidek and Marshall Kirk McKusick. Porting the Solaris ZFS file system to the FreeBSD operating system. ;login: the USENIX Association newsletter, 32(3):19–24, June 2007. CODEN LOGNEM. ISSN 1044-6397. URL

https://www.usenix.org/
publications/login/june-
2007-volume-32-number-3/
porting-solaris-zfs-file-
system-freebsd-operating.

Dolotta:2006:UNL

[Dol06]  Ted Dolotta.  USENIX

notes: Letter to the Editor. ;login: the USENIX Association newsletter, 31(2): 80, April 2006. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2006-04/openpdfs/letter.
pdf.

Dhamija:2000:DVU

[DP00] Rachna Dhamija and Adrian Perrig. Déjà Vu — a user study: Using images for authentication. In USE-NIX [USE00i], page ?? ISBN 1-880446-18-9. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/sec2000/
dhamija.html.

Dreyfus:2000:PUC

[Dre00] Suelette Dreyfus. The practical use of cryptography in human rights groups,

2000. URL http://www.
usenix.org/publications/
library/proceedings/sec2000/
invitedtalks.html. Unpub-

lished invited talk at Ninth USENIX Security Symposium.

Drzyzgula:2000:DDC

[Drz00]  Robert Drzyzgula.  De-

signing a data center instrumentation system. In USENIX [USE00g], page ?? ISBN 1-880446-13-8. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/lisa2000/
                                        drzyzgula.html.

Dean:2001:UCP

[DS01] Drew Dean and Adam Stubblefield. Using client puzzles to protect TLS. In USENIX [USE01h], page ?? ISBN 1-880446-18-9, 1-880446-07-

3. LCCN QA76.9.A25 U565

                                        2001. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/sec01/
                                        dean.html.

Duffin:2000:F

[Duf00] Paul Duffin. Feather. In USENIX [USE00e], page ?? ISBN 1-880446-24-3. LCCN

                                        ????  URL http://db.
                                        usenix.org/publications/
                                        library/proceedings/tcl2k/
                                        duffin.html.

Dusek:2008:CPP

[Dus08] Jason Dusek. Concurrent patterns: Parallels in system and language design. ;login: the USENIX Association newsletter, 33(5): ??, October 2008. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        october-2008-volume-33-

number-5/concurrent-patterns-
parallels-system-and-language.

Diklic:2001:ROS

[DVWW01] Dejan Diklic, Venkatesh Velayutham, Steve Welch, and Roger Williams. Remote outsourcing services for multiple branch offices and small businesses via the Internet. In USENIX [USE01a], page ?? ISBN 1-880446-05-7. LCCN

????  URL http://db.
usenix.org/publications/
library/proceedings/lisa2001/
tech/diklic.html.

Dzonsons:2009:FSL

[Dzo09]  Kristaps Dzonsons.  Fix-

ing on a standard language for UNIX manuals. ;login: the USENIX Association newsletter, 34(5):??, October 2009. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
october-2009-volume-34-
number-5/fixing-standard-
language-unix-manuals.

Edwards:2001:SLP

[EBC01] Nigel Edwards, Joubert Berger, and Tse Huong Choo. A secure Linux platform. In USENIX [USE01d], page ?? ISBN ???? LCCN ???? URL http:

//www.linuxshowcase.org/
edwards.html.

Engler:2000:CSR

[ECCH00] Dawson Engler, Benjamin Chelf, Andy Chou, and Seth Hallem. Checking system rules using systemspecific, programmer-written compiler extensions. In USENIX [USE00h], page ?? ISBN 1-880446-16-2. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/osdi2000/
engler.html.

Evard:2000:CAB

[ECF^{+}00] Remy Evard, Susan Coghlan, Richard Ferri, Brian Finley, Greg Lindahl, John-Paul Navarro, Lee Ward, and Stephen L. Scott. Cluster administration has been solved, or has it?, 2000. ISBN 1-880446-17-0. URL http://

www.usenix.org/publications/
library/proceedings/als2000/

. Unpublished panel talk at Annual Linux Showcase 2000.

Eckenwiler:2001:MTL

[Eck01] Mark Eckenwiler. A maze of twisty little statutes, all alike: The Electronic Communications Privacy Act of 1986,

2001. URL http://www.
usenix.org/publications/
library/proceedings/sec01/
eckenwiler/index.htm. Un-

published invited talk, Tenth USENIX Security Symposium, August 13–17, 2001, Washington, DC, USA.

Eugster:2001:CBP

[EG01] Patrick Th. Eugster and Rachid Guerraoui. Contentbased publish/subscribe with structural reflection. In USENIX [USE01e], page ?? ISBN 1-880446-12-X. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/coots01/
eugster.html.

Egelman:2004:SSF

[Ege04] Serge Egelman. Suing spammers for fun and profit. ;login: the USENIX Association newsletter, 29(2): ??, April 2004. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2004-04/pdfs/Egelman.pdf.

El-Kadi:2005:AID

[EKNG^{+}05] Amr El-Kadi, Ahmed Nashed, Kareem El Gebaly, Mahmoud Abo Daoud, Noha El Sharawy, Rania Nazmi, and Mostafa Mazen. Architecture and internal design of the AUC-Abyss Web server. ;login: the USENIX Association newsletter, 30(3): ??, June 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-06/pdfs/el-kadi0506.
pdf.

Engelschall:2000:PMS

[Eng00] Ralf S. Engelschall. Portable multithreading — the signal stack trick for user-space thread creation. In USE-NIX [USE00a], pages 239– 249. ISBN 1-880446-22-7.

LCCN ???? URL http:

                                        //www.usenix.org/events/
                                        usenix2000/general/engelschall.
                                        html.

Ely:2001:AUL

[ESW01] David Ely, Stefan Savage, and David Wetherall. Alpine: a user-level infrastructure for network protocol development. In USENIX [USE01c], page ?? ISBN 1-880446-14-6. LCCN TK5105.875.I57

                                        U96 2001. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/usits01/
                                        ely.html.

Enright:2008:SWR

[EVS^{+}08] Brandon Enright, Geoff Voelker, Stefan Savage, Chris Kanich, and Kiril Levchenko. Storm: When researchers collide. ;login: the USENIX Association newsletter, 33(4): ??, August 2008. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        august-2008-volume-33-number-
                                        4/storm-when-researchers-
                                        collide.

Ezolt:2001:SMC

[Ezo01] Phillip Ezolt. A study in malloc: a case of excessive minor faults. In USENIX [USE01d], page ?? ISBN ???? LCCN

                                        ????  URL http://www.
                                        linuxshowcase.org/ezolt.
                                        html.

Forney:2002:SAC

[FADAD02] Brian C. Forney, Andrea C.

Arpaci-Dusseau, and Remzi H. Arpaci-Dusseau. Storageaware caching: Revisiting caching for heterogeneous storage systems. In USENIX [USE02c], page ?? ISBN 1-880446-03-0. LCCN QA76.9.F5 C65 2002. URL

http://www.usenix.org/publications/
library/proceedings/fast02/
forney.html.

Farrow:2000:I

[Far00a] Rik Farrow. In this issue.

;login: the USENIX Association newsletter, 25(7):??, November 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-11/inthisissue.pdf.

Farrow:2000:Ma

[Far00b] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 25(1):??, February 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-2/features/musings.
html.

Farrow:2000:Mb

[Far00c] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 25(2): ??, April 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/

                                        2000-4/features/musings.
                                        html.

Farrow:2000:Mc

[Far00d] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 25(3): ??, June 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-6/features/musings.
                                        html.

Farrow:2000:Md

[Far00e] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 25(5):??, August 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-8/features/musings.
                                        html.

Farrow:2000:Me

[Far00f] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 25(6):??, October 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-10/pdfs/musings.pdf.

Farrow:2000:Mf

[Far00g] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 25(8):??, December 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-12/pdfs/musings.pdf.

Farrow:2000:MES

[Far00h] Rik Farrow. Musings on embedded systems. ;login: the USENIX Association newsletter, 25(4): ??, July 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-7/features/musings.
html.

Farrow:2000:SSR

[Far00i] Rik Farrow. Smart-space researchers: An interview with Kevin Mills and Alden Dima. ;login: the USE-NIX Association newsletter, 25(4):??, July 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-7/features/smart.html.

Farrow:2001:I

[Far01a] Rik Farrow. In this issue.

;login: the USENIX Association newsletter, 26(7):??, November 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-11/openpdfs/farrow.
pdf.

Farrow:2001:Ma

[Far01b] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 26(1):??, February 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.

                                        org/publications/login/
                                        2001-02/pdfs/farrow.pdf.

Farrow:2001:Mb

[Far01c] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 26(2): ??, April 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-04/pdfs/farrow.pdf.

Farrow:2001:Mc

[Far01d] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 26(3): ??, June 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-06/pdfs/farrow.pdf.

Farrow:2001:Md

[Far01e] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 26(4): ??, July 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-07/pdfs/farrow.pdf.

Farrow:2001:Me

[Far01f] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 26(6):??, October 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-10/pdfs/farrow.pdf.

Farrow:2001:Mf

[Far01g] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 26(8):??, December 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-12/pdfs/farrow.pdf.

Farrow:2001:RLK

[Far01h] Rik Farrow. A report on the Linux 2.5 Kernel Developers Summit. ;login: the USENIX Association newsletter, 26(3):??, June 2001. CODEN LOGNEM. ISSN 1044-6397. URL http:/

/www.usenix.org/events/
kernel01/summit.pdf.

Farrow:2002:I

[Far02a] Rik Farrow. In this issue. ;login: the USENIX Association newsletter, 27(6): ??, December 2002. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2002-12/openpdfs/
inthisissue.pdf.

Farrow:2002:Ma

[Far02b] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 27(1):??, February 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-02/pdfs/farrow.pdf.

Farrow:2002:Mb

[Far02c] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 27(2): ??, April 2002. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2002-04/pdfs/farrow.pdf.

Farrow:2002:Mc

[Far02d] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 27(3): ??, June 2002. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2002-06/pdfs/farrow.pdf.

Farrow:2002:Md

[Far02e] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 27(4):??, August 2002. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2002-08/pdfs/farrow.pdf.

Farrow:2002:MWD

[Far02f] Rik Farrow. Musings, or what I did on my summer vacation. ;login: the USENIX Association newsletter, 27(5): ??, October 2002. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2002-10/pdfs/farrow.pdf.

Farrow:2003:I

[Far03a] Rik Farrow. In this issue.

;login: the USENIX Association newsletter, 28(6):??, December 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-12/openpdfs/fm.pdf.

Farrow:2003:Ma

[Far03b] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 28(1):??, February 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-02/pdfs/farrow.pdf.

Farrow:2003:Mb

[Far03c] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 28(2): ??, April 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-04/pdfs/farrow.pdf.

Farrow:2003:Mc

[Far03d] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 28(3): ??, June 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-06/pdfs/farrow.pdf.

Farrow:2003:Md

[Far03e] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 28(4):??, August 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-08/pdfs/farrow.pdf.

Farrow:2003:Me

[Far03f] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 28(5):??, October 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-10/pdfs/farrow.pdf.

Farrow:2004:BRC

[Far04a] Rik Farrow. Book review: Cyrus Peikari and Anton Chuvakin, Security Warrior. ;login: the USENIX Association newsletter, 29(2): ??, April 2004. CODEN LOGNEM. ISSN 1044-6397.

Farrow:2004:BRN

[Far04b] Rik Farrow. Book review: Neal Stephenson, Quicksilver. ;login: the USENIX Association newsletter, 29(2): ??, April 2004. CODEN LOGNEM. ISSN 1044-6397.

Farrow:2004:BRa

[Far04c] Rik Farrow. Book reviews. ;login: the USENIX Association newsletter, 29(4): ??, August 2004. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2004-08/openpdfs/
rikreviews.pdf.

Farrow:2004:BRb

[Far04d] Rik Farrow. Book reviews.

;login: the USENIX Association newsletter, 29(6): ??, December 2004. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2004-12/openpdfs/
reviews.pdf.

Farrow:2004:ETK

[Far04e]  Rik Farrow.  Editor’s

thoughts (a.k.a. musings). ;login: the USENIX Association newsletter, 29(6): ??, December 2004. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2004-12/openpdfs/
musings.pdf.

Farrow:2004:Ma

[Far04f] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 29(1):??, February 2004. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2004-02/openpdfs/
musings.pdf.

Farrow:2004:Mb

[Far04g] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 29(2): ??, April 2004. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2004-04/pdfs/Farrow.pdf.

Farrow:2004:Mc

[Far04h] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 29(3): ??, June 2004. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2004-06/pdfs/musings.pdf.

Farrow:2004:Md

[Far04i] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 29(4):??, August 2004. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2004-08/pdfs/musings.pdf.

Farrow:2004:Me

[Far04j] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 29(5):??, October 2004. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2004-10/pdfs/musings.pdf.

Farrow:2005:BRb

[Far05a] Rik Farrow. Book reviews. ;login: the USENIX Association newsletter, 30(4): ??, August 2005. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2005-08/openpdfs/
                                        reviews.pdf.

Farrow:2005:Ma

[Far05b] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 30(1):??, February 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-02/pdfs/musings.pdf.

Farrow:2005:Mb

[Far05c] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 30(2): ??, April 2005. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2005-04/openpdfs/
musings0504.pdf.

Farrow:2005:Mc

[Far05d] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 30(3): ??, June 2005. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2005-06/openpdfs/
musings0506.pdf.

Farrow:2005:Md

[Far05e] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 30(4):??, August 2005. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2005-08/openpdfs/
musings.pdf.

Farrow:2005:Me

[Far05f] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 30(5):??, October 2005. CODEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2005-10/openpdfs/
                                        musings.pdf.

Farrow:2005:Mf

[Far05g] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 30(6):??, December 2005. CODEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2005-12/openpdfs/
                                        musings.pdf.

Farrow:2005:UNS

[Far05h] Rik Farrow. USENIX notes: Short topics booklets. ;login: the USENIX Association newsletter, 30(2): ??, April 2005. CODEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2005-04/openpdfs/
                                        shorttopics.pdf.

Farrow:2006:Ma

[Far06a] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 31(1):??, February 2006. CODEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/

login/2006-02/openpdfs/
musings.pdf.

Farrow:2006:Mb

[Far06b] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 31(2):2– 4, April 2006. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2006-04/openpdfs/
musings.pdf.

Farrow:2006:Mc

[Far06c] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 31(3):??, June 2006. CODEN LOGNEM. ISSN 1044-6397. URL

https://www.usenix.org/
publications/login/june-
2006-volume-31-number-3/
musings.

Farrow:2006:Md

[Far06d] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 31(4):??, August 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2006-volume-31-number-
4/musings.

Farrow:2006:Me

[Far06e] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 31(5):??, October 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
october-2006-volume-31-
number-5/musings.

Farrow:2006:Mf

[Far06f] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 31(6):??, December 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
december-2006-volume-31-
number-6/musings.

Farrow:2006:WOS

[Far06g] Rik Farrow. WISPER: Open source, long-distance wireless. ;login: the USENIX Association newsletter, 31(5): ??, October 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
october-2006-volume-31-
number-5/wisper-open-source-
long-distance-wireless.

Farrow:2007:Ma

[Far07a] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 32(1):??, February 2007. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
february-2007-volume-32-
number-1/musings.

Farrow:2007:Mb

[Far07b] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 32(2): ??, April 2007. CODEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/april-2007-volume-
32-number-2/musings.

Farrow:2007:Mc

[Far07c] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 32(3):??, June 2007. CODEN LOGNEM. ISSN 1044-6397. URL

https://www.usenix.org/
publications/login/june-
2007-volume-32-number-3/
musings.

Farrow:2007:Md

[Far07d] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 32(4):??, August 2007. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2007-volume-32-number-
4/musings.

Farrow:2007:Me

[Far07e] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 32(5):??, October 2007. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
october-2007-volume-32-
number-5/musings.

Farrow:2007:Mf

[Far07f] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 32(6):??, December 2007. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        december-2007-volume-32-
                                        number-6/musings.

Farrow:2007:PPI

[Far07g]  Rik Farrow.  Promises,

promises: An interview with Mark Burgess. ;login: the USENIX Association newsletter, 32(2): ??, April 2007. CODEN LOGNEM. ISSN 1044-

                                        6397. URL https://www.
                                        usenix.org/publications/
                                        login/april-2007-volume-
                                        32-number-2/promises-promises-
                                        interview-mark-burgess.

Farrow:2008:Ma

[Far08a] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 33(1):??, February 2008. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        february-2008-volume-33-
                                        number-1/musings.

Farrow:2008:Mb

[Far08b] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 33(2): ??, April 2008. CODEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/april-2008-volume-
33-number-2/musings.

Farrow:2008:Mc

[Far08c] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 33(3):??, June 2008. CODEN LOGNEM. ISSN 1044-6397. URL

https://www.usenix.org/
publications/login/june-
2008-volume-33-number-3/
musings.

Farrow:2008:Md

[Far08d] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 33(4):??, August 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2008-volume-33-number-
4/musings.

Farrow:2008:Me

[Far08e] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 33(5):??, October 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
october-2008-volume-33-
number-5/musings.

Farrow:2008:Mf

[Far08f] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 33(6):??, December 2008. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        december-2008-volume-33-
                                        number-6/musings.

Farrow:2009:Ma

[Far09a] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 34(1):??, February 2009. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        february-2009-volume-34-
                                        number-1/musings.

Farrow:2009:Mb

[Far09b] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 34(2): ??, April 2009. CODEN LOGNEM. ISSN 1044-

                                        6397. URL https://www.
                                        usenix.org/publications/
                                        login/april-2009-volume-
                                        34-number-2/musings.

Farrow:2009:Mc

[Far09c] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 34(3):??, June 2009. CODEN LOGNEM. ISSN 1044-6397. URL

                                        https://www.usenix.org/
                                        publications/login/june-
                                        2009-volume-34-number-3/
                                        musings.

Farrow:2009:Md

[Far09d] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 34(4):??, August 2009. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2009-volume-34-number-
4/musings.

Farrow:2009:Me

[Far09e] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 34(5):??, October 2009. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
october-2009-volume-34-
number-5/musings.

Farrow:2009:Mf

[Far09f] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 34(6):??, December 2009. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
december-2009-volume-34-
number-6/musings.

Feamster:2002:ICW

[FBH^{+}02] Nick Feamster, Magdalena Balazinska, Greg Harfst, Hari Balakrishnan, and David Karger. Infranet: Circumventing Web censorship and surveillance. In USENIX [USE02b], pages 237–262. ISBN 1-931971-00-5. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/sec02/
feamster.html.

Forte:2004:SWI

[FBMZ04] Dario Forte, Davide Bertoletti, Cristiano Maruti, and Michele Zambelli. Sebek Web “ITA” interface. ;login: the USENIX Association newsletter, 29(6):??, December 2004. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2004-12/pdfs/sebek.pdf.

Fong:2001:PLD

[FC01] Philip W. L. Fong and Robert D. Cameron. Proof linking: Distributed verification of Java classfiles in the presence of multiple classloaders. In USE-NIX [USE01g], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38 J42 2001. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/jvm01/
                                        fong.html.

Fu:2002:EPE

[FCTV02] Yun Fu, Ludmila Cherkasova, Wenting Tang, and Amin Vahdat. EtE: Passive end-to-end Internet service performance monitoring. In USENIX [USE02e], page ?? ISBN 1-880446-00-6. LCCN QA76.8.U65

                                        U84 2002. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/usenix02/
                                        fu.html.

Flood:2001:PGC

[FDSZ01] Christine H. Flood, David Detlefs, Nir Shavit, and Xiolan Zhang. Parallel garbage collection for shared memory multiprocessors. In USENIX [USE01g], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38

J42 2001. URL http://www.
usenix.org/publications/
library/proceedings/jvm01/
flood.html.

Felten:2002:FT

[Fel02] Ed Felten. Freedom to tinker, 2002. URL http://www.

usenix.org/publications/
library/proceedings/sec02/

tech.html. Unpublished.

Fennely:2000:IBB

[Fen00] Carole Fennely. An interview with Blaine Burnham. ;login: the USENIX Association newsletter, 25(7):??, November 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-11/pdfs/burnham.pdf.

Ferrell:2006:DRa

[Fer06a] Robert G. Ferrell. /dev/random.

;login: the USENIX Association newsletter, 31(2): 70–71, April 2006. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2006-04/openpdfs/
ferrell.pdf.

Ferrell:2006:DRb

[Fer06b] Robert G. Ferrell. /dev/random. ;login: the USENIX Association newsletter, 31(3): ??, June 2006. CODEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/june-2006-volume-
31-number-3/devrandom.

Ferrell:2006:DRc

[Fer06c] Robert G. Ferrell. /dev/random. ;login: the USENIX Association newsletter, 31(5): ??, October 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
october-2006-volume-31-
number-5/devrandom.

Ferrell:2006:DRd

[Fer06d] Robert G. Ferrell. /dev/random. ;login: the USENIX Association newsletter, 31(6):??, December 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
december-2006-volume-31-
number-6/devrandom.

Ferrell:2007:DRa

[Fer07a] Robert G. Ferrell. /dev/random. ;login: the USENIX Association newsletter, 32(1):??, February 2007. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
february-2007-volume-32-
number-1/devrandom.

Ferrell:2007:DRb

[Fer07b] Robert G. Ferrell. /dev/random.

;login: the USENIX Association newsletter, 32(2): ??, April 2007. CODEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/april-2007-volume-
32-number-2/devrandom.

Ferrell:2007:DRc

[Fer07c] Robert G. Ferrell. /dev/random.

;login: the USENIX Association newsletter, 32(3): ??, June 2007. CODEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/june-2007-volume-
32-number-3/devrandom.

Ferrell:2007:DRd

[Fer07d] Robert G. Ferrell. /dev/random.

;login: the USENIX Association newsletter, 32(4): ??, August 2007. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2007-volume-32-number-
4/devrandom.

Ferrell:2007:DRe

[Fer07e] Robert G. Ferrell. /dev/random.

;login: the USENIX Association newsletter, 32(5): ??, October 2007. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
october-2007-volume-32-
number-5/devrandom.

Ferrell:2007:DRf

[Fer07f] Robert G. Ferrell. /dev/random. ;login: the USENIX Association newsletter, 32(6):??, December 2007. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        december-2007-volume-32-
                                        number-6/devrandom.

Ferrell:2008:DRa

[Fer08a] Robert G. Ferrell. /dev/random. ;login: the USENIX Association newsletter, 33(1):??, February 2008. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        february-2008-volume-33-
                                        number-1/devrandom.

Ferrell:2008:DRb

[Fer08b] Robert G. Ferrell. /dev/random. ;login: the USENIX Association newsletter, 33(2): ??, April 2008. CODEN LOGNEM. ISSN 1044-

                                        6397. URL https://www.
                                        usenix.org/publications/
                                        login/april-2008-volume-
                                        33-number-2/devrandom.

Ferrell:2008:DRc

[Fer08c] Robert G. Ferrell. /dev/random. ;login: the USENIX Association newsletter, 33(3): ??, June 2008. CODEN LOGNEM. ISSN 1044-

                                        6397. URL https://www.
                                        usenix.org/publications/
                                        login/june-2008-volume-
                                        33-number-3/devrandom.

Ferrell:2008:DRd

[Fer08d] Robert G. Ferrell. /dev/random.

;login: the USENIX Association newsletter, 33(4): ??, August 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2008-volume-33-number-
4/devrandom.

Ferrell:2008:DRe

[Fer08e] Robert G. Ferrell. /dev/random.

;login: the USENIX Association newsletter, 33(5): ??, October 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
october-2008-volume-33-
number-5/devrandom.

Ferrell:2008:DRf

[Fer08f] Robert G. Ferrell. /dev/random.

;login: the USENIX Association newsletter, 33(6):??, December 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
december-2008-volume-33-
number-6/devrandom.

Ferrell:2009:DRa

[Fer09a] Robert G. Ferrell. /dev/random.

;login: the USENIX Association newsletter, 34(1):??, February 2009. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
february-2009-volume-34-
number-1/devrandom.

Ferrell:2009:DRb

[Fer09b] Robert G. Ferrell. /dev/random. ;login: the USENIX Association newsletter, 34(2): ??, April 2009. CODEN LOGNEM. ISSN 1044-

                                        6397. URL https://www.
                                        usenix.org/publications/
                                        login/april-2009-volume-
                                        34-number-2/devrandom.

Ferrell:2009:DRc

[Fer09c] Robert G. Ferrell. /dev/random. ;login: the USENIX Association newsletter, 34(3): ??, June 2009. CODEN LOGNEM. ISSN 1044-

                                        6397. URL https://www.
                                        usenix.org/publications/
                                        login/june-2009-volume-
                                        34-number-3/devrandom.

Ferrell:2009:DRd

[Fer09d] Robert G. Ferrell. /dev/random. ;login: the USENIX Association newsletter, 34(4): ??, August 2009. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        august-2009-volume-34-number-
                                        4/devrandom.

Ferrell:2009:DRe

[Fer09e] Robert G. Ferrell. /dev/random. ;login: the USENIX Association newsletter, 34(5): ??, October 2009. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        october-2009-volume-34-
                                        number-5/devrandom.

Ferrell:2009:DRR

[Fer09f] Robert G. Ferrell. /dev/random:

a Realist’s glossary of terms widely employed in the information security arena. ;login: the USENIX Association newsletter, 34(6):??, December 2009. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
december-2009-volume-34-
number-6/devrandom-realists-
glossary-terms-widely.

Fink:2007:ACF

[FF07] Glenn Fink and Deb Frincke.

Autonomic computing: Freedom or threat? ;login: the USENIX Association newsletter, 32(2): ??, April 2007. CODEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/april-2007-volume-
32-number-2/autonomic-computing-
freedom-or-threat.

Farrow:2005:BRa

[FH05] Rik Farrow and Chuck Hardin. Book reviews. ;login: the USENIX Association newsletter, 30(1):??, February 2005. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2005-02/openpdfs/
reviews.pdf.

Finke:2000:IAG

[Fin00a] Jon Finke. An improved approach for generating configuration files from a database. In USENIX [USE00g], page ?? ISBN 1-880446-13-8. LCCN ????

URL http://www.usenix.
org/publications/library/
proceedings/lisa2000/finke.
html.

Finley:2000:VS

[Fin00b] Brian Elliot Finley. VA SystemImager. In USE-NIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/als2000/
finley.html.

Finke:2005:WWC

[Fin05] Jon Finke. When worlds collide, 2: The two-sided sword of technology integration. ;login: the USENIX Association newsletter, 30(3): ??, June 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-06/pdfs/finke0506.
pdf.

Fiuczynski:2005:BTK

[Fiu05] Marc E. Fiuczynski. Better tools for kernel evolution, please! ;login: the USENIX Association newsletter, 30(5): ??, October 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-10/pdfs/fiuczynski.
pdf.

Fu:2000:FSD

[FKM00] Kevin Fu, M. Frans Kaashoek, and David Mazières. Fast and secure distributed read-only file system. In USENIX [USE00h], page ?? ISBN 1-880446-16-2. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/osdi2000/
fu.html.

Fu:2002:USS

[FKM02] Kevin Fu, Michael Kaminsky, and David Mazières. Using SFS for a secure network file system. ;login: the USENIX Association newsletter, 27(6): ??, December 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-12/pdfs/Fu.pdf.

Frisch:2008:BR

[FKS08] Æleen Frisch, Brad Knowles, and Sam Stover. Book reviews. ;login: the USENIX Association newsletter, 33(1): ??, February 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
february-2008-volume-33-
number-1/book-reviews.

Fowler:2000:EDF

[FKV00] Glenn S. Fowler, David G.

Korn, and Kiem-Phong Vo. Extended data formatting using Sfio. In USE-NIX [USE00a], pages 107–

116. ISBN 1-880446-22-7.

LCCN ???? URL http:

                                        //www.usenix.org/events/
                                        usenix2000/general/fowler.
                                        html.

Farrow:2006:WL

[FL06] Rik Farrow and Jane-Ellen Long. Writing for ;login;. ;login: the USENIX Association newsletter, 31(5): ??, October 2006. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        october-2006-volume-31-
                                        number-5/writing-login.

Flynt:2000:TSa

[Fly00a] Clif Flynt. The Tclsh spot. ;login: the USENIX Association newsletter, 25(2): ??, April 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-4/features/tclsh.html.

Flynt:2000:TSb

[Fly00b] Clif Flynt. The Tclsh spot. ;login: the USENIX Association newsletter, 25(3): ??, June 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-6/features/tclsh.html.

Flynt:2000:TSc

[Fly00c] Clif Flynt. The Tclsh spot. ;login: the USENIX Association newsletter, 25(5): ??, August 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-8/features/tclsh.html.

Flynt:2000:TSd

[Fly00d] Clif Flynt. The Tclsh spot.

;login: the USENIX Association newsletter, 25(8):??, December 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-12/pdfs/flynt.pdf.

Flynt:2000:BRA

[Fly00e] Clyf Flynt. Book review:

Applying UML and Patterns. ;login: the USENIX Association newsletter, 25(1):??, February 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-2/patterns.html.

Flynt:2001:ID

[Fly01a] Clif Flynt. Intrusion detection. ;login: the USE-NIX Association newsletter, 26(3):??, June 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-06/pdfs/flynt.pdf.

Flynt:2001:TSa

[Fly01b] Clif Flynt. The Tclsh spot.

;login: the USENIX Association newsletter, 26(1):??, February 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.

                                        org/publications/login/
                                        2001-02/pdfs/flynt.pdf.

Flynt:2001:TSb

[Fly01c] Clif Flynt. The Tclsh spot. ;login: the USENIX Association newsletter, 26(2): ??, April 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-04/pdfs/flynt.pdf.

Flynt:2001:TSc

[Fly01d] Clif Flynt. The Tclsh spot. ;login: the USENIX Association newsletter, 26(4): ??, July 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-07/pdfs/flynt.pdf.

Flynt:2001:TSd

[Fly01e] Clif Flynt. The Tclsh spot. ;login: the USENIX Association newsletter, 26(6): ??, October 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-10/pdfs/flynt.pdf.

Flynt:2001:TSe

[Fly01f] Clif Flynt. The Tclsh spot. ;login: the USENIX Association newsletter, 26(8):??, December 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-12/pdfs/flynt.pdf.

Flynt:2002:GEP

[Fly02a]  Clif Flynt.  Generating

Ethernet packets. ;login: the USENIX Association newsletter, 27(5):??, October 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-10/pdfs/flynt.pdf.

Flynt:2002:TSa

[Fly02b] Clif Flynt. The Tclsh spot.

;login: the USENIX Association newsletter, 27(1):??, February 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-02/pdfs/flynt.pdf.

Flynt:2002:TSb

[Fly02c] Clif Flynt. The Tclsh spot.

;login: the USENIX Association newsletter, 27(2): ??, April 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-04/pdfs/flynt.pdf.

Flynt:2002:TSc

[Fly02d] Clif Flynt. The Tclsh spot.

;login: the USENIX Association newsletter, 27(3): ??, June 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-06/pdfs/flynt.pdf.

Flynt:2002:TSd

[Fly02e] Clif Flynt. The Tclsh spot. ;login: the USENIX Association newsletter, 27(4): ??, August 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-08/pdfs/flynt.pdf.

Flynt:2003:TSa

[Fly03a] Clif Flynt. The Tclsh spot. ;login: the USENIX Association newsletter, 28(1):??, February 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-02/pdfs/flynt.pdf.

Flynt:2003:TSb

[Fly03b] Clif Flynt. The Tclsh spot. ;login: the USENIX Association newsletter, 28(2): ??, April 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-04/pdfs/flynt.pdf.

Flynt:2003:TSCa

[Fly03c] Clif Flynt. The Tclsh spot: Capturing and analyzing IP packets on the AX-4000 using AxTcl. ;login: the USE-NIX Association newsletter, 28(3):??, June 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-06/pdfs/flynt.pdf.

Flynt:2003:TSCb

[Fly03d] Clif Flynt. The Tclsh spot:

Client server sockets. ;login: the USENIX Association newsletter, 28(5):??, October 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-10/pdfs/flynt.pdf.

Flynt:2003:TSM

[Fly03e] Clif Flynt. The Tclsh spot:

Monitor your system with expect. ;login: the USENIX Association newsletter, 28(4): ??, August 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-08/pdfs/flynt.pdf.

Flynt:2004:TSa

[Fly04a] Clif Flynt. The Tclsh spot.

;login: the USENIX Association newsletter, 29(1):??, February 2004. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2004-02/pdfs/flynt.pdf.

Flynt:2004:TSc

[Fly04b] Clif Flynt. The Tclsh spot.

;login: the USENIX Association newsletter, 29(4): ??, August 2004. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2004-08/pdfs/flynt.pdf.

Flynt:2004:TSd

[Fly04c] Clif Flynt. The Tclsh spot. ;login: the USENIX Association newsletter, 29(5): ??, October 2004. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2004-10/pdfs/flynt.pdf.

Flynt:2004:TSb

[Fly04d] Cliff Flynt. The Tclsh spot. ;login: the USENIX Association newsletter, 29(3): ??, June 2004. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2004-06/pdfs/flynt.pdf.

Flynt:2004:TSM

[Fly04e]  Cliff Flynt.  The Tclsh

spot: Mobile agents in Tcl. ;login: the USENIX Association newsletter, 29(2): ??, April 2004. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2004-04/pdfs/Flynt.pdf.

Flynt:2005:TS

[Fly05] Clif Flynt. The Tclsh spot. ;login: the USENIX Association newsletter, 30(1):??, February 2005. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2005-02/pdfs/flynt.pdf.

Forrester:2000:ESR

[FM00] Justin E. Forrester and Barton P. Miller. An empirical study of the robustness of Windows NT applications using random testing. In USENIX [USE00d], page ?? ISBN 1-880446-20-0. LCCN

????  URL http://db.
usenix.org/publications/
library/proceedings/usenix-
win2000/forrester.html.

Franke:2001:PSL

[FNKR01] Hubertus Franke, Shailabh Nagar, Mike Kravetz, and Rajan Ravindran. PMQS: Scalable Linux scheduling for high end servers. In USE-NIX [USE01d], page ?? ISBN ???? LCCN ???? URL http:

//www.linuxshowcase.org/
franke.html.

Forte:2000:UTS

[For00] Dario Forte. Using Trinux for security testing. ;login: the USENIX Association newsletter, 25(8):??, December 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-12/pdfs/forte.pdf.

Forte:2001:UTS

[For01] Dario Forte. Using tcpdump and Sanitize for system security. ;login: the USENIX Association newsletter, 26(3): 57–59, June 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-06/pdfs/forte.pdf.

Forte:2003:CIR

[For03] Dario Forte. Coordinated incident response procedures. ;login: the USENIX Association newsletter, 28(6):??, December 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-12/pdfs/forte.pdf.

Fourdan:2000:XLD

[Fou00]  Olivier Fourdan.  Xfce:

a lightweight desktop environment. In USENIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN ????

URL http://www.usenix.
org/publications/library/
proceedings/als2000/fourdan.
html.

Foust:2002:ITU

[Fou02] Robert Foust. Identifying and tracking unauthorized 802.11 cards and access points. ;login: the USENIX Association newsletter, 27(4):32– 43, August 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-08/pdfs/foust.pdf.

Fraser:2001:LMY

[Fra01]  Timothy Fraser.  LO-

MAC: MAC you can live with. In USENIX [USE01f], page ?? ISBN 1-880446-10-3. LCCN QA76.8.U65

U84 2001. URL http://www.
usenix.org/publications/
library/proceedings/usenix01/
freenix01/fraser.html.

Freedman:2006:ASS

[Fre06] Michael J. Freedman. Automating server selection with OASIS. ;login: the USE-NIX Association newsletter, 31(5):??, October 2006. CO-DEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/october-2006-volume-
31-number-5/automating-
server-selection-oasis.

Frisch:2005:BRBa

[Fri05a]  Æleen Frisch.  Book re-

views: The bookworm. ;login: the USENIX Association newsletter, 30(2): ??, April 2005. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2005-04/openpdfs/
bookworm0504.pdf.

Frisch:2005:BRBb

[Fri05b]  Æleen Frisch.  Book re-

views: The bookworm. ;login: the USENIX Association newsletter, 30(3): ??, June 2005. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2005-06/openpdfs/
bookworm0506.pdf.

FitzGibbon:2000:DCM

[FS00] James FitzGibbon and Tim Strike. Distributed computing: Moving from CGI to CORBA. In USE-NIX [USE00a], pages 139–

148. ISBN 1-880446-22-7. LCCN ???? URL http:

                                        //www.usenix.org/events/
                                        usenix2000/general/fitzgibbon.
                                        html.

Frantzen:2001:SHF

[FS01] Mike Frantzen and Mike Shuey. StackGhost: Hardware facilitated stack protection. In USENIX [USE01h], page ?? ISBN 1-880446-18-9, 1-880446-07-3. LCCN QA76.9.A25 U565

                                        2001. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/sec01/
                                        frantzen.html.

Fu:2001:CCA

[FSAK01] Xiaodong Fu, Weisong Shi, Anatoly Akkerman, and Vijay Karamcheti. CANS: Composable, adaptive network services infrastructure. In USENIX [USE01c], page ?? ISBN 1-880446-14-6. LCCN TK5105.875.I57

                                        U96 2001. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/usits01/
                                        fu.html.

Fassino:2002:TSF

[FSLM02] Jean-Philippe Fassino, Jean-Bernard Stefani, Julia Lawall, and Gilles Muller. Think: a software framework for component-based operating system kernels. In USE-NIX [USE02e], page ?? ISBN 1-880446-00-6. LCCN QA76.8.U65 U84 2002. URL

http://www.usenix.org/publications/
library/proceedings/usenix02/
fassino.html.

Fu:2001:DCA

[FSSF01] Kevin Fu, Emil Sit, Kendra Smith, and Nick Feamster. The dos and don’ts of client authentication on the Web. In USENIX [USE01h], page ?? ISBN 1-880446-18-9, 1-880446-07-

3. LCCN QA76.9.A25 U565

2001. URL http://www.
usenix.org/publications/
library/proceedings/sec01/
fu.html.

Fuhriman:2006:IMI

[Fuh06]  Darrell Fuhriman.  Iden-

tity management identified. ;login: the USENIX Association newsletter, 31(4): ??, August 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2006-volume-31-number-
4/identity-management-identified.

Filderman:2001:A

[FZ01] Esther Filderman and Garry Zacheiss. AFS, 2001.

URL http://db.usenix.
org/publications/library/
proceedings/lisa2001/tech/

. Unpublished invited talk, LISA 2001: 15th Systems Administration Conference, December 2–7, 2001, Town and Country Resort Hotel, San Diego, CA.

Galvin:2005:SC

[Gal05]  Peter Baer Galvin.  So-

laris 10 Containers. ;login: the USENIX Association newsletter, 30(5):??, October 2005. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2005-10/pdfs/galvin.pdf.

Galvin:2008:PATa

[Gal08a] Peter Baer Galvin. Pete’s all things sun (PATS): The future of Sun. ;login: the USENIX Association newsletter, 33(1):??, February 2008. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        february-2008-volume-33-
                                        number-1/petes-all-things-
                                        sun-pats-future-sun.

Galvin:2008:PATb

[Gal08b] Peter Baer Galvin. Pete’s all things sun (PATS): The security sheriff. ;login: the USE-NIX Association newsletter, 33(2):??, April 2008. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL https://www.
                                        usenix.org/publications/
                                        login/april-2008-volume-
                                        33-number-2/petes-all-things-
                                        sun-pats-security-sheriff.

Galvin:2008:PATc

[Gal08c] Peter Baer Galvin. Pete’s all things sun (PATS): The state of ZFS. ;login: the USENIX Association newsletter, 33(3):??, June 2008. CODEN LOGNEM. ISSN 1044-6397. URL

https://www.usenix.org/
publications/login/june-
2008-volume-33-number-3/
petes-all-things-sun-pats-
state-zfs.

Galvin:2008:PATd

[Gal08d] Peter Baer Galvin. Pete’s all things Sun: Solaris System Analysis 101. ;login: the USENIX Association newsletter, 33(4):??, August 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2008-volume-33-number-
4/petes-all-things-sun-
solaris-system-analysis-
101.

Galvin:2008:PATe

[Gal08e] Peter Baer Galvin. Pete’s all things Sun: Solaris System Analysis 102. ;login: the USENIX Association newsletter, 33(5):??, October 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
october-2008-volume-33-
number-5/petes-all-things-
sun-solaris-system-analysis-

Galvin:2008:PATf

[Gal08f] Peter Baer Galvin. Pete’s all things Sun: The death of Solaris. ;login: the USENIX Association newsletter, 33(6):51–56, December 2008. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        december-2008-volume-33-
                                        number-6/petes-all-things-
                                        sun-death-solaris.

Galvin:2009:PATa

[Gal09a] Peter Baer Galvin. Pete’s all things Sun: Analytics for the masses (of storage administrators). ;login: the USENIX Association newsletter, 34(1): ??, February 2009. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        february-2009-volume-34-
                                        number-1/petes-all-things-
                                        sun-analytics-masses-storage.

Galvin:2009:PATc

[Gal09b] Peter Baer Galvin. Pete’s all things sun (PATS): AMPing up your Web environment. ;login: the USE-NIX Association newsletter, 34(3):??, June 2009. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL https://www.
                                        usenix.org/publications/
                                        login/june-2009-volume-
                                        34-number-3/petes-all-things-
                                        sun-pats-amping-your-web.

Galvin:2009:PATf

[Gal09c] Peter Baer Galvin. Pete’s all things Sun: Swaddling applications in a security blanket. ;login: the USENIX Association newsletter, 34(6): ??, December 2009. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
december-2009-volume-34-
number-6/petes-all-things-
sun-swaddling-applications.

Galvin:2009:PATd

[Gal09d] Peter Baer Galvin. Pete’s all things Sun: T servers

— why, and why not. ;login: the USENIX Association newsletter, 34(4):??, August 2009. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2009-volume-34-number-
4/petes-all-things-sun-
t-servers%E2%80%94why-and-
why-not.

Galvin:2009:PATb

[Gal09e] Peter Baer Galvin. Pete’s all things Sun: The Sun virtualization guide. ;login: the USENIX Association newsletter, 34(2): ??, April 2009. CODEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/april-2009-volume-
34-number-2/petes-all-things-
sun-sun-virtualization-
guide.

Galvin:2009:PATe

[Gal09f] Peter Baer Galvin. Pete’s all things Sun: VMware vSphere 4 vs. Microsoft Hyper-V R2. ;login: the USENIX Association newsletter, 34(5):58– 67, October 2009. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        october-2009-volume-34-
                                        number-5/petes-all-things-
                                        sun-vmware-vsphere-4-vs.

Garbee:2002:LLP

[Gar02] Bdale Garbee. Linux on laptop/PDA. In USE-NIX [USE02e], page 355. ISBN 1-880446-00-6. LCCN QA76.8.U65 U84 2002. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/usenix02/
                                        tech/techonefile.html. Un-

published guru session talk, 2002 USENIX Annual Technical Conference, June 10-15, 2002, Monterey Conference Center, Monterey, CA.

Garfinkel:2007:CGC

[Gar07] Simson Garfinkel. Commodity grid computing with Amazon’s S3 and EC2. ;login: the USENIX Association newsletter, 32(1):??, February 2007. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        february-2007-volume-32-
                                        number-1/commodity-grid-
                                        computing-amazons-s3-and-
                                        ec2.

Gleitz:2001:TAR

[GB01] Peter M. Gleitz and Steven M.

Bellovin. Transient addressing for related processes: Improved firewalling by using IPV6 and multiple addresses per host. In USENIX [USE01h], page ?? ISBN 1-880446-18-9, 1-880446-07-

3. LCCN QA76.9.A25 U565

2001. URL http://www.
usenix.org/publications/
library/proceedings/sec01/
gleitz.html.

Goel:2004:BMS

[GB04] Sanjay Goel and Stephen F.

Bush. Biological models of security for virus propagation in computer networks. ;login: the USENIX Association newsletter, 29(6):??, December 2004. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2004-12/pdfs/propagation.
pdf.

Gribble:2000:SDD

[GBHC00] Steven D. Gribble, Eric A.

Brewer, Joseph M. Hellerstein, and David Culler. Scalable, distributed data structures for Internet service construction. In USE-NIX [USE00h], page ?? ISBN 1-880446-16-2. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/osdi2000/
gribble.html.

Gilfix:2000:PNA

[GC00] Michael Gilfix and Alva L. Couch. Peep (the network auralizer): Monitoring your network with sound. In USENIX [USE00g], page ?? ISBN 1-880446-13-8. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/lisa2000/
                                        gilfix.html.

Gritter:2001:ACR

[GC01] Mark Gritter and David R. Cheriton. An architecture for content routing support in the Internet. In USENIX [USE01c], page ?? ISBN 1-880446-14-6. LCCN TK5105.875.I57 U96 2001.

                                        URL http://www.usenix.
                                        org/publications/library/
                                        proceedings/usits01/gritter.
                                        html.

Geer:2000:P

[Gee00a] Daniel Geer. From the President. ;login: the USE-NIX Association newsletter, 25(5):??, August 2000. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2000-8/usenixnews.
                                        html#president.

Geer:2000:MP

[Gee00b]  Daniel Geer.  Message

from the President. ;login: the USENIX Association newsletter, 25(4): ??, July 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-7/usenixnews.html#
message.

Geer:2003:CNS

[Gee03a] Dan Geer. Comments of the National Strategy to Secure Cyberspace. ;login: the USE-NIX Association newsletter, 28(2):??, April 2003. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2003-04/pdfs/geer.
pdf.

Geer:2003:GPS

[Gee03b] Dan Geer. Getting the problem statement right. ;login: the USENIX Association newsletter, 28(3): ??, June 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-06/pdfs/geer.pdf.

Geer:2003:PW

[Gee03c]  Dan Geer.  Patch work.

;login: the USENIX Association newsletter, 28(4): ??, August 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-08/pdfs/geer.pdf.

Geer:2005:MBE

[Gee05]  Dan Geer.  Monoculture

on the back of the envelope. ;login: the USENIX Association newsletter, 30(6):

??, December 2005. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2005-12/openpdfs/
geer.pdf.

Geer:2007:QLF

[Gee07] Dan Geer. A quant looks at the future. ;login: the USE-NIX Association newsletter, 32(3):??, June 2007. CO-DEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/june-2007-volume-
32-number-3/quant-looks-
future.

Gerbarg:2002:ASM

[Ger02] Louis G. Gerbarg. Advanced synchronization in Mac OS X: Extending UNIX to SMP and real-time. In USE-NIX [USE02a], pages 37–

45. ISBN 1-880446-02-2. LCCN QA76.76.O63 B736

2002. URL http://www.
usenix.org/publications/
library/proceedings/bsdcon02/
gerbarg.html.

Gettys:2002:FCW

[Get02a]  James Gettys.  The fu-

ture is coming: Where the X Window System should go. In USENIX [USE02d], page ?? ISBN 1-880446-01-

4. LCCN QA76.8.U65 P765

2002. URL http://www.
usenix.org/publications/
library/proceedings/usenix02/
tech/freenix/gettys.html.

Gettys:2002:GR

[Get02b] Jim Gettys. General/random.

In USENIX [USE02e], page

355. ISBN 1-880446-00-

6. LCCN QA76.8.U65 U84

2002. URL http://www.
usenix.org/publications/
library/proceedings/usenix02/
tech/techonefile.html. Un-

published guru session talk, 2002 USENIX Annual Technical Conference, June 10-15, 2002, Monterey Conference Center, Monterey, CA.

Golm:2002:JOS

[GFWK02] Michael Golm, Meik Felser, Christian Wawersich, and Jürgen Kleinoeder. The JX operating system. In USENIX [USE02e], page ?? ISBN 1-880446-00-6. LCCN QA76.8.U65 U84 2002. URL

http://www.usenix.org/publications/
library/proceedings/usenix02/
golm.html.

Gagnon:2001:SRF

[GH01] Etienne M. Gagnon and Laurie J. Hendren. SableVM: a research framework for the efficient execution of Java bytecode. In USENIX [USE01g], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38

J42 2001. URL http://www.
usenix.org/publications/
library/proceedings/jvm01/
gagnon.html.

Gobel:2006:AHB

[GHH06] Jan Göbel, Jens Hektor, and Thorsten Holz. Advanced honeypot-based intrusion detection. ;login: the USENIX Association newsletter, 31(6): ??, December 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
december-2006-volume-31-
number-6/advanced-honeypot-
based-intrusion-detection.

Gibbs:2000:IT

[Gib00]  Simon Gibbs.  Interac-

tion with TV in 2003,

2000. URL http://www.
usenix.org/publications/
library/proceedings/osdi2000/
wiess2000/invitedtalks/

gibbs.pdf. Unpublished invited talk at First Workshop on Industrial Experiences with Systems Software (WIESS 2000), October 22, 2000, Paradise Point Resort, San Diego, California, USA.

Gilfix:2001:IPN

[Gil01] Michael Gilfix. Introducing Peep: The network auralizer. ;login: the USENIX Association newsletter, 26(2): ??, April 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-04/pdfs/gilfix.pdf.

Giffin:2002:DMR

[GJM02] Jonathon T. Giffin, Somesh Jha, and Barton P. Miller. Detecting manipulated remote call streams. In USE-NIX [USE02b], pages 61–79.

ISBN 1-931971-00-5. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/sec02/
giffin.html.

Goldsmith:2004:CAI

[GK04] Clair W. Goldsmith and Rob Kolstad. A conversation about identity management. ;login: the USENIX Association newsletter, 29(5): ??, October 2004. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2004-10/pdfs/interview.
pdf.

Glass:2002:LMB

[Gla02] Brett Glass. Log monitors in BSD UNIX. In USE-NIX [USE02a], pages 131–

142. ISBN 1-880446-02-2. LCCN QA76.76.O63 B736

2002. URL http://www.
usenix.org/publications/
library/proceedings/bsdcon02/
glass.html.

Gleim:2002:JPI

[Gle02]  Urs Gleim.  JaRTS: a

portable implementation of real-time core extensions for Java. In USENIX [USE02f], page ?? ISBN 1-931971-01-3. LCCN QA76.73 .J38

J42 2002. URL http://www.
usenix.org/publications/
library/proceedings/javavm02/
gleim.html.

Grunwald:2000:PDC

[GLF^{+}00] Dirk Grunwald, Philip Levis, Keith I. Farkas, Charles B. Morrey III, and Michael Neufeld. Policies for dynamic clock scheduling. In USENIX [USE00h], page ?? ISBN 1-880446-16-2. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/osdi2000/
grunwald.html.

Gupta:2009:DE

[GLV^{+}09] Diwaker Gupta, Sangmin Lee, Michael Vrable, Stefan Savage, Alex C. Snoeren, George Varghese, Geoffrey M. Voelker, and Amin Vahdat. Difference engine. ;login: the USENIX Association newsletter, 34(2): 24–31, April 2009. CO-DEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/april-2009-volume-
34-number-2/difference-
engine.

Garg:2001:ASG

[GMAM01] Rahul Garg, Parul Mittal, Vikas Agarwal, and Natwar Modani. An architecture for secure generation and verification of electronic coupons. In USE-NIX [USE01b], page ?? ISBN 1-880446-09-X. LCCN QA76.8.U65 U84 2001. URL

http://www.usenix.org/publications/
library/proceedings/usenix01/
garg.html.

Goldberg:2000:PDT

[Gol00]  Ian Goldberg.  Privacy-

degrading technologies: How not to build the future,

2000. URL http://www.
usenix.org/publications/
library/proceedings/sec2000/
invitedtalks.html. Unpub-

lished invited talk at Ninth USENIX Security Symposium.

Golubitsky:2006:SSF

[Gol06]  Chaos Golubitsky.  Sim-

ple software flow analysis using GNU cflow. ;login: the USENIX Association newsletter, 31(2):37– 41, April 2006. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2006-04/pdfs/golubitsky.
pdf.

Gosling:2000:JLR

[Gos00] James Gosling. JAVA: a language for the real world,

2000. URL http://www.
usenix.org/publications/
library/proceedings/osdi2000/
wiess2000/invitedtalks/
gosling_html. Unpublished

invited talk at First Workshop on Industrial Experiences with Systems Software (WIESS 2000), October 22, 2000, Paradise Point Resort, San Diego, California, USA.

Gettys:2001:XRR

[GP01a] Jim Gettys and Keith Packard.

The X Resize and Rotate Extension — RandR. In USENIX [USE01f], page ?? ISBN 1-880446-10-3. LCCN QA76.8.U65 U84 2001. URL

http://www.usenix.org/publications/
library/proceedings/usenix01/
freenix01/gettys.html.

Gil:2001:MDS

[GP01b] Thomer M. Gil and Massimiliano Poletto. MULTOPS: a data-structure for bandwidth attack detection. In USENIX [USE01h], page ?? ISBN 1-880446-18-9, 1-880446-07-

3. LCCN QA76.9.A25 U565

2001. URL http://www.
usenix.org/publications/
library/proceedings/sec01/
gil.html.

Griffin:2007:XIC

[GR07] Leigh Griffin and John Ronan. Xen installation and configuration. ;login: the USENIX Association newsletter, 32(1):??, February 2007. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
february-2007-volume-32-
number-1/xen-installation-
and-configuration.

Granroth:2000:UKC

[Gra00a]  Kurt Granroth.  Using

KDE components (KParts),

2000. URL http://www.
usenix.org/publications/
library/proceedings/als2000/

. Unpublished invited talk at Annual Linux Showcase 2000.

Gray:2000:PTS

[Gra00b] Bob Gray. Performance tuning with source code UNIX. ;login: the USENIX Association newsletter, 25(2): ??, April 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-4/features/performance.
html.

Gray:2000:SCU

[Gra00c]  Bob Gray.  Source code

UNIX. ;login: the USENIX Association newsletter, 25(1): ??, February 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-2/features/sourcecode.
html.

Gray:2000:TOS

[Gra00d] Bob Gray. Teaching operating systems with source code UNIX. ;login: the USE-NIX Association newsletter, 25(4):??, July 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-7/features/teaching.
html.

Gray:2001:LTF

[Gra01a] Bob Gray. A logging and tracing facility for an embedded source code UNIX product. ;login: the USE-NIX Association newsletter, 26(2):??, April 2001. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2001-04/pdfs/gray.
                                        pdf.

Gray:2001:STY

[Gra01b] Bob Gray. Searching through your files with Glimpse. ;login: the USENIX Association newsletter, 26(8):??, December 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-12/pdfs/gray.pdf.

Gray:2001:SOR

[Gra01c]  Peter D. Gray.  SUS,

an object reference model for distributing UNIX super user privileges. In USENIX [USE01a], page ?? ISBN 1-880446-05-7. LCCN

                                        ????  URL http://db.
                                        usenix.org/publications/
                                        library/proceedings/lisa2001/
                                        tech/gray.html.

Gray:2002:SBA

[Gra02] Jim Gray. Storage bricks have arrived. In USE-NIX [USE02c], page ?? ISBN 1-880446-03-0. LCCN QA76.9.F5 C65 2002. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/fast02/

.

Granick:2004:SB

[Gra04] Jennifer S. Granick. Strike back. ;login: the USENIX Association newsletter, 29(6): ??, December 2004. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2004-12/pdfs/strike.pdf.

Grossman:2009:TWH

[Gro09] Jeremiah Grossman. Top 10 Web hacking techniques: ‘what’s possible, not probable’. ;login: the USENIX Association newsletter, 34(6): ??, December 2009. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
december-2009-volume-34-
number-6/top-10-web-hacking-
techniques-whats-possible-
not.

Gallesio:2002:BWL

[GS02] Erick Gallesio and Manuel Serrano. Biglook: a widget library for the Scheme programming language. In USENIX [USE02d], page ?? ISBN 1-880446-01-4. LCCN QA76.8.U65 P765 2002.

URL http://www.usenix.
org/publications/library/
proceedings/usenix02/tech/
freenix/gallesio.html.

Gogoulos:2006:ULL

[GS06] Markos Gogoulos and Diomidis Spinellis. Using Linux Live CDs for penetration testing. ;login: the USE-NIX Association newsletter, 31(3):??, June 2006. CO-DEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/

                                        login/june-2006-volume-
                                        31-number-3/using-linux-
                                        live-cds-penetration-testing.

Gupta:2004:MCB

[GSAA04] Abhishek Gupta, Ozgur D. Sahin, Divyakant Agrawal, and Amr El Abbadi. Meghdoot: Content-based publish/subscribe over P2P networks. In Jacobsen [Jac04], pages 254–?? CODEN LNCSD9. ISBN 3-540-23428-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.M54 I33 2004. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.springerlink.
                                        com/openurl.asp?genre=issue&
                                        issn=0302-9743&volume=3231;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=volume&
                                        id=doi:10.1007/b101561.

Gschwind:2001:PCP

[Gsc01] Thomas Gschwind. PSTL — a C++ Persistent Standard Template Library. In USENIX [USE01e], page ?? ISBN 1-880446-12-X. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/coots01/
                                        gschwind.html.

Griffin:2000:OSM

[GSGN00] John Linwood Griffin, Steven W. Schlosser, Gregory R. Ganger, and David F. Nagle. Operating system management of MEMS-based storage devices. In USE-NIX [USE00h], page ?? ISBN 1-880446-16-2. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/osdi2000/
griffin.html.

Griffin:2002:TAS

[GSS^{+}02] John Linwood Griffin, Jiri Schindler, Steven W. Schlosser, John C. Bucy, and Gregory R. Ganger. Timingaccurate storage emulation. In USENIX [USE02c], page ?? ISBN 1-880446-03-

0. LCCN QA76.9.F5 C65

2002. URL http://www.
usenix.org/publications/
library/proceedings/fast02/
griffin.html.

Grier:2008:BMS

[GTK08] Chris Grier, Shuo Tang, and Samuel T. King. Building a more secure Web browser. ;login: the USENIX Association newsletter, 33(4): ??, August 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2008-volume-33-number-
4/building-more-secure-
web-browser.

Gutmann:2000:OSC

[Gut00] Peter Gutmann. An opensource cryptographic coprocessor. In USENIX [USE00i], page ?? ISBN 1-880446-18-9. LCCN ????

URL http://www.usenix.

                                        org/publications/library/
                                        proceedings/sec2000/gutmann.
                                        html.

Gutmann:2001:DRS

[Gut01]  Peter Gutmann.  Data

remanence in semiconductor devices. In USENIX [USE01h], page ?? ISBN 1-880446-18-9, 1-880446-07-3. LCCN QA76.9.A25 U565

                                        2001. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/sec01/
                                        gutmann.html.

Gutmann:2002:LLI

[Gut02] Peter Gutmann. Lessons learned in implementing and deploying crypto software. In USENIX [USE02b], pages 315–325. ISBN 1-931971-00-5. LCCN ????

                                        URL http://www.usenix.
                                        org/publications/library/
                                        proceedings/sec02/gutmann.
                                        html.

Guy:2001:SQB

[Guy01] Alexander Guy. Salamander: The quest to build a useful handheld computing environment. In USENIX [USE01d], page ?? ISBN ???? LCCN ???? URL http:

                                        //www.linuxshowcase.org/
                                        guy.html.

Goel:2002:GVT

[GW02] Ashvin Goel and Jonathan Walpole. Gscope: a visualization tool for timesensitive software. In USE-NIX [USE02d], page ??

ISBN 1-880446-01-4. LCCN QA76.8.U65 P765 2002.

URL http://www.usenix.
org/publications/library/
proceedings/usenix02/tech/
freenix/goel.html.

Garfinkel:2007:WVC

[GW07] Tal Garfinkel and Andrew Warfield. What virtualization can do for security. ;login: the USENIX Association newsletter, 32(6):??, December 2007. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
december-2007-volume-32-
number-6/what-virtualization-
can-do-security.

Gwinn:2000:PSS

[Gwi00] Joe Gwinn. POSIX System Services Working Group

— Realtime (SSWG-RT). ;login: the USENIX Association newsletter, 25(5): ??, August 2000. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/standards/47.SSWG-
RT.html.

Hardt:2000:PPZ

[HAE00] Dick Hardt, Gisle Aas, and Paul Everitt. Perl, Python and Zope. In USENIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/als2000/
aas.html.

Hall:2000:EPPc

[Hal00a]  Joseph Hall.  Effective

Perl programming. ;login: the USENIX Association newsletter, 25(4): ??, July 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-7/features/effective.
                                        html.

Hall:2000:EPPa

[Hal00b]  Joseph N. Hall.  Effec-

tive Perl programming. ;login: the USENIX Association newsletter, 25(1):??, February 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-2/features/effectiveperl.
                                        html.

Hall:2000:EPPb

[Hal00c]  Joseph N. Hall.  Effec-

tive Perl programming. ;login: the USENIX Association newsletter, 25(2): ??, April 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-4/features/perl.html.

Halprin:2000:ORM

[Hal00d]  Geoff Halprin.  An ode

to risk management. ;login: the USENIX Association newsletter, 25(6):??, October 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-10/pdfs/odetoriskmanagement.
pdf.

Hammel:2000:SAL

[Ham00] Michael J. Hammel. The state of the arts — Linux tools for the graphic artist. In USENIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/als2000/
hammel.html.

Hamilton:2001:LIC

[Ham01]  Jennifer Hamilton.  Lan-

guage integration in the Common Language Runtime,

2001.  URL http://db.
usenix.org/publications/
library/proceedings/coots01/
invited_talks/hamilton_

html/. Unpublished invited talk, 6th USENIX Conference on Object-Oriented Technologies and Systems, January 29–February 2, 2001, San Antonio, Texas, USA.

Hankins:2003:IBG

[Han03] Greg Hankins. Introduction to the Border Gateway Protocol. ;login: the USENIX Association newsletter, 28(1): ??, February 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-02/pdfs/hankins.pdf.

Hardin:2002:BRB

[Har02a] Chuck Hardin. Book review: Brian Shea: Have You Locked the Castle Gates? ;login: the USENIX Association newsletter, 27(4): ??, August 2002. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2002-08/openpdfs/books.
                                        pdf.

Harris:2002:AMK

[Har02b] Trey Harris. Ask Mr. Know-It-All. ;login: the USENIX Association newsletter, 27(1): ??, February 2002. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2002-02/pdfs/harris.pdf.

Hartmeier:2002:DPO

[Har02c] Daniel Hartmeier. Design and performance of the OpenBSD stateful packet filter (pf). In USENIX [USE02d], page ?? ISBN 1-880446-01-4. LCCN QA76.8.U65 P765 2002.

                                        URL http://www.usenix.
                                        org/publications/library/
                                        proceedings/usenix02/tech/
                                        freenix/hartmeier.html.

Haskins:2000:MMM

[Has00]  Robert Haskins.  Mail

message metering. ;login: the USENIX Association newsletter, 25(8):??, December 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-12/pdfs/haskins.pdf.

Haskins:2001:Ia

[Has01a] Robert Haskins. ISPAdmin.

;login: the USENIX Association newsletter, 26(1):??, February 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-02/pdfs/haskins.pdf.

Haskins:2001:Ib

[Has01b]  Robert Haskins.  ISPAd-

min. ;login: the USENIX Association newsletter, 26(2): ??, April 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-04/pdfs/haskins.pdf.

Haskins:2001:Ic

[Has01c]  Robert Haskins.  ISPAd-

min. ;login: the USENIX Association newsletter, 26(3): ??, June 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-06/pdfs/haskins.pdf.

Haskins:2001:Id

[Has01d]  Robert Haskins.  ISPAd-

min. ;login: the USENIX Association newsletter, 26 (4):??, July 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-07/pdfs/haskins.pdf.

Haskins:2001:Ie

[Has01e]  Robert Haskins.  ISPAd-

min. ;login: the USENIX Association newsletter, 26(6): ??, October 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-10/pdfs/haskins.pdf.

Haskins:2001:If

[Has01f] Robert Haskins. ISPAdmin. ;login: the USENIX Association newsletter, 26(8):??, December 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-12/pdfs/haskins.pdf.

Haskins:2002:Ia

[Has02a]  Robert Haskins.  ISPAd-

min. ;login: the USENIX Association newsletter, 27(2): ??, April 2002. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2002-04/pdfs/haskins.pdf.

Haskins:2002:Ib

[Has02b]  Robert Haskins.  ISPAd-

min. ;login: the USENIX Association newsletter, 27(3): ??, June 2002. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2002-06/pdfs/haskins.pdf.

Haskins:2002:Ic

[Has02c]  Robert Haskins.  ISPAd-

min. ;login: the USENIX Association newsletter, 27(4): ??, August 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-08/pdfs/haskins.pdf.

Haskins:2002:Id

[Has02d]  Robert Haskins.  ISPAd-

min. ;login: the USENIX Association newsletter, 27(5): ??, October 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-10/pdfs/haskins.pdf.

Haskins:2003:ISPa

[Has03a] Robert Haskins. ISPadmin:

Service provider book reviews. ;login: the USENIX Association newsletter, 28(1): ??, February 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-02/pdfs/haskins.pdf.

Haskins:2003:ISPb

[Has03b] Robert Haskins. ISPadmin:

Service provider book reviews II. ;login: the USENIX Association newsletter, 28(4): ??, August 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-08/pdfs/haskins.pdf.

Haskins:2003:IMP

[Has03c] Robert D. Haskins. ISPadmin: Managing and providing ISP services. ;login: the USENIX Association newsletter, 28(5):??, October 2003. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2003-10/pdfs/haskins.pdf.

Haskins:2004:I

[Has04a]  Robert Haskins.  ISPad-

min. ;login: the USENIX Association newsletter, 29(3): ??, June 2004. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2004-06/pdfs/haskins.pdf.

Haskins:2004:IIP

[Has04b] Robert Haskins. ISPadmin: An interview with Paul Graham. ;login: the USENIX Association newsletter, 29(1): ??, February 2004. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2004-02/pdfs/haskins.pdf.

Haskins:2004:IRS

[Has04c] Robert Haskins. ISPadmin: Recent spam-fighting developments. ;login: the USENIX Association newsletter, 29(4): ??, August 2004. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2004-08/pdfs/haskins.pdf.

Haskins:2005:I

[Has05a] Robert Haskins. ISPadmin. ;login: the USENIX Association newsletter, 30(1):??, February 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-02/pdfs/haskins.pdf.

Haskins:2005:IEH

[Has05b]  Robert Haskins.  ISPad-

min: Embedded hardware. ;login: the USENIX Association newsletter, 30(5): ??, October 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-10/pdfs/haskins.pdf.

Haskins:2005:IIV

[Has05c] Robert Haskins. ISPadmin:

Interview with Vipul Ved Prakash. ;login: the USE-NIX Association newsletter, 30(2):??, April 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-04/pdfs/haskins0504.
pdf.

Haskins:2005:IUM

[Has05d]  Robert Haskins.  ISPad-

min: Understanding and mitigating DDoS attacks. ;login: the USENIX Association newsletter, 30(4): ??, August 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-08/pdfs/haskins.pdf.

Haskins:2006:IAS

[Has06a]  Robert Haskins.  ISPad-

min: Anti-spam roundup. ;login: the USENIX Association newsletter, 31(5): ??, October 2006. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        october-2006-volume-31-
                                        number-5/ispadmin-anti-
                                        spam-roundup.

Haskins:2006:IBN

[Has06b] Robert Haskins. ISPadmin: Blocking non-email spam. ;login: the USENIX Association newsletter, 31(1):??, February 2006. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2006-02/pdfs/haskins.pdf.

Haskins:2006:IPE

[Has06c]  Robert Haskins.  ISPad-

min: Policy enforcement. ;login: the USENIX Association newsletter, 31(3):??, June 2006. CODEN LOGNEM. ISSN 1044-6397. URL

                                        https://www.usenix.org/
                                        publications/login/june-
                                        2006-volume-31-number-3/
                                        ispadmin-policy-enforcement.

Haskins:2006:ITS

[Has06d]  Robert Haskins.  ISPad-

min: Traffic shaping. ;login: the USENIX Association newsletter, 31(4):??, August 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2006-volume-31-number-
4/ispadmin-traffic-shaping.

Haskins:2006:IW

[Has06e] Robert Haskins. ISPadmin:

Wireless. ;login: the USENIX Association newsletter, 31(6): ??, December 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
december-2006-volume-31-
number-6/ispadmin-wireless.

Haskins:2007:IDS

[Has07a]  Robert Haskins.  ISPad-

min: DHCP services. ;login: the USENIX Association newsletter, 32(1):??, February 2007. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
february-2007-volume-32-
number-1/ispadmin-dhcp-
services.

Haskins:2007:VIS

[Has07b] Robert Haskins. VoIP in an Internet Service Provider network. ;login: the USE-NIX Association newsletter, 32(2):??, April 2007. CO-DEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/april-2007-volume-
32-number-2/voip-internet-
service-provider-network.

Haynes:2006:IZ

[Hay06a]  Tom Haynes.  Introduc-

tion to ZFS. ;login: the USENIX Association newsletter, 31(3):??, June 2006. CODEN LOGNEM. ISSN 1044-6397. URL

                                        https://www.usenix.org/
                                        publications/login/june-
                                        2006-volume-31-number-3/
                                        introduction-zfs.

Haynes:2006:OM

[Hay06b]  Tom Haynes.  OpenSo-

laris: The model. ;login: the USENIX Association newsletter, 31(2):5– 8, April 2006. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2006-04/pdfs/haynes.pdf.

Hess:2001:ADO

[HBC^{+}01] Christopher K. Hess, Francisco Ballesteros, Rey Juan Carlos, Roy H. Campbell, and M. Dennis Mickunas. An adaptive data object service for pervasive computing environments. In USENIX [USE01e], page ?? ISBN 1-880446-12-X. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/coots01/
                                        hess.html.

Herder:2006:MSP

[HBG^{+}06] Jorrit N. Herder, Herbert Bos, Ben Gras, Philip Homburg, and Andrew S. Tanenbaum. Modular system programming in MINIX 3. ;login: the USENIX Association newsletter, 31(2):19–28, April 2006. CODEN LOGNEM. ISSN 1044-6397. URL

http://www.minix3.org/;
 http://www.usenix.org/
publications/login/2006-
04/openpdfs/herder.pdf.

Herder:2007:RFR

[HBG^{+}07] Jorrit N. Herder, Herbert Bos, Ben Gras, Philip Homburg, and Andrew S. Tanenbaum. Roadmap to a failureresilient operating system. ;login: the USENIX Association newsletter, 32(1):??, February 2007. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
february-2007-volume-32-
number-1/roadmap-failure-
resilient-operating-system.

Hohndel:2000:DDE

[HC00] Dirk Hohndel and Robin Cutshaw. Developing drivers and extensions for XFree86-4.x. In USENIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN ????

URL http://www.usenix.
org/publications/library/
proceedings/als2000/hohndel.
html.

Huang:2001:CDE

[HcC01] Lan Huang and Tzi cker Chiueh. Charm: An I/Odriven execution strategy for high-performance transaction processing. In USE-NIX [USE01b], page ?? ISBN 1-880446-09-X. LCCN QA76.8.U65 U84 2001. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/usenix01/
                                        huang.html.

Hay:2001:CNT

[HcFG01] Jeffrey R. Hay, Wu chun Feng, and Mark K. Gardner. Capturing network traffic with a MAGNeT. In USE-NIX [USE01d], page ?? ISBN ???? LCCN ???? URL http:

                                        //www.linuxshowcase.org/
                                        hay.html.

Hume:2002:NLC

[HD02] Andrew Hume and Scott Daniels. Ningaui: a Linux cluster for business. In USENIX [USE02d], page ?? ISBN 1-880446-01-

4. LCCN QA76.8.U65 P765

                                        2002. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/usenix02/
                                        tech/freenix/hume.html.

Hume:2000:GTV

[HDM00] Andrew Hume, Scott Daniels, and Angus MacLellan. Gecko: Tracking a very large billing system. In USENIX [USE00a], pages 93–105. ISBN 1-880446-22-7. LCCN ????

                                        URL http://www.usenix.
                                        org/events/usenix2000/general/
                                        hume.html.

Hu:2008:SVO

[HDM08] Wenjin Hu, Todd Deshane, and Jeanna Matthews. Solaris virtualization options. ;login: the USENIX Association newsletter, 33(5): ??, October 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
october-2008-volume-33-
number-5/solaris-virtualization-
options.

Hsieh:2001:REI

[HEB01] Wilson C. Hsieh, Dawson Engler, and Godmar Back. Reverse-engineering instruction encodings. In USE-NIX [USE01b], page ?? ISBN 1-880446-09-X. LCCN QA76.8.U65 U84 2001. URL

http://www.usenix.org/publications/
library/proceedings/usenix01/
hsieh.html.

Heiser:2005:SES

[Hei05] Gernot Heiser. Secure embedded systems need microkernels. ;login: the USENIX Association newsletter, 30(6): ??, December 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-12/pdfs/heiser.pdf.

Heiser:2007:YSS

[Hei07] Gernot Heiser. Your system is secure? Prove it! ;login: the USENIX Association newsletter, 32(6):??, December 2007. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/

                                        december-2007-volume-32-
                                        number-6/your-system-secure-
                                        prove-it.

Hellwig:2009:XBS

[Hel09] Christoph Hellwig. XFS: The big storage file system for Linux. ;login: the USENIX Association newsletter, 34(5):10–18, October 2009. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        october-2009-volume-34-
                                        number-5/xfs-big-storage-
                                        file-system-linux.

Hemmerich:2000:ARB

[Hem00] Christopher Hemmerich. Automating request-based software distribution. In USE-NIX [USE00g], page ?? ISBN 1-880446-13-8. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/lisa2000/
                                        hemmerich.html.

Henon:2006:EM

[Hén06]  Alain Hénon.  Extreme

makeover. ;login: the USE-NIX Association newsletter, 31(1):??, February 2006. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2006-02/pdfs/henon.
                                        pdf.

Herzog:2006:PH

[Her06] Pete Herzog. The protocol historian. ;login: the USE-NIX Association newsletter, 31(4):??, August 2006. CO-DEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/august-2006-volume-
31-number-4/protocol-historian.

Hohmuth:2001:PNS

[HH01] Michael Hohmuth and Hermann Härtig. Pragmatic nonblocking synchronization for real-time systems. In USENIX [USE01b], page ?? ISBN 1-880446-09-X. LCCN QA76.8.U65 U84 2001. URL

http://www.usenix.org/publications/
library/proceedings/usenix01/
hohmuth.html.

Hill:2000:KII

[Hil00] Paul B. Hill. Kerberos interoperability issues. In USENIX [USE00b], page ?? ISBN 1-880446-19-7. LCCN

????  URL http://db.
usenix.org/publications/
library/proceedings/lisa-
nt2000/hill.html.

Hallyn:2000:DTE

[HK00a] Serge Hallyn and Phil Kearns.

Domain and type enforcement for Linux. In USE-NIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/als2000/
hallyn.html.

Howell:2000:EEA

[HK00b] Jon Howell and David Kotz.

End-to-end authorization. In USENIX [USE00h], page ?? ISBN 1-880446-16-2. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/osdi2000/
howell.html.

Haskins:2003:IBS

[HK03] Robert Haskins and Rob Kolstad. ISPadmin: Bayesian spam-filtering techniques. ;login: the USENIX Association newsletter, 28(3): ??, June 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-06/pdfs/haskins.pdf.

Haeberlen:2000:SCP

[HLU^{+}00] Andreas Haeberlen, Jochen Liedtke, Karlsruhe University, Yoonho Park, Lars Reuther, and Volkmar Uhlig. Stub-code performance is becoming important. In USENIX [USE00f], page ?? ISBN 1-880446-15-

4. LCCN QA76.76.S95 W67

2000. URL http://www.
usenix.org/publications/
library/proceedings/osdi2000/
wiess2000/haeberlen.html.

Hobbit:2002:HA

[Hob02] Hobbit. High availability. ;login: the USENIX Association newsletter, 27(4): ??, August 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-08/pdfs/hobbit.pdf.

Hobbit:2005:DBS

[Hob05] Hobbit. DNS-based spam rejection. ;login: the USENIX Association newsletter, 30(5): ??, October 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-10/openpdfs/hobbit.
pdf.

Hohndel:2001:RDT

[Hoh01] Dirk Hohndel. Reviving dead technology: How XFree86 and open source prevailed against the X Consortium,

2001. URL http://www.
linuxshowcase.org/tech.

html. Unpublished invited talk, 5th Annual Linux Showcase and Conference, November 5–10, Oakland, CA.

Holz:2005:SB

[Hol05] Thorsten Holz. Spying with bots. ;login: the USENIX Association newsletter, 30(6): ??, December 2005. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2005-12/pdfs/holz.
pdf.

Hope:2001:CR

[Hop01] Paco Hope. Certification revocation. ;login: the USENIX Association newsletter, 26(8): ??, December 2001. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/

                                        login/2001-12/pdfs/hope.
                                        pdf.

Hope:2002:UJF

[Hop02] Paco Hope. Using jails in FreeBSD for fun and profit. ;login: the USENIX Association newsletter, 27(3): ??, June 2002. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2002-06/pdfs/hope.pdf.

Hoskins:2003:OHB

[Hos03] Matthew Hoskins. Oracle hot backup. ;login: the USENIX Association newsletter, 28(4): ??, August 2003. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2003-08/pdfs/hoskins.pdf.

Howard:2003:P

[How03] Travis Howard. .profile. ;login: the USENIX Association newsletter, 28(3): ??, June 2003. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2003-06/pdfs/howard.pdf.

Howard:2004:CT

[How04] Travis Howard. CFS travails. ;login: the USENIX Association newsletter, 29(4): ??, August 2004. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2004-08/pdfs/howard.pdf.

Howe:2005:SMS

[How05] Anthony Howe. Shoot the messenger: Some techniques for spam control. ;login: the USENIX Association newsletter, 30(3): ??, June 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-06/pdfs/howe0506.pdf.

Howard:2006:HOS

[How06a] Mike Howard. How often should you change your password? ;login: the USENIX Association newsletter, 31(6): ??, December 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
december-2006-volume-31-
number-6/how-often-should-
you-change-your-password.

Howard:2006:MYS

[How06b] Mike Howard. Maybe you should use Python. ;login: the USENIX Association newsletter, 31(5):??, October 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
october-2006-volume-31-
number-5/maybe-you-should-
use-python.

Hovemeyer:2001:MEN

[HP01] David Hovemeyer and William Pugh. More efficient network class loading through bundling. In USENIX [USE01g], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38

J42 2001. URL http://www.
usenix.org/publications/
library/proceedings/jvm01/
hovemeyer.html.

Handley:2001:NID

[HPK01] Mark Handley, Vern Paxson, and Christian Kreibich. Network intrusion detection: Evasion, traffic normalization, and end-to-end protocol semantics. In USENIX [USE01h], page ?? ISBN 1-880446-18-9, 1-880446-07-

3. LCCN QA76.9.A25 U565

2001. URL http://www.
usenix.org/publications/
library/proceedings/sec01/
handley.html.

Huang:2002:IWF

[HPS02] Hai Huang, Padmanabhan Pillai, and Kang G. Shin. Improving waitfree algorithms for interprocess communication in embedded real-time systems. In USENIX [USE02e], page ?? ISBN 1-880446-00-6. LCCN QA76.8.U65

U84 2002. URL http://www.
usenix.org/publications/
library/proceedings/usenix02/
huang.html.

Heursch:2001:RRL

[HR01a] Arnd Christian Heursch and Helmut Rzehak. Rapid reaction Linux: Linux with low latency and high timing accuracy. In USENIX [USE01d], page ?? ISBN ???? LCCN ???? URL http:

//www.linuxshowcase.org/
heursch.html.

Holmgren:2001:LCT

[HR01b] Don Holmgren and Ron Rechenmacher. Large clusters for theoretical physics at Fermi Lab. ;login: the USE-NIX Association newsletter, 26(5):??, August 2001. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2001-08/pdfs/holmgren.
pdf.

Holmwood:2000:PSA

[HRF00] John Holmwood, Kevin Reichert, and Blaine Feniak. Providing secure access to information using the Internet. In USE-NIX [USE00b], page ?? ISBN 1-880446-19-7. LCCN

????  URL http://db.
usenix.org/publications/
library/proceedings/lisa-
nt2000/holmwood.html.

Harris:2000:FTD

[HS00a] David K. Z. Harris and Bryan Stansell. Finding time to do it all. ;login: the USENIX Association newsletter, 25(6): ??, October 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-10/pdfs/findingtimetodoitall.
pdf.

Heidemann:2000:LAS

[HS00b] John Heidemann and Dhaval Shah. Location-aware scheduling with minimal infrastructure. In USENIX [USE00a], pages 131–138. ISBN 1-880446-22-7. LCCN ???? URL http:/

                                        /www.usenix.org/events/
                                        usenix2000/general/heidemann.
                                        html.

Haugerud:2001:SUD

[HS01] Hårek Haugerud and Sigmund Straumsnes. Simulation of user-driven computer behaviour. In USE-NIX [USE01a], page ?? ISBN 1-880446-05-7. LCCN

                                        ????  URL http://db.
                                        usenix.org/publications/
                                        library/proceedings/lisa2001/
                                        tech/haugerud.html.

Hasan:2009:SPP

[HSW09] Ragib Hasan, Radu Sion, and Marianne Winslett. Secure provenance: Protecting the genealogy of bits. ;login: the USENIX Association newsletter, 34(3):??, June 2009. CODEN LOGNEM. ISSN 1044-6397. URL

                                        https://www.usenix.org/
                                        publications/login/june-
                                        2009-volume-34-number-3/
                                        secure-provenance-protecting-
                                        genealogy-bits.

Hughes:2000:UCA

[Hug00]  Douglas Hughes.  User-

centric account management with heterogeneous password changing. In USE-NIX [USE00g], page ?? ISBN 1-880446-13-8. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/lisa2000/
hughes.html.

Hume:2000:LTA

[Hum00a] Andrew Hume. So long, and thanks for all the fish. ;login: the USENIX Association newsletter, 25(3): ??, June 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-6/usenixnews.html#
fish.

Hume:2000:TTA

[Hum00b] Andrew Hume. The times, they are A’changin’. ;login: the USENIX Association newsletter, 25(1):??, February 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-2/times.html.

Hume:2002:LDC

[Hum02] Andrew Hume. Large data/ clusters/resilient computing. In USENIX [USE02e], page

355. ISBN 1-880446-00-

6. LCCN QA76.8.U65 U84

2002. URL http://www.
usenix.org/publications/
library/proceedings/usenix02/
tech/techonefile.html. Un-

published guru session talk, 2002 USENIX Annual Technical Conference, June 10-15, 2002, Monterey Conference Center, Monterey, CA.

Hume:2005:HYD

[Hum05] Andrew Hume. How’s your OS these days? ;login: the USENIX Association newsletter, 30(3): ??, June 2005. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2005-06/pdfs/hume0506.pdf.

Hundt:2000:HCA

[Hun00] Robert Hundt. HP Caliper: An architecture for performance analysis tools. In USENIX [USE00f], page ?? ISBN 1-880446-15-4. LCCN QA76.76.S95 W67 2000.

                                        URL http://www.usenix.
                                        org/publications/library/
                                        proceedings/osdi2000/wiess2000/
                                        hundt.html.

Hand:2007:HVX

[HWF07] Steven Hand, Andrew Warfield, and Keir Fraser. Hardware virtualization with Xen. ;login: the USENIX Association newsletter, 32(1):??, February 2007. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        february-2007-volume-32-
                                        number-1/hardware-virtualization-
                                        xen.

Iams:2001:CAL

[Iam01] Tony Iams. A competitive assessment of Linux in the enterprise. In USENIX [USE01d], page ???? ISBN ???? LCCN ???? URL http:

//www.linuxshowcase.org/

tech.html. Unpublished invited talk, 5th Annual Linux Showcase and Conference, November 5–10, Oakland, CA.

Ioannidis:2001:BSW

[IB01] Sotiris Ioannidis and Steven M.

Bellovin. Building a secure Web browser. In USENIX [USE01f], page ?? ISBN 1-880446-10-3. LCCN QA76.8.U65 U84 2001. URL

http://www.usenix.org/publications/
library/proceedings/usenix01/
freenix01/ioannidis.html.

itojunHagino:2001:CPT

[iiH01] Jun ichiro itojun Hagino. Citrus Project: True multilingual support for BSD operating systems. In USE-NIX [USE01f], page ?? ISBN 1-880446-10-3. LCCN QA76.8.U65 U84 2001. URL

http://www.usenix.org/publications/
library/proceedings/usenix01/
freenix01/hagino.html.

Davis:2001:PTS

[III01] Gervaise (Gerry) Davis III.

Protecting trademarks on software in an open source environment: The Linux mark experience, 2001. URL http:

//www.linuxshowcase.org/

tech.html. Unpublished invited talk, 5th Annual Linux Showcase and Conference, November 5–10, Oakland, CA.

Ingham:2000:GIT

[Ing00] James Ingham. GDBTk: Integrating Tk into a recalcitrant command-line application. In USENIX [USE00e], page ?? ISBN 1-880446-24-3. LCCN ????

                                        URL http://db.usenix.
                                        org/publications/library/
                                        proceedings/tcl2k/ingham.
                                        html.

Ioannidis:2002:HCW

[Ioa02] John Ioannidis. How come we still don’t have IPSec, dammit?, 2002. URL http:/

                                        /www.usenix.org/publications/
                                        library/proceedings/sec02/

tech.html. Unpublished.

Itoi:2000:SCI

[Ito00] Naomaru Itoi. Secure coprocessor integration with Kerberos V5. In USE-NIX [USE00i], page ?? ISBN 1-880446-18-9. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/sec2000/
                                        itoi.html.

Itoi:2001:SCS

[Ito01]  Naomaru Itoi.  SC-CFS:

Smartcard secured cryptographic file system. In USE-NIX [USE01h], page ?? ISBN 1-880446-18-9, 1-880446-07-

3. LCCN QA76.9.A25 U565

2001. URL http://www.
usenix.org/publications/
library/proceedings/sec01/
itoi.html.

Jacobsen:2004:MAI

[Jac04] Hans-Arno Jacobsen, editor. Middleware 2004: ACM/IFIP/USENIX International Middleware Conference, Toronto, Canada, October 18–22, 2004: proceedings, volume 3231 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-23428-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.M54 I33 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.springerlink.
com/openurl.asp?genre=issue&
issn=0302-9743&volume=3231;
 http://www.springerlink.
com/openurl.asp?genre=volume&
id=doi:10.1007/b101561.

Jones:2006:ATP

[JADAD06] Stephen T. Jones, Andrea C. Arpaci-Dusseau, and Remzi H. Arpaci-Dusseau. Antfarm: Tracking processes in a virtual machine environment. In USENIX [USE06], page ?? ISBN 1-931971-44-

7. LCCN ???? URL http:

//www.usenix.org/events/
usenix06/tech/jones.html.

Jaeger:2000:IAN

[JBH^{+}00] R. Jaeger, S. Bhattachargee, J. K. Hollingsworth, R. Duncan, T. Lavian, and F. Travostino. Integrating active networking and commercial-grade routing platforms. In USE-NIX [USE00j], page ?? ISBN ???? LCCN

                                        ????  URL http://db.
                                        usenix.org/publications/
                                        library/proceedings/ine2000/
                                        jaeger.html.

Jordan:2004:EJT

[JCKS04] Mick Jordan, Grzegorz Czajkowski, Kirill Kouklinski, and Glenn Skinner. Extending a J2EE server with dynamic and flexible resource management. In Jacobsen [Jac04], pages 439– ?? CODEN LNCSD9. ISBN 3-540-23428-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.M54 I33 2004. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;
                                        http://www.springerlink.
                                        com/openurl.asp?genre=issue&
                                        issn=0302-9743&volume=3231;
                                         http://www.springerlink.
                                        com/openurl.asp?genre=volume&
                                        id=doi:10.1007/b101561.

Jenkins:2004:OPG

[Jen04] Stephen B. Jenkins. Offline programmatic generation of Web pages. ;login: the USENIX Association newsletter, 29(5):??, October 2004. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2004-10/pdfs/jenkins.pdf.

Ji:2000:AIB

[JFWS00] Minwen Ji, Edward W.

Felten, Randolph Wang, and Jaswinder Pal Singh. Archipelago: An islandbased file system for highly available and scalable Internet services. In USE-NIX [USE00d], page ?? ISBN 1-880446-20-0. LCCN

????  URL http://db.
usenix.org/publications/
library/proceedings/usenix-
win2000/ji.html.

Jones:2008:UNC

[JG08] Michael B. Jones and Dan Geer. 2008 USENIX Nominating Committee report. ;login: the USENIX Association newsletter, 33(1):??, February 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
february-2008-volume-33-
number-1/2008-usenix-nominating-
committee-report.

Jannotti:2000:ORM

[JGJ^{+}00] John Jannotti, David K.

Gifford, Kirk L. Johnson, M. Frans Kaashoek, and James W. O’Toole, Jr. Overcast: Reliable multicasting with an overlay network. In USENIX [USE00h], page ??

ISBN 1-880446-16-2. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/osdi2000/
jannotti.html.

Jelasity:2004:PSS

[JGKvS04] Márk Jelasity, Rachid Guerraoui, Anne-Marie Kermarrec, and Maarten van Steen. The Peer Sampling Service: Experimental evaluation of unstructured Gossip-based implementations. In Jacobsen [Jac04], pages 79–?? CODEN LNCSD9. ISBN 3-540-23428-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.M54 I33 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.springerlink.
com/openurl.asp?genre=issue&
issn=0302-9743&volume=3231;
 http://www.springerlink.
com/openurl.asp?genre=volume&
id=doi:10.1007/b101561.

Jiang:2009:SSP

[JHP^{+}09] Weihang Jiang, Chongfeng Hu, Shankar Pasupathy, Arkady Kanevsky, and Yuanyuan Zhou. Storage system problem troubleshooting and system logs. ;login: the USE-NIX Association newsletter, 34(3):??, June 2009. CO-DEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/june-2009-volume-
34-number-3/storage-system-

problem-troubleshooting-
and-system.

Jiang:2008:DBD

[JHZK08] Weihang Jiang, Chongfeng Hu, Yuanyuan Zhou, and Arkady Kanevsky. Don’t blame disks for every storage subsystem failure. ;login: the USENIX Association newsletter, 33(3):??, June 2008. CODEN LOGNEM. ISSN 1044-6397. URL

https://www.usenix.org/
publications/login/june-
2008-volume-33-number-3/
dont-blame-disks-every-
storage-subsystem-failure.

Jakobsson:2002:MMN

[JJR02] Markus Jakobsson, Ari Juels, and Ronald L. Rivest. Making mix nets robust for electronic voting by randomized partial checking. In USENIX [USE02b], pages 339–353. ISBN 1-931971-00-5. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/sec02/
jakobsson.html.

Joubert:2001:HPM

[JKN^{+}01] Philippe Joubert, Robert King, Richard Neves, Mark Russinovich, and John Tracey. High-performance memorybased Web servers: Kernel and user-space performance. In USENIX [USE01b], page ?? ISBN 1-880446-09-X. LCCN QA76.8.U65 U84 2001. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/usenix01/
                                        joubert.html.

Jakobsson:2003:UEC

[JM03] Markus Jakobsson and Filippo Menczer. Untraceable email cluster bombs. ;login: the USENIX Association newsletter, 28(6):??, December 2003. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2003-12/pdfs/jacobsson.
                                        pdf.

Jim:2002:CSD

[JMG^{+}02] Trevor Jim, Greg Morrisett, Dan Grossman, Michael Hicks, James Cheney, and Yanling Wang. Cyclone: a safe dialect of C. In USENIX [USE02e], page ?? ISBN 1-880446-00-6. LCCN QA76.8.U65 U84 2002. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/usenix02/
                                        jim.html.

Jeon:2000:DBC

[JN00] H. Seok Jeon and Sam H. Noh. Dynamic buffer cache management scheme based on simple and aggressive prefetching. In USENIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/als2000/
                                        jeon.html.

Jin:2004:EPO

[JN04] Jingwen Jin and Klara Nahrstedt. On exploring performance optimizations in Web service composition. In Jacobsen [Jac04], pages 115–?? CODEN LNCSD9. ISBN 3-540-23428-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.M54 I33 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.springerlink.
com/openurl.asp?genre=issue&
issn=0302-9743&volume=3231;
 http://www.springerlink.
com/openurl.asp?genre=volume&
id=doi:10.1007/b101561.

Johnson:2002:BRJ

[Joh02] Steve Johnson. Book review: Jim and Michele Mc-Carty: Software for Your Head: Core Protocols for Creating and Maintaining Shared Vision. ;login: the USE-NIX Association newsletter, 27(3):??, June 2002. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2002-06/openpdfs/
book.pdf.

Johnson:2006:AC

[Joh06] Stephen C. Johnson. Algorithms for the 21st century. ;login: the USENIX Association newsletter, 31(5): ??, October 2006. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        october-2006-volume-31-
                                        number-5/algorithms-21st-
                                        century.

Jones:2001:IR

[Jon01a] Keith J. Jones. Incident response. ;login: the USENIX Association newsletter, 26(7): ??, November 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-11/pdfs/jones1.pdf.

Jones:2001:LKM

[Jon01b]  Keith J. Jones.  Load-

able kernel modules. ;login: the USENIX Association newsletter, 26(7):??, November 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-11/pdfs/jones2.pdf.

Jones:2002:CNI

[Jon02] George M. Jones. The case for network infrastructure security. ;login: the USENIX Association newsletter, 27(6): ??, December 2002. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2002-12/pdfs/jones.pdf.

Josey:2000:AGS

[Jos00a]  Andrew Josey.  Austin

Group status update. ;login: the USENIX Association newsletter, 25(5):??, August 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
standards/43.josey.html.

Josey:2000:PRP

[Jos00b] Andrew Josey. POSIX revision project. ;login: the USE-NIX Association newsletter, 25(1):??, February 2000. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/standards/33.posixrevision.
html.

Josey:2000:SRA

[Jos00c] Andrew Josey. Standards report: Austin Group update. ;login: the USENIX Association newsletter, 25(2): ??, April 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
standards/35.austin.html.

Josey:2000:SRR

[Jos00d] Andrew Josey. Standards report: Recent publications from The Open Group. ;login: the USENIX Association newsletter, 25(2): ??, April 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
standards/36.opengroup.
html.

Josey:2001:SRA

[Jos01a] Andrew Josey. Standard report: Austin Group status update. ;login: the USENIX Association newsletter, 26(5): ??, August 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
standards/51.austin.pdf.

Josey:2001:OGB

[Jos01b] Andrew Josey. The Open Group Base Working Group update. ;login: the USENIX Association newsletter, 26(1): ??, February 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
standards/40.opengrp.html.

Josey:2002:SRR

[Jos02] Andrew Josey. Standard report: Report received August 2002 containing Austin Group (POSIX) status update (May 14, 2002) and Committee Maintenance Procedures for the Approved Standard (POSIX) (revised May 21, 2002). ;login: the USENIX Association newsletter, 27(5):??, October 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
standards/54.standards.
pdf.

Josey:2003:SRU

[Jos03] Andrew Josey. Standard report: Updates. ;login: the USENIX Association newsletter, 28(4):??, August 2003. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2003-08/openpdfs/
standards.pdf.

Josephsen:2006:HVS

[Jos06a] Dave Josephsen. Homeless Vikings: Short-lived BGP session hijacking — a new chapter in the spam wars. ;login: the USENIX Association newsletter, 31(6):??, December 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
december-2006-volume-31-
number-6/homeless-vikings-
short-lived-bgp-session-
hijacki.

Josey:2006:WGI

[Jos06b] Andrew Josey. Why get involved in POSIX? ;login: the USENIX Association newsletter, 31(5):??, October 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
october-2006-volume-31-
number-5/why-get-involved-
posix.

Josephsen:2007:IVS

[Jos07a] David Josephsen. iVoyeur: a view from someplace nearby. ;login: the USENIX Association newsletter, 32(4): ??, August 2007. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        august-2007-volume-32-number-
                                        4/ivoyeur-view-someplace-
                                        nearby.

Josephsen:2007:IMF

[Jos07b] David Josephsen. iVoyeur: Mystical flows. ;login: the USENIX Association newsletter, 32(6):??, December 2007. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        december-2007-volume-32-
                                        number-6/ivoyeur-mystical-
                                        flows.

Josephsen:2007:IOB

[Jos07c] David Josephsen. iVoyeur: Opaque brews. ;login: the USENIX Association newsletter, 32(5):??, October 2007. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        october-2007-volume-32-
                                        number-5/ivoyeur-opaque-
                                        brews.

Josephsen:2008:IAR

[Jos08a] David Josephsen. iVoyeur: Admin, root thyself. ;login: the USENIX Association newsletter, 33(3):77–62, June 2008. CODEN LOGNEM. ISSN 1044-6397. URL

                                        https://www.usenix.org/
                                        publications/login/june-
                                        2008-volume-33-number-3/
                                        ivoyeur-admin-root-thyself.

Josephsen:2008:IC

[Jos08b] David Josephsen. iVoyeur:

Comply. ;login: the USE-NIX Association newsletter, 33(2):??, April 2008. CO-DEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/april-2008-volume-
33-number-2/ivoyeur-comply.

Josephsen:2008:IHP

[Jos08c] David Josephsen. iVoyeur:

Hold the pixels. ;login: the USENIX Association newsletter, 33(4):??, August 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2008-volume-33-number-
4/ivoyeur-hold-pixels.

Josephsen:2008:IPP

[Jos08d] David Josephsen. iVoyeur:

Permission to parse. ;login: the USENIX Association newsletter, 33(1):65– 71, February 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
february-2008-volume-33-
number-1/ivoyeur-permission-
parse.

Josephsen:2008:IYSa

[Jos08e] David Josephsen. iVoyeur:

You should write an NEB module. ;login: the USENIX Association newsletter, 33(5): ??, October 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
october-2008-volume-33-
number-5/ivoyeur-you-should-
write-neb-module.

Josephsen:2008:IYSb

[Jos08f] David Josephsen. iVoyeur: You should write an NEB module, revisited. ;login: the USENIX Association newsletter, 33(6):??, December 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
december-2008-volume-33-
number-6/ivoyeur-you-should-
write-neb-module-revisited.

Josephsen:2009:IHH

[Jos09a] Dave Josephsen. iVoyeur: 7 habits of highly effective Nagios implementations. ;login: the USENIX Association newsletter, 34(6):??, December 2009. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
december-2009-volume-34-
number-6/ivoyeur-7-habits-
highly-effective-nagios.

Josephsen:2009:IMB

[Jos09b] Dave Josephsen. iVoyeur: Message in a bottle — replacing email warnings with SMS. ;login: the USENIX Association newsletter, 34(1): ??, February 2009. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
february-2009-volume-34-
number-1/ivoyeur-message-
bottle%E2%80%94replacing-
email-warnings.

Josephsen:2009:IPL

[Jos09c] Dave Josephsen. iVoyeur:

Packet-level, per-user network access control and monitoring. ;login: the USENIX Association newsletter, 34(5): ??, October 2009. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
october-2009-volume-34-
number-5/ivoyeur-packet-
level-user-network-access-
control.

Josephsen:2009:IDM

[Jos09d] Dave Josephsen. iVoyeur:

The dangers of marketing. ;login: the USENIX Association newsletter, 34(3): ??, June 2009. CODEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/june-2009-volume-
34-number-3/ivoyeur-dangers-
marketing.

Josephsen:2009:IT

[Jos09e] Dave Josephsen. iVoyeur:

Top 5 2008. ;login: the USE-NIX Association newsletter, 34(2):??, April 2009. CO-DEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/

                                        login/april-2009-volume-
                                        34-number-2/ivoyeur-top-
                                        5-2008.

Josephsen:2009:IWI

[Jos09f] Dave Josephsen. iVoyeur: Who invited the salesmen? ;login: the USENIX Association newsletter, 34(4): ??, August 2009. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        august-2009-volume-34-number-
                                        4/ivoyeur-who-invited-salesmen.

Jacob:2002:CAP

[JR02] Matthias Jacob and Keith Randall. Cross-architectural performance portability of a Java Virtual Machine implementation. In USE-NIX [USE02f], page ?? ISBN 1-931971-01-3. LCCN QA76.73 .J38 J42 2002.

                                        URL http://www.usenix.
                                        org/publications/library/
                                        proceedings/javavm02/jacob.
                                        html.

Johnson:2000:C

[JW00a] Steve Johnson and Dusty White. Chunks. ;login: the USENIX Association newsletter, 25(1):??, February 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-2/features/chunks.
                                        html.

Johnson:2000:FMM

[JW00b] Steve Johnson and Dusty White. The first miracle of management. ;login: the USENIX Association newsletter, 25(8):??, December 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-12/pdfs/johnson.pdf.

Johnson:2000:DR

[JW00c] Steve Johnson and Dusty White. I don’t have the resources. ;login: the USE-NIX Association newsletter, 25(4):??, July 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-7/features/resources.
html.

Johnson:2000:JP

[JW00d] Steve Johnson and Dusty White. Just presuppose. . .. ;login: the USENIX Association newsletter, 25(6): ??, October 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-10/pdfs/justpresuppose.
pdf.

Johnson:2000:CD

[JW00e] Steve Johnson and Dusty White. “this company doesn’t… ”. ;login: the USE-NIX Association newsletter, 25(2):??, April 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-4/features/company.
                                        html.

Johnson:2000:L

[JW00f] Steve Johnson and Dusty White. To the limit. ;login: the USENIX Association newsletter, 25(5):??, August 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-8/features/tothelimit.
                                        html.

Johnson:2000:Y

[JW00g] Steve Johnson and Dusty White. Yep! ;login: the USE-NIX Association newsletter, 25(3):??, June 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-6/features/yep.html.

Johnson:2001:DYB

[JW01a] Steve Johnson and Dusty White. Do you have both ORs in the water? ;login: the USENIX Association newsletter, 26(2): ??, April 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-04/pdfs/johnson.pdf.

Johnson:2001:GMG

[JW01b] Steve Johnson and Dusty White. Get my goat. ;login: the USENIX Association newsletter, 26(4):

??, July 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-07/pdfs/johnson.pdf.

Johnson:2001:HCG

[JW01c] Steve Johnson and Dusty White. Here comes the grooming. ;login: the USE-NIX Association newsletter, 26(8):??, December 2001. CODEN LOGNEM. ISSN 1044-6397. URL http://

www.usenix.org/publications/
login/2001-12/pdfs/johnson.
pdf.

Johnson:2001:JVP

[JW01d] Steve Johnson and Dusty White. Jobs vs. people. ;login: the USENIX Association newsletter, 26(1):??, February 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-02/pdfs/johnson.pdf.

Johnson:2001:KB

[JW01e] Steve Johnson and Dusty White. Kick those BUTS. ;login: the USENIX Association newsletter, 26(3): ??, June 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-06/pdfs/johnson.pdf.

Johnson:2001:T

[JW01f] Steve Johnson and Dusty

White.  Turf.  ;login:

the USENIX Association newsletter, 26(6):??, October 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-07/pdfs/johnson.pdf.

Johnson:2002:J

[JW02] Steve Johnson and Dusty White. Judgment. ;login: the USENIX Association newsletter, 27(1):??, February 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-02/pdfs/johnson.pdf.

Jiang:2001:APR

[JZ01] Song Jiang and Xiaodong Zhang. Adaptive page replacement to protect thrashing in Linux. In USENIX [USE01d], page ?? ISBN ???? LCCN ???? URL http:

//www.linuxshowcase.org/
jiang.html.

Krishnamurthy:2001:PCP

[KA01] Balachander Krishnamurthy and Martin Arlitt. PRO-COW: Protocol compliance on the Web — a longitudinal study. In USENIX [USE01c], page ?? ISBN 1-880446-14-6. LCCN TK5105.875.I57

U96 2001. URL http://www.
usenix.org/publications/
library/proceedings/usits01/
krishnamurthy.html.

Kaldewey:2009:PVC

[Kal09] Tim Kaldewey. Programming video cards for database applications. ;login: the USE-NIX Association newsletter, 34(4):??, August 2009. CO-DEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/august-2009-volume-
34-number-4/programming-
video-cards-database-applications.

Kamp:2002:RDD

[Kam02] Poul-Henning Kamp. Rethinking /dev and devices in the UNIX kernel. In USE-NIX [USE02a], pages 77–

88. ISBN 1-880446-02-2. LCCN QA76.76.O63 B736

2002. URL http://www.
usenix.org/publications/
library/proceedings/bsdcon02/
kamp.html.

Kamath:2004:CL

[Kam04] Sean Kamath. Commentary:

Lisa ’03. ;login: the USE-NIX Association newsletter, 29(1):??, February 2004. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2004-02/openpdfs/
opinion.pdf.

Kandula:2005:SDA

[Kan05a] Srikanth Kandula. Surviving DDoS attacks. ;login: the USENIX Association newsletter, 30(5):??, October 2005. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2005-10/pdfs/kandula.pdf.

Kanies:2005:UVC

[Kan05b] Luke Kanies. Using version control in system administration. ;login: the USENIX Association newsletter, 30(6): ??, December 2005. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2005-12/pdfs/kanies.pdf.

Kanies:2006:PNG

[Kan06a]  Luke Kanies.  Puppet:

Next-generation configuration management. ;login: the USENIX Association newsletter, 31(1):??, February 2006. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2006-02/pdfs/kanies.pdf.

Kanies:2006:WYS

[Kan06b]  Luke Kanies.  Why you

should use Ruby. ;login: the USENIX Association newsletter, 31(2):42– 48, April 2006. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2006-04/pdfs/kanies.pdf.

Karhuluoma:2002:BRJ

[Kar02] Nina Karhuluoma. Book review: John Viega and Gary McGraw: Building Secure Software. ;login: the USE-NIX Association newsletter, 27(4):??, August 2002. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2002-08/openpdfs/
books.pdf.

Kiriansky:2002:SEP

[KBA02] Vladimir Kiriansky, Derek Bruening, and Saman Amarasinghe. Secure execution via program shepherding. In USENIX [USE02b], pages 191–206. ISBN 1-931971-00-5. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/sec02/
kiriansky.html.

Kent:2000:NIM

[KC00] Alexander Kent and James Clifford. Network information management and distribution in a heterogeneous and decentralized enterprise environment. In USENIX [USE00g], page ?? ISBN 1-880446-13-8. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/lisa2000/
kent.html.

claffy:2001:IMM

[kc01] kc claffy. Internet measurement: Myths about Internet data, 2001. URL http://db.

usenix.org/publications/
library/proceedings/lisa2001/

tech/. Unpublished invited talk, LISA 2001: 15th Systems Administration Conference, December 2–7, 2001, Town and Country Resort Hotel, San Diego, CA.

Kim:2000:LOH

[KCK^{+}00] Jong Min Kim, Jongmoo Choi, Jesung Kim, Sam H. Noh, Sang Lyul Min, Yookun Cho, and Chong Sang Kim. A low-overhead, highperformance unified buffer management scheme that exploits sequential and looping references. In USE-NIX [USE00h], page ?? ISBN 1-880446-16-2. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/osdi2000/
                                        kim.html.

Kim:2002:SVP

[KCN02] Minkyong Kim, Landon Cox, and Brian Noble. Safety, visibility, and performance in a wide-area file system. In USENIX [USE02c], page ?? ISBN 1-880446-03-

0. LCCN QA76.9.F5 C65

                                        2002. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/fast02/
                                        kim.html.

Kersting:2002:IGA

[KD02] Oliver Kersting and Jürgen Döllner. Interactive 3D graphics applications for Tcl. In USENIX [USE02e], page ?? ISBN 1-880446-00-6. LCCN QA76.8.U65

U84 2002. URL http://www.
usenix.org/publications/
library/proceedings/usenix02/
tech/freenix/kersting.html.

Kenneally:2001:SDS

[Ken01] Erin Kenneally. Stepping on the digital scale. ;login: the USENIX Association newsletter, 26(8):??, December 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-12/pdfs/kenneally.
pdf.

Kenneally:2002:CF

[Ken02a] Erin Kenneally. Computer forensics. ;login: the USENIX Association newsletter, 27(4): ??, August 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-08/pdfs/kenneally.
pdf.

Kenneally:2002:EBB

[Ken02b] Erin Kenneally. Evolving behavioral boundaries in cyberspace. ;login: the USE-NIX Association newsletter, 27(3):??, June 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-06/pdfs/kenneally.
pdf.

Kenneally:2002:IDD

[Ken02c]  Erin Kenneally.  “it de-

pends”: Defining legal values for network behavior. ;login: the USENIX Association newsletter, 27(6):??, December 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-12/pdfs/kenneally.
pdf.

Kenneally:2003:EET

[Ken03a] Erin Kenneally. Evidence enhancing technology. ;login: the USENIX Association newsletter, 28(6):??, December 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-12/pdfs/kenneally.
pdf.

Kenneally:2003:PYD

[Ken03b] Erin Kenneally. Put your data where your mouth is. ;login: the USENIX Association newsletter, 28(3): ??, June 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-06/pdfs/kenneally.
pdf.

Kereliuk:2003:IVI

[Ker03] Ryan Kereliuk. Introducing Voice Over Internet Protocol technology. ;login: the USENIX Association newsletter, 28(5):??, October 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.

org/publications/login/
2003-10/pdfs/kereliuk.pdf.

Kernighan:2006:CTR

[Ker06] Brian Kernighan. Code testing and its role in teaching. ;login: the USENIX Association newsletter, 31(2): 9–18, April 2006. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2006-04/pdfs/kernighan.
pdf.

Ko:2000:DCS

[KFBK00] Calvin Ko, Timothy Fraser, Lee Badger, and Douglas Kilpatrick. Detecting and countering system intrusions using software wrappers. In USENIX [USE00i], page ?? ISBN 1-880446-18-9. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/sec2000/
ko.html.

Kornievskaia:2001:KCT

[KHDC01] Olga Kornievskaia, Peter Honeyman, Bill Doster, and Kevin Coffman. Kerberized credential translation: a solution to Web access control. In USENIX [USE01h], page ?? ISBN 1-880446-18-9, 1-880446-07-3. LCCN QA76.9.A25 U565

2001. URL http://www.
usenix.org/publications/
library/proceedings/sec01/
kornievskaia.html.

Konerding:2001:CAD

[KHF01] David Konerding, Conrad Huang, and Thomas Ferrin. Chimera: Affordable desktop molecular modeling on Linux workstations. In USENIX [USE01d], page ?? ISBN ???? LCCN ???? URL http:

                                        //www.linuxshowcase.org/
                                        konerding.html.

Kim:2001:OSI

[Kim01] Gene Kim. Open source issues and opportunities for security tool vendors. In USENIX [USE01d], page ???? ISBN ???? LCCN

                                        ????  URL http://www.
                                        linuxshowcase.org/tech.

html. Unpublished invited talk, 5th Annual Linux Showcase and Conference, November 5–10, Oakland, CA.

Kiselyov:2000:SH

[Kis00]  Oleg Kiselyov.  Speaking

HTTP. ;login: the USE-NIX Association newsletter, 25(2):??, April 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-4/features/speaking.
                                        html.

Kosters:2008:ITM

[KK08] Mark Kosters and Megan Kruse. IPv6: It’s time to make the move. ;login: the USENIX Association newsletter, 33(2): ??, April 2008. CODEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/april-2008-volume-
33-number-2/ipv6-its-time-
make-move.

Kroeger:2001:DIP

[KL01] Thomas M. Kroeger and Darrell D. E. Long. Design and implementation of a predictive file prefetching algorithm. In USENIX [USE01b], page ?? ISBN 1-880446-09-X. LCCN QA76.8.U65

U84 2001. URL http://www.
usenix.org/publications/
library/proceedings/usenix01/
kroeger.html.

Kleiman:2000:STC

[Kle00]  Steve Kleiman.  Surfing

technology curves, 2000.

URL http://www.usenix.
org/publications/library/
proceedings/osdi2000/wiess2000/
invitedtalks/kleiman_html.

Unpublished invited talk at First Workshop on Industrial Experiences with Systems Software (WIESS 2000), October 22, 2000, Paradise Point Resort, San Diego, California, USA.

Klein:2009:CKP

[Kle09] Gerwin Klein. Correct OS kernel? Proof? Done! ;login: the USENIX Association newsletter, 34(6):??, December 2009. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.

                                        org/publications/login/
                                        december-2009-volume-34-
                                        number-6/correct-os-kernel-
                                        proof-done.

Kingpin:2001:SAP

[KM01] Kingpin and Mudge. Security analysis of the Palm operating system and its weaknesses against malicious code threats. In USENIX [USE01h], page ?? ISBN 1-880446-18-9, 1-880446-07-3. LCCN QA76.9.A25 U565

                                        2001. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/sec01/
                                        kingpin.html.

Knell:2001:LLD

[Kne01] Mike Knell. The loneliness of the long-distance SysAdmin, or where geeks gather. ;login: the USENIX Association newsletter, 26(3): ??, June 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-06/pdfs/knell.pdf.

Knopper:2000:BSC

[Kno00] Klaus Knopper. Building a self-contained autoconfiguring Linux system on an ISO9660 file system. In USENIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/als2000/
                                        knopper.html.

Knowles:2006:AT

[Kno06] Brad Knowles. It’s about time . . .. ;login: the USENIX Association newsletter, 31(5): ??, October 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
october-2006-volume-31-
number-5/its-about-time.

Knowles:2008:BSN

[Kno08] Brad Knowles. Building scalable NTP server infrastructures. ;login: the USENIX Association newsletter, 33(5): ??, October 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
october-2008-volume-33-
number-5/building-scalable-
ntp-server-infrastructures.

Krishnan:2000:SCO

[KNRS00] P. Krishnan, Tejas Naik, Ganesan Ramu, and Roshan Sequeira. Sequencing of configuration operations for IP networks. In USE-NIX [USE00g], page ?? ISBN 1-880446-13-8. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/lisa2000/
krishnan.html.

Kocher:2002:IS

[Koc02] Paul Kocher. Illusions of security, 2002. URL http://

www.usenix.org/publications/
library/proceedings/sec02/

tech.html. Unpublished.

Kolstad:2000:UR

[Kol00a] Rob Kolstad. The 1999– 2000 USACO report. ;login: the USENIX Association newsletter, 25(5):??, August 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-8/usenixnews.html#
                                        usaco.

Kolstad:2000:IGH

[Kol00b] Rob Kolstad. An interview with Geoff Halprin. ;login: the USENIX Association newsletter, 25(6):??, October 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-10/pdfs/interviewwithhalprin.
                                        pdf.

Kolstad:2000:IM

[Kol00c] Rob Kolstad. An interview with Mudge. ;login: the USE-NIX Association newsletter, 25(2):??, April 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-4/features/mudge.html.

Kolstad:2000:Ma

[Kol00d] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 25(1):??, February 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-2/motd.html.

Kolstad:2000:Mb

[Kol00e] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 25(2): ??, April 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-4/motd.html.

Kolstad:2000:Mc

[Kol00f] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 25(3): ??, June 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-6/motd.html.

Kolstad:2000:Md

[Kol00g] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 25(4): ??, July 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-7/motd.html.

Kolstad:2000:Me

[Kol00h] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 25(5):??, August 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-8/motd.html.

Kolstad:2000:Mf

[Kol00i] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 25(6):??, October 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-10/MOTD.pdf.

Kolstad:2000:Mg

[Kol00j] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 25(8):??, December 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-12/motd.pdf.

Kolstad:2001:Ma

[Kol01a] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 26(1):??, February 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-02/motd.pdf.

Kolstad:2001:Mb

[Kol01b] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 26(2): ??, April 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-04/motd.pdf.

Kolstad:2001:Mc

[Kol01c] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 26(3): ??, June 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-06/motd.pdf.

Kolstad:2001:Md

[Kol01d] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 26(4): ??, July 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-07/motd.pdf.

Kolstad:2001:Me

[Kol01e] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 26(5):??, August 2001. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2001-08/openpdfs/
motd.pdf.

Kolstad:2001:Mf

[Kol01f] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 26(6):??, October 2001. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2001-10/openpdfs/
motd.pdf.

Kolstad:2001:Mg

[Kol01g] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 26(8):??, December 2001. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.

                                        usenix.org/publications/
                                        login/2001-12/openpdfs/
                                        motd.pdf.

Kolstad:2002:Ma

[Kol02a] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 27(1):??, February 2002. CODEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2002-02/openpdfs/
                                        motd.pdf.

Kolstad:2002:Mb

[Kol02b] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 27(2): ??, April 2002. CODEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2002-04/openpdfs/
                                        motd.pdf.

Kolstad:2002:Mc

[Kol02c] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 27(3): ??, June 2002. CODEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2002-06/openpdfs/
                                        motd.pdf.

Kolstad:2002:Md

[Kol02d] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 27(4):??, August 2002. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2002-08/openpdfs/
motd.pdf.

Kolstad:2002:Me

[Kol02e] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 27(5):??, October 2002. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2002-10/openpdfs/
motd.pdf.

Kolstad:2002:Mf

[Kol02f] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 27(6):??, December 2002. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2002-12/openpdfs/
motd.pdf.

Kolstad:2003:Ma

[Kol03a] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 28(1):??, February 2003. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2003-02/openpdfs/
motd.pdf.

Kolstad:2003:Mb

[Kol03b] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 28(2):

??, April 2003. CODEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2003-04/openpdfs/
                                        motd.pdf.

Kolstad:2003:Mc

[Kol03c] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 28(3): ??, June 2003. CODEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2003-06/openpdfs/
                                        motd.pdf.

Kolstad:2003:Md

[Kol03d] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 28(4):??, August 2003. CODEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2003-08/openpdfs/
                                        motd.pdf.

Kolstad:2003:Me

[Kol03e] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 28(5):??, October 2003. CODEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2003-10/openpdfs/
                                        motd.pdf.

Kolstad:2004:AHR

[Kol04a] Rob Kolstad. Alain Hénon retires. ;login: the USENIX Association newsletter, 29(4): ??, August 2004. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2004-08/openpdfs/unotes.
pdf.

Kolstad:2004:Ma

[Kol04b] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 29(2): ??, April 2004. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2004-04/openpdfs/
MOTD.pdf.

Kolstad:2004:Mb

[Kol04c] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 29(3): ??, June 2004. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2004-06/openpdfs/
motd.pdf.

Kolstad:2004:Mc

[Kol04d] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 29(4):??, August 2004. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2004-08/openpdfs/
motd.pdf.

Kolstad:2004:Md

[Kol04e] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 29(5):??, October 2004. CODEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2004-10/openpdfs/
                                        motd.pdf.

Kolstad:2004:UIF

[Kol04f] Rob Kolstad. USACO: The international finals. ;login: the USENIX Association newsletter, 29(6):??, December 2004. CODEN LOGNEM. ISSN 1044-6397.

Kolstad:2004:UUC

[Kol04g] Rob Kolstad. USACO: The USA Computing Olympiad. ;login: the USENIX Association newsletter, 29(4): ??, August 2004. CODEN LOGNEM. ISSN 1044-6397.

Kolstad:2004:UHC

[Kol04h] Rob Kolstad. USENIX haiku contest. ;login: the USE-NIX Association newsletter, 29(5):??, October 2004. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2004-10/openpdfs/
                                        haiku.pdf.

Kolstad:2005:Ma

[Kol05a] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 30(1):??, February 2005. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2005-02/openpdfs/
motd.pdf.

Kolstad:2005:Mb

[Kol05b] Rob Kolstad. motd. ;login: the USENIX Association newsletter, 30(3): ??, June 2005. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2005-06/openpdfs/
motd0506.pdf.

Kolstad:2005:ML

[Kol05c] Rob Kolstad. motd: Leverage. ;login: the USENIX Association newsletter, 30 (2):??, April 2005. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2005-04/openpdfs/
motd0504.pdf.

Kolstad:2005:SNA

[Kol05d] Rob Kolstad. SAGE-news article on SCO defacement. ;login: the USENIX Association newsletter, 30(1): ??, February 2005. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2005-02/openpdfs/
sco.pdf.

Kolstad:2005:UNUb

[Kol05e] Rob Kolstad. USENIX notes:

USACO team brings home the gold. ;login: the USENIX Association newsletter, 30(6): ??, December 2005. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2005-12/openpdfs/
usaco.pdf.

Kolstad:2005:UNUa

[Kol05f] Rob Kolstad. USENIX notes: USACO, the USA Computing Olympiad. ;login: the USE-NIX Association newsletter, 30(5):??, October 2005. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2005-10/openpdfs/
usaco05.pdf.

Kolstad:2006:UN

[Kol06] Rob Kolstad. USACO news. ;login: the USENIX Association newsletter, 31(5): ??, October 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
october-2006-volume-31-
number-5/usaco-news.

Kolstad:2007:UUa

[Kol07a] Rob Kolstad. USACO update. ;login: the USENIX Association newsletter, 32(4): ??, August 2007. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2007-volume-32-number-
4/usaco-update.

Kolstad:2007:UUb

[Kol07b] Rob Kolstad. USACO update. ;login: the USENIX Association newsletter, 32(6): ??, December 2007. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
december-2007-volume-32-
number-6/usaco-update.

Kolstad:2008:RU

[Kol08] Rob Kolstad. Report on US-ACO. ;login: the USENIX Association newsletter, 33(5): ??, October 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
october-2008-volume-33-
number-5/report-usaco.

Kolstad:2009:IOI

[Kol09a] Rob Kolstad. The International Olympiad on Informatics. ;login: the USENIX Association newsletter, 34(1): ??, February 2009. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
february-2009-volume-34-
number-1/international-
olympiad-informatics.

Kolstad:2009:UTS

[Kol09b] Rob Kolstad. USACO teams shine. ;login: the USENIX Association newsletter, 34(6): ??, December 2009. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.

org/publications/login/
december-2009-volume-34-
number-6/usaco-teams-shine.

Kaufman:2001:PNS

[KP01] Charlie Kaufman and Radia Perlman. PDM: a new strong password-based protocol. In USENIX [USE01h], page ?? ISBN 1-880446-18-9, 1-880446-07-3. LCCN QA76.9.A25 U565

2001. URL http://www.
usenix.org/publications/
library/proceedings/sec01/
kaufman.html.

Keskin:2000:TTS

[KPv00] Ahmet C. Keskin, Till I. Patzchke, and Ernst von-Voight. Tcl/Tk: a strong basis for complex load testing systems. In USE-NIX [USE00e], page ?? ISBN 1-880446-24-3. LCCN

????  URL http://db.
usenix.org/publications/
library/proceedings/tcl2k/
keskin.html.

Kriftcher:2000:TSU

[Kri00a] Noel N. Kriftcher. Tomorrow’s scientist and urban planners create the city of the future. ;login: the USE-NIX Association newsletter, 25(2):??, April 2000. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2000-4/scientists.
html.

Krishnaswamy:2000:APH

[Kri00b] Tara Krishnaswamy. Automatic precompiled headers: Speeding up C++ application build times. In USENIX [USE00f], page ?? ISBN 1-880446-15-4. LCCN QA76.76.S95 W67 2000.

URL http://www.usenix.
org/publications/library/
proceedings/osdi2000/wiess2000/
krishnaswamy.html.

Kryder:2002:FMR

[Kry02] Mark H. Kryder. Future magnetic recording technologies. In USENIX [USE02c], page ?? ISBN 1-880446-03-

0. LCCN QA76.9.F5 C65

2002. URL http://www.
usenix.org/publications/
library/proceedings/fast02/
kryder.pdf.

Krzywinski:2001:PCP

[Krz01] Martin Krzywinski. Picking cluster parts: Cluster construction at the Genome Sequence Centre. ;login: the USENIX Association newsletter, 26(5):??, August 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-08/pdfs/krzy.pdf.

Katz:2000:CCA

[KS00] Jonathan Katz and Bruce Schneier. A chosen ciphertext attack against several Email encryption protocols. In USENIX [USE00i], page ?? ISBN 1-880446-18-9. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/sec2000/
katz.html.

Khan:2001:PPF

[KT01] Javed I. Khan and Qingping Tao. Partial prefetch for faster surfing in composite hypermedia. In USE-NIX [USE01c], page ?? ISBN 1-880446-14-6. LCCN TK5105.875.I57 U96 2001.

URL http://www.usenix.
org/publications/library/
proceedings/usits01/khan.
html.

Kuhn:2000:PCL

[Kuh00] Markus G. Kuhn. Probabilistic counting of large digital signature collections. In USENIX [USE00i], page ?? ISBN 1-880446-18-9. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/sec2000/
kuhn.html.

Koster:2002:AGW

[KV02] Cornelis H. A. Koster and Erik Verbruggen. The AGFL grammar work lab. In USENIX [USE02d], page ?? ISBN 1-880446-01-4. LCCN QA76.8.U65 P765 2002.

URL http://www.usenix.
org/publications/library/
proceedings/usenix02/tech/
freenix/koster.html.

Kumar:2000:HSC

[KW00] Arun Kumar and Randy Wright. HP Scalable Computing Architecture. In USENIX [USE00f], page ?? ISBN 1-880446-15-4. LCCN QA76.76.S95 W67 2000.

URL http://www.usenix.
org/publications/library/
proceedings/osdi2000/wiess2000/
wright.html.

Kon:2001:DRM

[KYH^{+}01] Fabio Kon, Tomonori Yamane, Christopher K. Hess, Roy H. Campbell, and M. Dennis Mickunas. Dynamic resource management and automatic configuration of distributed component systems. In USE-NIX [USE01e], page ?? ISBN 1-880446-12-X. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/coots01/
kon.html.

Laffitte:2004:DIP

[LA04] Hernan Laffitte and Eric Anderson. Document it with a picture. ;login: the USE-NIX Association newsletter, 29(2):??, April 2004. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2004-04/pdfs/Laffitte.pdf.

LaBonte:2000:FII

[LaB00] Alain LaBonté. FAQ: ISO/ IEC 106461 versus Unicode.

;login: the USENIX Association newsletter, 25(3): ??, June 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
standards/38.labonte.html.

Lambright:2000:EMR

[Lam00] Dan Lambright. Experiences in measuring the reliability of a cache-based storage system. In USENIX [USE00f], page ?? ISBN 1-880446-15-

4. LCCN QA76.76.S95 W67

2000. URL http://www.
usenix.org/publications/
library/proceedings/osdi2000/
wiess2000/lambright.html.

Langley:2006:ASU

[Lan06] Randolph Langley. Auditing superuser usage. ;login: the USENIX Association newsletter, 31(2):56– 59, April 2006. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2006-04/pdfs/langley.pdf.

Langheinrich:2008:HBT

[Lan08] Eric Langheinrich. http:BL: Taking DNSBL beyond SMTP. ;login: the USENIX Association newsletter, 33(1):??, February 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
february-2008-volume-33-
number-1/httpbl-taking-
dnsbl-beyond-smtp.

Lasser:2001:ISP

[Las01] Jon Lasser. Implementing Snort in a production environment. ;login: the USENIX Association newsletter, 26(7): ??, November 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-11/pdfs/lasser.pdf.

Lawton:2000:PVM

[Law00] Kevin P. Lawton. Plex86: An 180x86 virtual machine. In USENIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/als2000/
lawton.html.

Lu:2002:AOD

[LAW02] Chenyang Lu, Guillermo A.

Alvarez, and John Wilkes. Aqueduct: Online data migration with performance guarantees. In USENIX [USE02c], page ?? ISBN 1-880446-03-0. LCCN QA76.9.F5

C65 2002. URL http://www.
usenix.org/publications/
library/proceedings/fast02/
lu.html.

Lavana:2000:CTR

[LB00] Hemang Lavana and Franc Brglez. CollabWiseTk: a toolkit for rendering standalone applications collaborative. In USENIX [USE00e], page ?? ISBN 1-880446-24-3. LCCN ????

URL http://db.usenix.
org/publications/library/
proceedings/tcl2k/lavana.
html.

Lai:2001:NTM

[LB01] Kevin Lai and Mary Baker. Nettimer: a tool for measuring bottleneck link bandwidth. In USENIX [USE01c], page ?? ISBN 1-880446-14-6. LCCN TK5105.875.I57

U96 2001. URL http://www.
usenix.org/publications/
library/proceedings/usits01/
lai.html.

Lampoudi:2002:SSE

[LB02] Sotiria Lampoudi and David M. Beazley. SWILL: a simple embedded Web server library. In USENIX [USE02d], page ?? ISBN 1-880446-01-

4. LCCN QA76.8.U65 P765

2002. URL http://www.
usenix.org/publications/
library/proceedings/usenix02/
tech/freenix/lampoudi.html.

Levy:2001:LMD

[LBN01] Lior Levy, Liad Blumrosen, and Noam Nisan. On-line markets for distributed object services: The MAJIC system. In USENIX [USE01c], page ?? ISBN 1-880446-14-6. LCCN TK5105.875.I57

U96 2001. URL http://www.
usenix.org/publications/
library/proceedings/usits01/
levy.html.

Long:2007:SUb

[LC07] Jane-Ellen Long and Alva Couch. SAGE update. ;login: the USENIX Association newsletter, 32(2): ??, April 2007. CODEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/april-2007-volume-
32-number-2/sage-update.

Lowell:2000:EFT

[LCC00] David E. Lowell, Subhachandra Chandra, and Peter Chen. Exploring failure transparency and the limits of generic recovery. In USENIX [USE00h], page ?? ISBN 1-880446-16-2. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/osdi2000/
lowell.html.

Larochelle:2001:SDL

[LE01] David Larochelle and David Evans. Statically detecting likely buffer overflow vulnerabilities. In USENIX [USE01h], page ?? ISBN 1-880446-18-9, 1-880446-07-

3. LCCN QA76.9.A25 U565

2001. URL http://www.
usenix.org/publications/
library/proceedings/sec01/
larochelle.html.

Lear:2004:BAK

[Lea04] Eliot Lear. Being awash in keys. ;login: the USENIX Association newsletter, 29(6):

??, December 2004. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2004-12/pdfs/keys.
pdf.

LeFebvre:2002:CCF

[LeF02] William LeFebvre. CNN.com: Facing a world crisis. In USENIX [USE02e], page

355. ISBN 1-880446-00-

6. LCCN QA76.8.U65 U84

2002. URL http://www.
usenix.org/publications/
library/proceedings/usenix02/
tech/techonefile.html. Un-

published invited talk, 2002 USENIX Annual Technical Conference, June 10-15, 2002, Monterey Conference Center, Monterey, CA.

Lehey:2001:IFS

[Leh01] Greg Lehey. Improving the FreeBSD SMP implementation. In USENIX [USE01f], page ?? ISBN 1-880446-10-3. LCCN QA76.8.U65

U84 2001. URL http://www.
usenix.org/publications/
library/proceedings/usenix01/
freenix01/lehey.html.

Leighton:2000:NSI

[Lei00] Luke K. C. Leighton. NT5 and Samba integration issues. In USENIX [USE00b], page ?? ISBN 1-880446-19-7. LCCN ????

URL http://db.usenix.
org/publications/library/

proceedings/lisa-nt2000/
leighton.html.

Lemberg:2000:TTC

[Lem00] Paul Lemberg. Ten things to consider. ;login: the USE-NIX Association newsletter, 25(3):??, June 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-6/features/ten.html.

Lemon:2001:KGS

[Lem01] Jonathan Lemon. Kqueue — a generic and scalable event notification facility. In USENIX [USE01f], page ?? ISBN 1-880446-10-3. LCCN QA76.8.U65

U84 2001. URL http://www.
usenix.org/publications/
library/proceedings/usenix01/
freenix01/lemon.html.

Lemon:2002:RSF

[Lem02] Jonathan Lemon. Resisting SYN flood DoS attacks with a SYN cache. In USE-NIX [USE02a], pages 89–

97. ISBN 1-880446-02-2. LCCN QA76.76.O63 B736

2002. URL http://www.
usenix.org/publications/
library/proceedings/bsdcon02/
lemon.html.

Lever:2000:LKH

[Lev00] Chuck Lever. Linux kernel hash table behavior: Analysis and improvements. In USENIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/als2000/
lever.html.

Levin:2005:SVN

[Lev05] Adam Levin. SAN vs. NAS for Oracle: a tale of two protocols. ;login: the USENIX Association newsletter, 30(4): ??, August 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-08/pdfs/levin.pdf.

Leyton:2001:QID

[Ley01] Richard Leyton. A quick introduction to database systems. ;login: the USENIX Association newsletter, 26(8): ??, December 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-12/pdfs/leyton.pdf.

LaMeyer:2000:DDI

[LGJ00] Amy LaMeyer, Shankaranarayanan Ganesan, and Jesper M. Johansson. On designing a database for integrated user management: Pitfalls and possibilities. In USENIX [USE00b], page ?? ISBN 1-880446-19-7. LCCN

????  URL http://db.
usenix.org/publications/
library/proceedings/lisa-
nt2000/lameyer.html.

Li:2004:ILM

[LGW04] Baochun Li, Jiang Guo, and Mea Wang. iOverlay: a lightweight middleware infrastructure for overlay application implementations. In Jacobsen [Jac04], pages 135–?? CODEN LNCSD9. ISBN 3-540-23428-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.M54 I33 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.springerlink.
com/openurl.asp?genre=issue&
issn=0302-9743&volume=3231;
http://www.springerlink.
com/openurl.asp?genre=volume&
id=doi:10.1007/b101561.

Lever:2002:LNC

[LH02] Chuck Lever and Peter Honeyman. Linux NFS client write performance. In USE-NIX [USE02d], page ?? ISBN 1-880446-01-4. LCCN QA76.8.U65 P765 2002.

URL http://www.usenix.
org/publications/library/
proceedings/usenix02/tech/
freenix/lever.html.

Liu:2006:HPV

[LHAP06] Jiuxing Liu, Wei Huang, Bulent Abali, and Dhabaleswar K. Panda. High performance VMM-bypass I/O in virtual machines. In USE-NIX [USE06], page ?? ISBN 1-931971-44-7. LCCN ????

Liebmann:2000:DCN

[Lie00] Lenny Liebmann. Dot-coms’ newest secret weapon doesn’t have a name. ;login: the USE-NIX Association newsletter, 25(2):??, April 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-4/features/weapon.
html.

Lifka:2001:SPC

[Lif01] Dave Lifka. Special purpose clusters at the Cornell Theory Center. ;login: the USENIX Association newsletter, 26(5): ??, August 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-08/pdfs/lifka.pdf.

Limoncelli:2001:DYK

[Lim01] Tom Limoncelli. Do you know what’s in your firewall? ;login: the USENIX Association newsletter, 26(8): ??, December 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-12/pdfs/limoncelli.
pdf.

Limoncelli:2006:DW

[Lim06]  Tom Limoncelli.  Dele-

gating to the Web. ;login: the USENIX Association newsletter, 31(1):??, February 2006. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2006-02/pdfs/limoncelli.
pdf.

Limoncelli:2009:HIM

[Lim09] Thomas A. Limoncelli. Hey!

I have to install and maintain this crap too, ya know! ;login: the USENIX Association newsletter, 34(5): ??, October 2009. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
october-2009-volume-34-
number-5/hey-i-have-install-
and-maintain-crap-too-ya-
know.

Lidl:2002:FPF

[LLB02] Kurt J. Lidl, Deborah G.

Lidl, and Paul R. Borman. Flexible packet filtering: Providing a rich toolbox. In USENIX [USE02a], pages 99–

110. ISBN 1-880446-02-2. LCCN QA76.76.O63 B736

2002. URL http://www.
usenix.org/publications/
library/proceedings/bsdcon02/
lidl.html.

Lloyd:2007:ISE

[Llo07] John Lloyd. Introducing system engineering to the system Admin. ;login: the USE-NIX Association newsletter, 32(2):??, April 2007. CO-DEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/april-2007-volume-
32-number-2/introducing-
system-engineering-system-
admin.

Lawrenz:2000:UTB

[LM00] Carsten H. Lawrenz and Rajkumar Madhuram. Using Tcl to build a buzzwordcompliant environment that glues together legacy analysis programs. In USE-NIX [USE00e], page ?? ISBN 1-880446-24-3. LCCN

                                        ????  URL http://db.
                                        usenix.org/publications/
                                        library/proceedings/tcl2k/
                                        lawrenz.html.

Long:2007:SUa

[Lon07] Jane-Ellen Long. SAGE update. ;login: the USENIX Association newsletter, 32(1): ??, February 2007. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        february-2007-volume-32-
                                        number-1/sage-update.

Long:2008:US

[Lon08] Jane-Ellen Long. Update on SAGE. ;login: the USENIX Association newsletter, 33(5): ??, October 2008. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        october-2008-volume-33-
                                        number-5/update-sage.

Long:2009:GM

[Lon09a] Jane-Ellen Long. Got a minute? ;login: the USENIX Association newsletter, 34(1): ??, February 2009. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
february-2009-volume-34-
number-1/got-minute.

Long:2009:UGG

[Lon09b] Jane-Ellen Long. USENIX goes greener. ;login: the USE-NIX Association newsletter, 34(3):??, June 2009. CO-DEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/june-2009-volume-
34-number-3/usenix-goes-
greener.

Losh:2002:PUF

[Los02] M. Warner Losh. Problems updating FreeBSD’s card system from ISA to PCI. In USENIX [USE02a], pages 11–

18. ISBN 1-880446-02-2. LCCN QA76.76.O63 B736

2002. URL http://www.
usenix.org/publications/
library/proceedings/bsdcon02/
losh.html.

Loza:2005:FTF

[Loz05a] Boris Loza. Finding trojans for fun and profit. ;login: the USENIX Association newsletter, 30(5):??, October 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-10/pdfs/loza.pdf.

Loza:2005:UAD

[Loz05b] Boris Loza. Under attack:

Dealing with missing UNIX files. ;login: the USENIX Association newsletter, 30(3): ??, June 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-06/pdfs/loza0506.pdf.

Larus:2002:UCS

[LP02] James R. Larus and Michael Parkes. Using cohort scheduling to enhance server performance. In USENIX [USE02e], page ?? ISBN 1-880446-00-6. LCCN QA76.8.U65

U84 2002. URL http://www.
usenix.org/publications/
library/proceedings/usenix02/
larus.html.

Lancaster:2001:MRW

[LR01] Larry Lancaster and Alan Rowe. Measuring realworld data availability. In USENIX [USE01a], page ?? ISBN 1-880446-05-7. LCCN

????  URL http://db.
usenix.org/publications/
library/proceedings/lisa2001/
tech/lancaster.html.

Leff:2004:AES

[LR04] Avraham Leff and James T. Rayfield. Alternative edgeserver architectures for Enterprise JavaBeans applications. In Jacobsen [Jac04], pages 195–?? CODEN LNCSD9. ISBN 3-540-23428-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.M54 I33 2004. URL ftp://

uiarchive.cso.uiuc.edu/

pub/etext/gutenberg/;
http://www.springerlink.
com/openurl.asp?genre=issue&
issn=0302-9743&volume=3231;
http://www.springerlink.
com/openurl.asp?genre=volume&
id=doi:10.1007/b101561.

Larson:2000:IAV

[LS00] Karl Larson and Todd Stansell. Improving availability in VERITAS environments. In USE-NIX [USE00g], page ?? ISBN 1-880446-13-8. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/lisa2000/
larson.html.

Lindemann:2001:ICT

[LS01a] Mark Lindemann and Sean W.

Smith. Improving DES coprocessor throughput for short operations. In USENIX [USE01h], page ?? ISBN 1-880446-18-9, 1-880446-07-

3. LCCN QA76.9.A25 U565

2001. URL http://www.
usenix.org/publications/
library/proceedings/sec01/
lindemann.html.

Loscocco:2001:IFS

[LS01b] Peter Loscocco and Stephen Smalley. Integrating flexible support for security policies into the Linux operating system. In USENIX [USE01f], page ?? ISBN 1-880446-10-3. LCCN QA76.8.U65

U84 2001. URL http://www.
usenix.org/publications/

library/proceedings/usenix01/
freenix01/loscocco.html.

Leung:2009:SMS

[LSB^{+}09] Andrew W. Leung, Minglong Shao, Timothy Bisson, Shankar Pasupathy, and Ethan L. Miller. Spyglass: Metadata search for largescale storage systems. ;login: the USENIX Association newsletter, 34(3):??, June 2009. CODEN LOGNEM. ISSN 1044-6397. URL

https://www.usenix.org/
publications/login/june-
2009-volume-34-number-3/
spyglass-metadata-search-
large-scale-storage-systems.

Lumb:2000:THD

[LSG^{+}00] Christopher Lumb, Jiri Schindler, Gregory R. Ganger, Erik Riedel, and David F. Nagle. Towards higher disk head utilization: Extracting “free” bandwidth from busy disk drives. In USE-NIX [USE00h], page ?? ISBN 1-880446-16-2. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/osdi2000/
lumb.html.

Lumb:2002:FSO

[LSG02] Christopher R. Lumb, Jiri Schindler, and Gregory R. Ganger. Freeblock scheduling outside of disk firmware. In USENIX [USE02c], page ?? ISBN 1-880446-03-0. LCCN QA76.9.F5 C65 2002. URL

http://www.usenix.org/publications/
library/proceedings/fast02/
lumb.html.

Lucovsky:2000:NWB

[Luc00] Mark Lucovsky. From NT OS/2 to Windows 2000 and beyond? A softwareengineering odyssey, 2000.

URL http://db.usenix.
org/events/usenix-win2000/
invitedtalks/lucovsky.ppt;
 http://db.usenix.org/events/
usenix-win2000/invitedtalks/
lucovsky_html.  Unpub-

lished talk at Fourth Usenix Windows Systems Symposium.

Lucas:2002:FDP

[Luc02] Michael Lucas. The FreeBSD documentation project. In USENIX [USE02a], pages viii + 151. ISBN 1-880446-02-

2. LCCN QA76.76.O63 B736

2002. URL http://www.
usenix.org/publications/
library/proceedings/bsdcon02/

lucas.pdf. Unpublished invited talk, BSDCON2002: Growing the BSD Community, February 11–14, 2002, Cathedral Hill Hotel, San Francisco, CA.

Lucas:2005:FS

[Luc05] Michael W. Lucas. FreeBSD 5 SMPng. ;login: the USENIX Association newsletter, 30(5): ??, October 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-10/pdfs/lucas.pdf.

Lumb:2001:LCH

[Lum01] Ian Lumb. Linux clustering for high-performance computing. ;login: the USE-NIX Association newsletter, 26(5):??, August 2001. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2001-08/pdfs/lumb.
                                        pdf.

Lumb:2003:PHR

[Lum03] Ian Lumb. Production HPC reinvented. ;login: the USE-NIX Association newsletter, 28(4):??, August 2003. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2003-08/pdfs/lumb.
                                        pdf.

Liao:2002:UTC

[LV02] Yihua Liao and V. Rao Vemuri. Using text categorization techniques for intrusion detection. In USE-NIX [USE02b], pages 51–59. ISBN 1-931971-00-5. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/sec02/
                                        liao.html.

Wong:2000:OSS

[lWS00] Alexander Ya li Wong and Margo I. Seltzer. Operating system support for multi-user, remote, graphical interaction. In USE-NIX [USE00a], pages 183– 196. ISBN 1-880446-22-7.

LCCN ???? URL http:

//www.usenix.org/events/
usenix2000/general/wong.
html.

Luo:2001:SSS

[LY01] Mon-Yen Luo and Chu-Sing Yang. System support for scalable, reliable, and highly manageable Web hosting service. In USENIX [USE01c], page ?? ISBN 1-880446-14-6. LCCN TK5105.875.I57

U96 2001. URL http://www.
usenix.org/publications/
library/proceedings/usits01/
luo.html.

Magoutis:2002:SPD

[MAF^{+}02] Kostas Magoutis, Salimah Addetia, Alexandra Fedorova, Margo Seltzer, Jeff Chase, Drew Gallatin, Richard Kisley, Rajiv Wickremesinghe, and Eran Gabber. Structure and performance of the Direct Access File System (DAFS). In USE-NIX [USE02d], page ?? ISBN 1-880446-01-4. LCCN QA76.8.U65 P765 2002.

URL http://www.usenix.
org/publications/library/
proceedings/usenix02/magoutis.
html.

Magoutis:2002:DID

[Mag02] Kostas Magoutis. Design and implementation of a Direct Access File System (DAFS) kernel server for FreeBSD. In USENIX [USE02a], pages 65–76. ISBN 1-880446-02-

2. LCCN QA76.76.O63 B736

                                        2002. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/bsdcon02/
                                        magoutis.html.

Mahmoud:2004:WJA

[Mah04] Qusay H. Mahmoud. Wireless Java applications development. ;login: the USE-NIX Association newsletter, 29(2):??, April 2004. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2004-04/pdfs/Mahmoud.pdf.

Malone:2005:STO

[Mal05]  David Malone.  Security

through obscurity: a review of a few of FreeBSD’s lesser-known security capabilities. ;login: the USENIX Association newsletter, 30(5): ??, October 2005. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2005-10/pdfs/malone.pdf.

Malone:2006:UHW

[Mal06] David Malone. Unwanted HTTP: Who has the time? ;login: the USENIX Association newsletter, 31(2):49– 54, April 2006. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2006-04/pdfs/malone.pdf.

Manning:2002:AD

[Man02] Bill Manning. Adventures in DNS. In USENIX [USE02e], page 355. ISBN 1-880446-00-6. LCCN QA76.8.U65

U84 2002. URL http://www.
usenix.org/publications/
library/proceedings/usenix02/
tech/techonefile.html. Un-

published invited talk, 2002 USENIX Annual Technical Conference, June 10-15, 2002, Monterey Conference Center, Monterey, CA.

Manzuik:2005:YDO

[Man05] Steve Manzuik. Your defense is offensive. ;login: the USENIX Association newsletter, 30(4):??, August 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-08/pdfs/manzuik.pdf.

Maone:2009:HWN

[Mao09] Giorgio Maone. Hardening the Web with NoScript. ;login: the USENIX Association newsletter, 34(6):??, December 2009. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
december-2009-volume-34-
number-6/hardening-web-
noscript.

Martin:2000:NSD

[Mar00] Brian Martin. A note on security disclosures. ;login: the USENIX Association newsletter, 25(8):??, December 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-12/pdfs/martin.pdf.

Marti:2001:ELS

[Mar01a] Don Marti. Embedded Linux success stories, or, it’s beginning to look a lot like a Linux Christmas. In USENIX [USE01d], page ???? ISBN ???? LCCN ???? URL http:

                                        //www.linuxshowcase.org/

tech.html. Unpublished talk, 5th Annual Linux Showcase and Conference, November 5–10, Oakland, CA.

Maruyama:2001:ODI

[Mar01b] Fuyuhiko Maruyama. Open-JIT 2: The design and implementation of application framework for JIT compilers. In USENIX [USE01g], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38

                                        J42 2001. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/jvm01/
                                        JVM_wips/S17.pdf.

Maroti:2004:DFR

[Mar04]  Miklós Maróti.  Directed

flood-routing framework for wireless sensor networks. In Jacobsen [Jac04], pages 99– ?? CODEN LNCSD9. ISBN 3-540-23428-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.M54 I33 2004. URL ftp://

                                        uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;

http://www.springerlink.
com/openurl.asp?genre=issue&
issn=0302-9743&volume=3231;
http://www.springerlink.
com/openurl.asp?genre=volume&
id=doi:10.1007/b101561.

Mason:2000:DTH

[Mas00] Dave Mason. DocBook: a tutorial for hackers and writers. In USENIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN ????

URL http://www.usenix.
org/publications/library/
proceedings/als2000/mason.
html.

Massey:2001:XXP

[Mas01] Bart Massey. XCB: An X protocol C binding. In USE-NIX [USE01d], page ?? ISBN ???? LCCN ???? URL http:

//www.linuxshowcase.org/
massey.html.

Mashey:2002:SST

[Mas02]  John R. Mashey.  Soft-

ware strategy from the 1980 time capsule. In USE-NIX [USE02a], page ?? ISBN 1-880446-02-2. LCCN QA76.76.O63 B736 2002.

URL http://www.usenix.
org/publications/library/
proceedings/bsdcon02/mashey_
army/;   http://www.usenix.
org/publications/library/
proceedings/bsdcon02/mashey_

small/. Keynote address.

Mazires:2001:TUL

[Maz01] David Mazières. A toolkit for user-level file systems. In USENIX [USE01b], page ?? ISBN 1-880446-09-X. LCCN QA76.8.U65 U84 2001. URL

http://www.usenix.org/publications/
library/proceedings/usenix01/
mazieres.html.

Martin:2000:AGO

[MB00] Joel D. Martin and Aaron D. Brooks. Automated generic operating system installation and maintenance. In USENIX [USE00b], page ?? ISBN 1-880446-19-7. LCCN

????  URL http://db.
usenix.org/publications/
library/proceedings/lisa-
nt2000/martin.html.

Miu:2001:DHC

[MB01] Allen Miu and Paramvir Bahl. Dynamic host configuration for managing mobility between public and private networks. In USE-NIX [USE01c], page ?? ISBN 1-880446-14-6. LCCN TK5105.875.I57 U96 2001.

URL http://www.usenix.
org/publications/library/
proceedings/usits01/miu.
html.

Maniatis:2002:EAS

[MB02a] Petros Maniatis and Mary Baker. Enabling the archival storage of signed documents. In USENIX [USE02c], page ?? ISBN 1-880446-03-

0. LCCN QA76.9.F5 C65

2002. URL http://www.
usenix.org/publications/
library/proceedings/fast02/
maniatis.html.

Maniatis:2002:SHP

[MB02b] Petros Maniatis and Mary Baker. Secure history preservation through timeline entanglement. In USENIX [USE02b], pages 297–312. ISBN 1-931971-00-5. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/sec02/
maniatis.html.

Massey:2002:XMZ

[MB02c] Bart Massey and Robert T.

Bauer. X meets Z: Verifying correctness in the presence of POSIX threads. In USENIX [USE02d], page ?? ISBN 1-880446-01-4. LCCN QA76.8.U65 P765 2002.

URL http://www.usenix.
org/publications/library/
proceedings/usenix02/tech/
freenix/massey.html.

McGarvey:2001:BCD

[MCB^{+}01] Brian McGarvey, Robert Cicconetti, Nathan Bushyager, Edan Dalton, and Manos Tentzeris. Beowulf cluster design for scientific PDE models. In USENIX [USE01d], page ?? ISBN ???? LCCN ???? URL http:

//www.linuxshowcase.org/
mcgarvey.html.

McCluskey:2000:JPa

[McC00a] Glen McCluskey. Java performance. ;login: the USENIX Association newsletter, 25(1): ??, February 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-2/features/javaperformance.
                                        html.

McCluskey:2000:JPb

[McC00b] Glen McCluskey. Java performance. ;login: the USE-NIX Association newsletter, 25(2):??, April 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-4/features/java.html.

McCluskey:2000:JPc

[McC00c] Glen McCluskey. Java performance. ;login: the USE-NIX Association newsletter, 25(3):??, June 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-6/features/java.html.

McCluskey:2000:JPd

[McC00d] Glen McCluskey. Java performance. ;login: the USENIX Association newsletter, 25(5): ??, August 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-8/features/java.html.

McCluskey:2000:JPe

[McC00e] Glen McCluskey. Java performance. ;login: the USENIX Association newsletter, 25(6): ??, October 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/

2000-10/pdfs/javaperformance.
pdf.

McCluskey:2000:JPf

[McC00f] Glen McCluskey. Java performance. ;login: the USENIX Association newsletter, 25(8): ??, December 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-12/pdfs/javaperf.pdf.

McCrickard:2000:ATE

[McC00g] D. Scott McCrickard. Agentk:

a toolkit for enhancing agent interfaces. ;login: the USE-NIX Association newsletter, 25(4):??, July 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-7/features/agentk.
html.

McCluskey:2001:DC

[McC01a]  Glen McCluskey.  Dec-

larations in C9X. ;login: the USENIX Association newsletter, 26(3): ??, June 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-06/pdfs/mccluskey.
pdf.

McCluskey:2001:FAM

[McC01b] Glen McCluskey. Flexible array members and designators in C9X. ;login: the USE-NIX Association newsletter, 26(4):??, July 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-07/pdfs/mccluskey.
                                        pdf.

McCluskey:2001:JPa

[McC01c] Glen McCluskey. Java performance. ;login: the USENIX Association newsletter, 26(1): ??, February 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-02/pdfs/mccluskey.
                                        pdf.

McCluskey:2001:JPb

[McC01d] Glen McCluskey. Java performance. ;login: the USE-NIX Association newsletter, 26(2):??, April 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-04/pdfs/mccluskey1.
                                        pdf.

McCluskey:2001:SNN

[McC01e]  Glen McCluskey.  Some

new numeric programming features in C9X. ;login: the USENIX Association newsletter, 26(8):??, December 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-12/pdfs/mccluskey.
                                        pdf.

McCluskey:2001:UNT

[McC01f] Glen McCluskey. Using new types and values in C9X. ;login: the USENIX Association newsletter, 26(2): ??, April 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-04/pdfs/mccluskey2.
pdf.

McCluskey:2001:VLA

[McC01g]  Glen McCluskey.  Vari-

able length arrays. ;login: the USENIX Association newsletter, 26(6):??, October 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-10/pdfs/mccluskey.
pdf.

McCluskey:2002:CC

[McC02a]  Glen McCluskey.  C99

and compatibility. ;login: the USENIX Association newsletter, 27(5):??, October 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-10/pdfs/mccluskey.
pdf.

McCluskey:2002:MLF

[McC02b] Glen McCluskey. Math library functions in C9X. ;login: the USENIX Association newsletter, 27(2):9– 13, April 2002. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2002-04/pdfs/mccluskey.
                                        pdf.

McCluskey:2002:NFC

[McC02c]  Glen McCluskey.  New

I/O features in C9X. ;login: the USENIX Association newsletter, 27(3):31– 35, June 2002. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2002-06/pdfs/mccluskey.
                                        pdf.

McCluskey:2002:NPF

[McC02d] Glen McCluskey. New preprocessor features in C9X. ;login: the USENIX Association newsletter, 27(1):14– 17, February 2002. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2002-02/openpdfs/
                                        mccluskey.pdf.

McCluskey:2002:WC

[McC02e] Glen McCluskey. Wide characters. ;login: the USENIX Association newsletter, 27(4): 12–15, August 2002. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2002-08/pdfs/mccluskey.
                                        pdf.

McCluskey:2003:CT

[McC03a] Glen McCluskey. C# types.

;login: the USENIX Association newsletter, 28(3): ??, June 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-06/pdfs/mccluskey.
pdf.

McCluskey:2003:ECH

[McC03b] Glen McCluskey. Examining the C# “hello world” program. ;login: the USE-NIX Association newsletter, 28(2):??, April 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-04/pdfs/mccluskey.
pdf.

McCluskey:2003:IC

[McC03c] Glen McCluskey. An introduction to C#. ;login: the USENIX Association newsletter, 28(1):??, February 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-02/pdfs/mccluskey.
pdf.

McCluskey:2003:UCP

[McC03d] Glen McCluskey. Using C# properties and static members. ;login: the USENIX Association newsletter, 28(5): ??, October 2003. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2003-10/pdfs/mccluskey.
                                        pdf.

McCluskey:2003:WCC

[McC03e]  Glen McCluskey.  Work-

ing with C# classes. ;login: the USENIX Association newsletter, 28(4):??, August 2003. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2003-08/pdfs/mccluskey.
                                        pdf.

McCluskey:2004:COO

[McC04a] Glen McCluskey. C# overloaded operators. ;login: the USENIX Association newsletter, 29(3): ??, June 2004. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2004-06/pdfs/mccluskey.
                                        pdf.

McCluskey:2004:MUC

[McC04b] Glen McCluskey. Making use of C# collections. ;login: the USENIX Association newsletter, 29(5):??, October 2004. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2004-10/pdfs/mccluskey.
                                        pdf.

McCluskey:2004:UCA

[McC04c]  Glen McCluskey.  Using

C# abstract classes. ;login: the USENIX Association newsletter, 29(2): ??, April 2004. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2004-04/pdfs/McCluskey.
pdf.

McCluskey:2004:UCR

[McC04d]  Glen McCluskey.  Us-

ing C# reflection. ;login: the USENIX Association newsletter, 29(4):??, August 2004. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2004-08/pdfs/mccluskey.
pdf.

McCluskey:2004:WCI

[McC04e] Glen McCluskey. Working with C# interfaces. ;login: the USENIX Association newsletter, 29(1):??, February 2004. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2004-02/pdfs/mccluskey.
pdf.

McCluskey:2005:WCS

[McC05a] Glen McCluskey. Working with C# serialization. ;login: the USENIX Association newsletter, 30(1):??, February 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-02/pdfs/mccluskey.
pdf.

McCullough:2005:SAF

[McC05b]  Mark McCullough.  Se-

cure automated file transfer. ;login: the USENIX Association newsletter, 30(4): ??, August 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-08/pdfs/mccullough.
pdf.

McCool:2008:AHP

[McC08] Michael D. McCool. Achieving high performance by targeting multiple parallelism mechanisms. ;login: the USE-NIX Association newsletter, 33(2):??, April 2008. CO-DEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/april-2008-volume-
33-number-2/achieving-high-
performance-targeting-multiple.

Mao:2002:PEE

[MCD^{+}02] Zhuoqing Mao, Charles Cranor, Fred Douglis, Michael Rabinovich, Oliver Spatscheck, and Jia Wang. A precise and efficient evaluation of the proximity between Web clients and their local DNS servers. In USENIX [USE02e], page ?? ISBN 1-880446-00-6. LCCN QA76.8.U65 U84 2002. URL

http://www.usenix.org/publications/
library/proceedings/usenix02/
mao.html.

Mathur:2007:ENG

[MCD07] Avantika Mathur, Mingming Cao, and Andreas Dilger. Ext4: The next generation of the Ext3 file system. ;login: the USENIX Association newsletter, 32(3):??, June 2007. CODEN LOGNEM. ISSN 1044-6397. URL

https://www.usenix.org/
publications/login/june-
2007-volume-32-number-3/
ext4-next-generation-ext3-
file-system.

McGuire:2000:EL

[McG00]  Nicholas McGuire.  Em-

bedded Linux. In USE-NIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/als2000/
mcguire.html.

McKusick:2002:RFB

[McK02] Marshall Kirk McKusick.

Running “fsck” in the background. In USE-NIX [USE02a], pages 55–

64. ISBN 1-880446-02-2. LCCN QA76.76.O63 B736

2002. URL http://www.
usenix.org/publications/
library/proceedings/bsdcon02/
mckusick.html.

McKusick:2004:JFF

[McK04a] Kirk McKusick. The jail facility in FreeBSD 5.2. ;login: the USENIX Association newsletter, 29(4):??, August 2004. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2004-08/pdfs/mckusick.pdf.

McKusick:2004:RSG

[McK04b] Kirk McKusick. Restructuring of SAGE governance. ;login: the USENIX Association newsletter, 29(4): ??, August 2004. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2004-08/openpdfs/unotes.
                                        pdf.

McKusick:2006:DPF

[McK06] Marshall Kirk McKusick. Disks from the perspective of a file system. ;login: the USENIX Association newsletter, 31(3):??, June 2006. CODEN LOGNEM. ISSN 1044-6397. URL

                                        https://www.usenix.org/
                                        publications/login/june-
                                        2006-volume-31-number-3/
                                        disks-perspective-file-
                                        system.

McKusick:2007:BHB

[McK07] Marshall Kirk McKusick. A brief history of the BSD Fast File System. ;login: the USENIX Association newsletter, 32(3):??, June 2007. CODEN LOGNEM. ISSN 1044-6397. URL

https://www.usenix.org/
publications/login/june-
2007-volume-32-number-3/
brief-history-bsd-fast-
file-system.

McQuillan:2000:LTS

[McQ00] Jim McQuillan. The Linux Terminal Server Project: Thin clients and Linux. In USENIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/als2000/
mcquillan.html.

Menon:2006:ONV

[MCZ06] Aravind Menon, Alan L. Cox, and Willy Zwaenepoel. Optimizing network virtualization in Xen. In USENIX [USE06], page ?? ISBN 1-931971-44-

7. LCCN ???? URL http:

//www.usenix.org/events/
usenix06/tech/menon.html.

Miller:2000:RBS

[MD00] Allan Miller and Alex Donnini. Relieving the burden of system administration through support automation. In USE-NIX [USE00g], page ?? ISBN 1-880446-13-8. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/lisa2000/
miller.html.

Mulligan:2005:UNN

[MD05] Tara Mulligan and Anne Dickison. USENIX notes: New membership benefits. ;login: the USENIX Association newsletter, 30(4): ??, August 2005. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2005-08/openpdfs/
                                        benefits.pdf.

Michalickova:2000:SAP

[MDH00] Katerina Michalickova, Moyez Dharsee, and Christopher W. V. Hogue. Sequence analysis on a 216-processor Beowulf cluster. In USE-NIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/als2000/
                                        michalickova.html.

Meek:2000:WAN

[Mee00]  Jon T. Meek.  Wide

area network packet capture and analysis. In USE-NIX [USE00g], page ?? ISBN 1-880446-13-8. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/lisa2000/
                                        meek.html.

Mewburn:2001:DIN

[Mew01] Luke Mewburn. The design and implementation of the NetBSD rc.d system. In USENIX [USE01f], page ?? ISBN 1-880446-10-3. LCCN QA76.8.U65

U84 2001. URL http://www.
usenix.org/publications/
library/proceedings/usenix01/
freenix01/mewburn.html.

Meyerovich:2009:RBP

[Mey09] Leo Meyerovich. Rethinking browser performance. ;login: the USENIX Association newsletter, 34(4): ??, August 2009. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2009-volume-34-number-
4/rethinking-browser-performance.

Milan-Franco:2004:AMD

[MFJPPMK04] Jesús M. Milan-Franco, Ricardo Jiménez-Peris, Marta Patiño-Martínez, and Bettina Kemme. Adaptive middleware for data replication. In Jacobsen [Jac04], pages 175–?? CODEN LNCSD9. ISBN 3-540-23428-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.M54 I33 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.springerlink.
com/openurl.asp?genre=issue&
issn=0302-9743&volume=3231;
 http://www.springerlink.
com/openurl.asp?genre=volume&
id=doi:10.1007/b101561.

Molloy:2001:SLS

[MH01] Stephen Molloy and Peter Honeyman. Scalable Linux scheduling. In USE-NIX [USE01f], page ?? ISBN 1-880446-10-3. LCCN QA76.8.U65 U84 2001. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/usenix01/
                                        freenix01/molloy.html.

Milojicic:2000:GMM

[MHM^{+}00] Dejan Milojicic, Steve Hoyle, Alan Messer, Albert Munoz, Lance Russell, Tom Wylegala, Vivekanand Vellanki, and Stephen Childs. Global memory management for a multi-computer system. In USENIX [USE00d], page ?? ISBN 1-880446-20-0. LCCN

                                        ????  URL http://db.
                                        usenix.org/publications/
                                        library/proceedings/usenix-
                                        win2000/milojicic.html.

Minnich:2000:LB

[MHW00] Ronald G. Minnich, James Hendricks, and Dale Webster. The Linux BIOS. In USENIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/als2000/
                                        minnichBIOS.html.

Mirtchovski:2007:WSD

[MI07] Andrey Mirtchovski and Latchesar Ionkov. Why some dead OSes still matter. ;login: the USENIX Association newsletter, 32(5):5– 12, October 2007. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.

org/publications/login/
october-2007-volume-32-
number-5/why-some-dead-
oses-still-matter.

Midkiff:2001:JCM

[Mid01] Sam Midkiff. A Java compiler for many memory models. In USENIX [USE01g], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38

J42 2001. URL http://www.
usenix.org/publications/
library/proceedings/jvm01/
JVM_wips/S11.pdf.

Miltchev:2002:SRC

[MIK02] Stefan Miltchev, Sotiris Ioannidis, and Angelos Keromytis. A study of the relative costs of network security protocols. In USENIX [USE02d], page ?? ISBN 1-880446-01-

4. LCCN QA76.8.U65 P765

2002. URL http://www.
usenix.org/publications/
library/proceedings/usenix02/
tech/freenix/miltchev.html.

Miller:2002:SMP

[Mil02] Frank Miller. Simple memory protection for embedded operating system kernels. In USENIX [USE02d], page ?? ISBN 1-880446-01-

4. LCCN QA76.8.U65 P765

2002. URL http://www.
usenix.org/publications/
library/proceedings/usenix02/
tech/freenix/miller.html.

Miller:2005:CRS

[Mil05] Hal Miller. Customer requirements specifications for system administrators. ;login: the USENIX Association newsletter, 30(2): ??, April 2005. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2005-04/pdfs/miller0504.
                                        pdf.

Minnich:2000:LLB

[Min00] Ronald Minnich. LOBOS (Linux OS Boots OS): Booting a kernel in 32-bit mode. In USENIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/als2000/
                                        minnichLOBOS.html.

Minnich:2001:SHP

[Min01] Ronald G. Minnich. Supermon: High-performance monitoring for Linux clusters. In USENIX [USE01d], page ?? ISBN ???? LCCN ???? URL http:

                                        //www.linuxshowcase.org/
                                        minnich.html.

Mitchell:2002:FMC

[Mit02] John C. Mitchell. Formal methods and computer security, 2002. URL http://www.

                                        usenix.org/publications/
                                        library/proceedings/sec02/

tech.html. Unpublished.

Marmorstein:2007:DFP

[MK07] Robert Marmorstein and Phil Kearns. Debugging a firewall policy with policy mapping. ;login: the USENIX Association newsletter, 32(1): ??, February 2007. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
february-2007-volume-32-
number-1/debugging-firewall-
policy-policy-mapping.

McDougall:2006:MCM

[ML06] Richard McDougall and James Laudon. Multi-core microprocessors are here. ;login: the USENIX Association newsletter, 31(5):??, October 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
october-2006-volume-31-
number-5/multi-core-microprocessors-
are-here.

Miller:2002:SSN

[MLFR02] Ethan Miller, Darrell Long, William Freeman, and Benjamin Reed. Strong security for network-attached storage. In USENIX [USE02c], page ?? ISBN 1-880446-03-

0. LCCN QA76.9.F5 C65

2002. URL http://www.
usenix.org/publications/
library/proceedings/fast02/
miller.html.

Miller:2000:ICS

[MM00] Robert C. Miller and Brad A.

Myers. Integrating a command shell into a Web browser. In USENIX [USE00a], pages 171–182. ISBN 1-880446-22-7. LCCN ???? URL http:/

                                        /www.usenix.org/events/
                                        usenix2000/general/miller.
                                        html.

Miller:2001:ISE

[MM01] Robert C. Miller and Brad A. Myers. Interactive simultaneous editing of multiple text regions. In USE-NIX [USE01b], page ?? ISBN 1-880446-09-X. LCCN QA76.8.U65 U84 2001. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/usenix01/
                                        miller.html.

Megiddo:2003:OL

[MM03] Nimrod Megiddo and Dharmendra S. Modha. One up on LRU. ;login: the USENIX Association newsletter, 28(4): ??, August 2003. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2003-08/pdfs/Megiddo.pdf.

Myers:2000:EWD

[MMBE00] Brad A. Myers, Robert C. Miller, Benjamin Bostwick, and Carl Evankovich. Extending the Windows desktop interface with connected handheld computers. In USENIX [USE00d], page ?? ISBN 1-880446-20-0. LCCN

                                        ????  URL http://db.
                                        usenix.org/publications/
                                        library/proceedings/usenix-
                                        win2000/myers.html.

Marvie:2001:TTC

[MMGL01] Raphaël Marvie, Philippe Merle, Jean-Marc Geib, and Sylvain Leblanc. TORBA: Trading contracts for CORBA. In USENIX [USE01e], page ?? ISBN 1-880446-12-X. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/coots01/
marvie.html.

McCarthy:2000:MPG

[MMS00] Cliff McCarthy, Michael Murphy, and Indira Subramanian. Meeting performance goals with the HP-UX workload manager. In USENIX [USE00f], page ?? ISBN 1-880446-15-4. LCCN QA76.76.S95 W67 2000.

URL http://www.usenix.
org/publications/library/
proceedings/osdi2000/wiess2000/
subramanian.html.

Mitchell:2000:NCN

[MNH00] Ellen L. Mitchell, Eric Nelson, and David Hess. ND: a comprehensive network administration and analysis tool. In USE-NIX [USE00g], page ?? ISBN 1-880446-13-8. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/lisa2000/
mitchell.html.

Memik:2002:EIF

[MoCA^{+}02] Gokhan Memik, University of California, Los Angeles, Mahmut Kandemir, and Alok Choudhary. Exploiting inter-file access patterns using multi-collective I/O. In USENIX [USE02c], page ?? ISBN 1-880446-03-

0. LCCN QA76.9.F5 C65

2002. URL http://www.
usenix.org/publications/
library/proceedings/fast02/
memik.html.

Moen:2002:RSL

[Moe02] Rick Moen. A recipe for a successful Linux user group. ;login: the USENIX Association newsletter, 27(2): ??, April 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-04/pdfs/moen.pdf.

Moore:2000:DPG

[Moo00] Richard J. Moore. Dynamic probes and generalised kernel hooks interface for Linux. In USENIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/als2000/
moore.html.

Moore:2001:UDT

[Moo01] Richard J. Moore. A universal dynamic trace for Linux and other operating systems. In USENIX [USE01f], page ?? ISBN 1-880446-10-3. LCCN QA76.8.U65

U84 2001. URL http://www.
usenix.org/publications/
library/proceedings/usenix01/
freenix01/moore.html.

Moore:2002:NTO

[Moo02]  David Moore.  Network

telescopes: Observing small or distant security events,

2002. URL http://www.
usenix.org/publications/
library/proceedings/sec02/

tech.html. Unpublished.

Morgan:2000:ILF

[Mor00] Richard R. Morgan. Introducing Linux file services into a Windows NT network. In USENIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/als2000/
morgan.html.

Morgan:2001:RTP

[Mor01] Kevin Morgan. Real-time performance in native Linux. In USENIX [USE01d], page ???? ISBN ???? LCCN

????  URL http://www.
linuxshowcase.org/tech.

html. Unpublished talk, 5th Annual Linux Showcase and Conference, November 5–10, Oakland, CA.

Morris:2002:SAP

[Mor02]  Robert Morris.  Stor-

age: From atoms to people. In USENIX [USE02c], page ?? ISBN 1-880446-03-0. LCCN QA76.9.F5 C65

2002. URL http://www.
usenix.org/publications/
library/proceedings/fast02/
morris.pdf.

Moskowitz:2003:EYO

[Mos03a]  Adam Moskowitz.  “eat

your own dog food”. ;login: the USENIX Association newsletter, 28(5):??, October 2003. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2003-10/pdfs/moskowitz.
                                        pdf.

Moskowitz:2003:CU

[Mos03b]  Adam Moskowitz.  On

choosing usernames. ;login: the USENIX Association newsletter, 28(4):??, August 2003. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2003-08/pdfs/moskowitz.
                                        pdf.

Massey:2001:NLP

[MP01a] Bart Massey and Keith Packard. Nickle: Language principles and pragmatics. In USENIX [USE01f], page ?? ISBN 1-880446-10-3. LCCN QA76.8.U65

                                        U84 2001. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/usenix01/
                                        freenix01/massey.html.

Moore:2001:EFJ

[MP01b] J. Strother Moore and George M. Porter. An executable formal Java Virtual Machine thread model. In USENIX [USE01g], page ??

ISBN 1-880446-11-1. LCCN QA76.73.J38 J42 2001. URL

http://www.usenix.org/publications/
library/proceedings/jvm01/
moore.html.

Murty:2008:DHP

[MPC^{+}08] Rohan Murty, Jitendra Padhye, Ranver Chandra, Alec Wolman, and Brian Zill. Designing high-performance enterprise Wi–Fi networks. ;login: the USENIX Association newsletter, 33(4): ??, August 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2008-volume-33-number-
4/designing-high-performance-
enterprise-wi-fi.

Mrillon:2000:DIH

[MRC^{+}00] Fabrice Mérillon, Laurent Réveillère, Charles Consel, Renaud Marlet, and Gilles Muller. Devil: An IDL for hardware programming. In USENIX [USE00h], page ?? ISBN 1-880446-16-2. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/osdi2000/
merillon.html.

Monrose:2002:TSG

[MRL^{+}02] Fabian Monrose, Michael Reiter, Qi Li, Daniel P. Lopresti, and Chilin Shih. Toward speech-generated cryptographic keys on resourceconstrained devices. In USE-NIX [USE02b], pages 283–

296. ISBN 1-931971-00-5.

LCCN ???? URL http://

                                        www.usenix.org/publications/
                                        library/proceedings/sec02/
                                        monrose.html.

Mitchell:2000:GSS

[MS00] Mark Mitchell and Alexander Samuel. GCC 3.0: The state of the source. In USENIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/als2000/
                                        mitchell.html.

Martin:2002:DUS

[MS02] David Martin and Andrew Schulman. Deanonymizing users of the SafeWeb anonymizing service. In USE-NIX [USE02b], pages 123– 137. ISBN 1-931971-00-5. LCCN ???? URL http://

                                        www.usenix.org/publications/
                                        library/proceedings/sec02/
                                        martin.html.

Melski:2000:SCX

[MSO00] Eric Melski, Scott Stanton, and John Ousterhout. Scriptics Connect: An XML integration server based on Tcl. In USE-NIX [USE00e], page ?? ISBN 1-880446-24-3. LCCN

                                        ????  URL http://db.
                                        usenix.org/publications/
                                        library/proceedings/tcl2k/
                                        melski.html.

Metzger:2009:ITS

[MSV09] Perry Metzger, William Allen Simpson, and Paul Vixie.

Improving TCP security with robust cookies. ;login: the USENIX Association newsletter, 34(6):??, December 2009. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
december-2009-volume-34-
number-6/improving-tcp-
security-robust-cookies.

Mudge:2000:MDA

[Mud00a] Mudge. Methods for detecting addressable promiscuous devices, 2000. URL http://

www.usenix.org/publications/
library/proceedings/sec2000/
invitedtalks/mudge_html.

Unpublished invited talk at Ninth USENIX Security Symposium.

Mudge:2000:SDM

[Mud00b]  Mudge.  Security devices

that might not be. ;login: the USENIX Association newsletter, 25(7):??, November 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-11/pdfs/mudge.pdf.

Mudge:2003:IT

[Mud03] Mudge. Insider threat. ;login: the USENIX Association newsletter, 28(6):??, December 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-12/pdfs/mudge.pdf.

Muentz:2007:HYS

[Mue07a] Alexander Muentz. Hardening your systems against litigation. ;login: the USENIX Association newsletter, 32(5): ??, October 2007. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        october-2007-volume-32-
                                        number-5/hardening-your-
                                        systems-against-litigation.

Muentz:2007:SKB

[Mue07b] Alexander Muentz. Script kiddies with briefcases: The legal system as threat. ;login: the USENIX Association newsletter, 32(3):??, June 2007. CODEN LOGNEM. ISSN 1044-6397. URL

                                        https://www.usenix.org/
                                        publications/login/june-
                                        2007-volume-32-number-3/
                                        script-kiddies-briefcases-
                                        legal-system-threat.

Mulligan:2004:MN

[Mul04a] Tara Mulligan. Membership news. ;login: the USENIX Association newsletter, 29(5): ??, October 2004. CODEN LOGNEM. ISSN 1044-6397.

Mulligan:2004:SUBb

[Mul04b] Tara Mulligan. Summary of the USENIX Board of Directors meetings. ;login: the USENIX Association newsletter, 29(5):??, October 2004. CODEN LOGNEM. ISSN 1044-6397.

Mulligan:2005:SUB

[Mul05a] Tara Mulligan. Summary of USENIX Board of Directors actions. ;login: the USE-NIX Association newsletter, 30(1):??, February 2005. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2005-02/openpdfs/
directors.pdf.

Mulligan:2005:UNS

[Mul05b] Tara Mulligan. USENIX notes: Summary of the USE-NIX Board of Directors actions. ;login: the USE-NIX Association newsletter, 30(3):??, June 2005. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2005-06/openpdfs/
minutes0506.pdf.

Munzner:2002:IVS

[Mun02] Tamara Munzner. Information visualization for systems people. In USE-NIX [USE02e], page 355. ISBN 1-880446-00-6. LCCN QA76.8.U65 U84 2002. URL

http://www.usenix.org/publications/
library/proceedings/usenix02/
tech/techonefile.html. Un-

published invited talk, 2002 USENIX Annual Technical Conference, June 10-15, 2002, Monterey Conference Center, Monterey, CA.

Murray:2001:CDC

[Mur01] Eric Murray. Changes in deployment of cryptography, and possible causes,

2001. URL http://www.
usenix.org/publications/
library/proceedings/sec01/
murray/index.htm. Unpub-

lished invited talk, Tenth USENIX Security Symposium, August 13–17, 2001, Washington, DC, USA.

Murray:2002:IYP

[Mur02] Mark R. V. Murray. An implementation of the Yarrow PRNG for FreeBSD. In USE-NIX [USE02a], pages 47–

53. ISBN 1-880446-02-2. LCCN QA76.76.O63 B736

2002. URL http://www.
usenix.org/publications/
library/proceedings/bsdcon02/
murray.html.

Mohapatra:2004:GTA

[MV04] Shivajit Mohapatra and Nalini Venkatasubramanian. A game theoretic approach for power aware middleware. In Jacobsen [Jac04], pages 417–?? CODEN LNCSD9. ISBN 3-540-23428-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.M54 I33 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.springerlink.
com/openurl.asp?genre=issue&
issn=0302-9743&volume=3231;
 http://www.springerlink.

com/openurl.asp?genre=volume&
id=doi:10.1007/b101561.

Maheshwari:2000:HBT

[MVS00] Umesh Maheshwari, Radek Vingralek, and Bill Shapiro. How to build a trusted database system on untrusted storage. In USE-NIX [USE00h], page ?? ISBN 1-880446-16-2. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/osdi2000/
maheshwari.html.

Moore:2001:IID

[MVS01] David Moore, Geoffrey M.

Voelker, and Stefan Savage. Inferring Internet denial-of-service activity. In USENIX [USE01h], page ?? ISBN 1-880446-18-9, 1-880446-07-

3. LCCN QA76.9.A25 U565

2001. URL http://www.
usenix.org/publications/
library/proceedings/sec01/
moore.html.

Mikulin:2000:ILH

[MVW00] Dmitry Mikulin, Murali Vijayasundaram, and Loreena Wong. Incremental linking on HP-UX. In USE-NIX [USE00f], page ?? ISBN 1-880446-15-4. LCCN QA76.76.S95 W67 2000.

URL http://www.usenix.
org/publications/library/
proceedings/osdi2000/wiess2000/
mikulin.html.

Mainland:2005:DAR

[MW05] Geoffrey Mainland and Matt Welsh. Distributed, adaptive resource allocation for sensor networks. ;login: the USENIX Association newsletter, 30(5):??, October 2005. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2005-10/pdfs/mainland.pdf.

Mulligan:2004:SUBa

[MY04] Tara Mulligan and Ellie Young. Summary of the USENIX Board of Directors actions. ;login: the USE-NIX Association newsletter, 29(1):??, February 2004. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2004-02/openpdfs/
                                        young.pdf.

Mulligan:2006:UNS

[MY06] Tara Mulligan and Ellie Young. USENIX notes: Summary of USENIX Board of Directors meetings and actions. ;login: the USENIX Association newsletter, 31(1): ??, February 2006. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2006-02/openpdfs/
                                        board.pdf.

McCrickard:2000:SIA

[MZ00] Scott McCrickard and Q. Alex Zhao. Supporting information awareness using animated widgets. In USE-NIX [USE00e], page ?? ISBN 1-880446-24-3. LCCN

????  URL http://db.
usenix.org/publications/
library/proceedings/tcl2k/
mccrickard.html.

Natt:2004:BRP

[Nat04] Joel E. Natt. Book review: Planet Broadband. ;login: the USENIX Association newsletter, 29(5):??, October 2004. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2004-10/openpdfs/review.
pdf.

Nayate:2004:TID

[NDI04] Amol Nayate, Mike Dahlin, and Arun Iyengar. Transparent information dissemination. In Jacobsen [Jac04], pages 212–?? CODEN LNCSD9. ISBN 3-540-23428-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.M54 I33 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.springerlink.
com/openurl.asp?genre=issue&
issn=0302-9743&volume=3231;
 http://www.springerlink.
com/openurl.asp?genre=volume&
id=doi:10.1007/b101561.

Nemeth:2000:RNC

[Nem00a] Evi Nemeth. Report of the Nominating Committee. ;login: the USENIX Association newsletter, 25(1):??, February 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-2/nominating.html.

Nemeth:2000:SD

[Nem00b] Evi Nemeth. Securing the DNS. ;login: the USENIX Association newsletter, 25(7): ??, November 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-11/pdfs/dns.pdf.

Nessett:2000:OSI

[Nes00] Dan Nessett. Open session on intelligence at the network edge. In USENIX [USE00j], page ?? ISBN ???? LCCN

                                        ????  URL http://db.
                                        usenix.org/publications/
                                        library/proceedings/ine2000/
                                        nessett_slides.html;   http:
                                        //db.usenix.org/publications/
                                        library/proceedings/ine2000/
                                        nessett_slides.ppt.

Novikov:2000:FMT

[NH00] Alexei Novikov and Martin Hamilton. FTP mirror tracker: a few steps towards URN. In USE-NIX [USE00g], page ?? ISBN 1-880446-13-8. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/lisa2000/
                                        novikov.html.

Nicholson:2000:EPW

[Nic00a] John Nicholson. Electronic privacy in the workplace. ;login: the USENIX Association newsletter, 25(2): ??, April 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-4/features/electronic.
html.

Nicholson:2000:HW

[Nic00b] John Nicholson. The hostile workplace. ;login: the USE-NIX Association newsletter, 25(3):??, June 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-6/features/hostile.
html.

Nicholson:2000:NLN

[Nic00c] John Nicholson. National laws and the Net. ;login: the USENIX Association newsletter, 25(8):??, December 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-12/pdfs/nicholson.
pdf.

Nicholson:2000:PC

[Nic00d]  John Nicholson.  Polite-

ness in computing. ;login: the USENIX Association newsletter, 25(1):??, February 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-2/features/politeness.
html.

Nicholson:2000:TWW

[Nic00e]  John Nicholson.  Tam-

ing the Wild West. ;login: the USENIX Association newsletter, 25(5):??, August 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-8/features/taming.
html.

Nicholson:2001:YBC

[Nic01a] John Nicholson. You’ve been cracked … and now you’re sued. ;login: the USENIX Association newsletter, 26(2): ??, April 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-04/pdfs/nicholson.
pdf.

Nicholsonp:2001:CPW

[Nic01b] John Nicholsonp. The corporate policy Web. ;login: the USENIX Association newsletter, 26(6):??, October 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-10/pdfs/nicholson.
pdf.

Nicholson:2002:DP

[Nic02a] John Nicholson. Data privacy. ;login: the USENIX Association newsletter, 27(2): ??, April 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-04/pdfs/nicholson.
pdf.

Nicholson:2002:SL

[Nic02b]  John Nicholson.  Soft-

ware Licensing 101. ;login: the USENIX Association newsletter, 27(5):??, October 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-10/pdfs/nicholson.
pdf.

Nicholson:2003:CPO

[Nic03a] John Nicholson. Common problems with outsourcing deals and how to avoid them. ;login: the USENIX Association newsletter, 28(1):??, February 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-02/pdfs/nicholson.
pdf.

Nicholson:2003:TOI

[Nic03b]  John Nicholson.  Trends

in the outsourcing industry. ;login: the USENIX Association newsletter, 28(3): ??, June 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/

                                        2003-06/pdfs/nicholson.
                                        pdf.

Nikander:2002:ACP

[Nik02] Pekka Nikander. Authorization and charging in public WLANs using FreeBSD and 802.1. In USENIX [USE02d], page ?? ISBN 1-880446-01-

4. LCCN QA76.8.U65 P765

                                        2002. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/usenix02/
                                        tech/freenix/nikander.html.

Nickless:2000:CCN

[NNW00] Bill Nickless, John-Paul Navarro, and Linda Winkler. Combining Cisco Net-Flow exports with relational database technology for usage statistics, intrusion detection, and network forensics. In USENIX [USE00g], page ?? ISBN 1-880446-13-8. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/lisa2000/
                                        navarro.html.

Norberg:2000:BWN

[Nor00] Stefan Norberg. Building a Windows NT bastion host. ;login: the USENIX Association newsletter, 25(1):??, February 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-2/features/bastionhost.
                                        html.

Ngo:2001:IJJ

[NS01] Ton Ngo and Steve Smith.

Implementing JNI in Java for Jalapeño. In USE-NIX [USE01g], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38 J42 2001. URL

http://www.usenix.org/publications/
library/proceedings/jvm01/
JVM_wips/S06.pdf.

Ng:2002:OHP

[NSH^{+}02] Wee Teck Ng, Hao Sun, Bruce Hillyer, Elizabeth Shriver, Eran Gabber, and Banu Ozden. Obtaining high performance for storage outsourcing. In USENIX [USE02c], page ?? ISBN 1-880446-03-

0. LCCN QA76.9.F5 C65

2002. URL http://www.
usenix.org/publications/
library/proceedings/fast02/
ng.html.

Naik:2004:ARS

[NSK04] Vijay K. Naik, Swaminathan Sivasubramanian, and Sriram Krishnan. Adaptive resource sharing in a Web services environment. In Jacobsen [Jac04], pages 311– ?? CODEN LNCSD9. ISBN 3-540-23428-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.M54 I33 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.springerlink.
com/openurl.asp?genre=issue&
issn=0302-9743&volume=3231;

                                         http://www.springerlink.
                                        com/openurl.asp?genre=volume&
                                        id=doi:10.1007/b101561.

Ng:2001:WSA

[NSZ01] T. S. Eugene Ng, Ion Stoica, and Hui Zhang. A waypoint service approach to connect heterogeneous Internet address spaces. In USENIX [USE01b], page ?? ISBN 1-880446-09-X. LCCN QA76.8.U65 U84 2001. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/usenix01/
                                        ng.html.

Nieh:2001:VTR

[NVZ01] Jason Nieh, Chris Vaill, and Hua Zhong. Virtualtime round-robin: An O(1) proportional share scheduler. In USENIX [USE01b], page ?? ISBN 1-880446-09-X. LCCN QA76.8.U65

                                        U84 2001. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/usenix01/
                                        nieh.html.

Nicholson:2004:COP

[NW04] John Nicholson and Rachel Wheeler. California Online Privacy Act has widespread effects. ;login: the USENIX Association newsletter, 29(4): ??, August 2004. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2004-08/pdfs/nicholson.
                                        pdf.

Neumann:2000:XOO

[NZ00] Gustav Neumann and Uwe Zdun. XOTcl — an object-oriented scripting language. In USENIX [USE00e], page ?? ISBN 1-880446-24-3. LCCN ????

URL http://db.usenix.
org/publications/library/
proceedings/tcl2k/neumann.
html.

Oberle:2004:DMS

[OESV04] Daniel Oberle, Andreas Eberhart, Steffen Staab, and Raphael Volz. Developing and managing software components in an ontology-based application server. In Jacobsen [Jac04], pages 459– ?? CODEN LNCSD9. ISBN 3-540-23428-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.M54 I33 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.springerlink.
com/openurl.asp?genre=issue&
issn=0302-9743&volume=3231;
 http://www.springerlink.
com/openurl.asp?genre=volume&
id=doi:10.1007/b101561.

Ong:2000:PCL

[OF00] Hong Ong and Paul A. Farrell. Performance comparison of LAM/MPI, MPICH, and MVICH on a Linux cluster connected by a Gigabit Ethernet network. In USENIX [USE00c], page ??

ISBN 1-880446-17-0. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/als2000/
                                        ong.html.

ORourke:2001:PEL

[OK01] Patrick O’Rourke and Mike Keefe. Performance evaluation of Linux Virtual Server. In USENIX [USE01a], page ?? ISBN 1-880446-05-7. LCCN ????

                                        URL http://db.usenix.
                                        org/publications/library/
                                        proceedings/lisa2001/tech/
                                        orourke.html.

Okita:2007:DI

[Oki07] Cat Okita. (digital) identity 2.0. ;login: the USENIX Association newsletter, 32(5): ??, October 2007. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        october-2007-volume-32-
                                        number-5/digital-identity-
                                        20.

Ogasawara:2002:OPO

[OKN02] Takeshi Ogasawara, Hideaki Komatsu, and Toshio Nakatani. Optimizing precision overhead for x86 processors. In USENIX [USE02f], page ?? ISBN 1-931971-01-3. LCCN QA76.73 .J38 J42 2002.

                                        URL http://www.usenix.
                                        org/publications/library/
                                        proceedings/javavm02/ogasawara.
                                        html.

Oliver:2001:CSF

[Oli01] Ross Oliver. Countering SYN flood denial-of-service (DoS) attacks, 2001. URL http:

//www.usenix.org/events/
sec01/invitedtalks/oliver.

pdf. Unpublished invited talk, Tenth USENIX Security Symposium, August 13–17, 2001, Washington, DC, USA.

Oppenheim:2000:DTP

[OM00] Kyle Oppenheim and Patrick McCormick. Deployme: Tellme’s package management and deployment system. In USENIX [USE00g], page ?? ISBN 1-880446-13-8. LCCN ????

URL http://www.usenix.
org/publications/library/
proceedings/lisa2000/oppenheim.
html.

Ott:2002:MTF

[OMP02] David Ott and Ketan Mayer-Patel. A mechanism for TCP-friendly transport-level protocol coordination. In USENIX [USE02e], page ?? ISBN 1-880446-00-6. LCCN QA76.8.U65 U84 2002. URL

http://www.usenix.org/publications/
library/proceedings/usenix02/
ott.html.

Opyrchal:2001:SDE

[OP01] Lukasz Opyrchal and Atul Prakash. Secure distribution of events in contentbased publish subscribe systems. In USENIX [USE01h], page ?? ISBN 1-880446-18-9, 1-880446-07-3. LCCN QA76.9.A25 U565

2001. URL http://www.
usenix.org/publications/
library/proceedings/sec01/
opyrchal.html.

Orgeron:2007:ILDa

[Org07a] Octave Orgeron. An introduction to logical domains, part 1. ;login: the USENIX Association newsletter, 32(4): ??, August 2007. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2007-volume-32-number-
4/introduction-logical-
domains-part-1.

Orgeron:2007:ILDb

[Org07b] Octave Orgeron. An introduction to logical domains, part 2. ;login: the USENIX Association newsletter, 32(5): ??, October 2007. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
october-2007-volume-32-
number-5/introduction-logical-
domains-part-2.

Orgeron:2008:ILDa

[Org08a] Octave Orgeron. An introduction to logical domains. ;login: the USENIX Association newsletter, 33(1):??, February 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/

february-2008-volume-33-
number-1/introduction-logical-
domains.

Orgeron:2008:ILDb

[Org08b] Octave Orgeron. An introduction to logical domains, part 4. ;login: the USE-NIX Association newsletter, 33(2):??, April 2008. CO-DEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/april-2008-volume-
33-number-2/introduction-
logical-domains-part-4.

Orman:2000:ANF

[Orm00a]  Hilarie Orman.  Active

networking fast path to QoS. In USENIX [USE00j], page ?? ISBN ???? LCCN

????  URL http://db.
usenix.org/publications/
library/proceedings/ine2000/
orman.p_slides.html;   http:
//db.usenix.org/publications/
library/proceedings/ine2000/
orman.p_slides.ppt.

Orman:2000:TPI

[Orm00b]  Hilarie Orman.  Towards

a platform for intelligent activity at the edge. In USENIX [USE00j], page ?? ISBN ???? LCCN

????  URL http://db.
usenix.org/publications/
library/proceedings/ine2000/
orman.html.

Ozment:2006:SOM

[OS06] Andy Ozment and Stuart E.

Schechter. The security of OpenBSD: Milk or wine? ;login: the USENIX Association newsletter, 31(6):??, December 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
december-2006-volume-31-
number-6/security-openbsd-
milk-or-wine.

Oleson:2000:OIS

[OSE^{+}00] Van Oleson, Karsten Schwan, Greg Eisenhouer, Beth Plale, Calton Pu, and Dick Amin. Operational information systems: An example from the airline industry. In USENIX [USE00f], page ?? ISBN 1-880446-15-4. LCCN QA76.76.S95 W67 2000.

URL http://www.usenix.
org/publications/library/
proceedings/osdi2000/wiess2000/
oleson.html.

Osterlund:2000:PPI

[Ost00] Robert Osterlund. PIKT: Problem Informant/Killer Tool. In USENIX [USE00g], page ?? ISBN 1-880446-13-8. LCCN ????

URL http://www.usenix.
org/publications/library/
proceedings/lisa2000/osterlund.
html.

Otto:2000:EMW

[Ott00] J. Nick Otto. Enterprise management of Windows NT services. In USE-NIX [USE00b], page ?? ISBN 1-880446-19-7. LCCN

????  URL http://db.

usenix.org/publications/
library/proceedings/lisa-
nt2000/otto.html.

Orr:2000:SPF

[OW00] Giles Orr and Jacob Wyatt. SSH port forwarding. In USENIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/als2000/
orr.html.

Owen:2002:PP

[Owe02] Howard Owen. The problem of PORCMOLSULB. ;login: the USENIX Association newsletter, 27(4): ??, August 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-08/pdfs/owen.pdf.

Plattner:2004:GSR

[PA04] Christian Plattner and Gustavo Alonso. Ganymed: Scalable replication for transactional Web applications. In Jacobsen [Jac04], pages 155–?? CODEN LNCSD9. ISBN 3-540-23428-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.M54 I33 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.springerlink.
com/openurl.asp?genre=issue&
issn=0302-9743&volume=3231;
 http://www.springerlink.

                                        com/openurl.asp?genre=volume&
                                        id=doi:10.1007/b101561.

Packard:2000:TWX

[Pac00]  Keith Packard.  Translu-

cent windows in X. In USENIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/als2000/
                                        packard.html.

Packard:2001:DIX

[Pac01a] Keith Packard. Design and implementation of the X Rendering Extension. In USENIX [USE01f], page ?? ISBN 1-880446-10-3. LCCN QA76.8.U65 U84 2001. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/usenix01/
                                        freenix01/packard.html.

Packard:2001:XFL

[Pac01b] Keith Packard. The Xft font library: Architecture and users guide. In USE-NIX [USE01d], page ?? ISBN ???? LCCN ???? URL http:

                                        //www.linuxshowcase.org/
                                        packard.html.

Palmer:2005:UNS

[Pal05]  Chris Palmer.  USENIX

notes: SAGE update. ;login: the USENIX Association newsletter, 30(5):??, October 2005. CODEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/

login/2005-10/openpdfs/
sage.pdf.

Parter:2000:SE

[Par00]  David Parter.  Sysadmin

ethics. ;login: the USENIX Association newsletter, 25(6): ??, October 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-10/pdfs/sysadminethics.
pdf.

Parseghian:2002:JBT

[Par02a] Pat Parseghian. The joy of breaking things. In USENIX [USE02e], page

355. ISBN 1-880446-00-

6. LCCN QA76.8.U65 U84

2002. URL http://www.
usenix.org/publications/
library/proceedings/usenix02/
tech/techonefile.html. Un-

published invited talk, 2002 USENIX Annual Technical Conference, June 10-15, 2002, Monterey Conference Center, Monterey, CA.

Parter:2002:SMG

[Par02b]  David Parter.  Sysadmin

management/general. In USENIX [USE02e], page

355. ISBN 1-880446-00-

6. LCCN QA76.8.U65 U84

2002. URL http://www.
usenix.org/publications/
library/proceedings/usenix02/
tech/techonefile.html. Un-

published guru session talk, 2002 USENIX Annual Technical Conference, June 10-15, 2002, Monterey Conference Center, Monterey, CA.

Parter:2005:UNS

[Par05]  David Parter.  USENIX

notes: SAGE update. ;login: the USENIX Association newsletter, 30(4):??, August 2005. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2005-08/openpdfs/parter.
                                        pdf.

Patterson:2002:AMP

[Pat02a]  David Patterson.  Avail-

ability and maintainability >> performance: New focus for a new century. In USENIX [USE02c], page ?? ISBN 1-880446-03-0. LCCN QA76.9.F5 C65 2002. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/fast02/
                                        patterson/sld001.htm.

Patterson:2002:ID

[Pat02b] David A. Patterson. An introduction to dependability. ;login: the USENIX Association newsletter, 27(4): ??, August 2002. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2002-08/pdfs/patterson.
                                        pdf.

Preslan:2000:SFR

[PBB^{+}00] Kenneth Preslan, Andrew Barry, Jonathan Brassow, Michael Declerck, A. J.

Lewis, Adam Manthei, Ben Marzinski, Erling Nygaard, Seth Van Oort, David Teigland, Mike Tilstra, Steve Whitehouse, and Matthew O’Keefe. Scalability and failure recovery in a Linux cluster file system. In USENIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/als2000/
preslan.html.

Peterson:2002:FCM

[PBP02] David S. Peterson, Matt Bishop, and Raju Pandey. A flexible containment mechanism for executing untrusted code. In USENIX [USE02b], pages 207–225. ISBN 1-931971-00-5. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/sec02/
peterson.html.

Peisert:2005:FSA

[Pei05]  Sean Peisert.  Forensics

for system administrators. ;login: the USENIX Association newsletter, 30(4): ??, August 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-08/pdfs/peisert.pdf.

Pelletier:2001:Z

[Pel01]  Michel Pelletier.  Zope,

2001.  URL http://db.
usenix.org/publications/

                                        library/proceedings/lisa2001/

tech/. Unpublished invited talk, LISA 2001: 15th Systems Administration Conference, December 2–7, 2001, Town and Country Resort Hotel, San Diego, CA.

Pennock:2004:ISC

[Pen04] Phil Pennock. On IMAP service for customers. ;login: the USENIX Association newsletter, 29(5):??, October 2004. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2004-10/pdfs/pennock.pdf.

Perrine:2001:CMS

[Per01] Tom Perrine. Cops are from Mars, sysadmins are from Pluto: Dealing with law enforcement, 2001. URL http:

                                        //www.usenix.org/events/
                                        sec01/invitedtalks/perrine.

pdf. Unpublished invited talk, Tenth USENIX Security Symposium, August 13–17, 2001, Washington, DC, USA.

Perrine:2002:KSO

[Per02] Tom Perrine. The kernelized secure operating system (KSOS). ;login: the USENIX Association newsletter, 27(6): ??, December 2002. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2002-12/pdfs/perrine.pdf.

Perrine:2003:ECP

[Per03]  Tom Perrine.  The end

of crypt() passwords … please? ;login: the USENIX Association newsletter, 28(6): ??, December 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-12/pdfs/perrine.pdf.

Ponnekanti:2004:IAI

[PF04] Shankar R. Ponnekanti and Armando Fox. Interoperability among independently evolving Web services. In Jacobsen [Jac04], pages 331–?? CODEN LNCSD9. ISBN 3-540-23428-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.M54 I33 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.springerlink.
com/openurl.asp?genre=issue&
issn=0302-9743&volume=3231;
 http://www.springerlink.
com/openurl.asp?genre=volume&
id=doi:10.1007/b101561.

P:2001:DIF

[PG01] Venkatesh P. and K. Gopinath. The design, implementation, and framework for a Linux-based temperature-sensitive storage system. In USENIX [USE01d], page ?? ISBN ???? LCCN ???? URL http:

//www.linuxshowcase.org/
venkatesh.html.

Palekar:2001:DIL

[PGCR01] Ashish Palekar, Narendran Ganapathym, Anshul Chadda, and Robert D. Russel. Design and implementation of a Linux SCSI target for storage area networks. In USE-NIX [USE01d], page ?? ISBN ???? LCCN ???? URL http:

                                        //www.linuxshowcase.org/
                                        palekar.html.

Provos:2001:SSI

[PH01] Niels Provos and Peter Honeyman. ScanSSH: Scanning the Internet for SSH servers. In USENIX [USE01a], page ?? ISBN 1-880446-05-7. LCCN

                                        ????  URL http://db.
                                        usenix.org/publications/
                                        library/proceedings/lisa2001/
                                        tech/provos.html.

Phillips:2000:TFF

[Phi00] Daniel Phillips. The Tux2 failsafe filesystem, 2000.

                                        URL http://www.usenix.
                                        org/publications/library/
                                        proceedings/als2000/. Un-

published talk at Annual

Linux Showcase 2000.

Phillips:2001:DIE

[Phi01] Daniel Phillips. A directory index for EXT2. In USE-NIX [USE01d], page ?? ISBN ???? LCCN ???? URL http:

                                        //www.linuxshowcase.org/
                                        phillips.html.

Piatek:2007:BBM

[PIA^{+}07] Michael Piatek, Tomas Isdal, Tom Anderson, Arvind Krishnamurthy, and Arun Venkataramani. Building BitTyrant, a (more) strategic BitTorrent client. ;login: the USENIX Association newsletter, 32(4):??, August 2007. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2007-volume-32-number-
4/building-bittyrant-more-
strategic-bittorrent.

Piele:2000:BOI

[Pie00] Don Piele. Balkan Olympiad in Informatics. ;login: the USENIX Association newsletter, 25(4):??, July 2000. CODEN LOGNEM. ISSN 1044-6397. URL http://

www.usenix.org/publications/
login/2000-7/usenixnews.
html#balkan.

Pierzchala:2003:CWO

[Pie03] Stephen Pierzchala. Compressing Web output using mod gzip and Apache. ;login: the USENIX Association newsletter, 28(2): ??, April 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-04/pdfs/pierzchala.
pdf.

Pike:2000:LFNb

[Pik00a] Rob Pike. Lexical file names in Plan 9, or getting dot-dot right. In USENIX [USE00a], pages 85–92. ISBN 1-880446-22-7. LCCN ???? URL http:

                                        //www.usenix.org/events/
                                        usenix2000/general/pikelex.
                                        html.

Pike:2000:POU

[Pik00b] Rob Pike. Plumbing and other utilities. In USE-NIX [USE00a], pages 159–

170. ISBN 1-880446-22-7. LCCN ???? URL http:

                                        //www.usenix.org/events/
                                        usenix2000/general/pikeplumb.
                                        html.

Piscitello:2008:CFR

[Pis08] David Piscitello. Are commercial firewalls ready for IP Version 6? ;login: the USE-NIX Association newsletter, 33(2):??, April 2008. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL https://www.
                                        usenix.org/publications/
                                        login/april-2008-volume-
                                        33-number-2/are-commercial-
                                        firewalls-ready-ip-version-
                                        6.

Plonka:2000:FNT

[Plo00] Dave Plonka. FlowScan: a network traffic flow reporting and visualization tool. In USENIX [USE00g], page ?? ISBN 1-880446-13-8. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/lisa2000/
                                        plonka.html.

Provos:2000:AOB

[PLT00] Niels Provos, Chuck Lever, and Stephen Tweedie. Analyzing the overload behavior of a simple Web server. In USENIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/als2000/
provos.html.

Patarin:2000:PFN

[PM00] Simon Patarin and Mesaac Makpangou. Pandora: a flexible network monitoring platform. In USENIX [USE00a], pages 27–40. ISBN 1-880446-22-7. LCCN ???? URL http:

//www.usenix.org/events/
usenix2000/general/patarin.
html.

Patterson:2002:SFS

[PMF^{+}02] R. Hugo Patterson, Stephen Manley, Mike Federwisch, Dave Hitz, Steve Kleiman, and Shane Owara. SnapMirror: File-system-based asynchronous mirroring for disaster recovery. In USE-NIX [USE02c], page ?? ISBN 1-880446-03-0. LCCN QA76.9.F5 C65 2002. URL

http://www.usenix.org/publications/
library/proceedings/fast02/
patterson.html.

Prasad:2004:SMT

[PMN^{+}04] Sushil K. Prasad, Vijay Madisetti, Shamkant B. Navathe, Raj Sunderraman, Erdogan Dogdu, Anu Bourgeois, Michael Weeks, Bing Liu, Janaka Balasooriya, Arthi Hariharan, Wanxia Xie, Praveen Madiraju, Srilaxmi Malladi, Raghupathy Sivakumar, Alex Zelikovsky, Yanqing Zhang, Yi Pan, and Saied Belkasim. SyD: a middleware testbed for collaborative applications over small heterogeneous devices and data stores. In Jacobsen [Jac04], pages 352–?? CODEN LNCSD9. ISBN 3-540-23428-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.M54 I33 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.springerlink.
com/openurl.asp?genre=issue&
issn=0302-9743&volume=3231;
 http://www.springerlink.
com/openurl.asp?genre=volume&
id=doi:10.1007/b101561.

Potter:2005:BTB

[PN05] Shaya Potter and Jason Nieh. Breaking the ties that bind: Application isolation and migration. ;login: the USENIX Association newsletter, 30(6): ??, December 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-12/pdfs/potter.pdf.

Powell:2001:FL

[Pow01] Brad Powell. Forensics lite. ;login: the USENIX Association newsletter, 26(7):??, November 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.

org/publications/login/
2001-11/pdfs/powell.pdf.

Park:2006:CSL

[PP06] KyoungSoo Park and Vivek S.

Pai. CoBlitz: a scalable large-file transfer service. ;login: the USENIX Association newsletter, 31(4): ??, August 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2006-volume-31-number-
4/coblitz-scalable-large-
file-transfer-service.

Prevelakis:2005:VF

[Pre05]  Vassilis Prevelakis.  The

virtual firewall. ;login: the USENIX Association newsletter, 30(6):??, December 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-12/pdfs/prevelakis.
pdf.

Prevelakis:2006:LE

[Pre06]  Vassilis Prevelakis.  Let-

ter to the Editor. ;login: the USENIX Association newsletter, 31(4):??, August 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2006-volume-31-number-
4/letter-editor.

Prevelakis:2007:SSL

[Pre07]  Vassilis Prevelakis.  Sup-

porting a security laboratory. ;login: the USENIX Association newsletter, 32 (3):??, June 2007. CO-DEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/june-2007-volume-
32-number-3/supporting-
security-laboratory.

Printezis:2001:HSB

[Pri01] Tony Printezis. Hot-swapping between a Mark&Sweep and a Mark&Compact garbage collector in a generational environment. In USE-NIX [USE01g], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38 J42 2001. URL

http://www.usenix.org/publications/
library/proceedings/jvm01/
printezis.html.

Provos:2000:EVM

[Pro00]  Niels Provos.  Encrypt-

ing virtual machine. In USENIX [USE00i], page ?? ISBN 1-880446-18-9. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/sec2000/
provos.html.

Provos:2001:DAS

[Pro01]  Niels Provos.  Defend-

ing against statistical steganalysis. In USENIX [USE01h], page ?? ISBN 1-880446-18-9, 1-880446-07-

3. LCCN QA76.9.A25 U565

2001. URL http://www.
usenix.org/publications/

library/proceedings/sec01/
provos.html.

Peterson:2006:PAP

[PRS06] Ryan Peterson, Venugopalan Ramasubramanian, and Emin Gün Sirer. A practical approach to peer-to-peer publishsubscribe. ;login: the USE-NIX Association newsletter, 31(4):??, August 2006. CO-DEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/august-2006-volume-
31-number-4/practical-approach-
peer-peer-publish-subscribe.

Prevelakis:2001:SA

[PS01] Vassilis Prevelakis and Diomidis Spinellis. Sandboxing applications. In USE-NIX [USE01f], page ?? ISBN 1-880446-10-3. LCCN QA76.8.U65 U84 2001. URL

http://www.usenix.org/publications/
library/proceedings/usenix01/
freenix01/prevelakis.html.

Pendarakis:2001:AAL

[PSVW01] Dimitrios Pendarakis, Sherlia Shi, Dinesh Verma, and Marcel Waldvogel. ALMI: An application level multicast infrastructure. In USENIX [USE01c], page ?? ISBN 1-880446-14-6. LCCN TK5105.875.I57 U96 2001.

URL http://www.usenix.
org/publications/library/
proceedings/usits01/shi.
html.

Perry:2001:ELT

[PSW^{+}01] Alexander Perry, John Sturtz, David Walsh, Vista Clara, and Brian Whitecotton. Embedding Linux to track concealed weapons. In USENIX [USE01d], page ?? ISBN ???? LCCN ???? URL http:

                                        //www.linuxshowcase.org/
                                        perry.html.

Puryear:2002:BRD

[Pur02] Dustin Puryear. Book review: Dave Sill: The qmail Handbook. ;login: the USENIX Association newsletter, 27(6): ??, December 2002. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2002-12/openpdfs/
                                        books.pdf.

Puryear:2006:CFP

[Pur06] Dustin Puryear. Consulting for fun and profit. ;login: the USENIX Association newsletter, 31(1):??, February 2006. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2006-02/pdfs/puryear.pdf.

Paleczny:2001:JHS

[PVC01] Michael Paleczny, Christopher Vick, and Cliff Click. The Java HotSpot^{T M} server compiler. In USENIX [USE01g], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38

                                        J42 2001. URL http://www.

usenix.org/publications/
library/proceedings/jvm01/
paleczny.html.

Poynor:2000:APS

[PW00] Todd Poynor and Tom Wylegala. Application programming on a shared memory multicomputer. In USENIX [USE00f], page ?? ISBN 1-880446-15-4. LCCN QA76.76.S95 W67 2000.

URL http://www.usenix.
org/publications/library/
proceedings/osdi2000/wiess2000/
poynor.html.

Quinlan:2002:VNA

[QD02] Sean Quinlan and Sean Dorward. Venti: a new approach to archival data storage. In USENIX [USE02c], page ?? ISBN 1-880446-03-

0. LCCN QA76.9.F5 C65

2002. URL http://www.
usenix.org/publications/
library/proceedings/fast02/
quinlan.html.

Quinn:2007:LLL

[Qui07] Brendan Quinn. Lessons learned from living with LDAP. ;login: the USE-NIX Association newsletter, 32(2):??, April 2007. CO-DEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/april-2007-volume-
32-number-2/lessons-learned-
living-ldap.

Ranganathan:2004:MML

[RAMC^{+}04] Anand Ranganathan, Jalal Al-Muhtadi, Shiva Chetan, Roy Campbell, and M. Dennis Mickunas. Middle-Where: a middleware for location awareness in ubiquitous computing applications. In Jacobsen [Jac04], pages 397–?? CODEN LNCSD9. ISBN 3-540-23428-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.M54 I33 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.springerlink.
com/openurl.asp?genre=issue&
issn=0302-9743&volume=3231;
 http://www.springerlink.
com/openurl.asp?genre=volume&
id=doi:10.1007/b101561.

Ranum:2000:NPBa

[Ran00a] Marcus J. Ranum. The network police blotter. ;login: the USENIX Association newsletter, 25(1):??, February 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-2/features/police.
html.

Ranum:2000:NPBb

[Ran00b] Marcus J. Ranum. The network police blotter. ;login: the USENIX Association newsletter, 25(2): ??, April 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-4/features/network.
html.

Ranum:2000:NPBc

[Ran00c] Marcus J. Ranum. The network police blotter. ;login: the USENIX Association newsletter, 25(3): ??, June 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-6/features/police.
html.

Ranum:2000:NPBd

[Ran00d] Marcus J. Ranum. The network police blotter. ;login: the USENIX Association newsletter, 25(5):??, August 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-8/features/blotter.
html.

Ranum:2000:NPBe

[Ran00e] Marcus J. Ranum. The network police blotter. ;login: the USENIX Association newsletter, 25(6):??, October 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-10/pdfs/networkpoliceblotter.
pdf.

Ranum:2000:NPBf

[Ran00f] Marcus J. Ranum. The network police blotter. ;login: the USENIX Association newsletter, 25(8):??, December 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-12/pdfs/ranum.pdf.

Raner:2002:LJV

[Ran02] Mirko Raner. A lightweight Java Virtual Machine for a stack-based microprocessor. In USENIX [USE02f], page ?? ISBN 1-931971-01-3. LCCN QA76.73 .J38

                                        J42 2002. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/javavm02/
                                        raner.html.

Ranum:2005:WNE

[Ran05] Marcus Ranum. Who needs an enemy when you can divide and conquer yourself? ;login: the USENIX Association newsletter, 30(1):??, February 2005. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2005-02/pdfs/ranum.pdf.

Rao:2000:UJa

[Rao00a] Prithvi Rao. Using Java. ;login: the USENIX Association newsletter, 25(1): ??, February 2000. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        java/usingjava20.html.

Rao:2000:UJb

[Rao00b] Prithvi Rao. Using Java.

;login: the USENIX Association newsletter, 25(2): ??, April 2000. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
java/usingjava21.html.

Rao:2000:UJc

[Rao00c] Prithvi Rao. Using Java.

;login: the USENIX Association newsletter, 25(3): ??, June 2000. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
java/usingjava22.html.

Rao:2000:UJd

[Rao00d] Prithvi Rao. Using Java.

;login: the USENIX Association newsletter, 25(5): ??, August 2000. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
java/usingjava22.html.

Rao:2000:UJf

[Rao00e] Prithvi Rao. Using Java.

;login: the USENIX Association newsletter, 25(6): ??, October 2000. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
java/usingjava24.pdf.

Rao:2000:UJg

[Rao00f] Prithvi Rao. Using Java.

;login: the USENIX Association newsletter, 25(8): ??, December 2000. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
java/usingjava25.pdf.

Rao:2001:UCJa

[Rao01a] Prithvi Rao. Using CORBA with Java. ;login: the USE-NIX Association newsletter, 26(1):??, February 2001. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
java/usingjava26.pdf.

Rao:2001:UCJb

[Rao01b] Prithvi Rao. Using CORBA with Java. ;login: the USE-NIX Association newsletter, 26(3):??, June 2001. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
java/usingjava27.pdf.

Rash:2004:CPK

[Ras04] Michael Rash. Combining port knocking and passive OS fingerprinting with fwknop. ;login: the USENIX Association newsletter, 29(6):??, December 2004. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2004-12/pdfs/fwknop.pdf.

Rash:2006:SPA

[Ras06] Michael Rash. Single packet authorization with Fwknop.

;login: the USENIX Association newsletter, 31(1):??, February 2006. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2006-02/pdfs/rash.pdf.

Rash:2007:ISM

[Ras07] Mike Rash. IDS signature matching with iptables,

psad, and fwsnort. ;lo-

gin: the USENIX Association newsletter, 32(6):??, December 2007. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
december-2007-volume-32-
number-6/ids-signature-
matching-iptables-psad-
and-fwsnort.

Russell:2001:HSA

[RB01] Kenneth Russell and Lars Bak. The HotSpot^{T M} serviceability agent: An outof-process high-level debugger for a Java^{T M} Virtual Machine. In USENIX [USE01g], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38

J42 2001. URL http://www.
usenix.org/publications/
library/proceedings/jvm01/
russell.html.

Ranganathan:2000:AMT

[RBPM00] M. Ranganathan, Mark Bednarek, Fernand Pors, and Doug Montgomery. AGNI: a multi-threaded middleware for distributed scripting. In USENIX [USE00e], page ??

ISBN 1-880446-24-3. LCCN

                                        ????  URL http://db.
                                        usenix.org/publications/
                                        library/proceedings/tcl2k/
                                        ranganathan.html.

Rezmierski:2000:SIC

[RCH00] Virginia Rezmierski, Adriana Carroll, and Jamie Hine. A study on incident costs and frequencies. ;login: the USENIX Association newsletter, 25(5): ??, August 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-8/features/incident.
                                        html.

Rescorla:2002:SCS

[RCK02] Eric Rescorla, Adam Cain, and Brian Korver. SS-LACC: a clustered SSL accelerator. In USENIX [USE02b], pages 229–246. ISBN 1-931971-00-5. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/sec02/
                                        rescorla.html.

Rajamony:2001:MCP

[RE01] Ramakrishnan Rajamony and Mootaz Elnozahy. Measuring client-perceived response time on the WWW. In USENIX [USE01c], page ?? ISBN 1-880446-14-6. LCCN TK5105.875.I57 U96 2001.

                                        URL http://www.usenix.
                                        org/publications/library/
                                        proceedings/usits01/rajamony.
                                        html.

Readers:2004:MH

[Rea04] Our Readers. More haikus!

;login: the USENIX Association newsletter, 29(1):??, February 2004. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2004-02/openpdfs/haikus.
pdf.

Regehr:2002:ISB

[Reg02]  John Regehr.  Inferring

scheduling behavior with Hourglass. In USENIX [USE02d], page ?? ISBN 1-880446-01-4. LCCN QA76.8.U65 P765 2002. URL http://

www.usenix.org/publications/
library/proceedings/usenix02/
tech/freenix/regehr.html.

Renfro:2002:VCP

[Ren02] Scott G. Renfro. VeriSign CZAG: Privacy leak in X.509 certificates. In USENIX [USE02b], pages 139–148. ISBN 1-931971-00-5. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/sec02/
renfro.html.

Romig:2000:OFT

[RFL00] Steve Romig, Mark Fullmer, and Ron Luman. The OSU Flow-tools package and CISCO NetFlow logs. In USENIX [USE00g], page ?? ISBN 1-880446-13-8. LCCN

????  URL http://www.
usenix.org/publications/

library/proceedings/lisa2000/
fullmer.html.

Reid:2000:KCC

[RFS^{+}00] Alastair Reid, Matthew Flatt, Leigh Stoller, Jay Lepreau, and Eric Eide. Knit: Component composition for systems software. In USENIX [USE00h], page ?? ISBN 1-880446-16-2. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/osdi2000/
reid.html.

Rangarajan:2000:SDS

[RI00a] Muralidharan Rangarajan and Liviu Iftode. Software distributed shared memory over virtual interface architecture: Implementation and performance. In USENIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/als2000/
rangarajan.html.

Robin:2000:AIP

[RI00b] John Scott Robin and Cynthia E. Irvine. Analysis of the Intel Pentium’s ability to support a secure virtual machine monitor. In USENIX [USE00i], page ?? ISBN 1-880446-18-9. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/sec2000/
robin.html.

Roman:2004:DPR

[RI04] Manuel Roman and Nayeem Islam. Dynamically programmable and reconfigurable middleware services. In Jacobsen [Jac04], pages 372–?? CODEN LNCSD9. ISBN 3-540-23428-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.M54 I33 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.springerlink.
com/openurl.asp?genre=issue&
issn=0302-9743&volume=3231;
http://www.springerlink.
com/openurl.asp?genre=volume&
id=doi:10.1007/b101561.

Rightnour:2002:SEH

[Rig02]  Tim Rightnour.  Sushi:

An extensible human interface for NetBSD. In USE-NIX [USE02a], pages 143–

151. ISBN 1-880446-02-2. LCCN QA76.76.O63 B736

2002. URL http://www.
usenix.org/publications/
library/proceedings/bsdcon02/
rightnour.html.

Regan:2001:CFN

[RJ01] Jude T. Regan and Christian D. Jensen. Capability file names: Separating authorisation from user management in an Internet file system. In USENIX [USE01h], page ?? ISBN 1-880446-18-9, 1-880446-07-

3. LCCN QA76.9.A25 U565

2001. URL http://www.
usenix.org/publications/
library/proceedings/sec01/
regan.html.

Riedel:2002:FES

[RKS02] Erik Riedel, Mahesh Kallahalla, and Ram Swaminathan. A framework for evaluating storage system security. In USENIX [USE02c], page ?? ISBN 1-880446-03-

0. LCCN QA76.9.F5 C65

2002. URL http://www.
usenix.org/publications/
library/proceedings/fast02/
riedel.html.

Roselli:2000:CFS

[RLA00] Drew Roselli, Jacob R. Lorch, and Thomas E. Anderson. A comparison of file system workloads. In USE-NIX [USE00a], pages 41–

54. ISBN 1-880446-22-7. LCCN ???? URL http:

//www.usenix.org/events/
usenix2000/general/roselli.
html.

Reisman:2002:IAT

[RMS02] Ron Reisman, James Murphy, and Rob Savoye. Introduction to air traffic management systems. In USE-NIX [USE02e], page ?? ISBN 1-880446-00-6. LCCN QA76.8.U65 U84 2002. URL

http://www.usenix.org/publications/
library/proceedings/usenix02/
tech/techonefile.html. Un-

published invited talk, 2002 USENIX Annual Technical Conference, June 10-15, 2002, Monterey Conference Center, Monterey, CA.

Reumann:2000:VSN

[RMSK00] John Reumann, Ashish Mehra, Kang G. Shin, and Dilip Kandlur. Virtual services: a new abstraction for server consolidation. In USE-NIX [USE00a], pages 117– 130. ISBN 1-880446-22-7. LCCN ???? URL http:

//www.usenix.org/events/
usenix2000/general/reumann.
html.

Robertson:2000:LHS

[Rob00] Alan Robertson. Linux-HA heartbeat system design. In USENIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/als2000/
robertson.html.

Romig:2000:CLF

[Rom00]  Steve Romig.  Correlat-

ing log file entries. ;login: the USENIX Association newsletter, 25(7):??, November 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-11/pdfs/log.pdf.

Roseman:2000:PTJ

[Ros00] Mark Roseman. Proxy Tk:

a Java applet user interface toolkit for Tcl. In USENIX [USE00e], page ?? ISBN 1-880446-24-3. LCCN

????  URL http://db.
usenix.org/publications/
library/proceedings/tcl2k/
roseman.html.

Rose:2001:WAC

[Ros01] Greg Rose. Where has all the crypto gone? Long time coming: a speculative and historical talk. In USENIX [USE01a], page ?? ISBN 1-880446-05-7. LCCN

????  URL http://db.
usenix.org/publications/
library/proceedings/lisa2001/
tech/rose/index.htm.

Roth:2001:DRS

[Rot01] Mark Roth. Defining the role of service manager: Sanity through organizational evolution. In USE-NIX [USE01a], page ?? ISBN 1-880446-05-7. LCCN

????  URL http://db.
usenix.org/publications/
library/proceedings/lisa2001/
tech/roth.html.

Rao:2000:CPR

[RR00] Josyula R. Rao and Pankaj Rohatgi. Can pseudonymity really guarantee privacy? In USENIX [USE00i], page ?? ISBN 1-880446-18-9. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/sec2000/
rao.html.

Rodgers:2001:MSN

[RS01] R. P. C. Rodgers and Ziying Sherwin. A management system for networksharable locally installed software: Merging RPM and the Depot scheme under Solaris. In USENIX [USE01a], pages 267–272. ISBN 1-880446-05-7. LCCN

????  URL http://db.
usenix.org/publications/
library/proceedings/lisa2001/
tech/rodgers.html.

Rubenstein:2001:WC

[Rub01a] Jennifer Rubenstein. Women in computing. ;login: the USENIX Association newsletter, 26(3):??, June 2001. CODEN LOGNEM. ISSN 1044-6397. URL http://

www.usenix.org/publications/
login/2001-06/pdfs/rubenstein.
pdf.

Rubin:2001:SCR

[Rub01b] Avi Rubin. Security considerations for remote electronic voting over the Internet. ;login: the USENIX Association newsletter, 26(1): ??, February 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-02/pdfs/rubin.pdf.

Rubin:2001:SVP

[Rub01c] Avi Rubin. Security for Evoting in public elections,

2001.  URL http://db.
usenix.org/publications/
library/proceedings/lisa2001/

tech/. Unpublished invited talk, LISA 2001: 15th Systems Administration Conference, December 2–7, 2001, Town and Country Resort Hotel, San Diego, CA.

Russo:2000:KEK

[Rus00a] Christopher M. Russo. Keeping employees by keeping them happy. ;login: the USENIX Association newsletter, 25(5):??, August 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-8/features/keeping.
                                        html.

Russo:2000:MAS

[Rus00b] Christopher M. Russo. The magic, art, and science of customer support. ;login: the USENIX Association newsletter, 25(2): ??, April 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-4/features/customersupport.
                                        html.

Russo:2000:RW

[Rus00c] Christopher M. Russo. Resumé writing. ;login: the USENIX Association newsletter, 25(4):??, July 2000. CODEN LOGNEM. ISSN 1044-6397. URL http://

                                        www.usenix.org/publications/
                                        login/2000-7/features/resume.
                                        html.

Russo:2002:WYP

[Rus02] Christopher M. Russo. What is your problem? That email was fine! ;login: the USENIX Association newsletter, 27(1): ??, February 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-02/pdfs/russo.pdf.

Russell:2003:HWG

[Rus03] Ryan Russell. How are we going to patch all these boxes? ;login: the USENIX Association newsletter, 28(6):??, December 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-12/pdfs/russell.pdf.

Russo:2004:KEKa

[Rus04a] Christopher M. Russo. Keeping employees by keeping them happy, Part II. ;login: the USENIX Association newsletter, 29(1):??, February 2004. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2004-02/pdfs/russo.pdf.

Russo:2004:KEKb

[Rus04b] Christopher M. Russo. Keeping employees by keeping them happy, Part III. ;login: the USENIX Association newsletter, 29(2): ??, April 2004. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2004-04/pdfs/Russo.pdf.

Ruwart:2000:LDS

[Ruw00] Thomas M. Ruwart. Linux disk subsystem I/O performance, 2000. URL http://

                                        www.usenix.org/publications/
                                        library/proceedings/als2000/

. Unpublished talk at Annual

Linux Showcase 2000.

Ressman:2000:UCA

[RV00] David Ressman and John Valdés. Use of Cfengine for automated, multi-platform software and patch distribution. In USENIX [USE00g], page ?? ISBN 1-880446-13-8. LCCN ????

                                        URL http://www.usenix.
                                        org/publications/library/
                                        proceedings/lisa2000/ressman.
                                        html.

Salus:2001:BRBg

[S^{+}01] Peter H. Salus et al. Book reviews: The bookworm. ;login: the USENIX Association newsletter, 26(7):??, November 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-11/openpdfs/bkrev.
                                        pdf.

Shapiro:2002:DEE

[SA02] Jonathan S. Shapiro and Jonathan Adams. Design evolution of the EROS single-level store. In USE-NIX [USE02e], page ?? ISBN 1-880446-00-6. LCCN QA76.8.U65 U84 2002. URL

http://www.usenix.org/publications/
library/proceedings/usenix02/
shapiro.html.

Sailer:2001:ITV

[SAB^{+}01a] Reiner Sailer, Arup Acharya, Mandis Beigi, Raymond Jennings, and Dinesh Verma. IPSECvalidate: a tool to validate IPSEC configurations. In USENIX [USE01a], page ?? ISBN 1-880446-05-7. LCCN ????

URL http://db.usenix.
org/publications/library/
proceedings/lisa2001/tech/
sailer.html.

Snoeren:2001:FGF

[SAB01b] Alex C. Snoeren, David G.

Andersen, and Hari Balakrishnan. Fine-grained failover using connection migration. In USENIX [USE01c], page ?? ISBN 1-880446-14-6. LCCN TK5105.875.I57

U96 2001. URL http://www.
usenix.org/publications/
library/proceedings/usits01/
snoeren.html.

Sacks:2008:LKR

[Sac08] Matthew Sacks. Linux kernel resource allocation in virtualized environments. ;login: the USENIX Association newsletter, 33(2): ??, April 2008. CODEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/april-2008-volume-
33-number-2/linux-kernel-
resource-allocation-virtualized.

Sahore:2008:RDS

[Sah08] Sandeep Sahore. Reclaim disk space by shrinking files. ;login: the USENIX Association newsletter, 33(5): ??, October 2008. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        october-2008-volume-33-
                                        number-5/reclaim-disk-space-
                                        shrinking-files.

Salus:2000:YAUb

[Sal00a] Peter Salus. 20 years ago in UNIX. ;login: the USE-NIX Association newsletter, 25(2):??, April 2000. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2000-4/20yearsago.
                                        html.

Salus:2000:YAUc

[Sal00b] Peter Salus. 20 years ago in UNIX. ;login: the USE-NIX Association newsletter, 25(3):??, June 2000. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2000-6/usenixnews.
                                        html#20.

Salus:2000:YAUa

[Sal00c]  Peter Salus.  20 years

ago in USENIX. ;login: the USENIX Association newsletter, 25(1):??, February 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-2/20yrsago.html.

Salus:2000:YA

[Sal00d] Peter Salus. N years ago.

;login: the USENIX Association newsletter, 25(5): ??, August 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-8/usenixnews.html#
salus.

Salus:2000:BRBa

[Sal00e] Peter H. Salus. Book review: The bookworm. ;login: the USENIX Association newsletter, 25(1):??, February 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-2/bookworm.html.

Salus:2000:BRBb

[Sal00f] Peter H. Salus. Book reviews: The bookworm. ;login: the USENIX Association newsletter, 25(2): ??, April 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-4/bookworm.html.

Salus:2000:BRBc

[Sal00g] Peter H. Salus. Book reviews: The bookworm. ;login: the USENIX Association newsletter, 25(3): ??, June 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-6/bookworm.html.

Salus:2000:BRBd

[Sal00h] Peter H. Salus. Book reviews: The bookworm. ;login: the USENIX Association newsletter, 25(5):??, August 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-8/bookworm.html.

Salus:2000:BRBe

[Sal00i] Peter H. Salus. Book reviews: The bookworm. ;login: the USENIX Association newsletter, 25(6):??, October 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-10/bookworm.pdf.

Salus:2000:BRBf

[Sal00j] Peter H. Salus. Book reviews: The bookworm. ;login: the USENIX Association newsletter, 25(7):??, November 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-11/bookworm.pdf.

Salus:2000:BRBg

[Sal00k] Peter H. Salus. Book reviews: The bookworm. ;login: the USENIX Association newsletter, 25(8):??, December 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-12/bookworm.pdf.

Salus:2001:CA

[Sal01a]  Peter Salus.  2001: a

communications anniversary,

2001.  URL http://db.
usenix.org/publications/
library/proceedings/lisa2001/

tech/. Unpublished invited talk, LISA 2001: 15th Systems Administration Conference, December 2–7, 2001, Town and Country Resort Hotel, San Diego, CA.

Salus:2001:CRC

[Sal01b]  Peter Salus.  Conference

report: COOTS 6. ;login: the USENIX Association newsletter, 26(2): ??, April 2001. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
library/proceedings/coots01/
saluscoots.pdf.

Salus:2001:BRBa

[Sal01c] Peter H. Salus. Book reviews: The bookworm. ;login: the USENIX Association newsletter, 26(1):??, February 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-02/bookworm.pdf.

Salus:2001:BRBb

[Sal01d] Peter H. Salus. Book reviews: The bookworm. ;login: the USENIX Association newsletter, 26(2): ??, April 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-04/bookworm.pdf.

Salus:2001:BRBc

[Sal01e] Peter H. Salus. Book reviews: The bookworm. ;login: the USENIX Association newsletter, 26(3): ??, June 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-06/bookworm.pdf.

Salus:2001:BRBd

[Sal01f] Peter H. Salus. Book reviews: The bookworm. ;login: the USENIX Association newsletter, 26(4): ??, July 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-07/bookworm.pdf.

Salus:2001:BRBe

[Sal01g] Peter H. Salus. Book reviews: The bookworm. ;login: the USENIX Association newsletter, 26(5):??, August 2001. CODEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/

login/2001-08/openpdfs/
bookworm.pdf.

Salus:2001:BRBf

[Sal01h] Peter H. Salus. Book reviews: The bookworm. ;login: the USENIX Association newsletter, 26(6):??, October 2001. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2001-10/openpdfs/
bookworm.pdf.

Salus:2001:BRBh

[Sal01i] Peter H. Salus. Book reviews: The bookworm. ;login: the USENIX Association newsletter, 26(8):??, December 2001. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2001-12/openpdfs/
bookworm.pdf.

Salus:2002:BRBb

[Sal02a] Peter H. Salus. Book review: The bookworm. ;login: the USENIX Association newsletter, 27(2): ??, April 2002. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2002-04/openpdfs/
book.pdf.

Salus:2002:BRBa

[Sal02b] Peter H. Salus. Book reviews: The bookworm. ;login: the USENIX Association newsletter, 27(1):??, February 2002. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2002-02/openpdfs/
bookreviews.pdf.

Salus:2002:BRBc

[Sal02c] Peter H. Salus. Book reviews: The bookworm. ;login: the USENIX Association newsletter, 27(3): ??, June 2002. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2002-06/openpdfs/
book.pdf.

Salus:2002:BRBd

[Sal02d] Peter H. Salus. Book reviews: The bookworm. ;login: the USENIX Association newsletter, 27(4):??, August 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-08/openpdfs/books.
pdf.

Salus:2002:BRBe

[Sal02e] Peter H. Salus. Book reviews: The bookworm. ;login: the USENIX Association newsletter, 27(5):??, October 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-10/openpdfs/books.
pdf.

Salus:2002:BRBf

[Sal02f] Peter H. Salus. Book reviews: The bookworm. ;login: the USENIX Association newsletter, 27(6):??, December 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-12/openpdfs/books.
pdf.

Salus:2002:SW

[Sal02g] Peter H. Salus. Secure from what? ;login: the USENIX Association newsletter, 27(6): ??, December 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-12/pdfs/salus.pdf.

Salus:2003:BRH

[Sal03a] Peter H. Salus. Book reviews and history: The bookworm: Networking anniversaries. ;login: the USE-NIX Association newsletter, 28(5):??, October 2003. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2003-10/openpdfs/
salus.pdf.

Salus:2003:BRBa

[Sal03b] Peter H. Salus. Book reviews: The bookworm. ;login: the USENIX Association newsletter, 28(1):??, February 2003. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2003-02/openpdfs/
bookworm.pdf.

Salus:2003:BRBe

[Sal03c] Peter H. Salus. Book reviews: The bookworm: Nearly 20 years ago in U[SE]NIX. ;login: the USENIX Association newsletter, 28(6):??, December 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-12/openpdfs/salus.
pdf.

Salus:2003:BRBb

[Sal03d] Peter H. Salus. Book reviews: The bookworm: Reviews of Pavol Cerven’s Crackproof Your Software and of Jenness and Cozens’s Extending and Embedding Perl. ;login: the USENIX Association newsletter, 28(2): ??, April 2003. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2003-04/openpdfs/
bookreviews.pdf.

Salus:2003:BRBd

[Sal03e] Peter H. Salus. Book reviews: The bookworm; twenty-five years ago in UNIX. ;login: the USENIX Association newsletter, 28(4):??, August 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.

org/publications/login/
2003-08/openpdfs/salus.
pdf.

Salus:2003:BRBc

[Sal03f]  Peter H. Salus.  Book

reviews: The bookworm: Twenty years ago in USE-NIX. ;login: the USE-NIX Association newsletter, 28(3):??, June 2003. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2003-06/openpdfs/
salus.pdf.

Salus:2004:SYA

[Sal04a] Peter Salus. STUG 20 years ago. ;login: the USENIX Association newsletter, 29 (2):??, April 2004. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2004-04/openpdfs/
History.pdf.

Salus:2004:BRI

[Sal04b] Peter H. Salus. Book review: Is it 10 years already? ;login: the USE-NIX Association newsletter, 29(1):??, February 2004. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2004-02/openpdfs/
history.pdf.

Salus:2004:BRBb

[Sal04c] Peter H. Salus. Book review: The bookworm. ;login: the USENIX Association newsletter, 29(2): ??, April 2004. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2004-04/openpdfs/Books.
pdf.

Salus:2004:BRBc

[Sal04d] Peter H. Salus. Book review: The bookworm. ;login: the USENIX Association newsletter, 29(3): ??, June 2004. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2004-06/openpdfs/
bookworm.pdf.

Salus:2004:BRBd

[Sal04e] Peter H. Salus. Book review: The bookworm. ;login: the USENIX Association newsletter, 29(5):??, October 2004. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2004-10/openpdfs/
bookworm.pdf.

Salus:2004:BRBe

[Sal04f] Peter H. Salus. Book review: The bookworm. ;login: the USENIX Association newsletter, 29(6):??, December 2004. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/

login/2004-12/openpdfs/
bookworm.pdf.

Salus:2004:BRUb

[Sal04g] Peter H. Salus. Book reviews and USENIX history: History and FUD. ;login: the USENIX Association newsletter, 29(4): ??, August 2004. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2004-08/openpdfs/
history.pdf.

Salus:2004:BRUa

[Sal04h] Peter H. Salus. Book reviews and USENIX history: The bookworm. ;login: the USE-NIX Association newsletter, 29(4):??, August 2004. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2004-08/openpdfs/
bookworm.pdf.

Salus:2004:BRUc

[Sal04i] Peter H. Salus. Book reviews and USENIX history: Twenty years ago … and more. ;login: the USENIX Association newsletter, 29(4): ??, August 2004. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2004-08/openpdfs/twenty.
pdf.

Salus:2004:BRBa

[Sal04j] Peter H. Salus. Book reviews: The bookworm. ;login: the USENIX Association newsletter, 29(1):??, February 2004. CODEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2004-02/openpdfs/
                                        bookworm.pdf.

Salus:2004:TYA

[Sal04k] Peter H. Salus. Twenty years ago in ;login:. ;login: the USENIX Association newsletter, 29(5):??, October 2004. CODEN LOGNEM. ISSN 1044-6397.

Salus:2005:BRc

[Sal05a] Peter H. Salus. Book reviews. ;login: the USENIX Association newsletter, 30(2): ??, April 2005. CODEN LOGNEM. ISSN 1044-6397.

Salus:2005:BRB

[Sal05b] Peter H. Salus. Book reviews: The bookworm. ;login: the USENIX Association newsletter, 30(1):??, February 2005. CODEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2005-02/openpdfs/
                                        bookworm.pdf.

Salus:2005:UNYd

[Sal05c] Peter H. Salus. USENIX notes: 20 years ago in ;login:. ;login: the USENIX Association newsletter, 30 (4):??, August 2005. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2005-08/openpdfs/
history.pdf.

Salus:2005:UNYa

[Sal05d] Peter H. Salus. USENIX notes: 20 years ago … and more. ;login: the USENIX Association newsletter, 30(1): ??, February 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-02/openpdfs/20.pdf.

Salus:2005:UNYe

[Sal05e] Peter H. Salus. USENIX notes: 25 years ago. ;login: the USENIX Association newsletter, 30(5):??, October 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-10/openpdfs/salus.
pdf.

Salus:2005:UNG

[Sal05f] Peter H. Salus. USENIX notes: Getting it wrong. ;login: the USENIX Association newsletter, 30(6):??, December 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-12/openpdfs/salus.
pdf.

Salus:2005:UNT

[Sal05g] Peter H. Salus. USENIX notes: Thirtieth anniversary, USENIX Association. ;login: the USENIX Association newsletter, 30(2): ??, April 2005. CODEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2005-04/openpdfs/
                                        thirtieth.pdf.

Salus:2005:UNYb

[Sal05h] Peter H. Salus. USENIX notes: Years and years ago. ;login: the USENIX Association newsletter, 30(2): ??, April 2005. CODEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2005-04/openpdfs/
                                        yearsandyears.pdf.

Salus:2005:UNYc

[Sal05i] Peter H. Salus. USENIX notes: Years and years ago. ;login: the USENIX Association newsletter, 30(3): ??, June 2005. CODEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2005-06/openpdfs/
                                        history0506.pdf.

Samuelson:2002:IDA

[Sam02] Pam Samuelson. Implications of the DMCA anticircumvention for security, research, and innovation,

2002. URL http://www.
usenix.org/publications/
library/proceedings/sec02/

tech.html. Unpublished.

Sandnes:2001:SPO

[San01] Frode Eika Sandnes. Scheduling partially ordered events in a randomised framework: Empirical results and implications for automatic configuration management. In USENIX [USE01a], page ?? ISBN 1-880446-05-7. LCCN

????  URL http://db.
usenix.org/publications/
library/proceedings/lisa2001/
tech/sandnes.html.

Samuel:2009:PMS

[SC09] Justin Samuel and Justin Cappos. Package managers still vulnerable: How to protect your systems. ;login: the USENIX Association newsletter, 34(1):??, February 2009. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
february-2009-volume-34-
number-1/package-managers-
still-vulnerable-how-protect.

Schafer:2000:PSS

[Sch00a] Peg Schafer. Perspectives on the SAGE salary survey. ;login: the USENIX Association newsletter, 25(3): ??, June 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-6/sagenews.html#perspectives.

Schrijvers:2000:AN

[Sch00b] Eljakim Schrijvers. Application note. ;login: the USE-NIX Association newsletter, 25(2):??, April 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-4/features/application.
                                        html.

Schuba:2000:SVA

[Sch00c] Christoph Schuba. Supernetworking: Virtual address resolution at the edge. In USENIX [USE00j], page ?? ISBN ???? LCCN

                                        ????  URL http://db.
                                        usenix.org/publications/
                                        library/proceedings/ine2000/
                                        schuba_slides.pdf.

Schilling:2001:AHS

[Sch01]  Richard Schilling.  Affil-

iated Health Services Information Systems clusters. ;login: the USENIX Association newsletter, 26(5): ??, August 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-08/pdfs/schilling.
                                        pdf.

Schadow:2002:FBL

[Sch02a]  Gunther Schadow.  A

FreeBSD-based low-cost broadband VPN router for a telemedicine application. In USENIX [USE02a], pages 111–122. ISBN 1-880446-02-2. LCCN QA76.76.O63 B736

2002. URL http://www.
usenix.org/publications/
library/proceedings/bsdcon02/
schadow.html.

Schneider:2002:BRA

[Sch02b] Raymond M. Schneider.

Book review: Alistair Cockburn: Agile Software Development. ;login: the USE-NIX Association newsletter, 27(2):??, April 2002. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2002-04/openpdfs/
book.pdf.

Schneider:2002:BRT

[Sch02c] Raymond M. Schneider.

Book review: Tom Clark: IPSans: A Guide to ICSI and IFCP and FCIP Protocols for Storage Area Networks. ;login: the USE-NIX Association newsletter, 27(3):??, June 2002. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2002-06/openpdfs/
book.pdf.

Schneider:2002:BRY

[Sch02d] Raymond M. Schneider.

Book review: Yukihiro Matsumoto with translated text by David L. Reynolds, Jr.: Ruby in a Nutshell. ;login: the USENIX Association newsletter, 27(6):??, December 2002. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2002-12/openpdfs/books.
                                        pdf.

Schneier:2002:FNS

[Sch02e]  Bruce Schneier.  Fixing

network security by hacking the business climate. In USENIX [USE02e], page

355. ISBN 1-880446-00-

6. LCCN QA76.8.U65 U84

                                        2002. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/usenix02/
                                        tech/techonefile.html. Un-

published invited talk, 2002 USENIX Annual Technical Conference, June 10-15, 2002, Monterey Conference Center, Monterey, CA.

Scher:2004:SRU

[Sch04] Michael B. Scher. SSO/ RSO/USO/oh no? ;login: the USENIX Association newsletter, 29(6):??, December 2004. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2004-12/pdfs/ohno.pdf.

Scher:2006:DSR

[Sch06] Michael B. Scher. On doing ‘being reasonable’. ;login: the USENIX Association newsletter, 31(6):??, December 2006. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/

december-2006-volume-31-
number-6/doing-being-reasonable.

Stover:2005:UMD

[SD05] Sam Stover and Matt Dickerson. Using memory dumps in digital forensics. ;login: the USENIX Association newsletter, 30(6):??, December 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-12/pdfs/stover.pdf.

Stover:2007:ASN

[SDHD07] Sam Stover, Dave Dittrich, John Hernandez, and Sven Dietrich. Analysis of the Storm and Nugache trojans: P2P is here. ;login: the USENIX Association newsletter, 32(6):??, December 2007. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
december-2007-volume-32-
number-6/analysis-storm-
and-nugache-trojans-p2p-
here.

Smetters:2008:AHG

[SDN^{+}08] Diana Smetters, Brinda Dalal, Les Nelson, Nathaniel Good, and Ame Elliott. Ad hoc guesting: When exceptions are the rule. ;login: the USENIX Association newsletter, 33(4):??, August 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.

                                        org/publications/login/
                                        august-2008-volume-33-number-
                                        4/ad-hoc-guesting-when-
                                        exceptions-are-rule.

Shriraman:2009:TPT

[SDS09] Arrvindh Shriraman, Sandhya Dwarkadas, and Michael L. Scott. Tapping into parallelism with transactional memory. ;login: the USE-NIX Association newsletter, 34(2):??, April 2009. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL https://www.
                                        usenix.org/publications/
                                        login/april-2009-volume-
                                        34-number-2/tapping-parallelism-
                                        transactional-memory.

Sears:2000:ECP

[Sea00] Chris Sears. The elements of cache programming style. In USENIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/als2000/
                                        sears.html.

Seeds:2001:AOP

[See01] Glen Seeds. Austin, October 2000 Plenary for D4 issue resolution. ;login: the USENIX Association newsletter, 26(1): ??, February 2001. CODEN LOGNEM. ISSN 1044-6397.

Seely:2006:SAC

[See06] Andy Seely. System administration of command and control systems. ;login: the USENIX Association newsletter, 31(4):??, August 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2006-volume-31-number-
4/system-administration-
command-and-control-systems.

Sellens:2000:SNM

[Sel00a] John Sellens. System and network monitoring. ;login: the USENIX Association newsletter, 25(5):??, August 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-8/features/monitoring.
html.

Sellens:2000:TDD

[Sel00b]  John Sellens.  Thresh

— a data-directed SNMP threshold poller. In USE-NIX [USE00g], page ?? ISBN 1-880446-13-8. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/lisa2000/
sellens.html.

Sellens:2001:DEI

[Sel01]  John Sellens.  Designing

an external infrastructure. ;login: the USENIX Association newsletter, 26(3): ??, June 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-06/pdfs/sellens.pdf.

Sellens:2002:TRM

[Sel02]  John Sellens.  Ten rea-

sons to monitor your system. ;login: the USENIX Association newsletter, 27(4): ??, August 2002. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2002-08/pdfs/sellens.pdf.

Seltzer:2008:PFS

[Sel08] Margo Seltzer. The present and future of SAN/NAS: Interview with Dave Hitz and Brian Pawlowsky of NetApp. ;login: the USE-NIX Association newsletter, 33(3):??, June 2008. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL https://www.
                                        usenix.org/publications/
                                        login/june-2008-volume-
                                        33-number-3/present-and-
                                        future-sannas-interview-
                                        dave-hitz-and.

Sengar:2007:IT

[Sen07] Hemant Sengar. IP telephony. ;login: the USENIX Association newsletter, 32(5): ??, October 2007. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        october-2007-volume-32-
                                        number-5/ip-telephony.

Schultz:2001:MME

[SEZ^{+}01] Matthew G. Schultz, Eleazar Eskin, Erez Zadok, Manasi Bhattacharyya, and Salvatore J. Stolfo. MEF, Malicious Email Filter — a UNIX mail filter that detects malicious Windows executables. In USENIX [USE01f], page ?? ISBN 1-880446-10-3. LCCN QA76.8.U65

U84 2001. URL http://www.
usenix.org/publications/
library/proceedings/usenix01/
freenix01/schultz.html.

Somayaji:2000:ARU

[SF00] Anil Somayaji and Stephanie Forrest. Automated response using system-call delays. In USENIX [USE00i], page ?? ISBN 1-880446-18-9. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/sec2000/
somayaji.html.

Sirer:2007:SLK

[SF07] Emin Gün Sirer and Rik Farrow. Some lesser-known laws of computer science. ;login: the USENIX Association newsletter, 32(4): ??, August 2007. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2007-volume-32-number-
4/some-lesser-known-laws-
computer-science.

Siddha:2001:PSD

[SG01] Suresh Siddha and K. Gopinath. A persistent snapshot device driver for Linux. In USENIX [USE01d], page ?? ISBN ???? LCCN ???? URL http:/

/www.linuxshowcase.org/
siddha.html.

Shriver:2001:SMW

[SGHS01] Elizabeth Shriver, Eran Gabber, Lan Huang, and Christopher A. Stein. Storage management for Web proxies. In USENIX [USE01b], page ?? ISBN 1-880446-09-X. LCCN QA76.8.U65

U84 2001. URL http://www.
usenix.org/publications/
library/proceedings/usenix01/
shriver.html.

Schindler:2002:TAE

[SGLG02] Jiri Schindler, John Linwood Griffin, Christopher R. Lumb, and Gregory R. Ganger. Track-aligned extents: Matching access patterns to disk drive characteristics. In USENIX [USE02c], page ?? ISBN 1-880446-03-

0. LCCN QA76.9.F5 C65

2002. URL http://www.
usenix.org/publications/
library/proceedings/fast02/
schindler.html.

Seltzer:2000:JVS

[SGM^{+}00] Margo I. Seltzer, Gregory R. Ganger, M. Kirk McKusick, Keith A. Smith, Craig A. N. Soules, and Christopher A. Stein. Journaling versus soft updates: Asynchronous meta-data protection in file systems. In USE-NIX [USE00a], pages 71–

84. ISBN 1-880446-22-7. LCCN ???? URL http:

//www.usenix.org/events/

usenix2000/general/seltzer.
html.

Storer:2008:PEE

[SGMV08] Mark W. Storer, Kevin M.

Greenan, Ethan L. Miller, and Kaladhar Voruganti. Pergamum: Energy-efficient archival storage with disk instead of tape. ;login: the USENIX Association newsletter, 33(3):??, June 2008. CODEN LOGNEM. ISSN 1044-6397. URL

https://www.usenix.org/
publications/login/june-
2008-volume-33-number-3/
pergamum-energy-efficient-
archival-storage-disk.

Strunk:2000:SSS

[SGS^{+}00] John D. Strunk, Garth R.

Goodson, Michael L. Scheinholtz, Craig A. N. Soules, and Gregory R. Ganger. Self-securing storage: Protecting data in compromised systems. In USE-NIX [USE00h], page ?? ISBN 1-880446-16-2. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/osdi2000/
strunk.html.

Sobti:2002:PMS

[SGZ^{+}02] Sumeet Sobti, Nitin Garg, Chi Zhang, Xiang Yu, Arvind Krishnamurthy, and Randolph Wang. PersonalRAID: Mobile storage for distributed and disconnected computers. In USENIX [USE02c], page ?? ISBN 1-880446-03-

0. LCCN QA76.9.F5 C65

                                        2002. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/fast02/
                                        sobti.html.

Schmuck:2002:GSD

[SH02] Frank Schmuck and Roger Haskin. GPFS: a shareddisk file system for large computing clusters. In USENIX [USE02c], page ?? ISBN 1-880446-03-0. LCCN QA76.9.F5 C65 2002. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/fast02/
                                        schmuck.html.

Shaffer:2000:LAC

[Sha00]  Michael W. Shaffer.  A

Linux appliance construction setv. In USENIX [USE00g], page ?? ISBN 1-880446-13-8. LCCN ????

                                        URL http://www.usenix.
                                        org/publications/library/
                                        proceedings/lisa2000/shaffer.
                                        html.

Shapiro:2002:CCM

[Sha02a] Jonathan S. Shapiro. CPCMS: a configuration management system based on cryptographic names. In USE-NIX [USE02d], page ?? ISBN 1-880446-01-4. LCCN QA76.8.U65 P765 2002.

                                        URL http://www.usenix.
                                        org/publications/library/
                                        proceedings/usenix02/tech/
                                        freenix/shapiro.html.

Shaylor:2002:JTC

[Sha02b]  Nik Shaylor.  A just-in-

time compiler for memoryconstrained low-power devices. In USENIX [USE02f], page ?? ISBN 1-931971-01-3. LCCN QA76.73 .J38

J42 2002. URL http://www.
usenix.org/publications/
library/proceedings/javavm02/
shaylor.html.

Shapiro:2004:ELM

[Sha04] Jonathan S. Shapiro. Extracting the lessons of Multics. ;login: the USENIX Association newsletter, 29(6): 57–59, December 2004. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2004-12/pdfs/multics.
pdf.

Shea:2000:MES

[She00] Linchi Shea. Monitor an enterprise of SQL servers

— automating management by exception with Perl. In USENIX [USE00b], page ?? ISBN 1-880446-19-7. LCCN

????  URL http://db.
usenix.org/publications/
library/proceedings/lisa-
nt2000/shea.html.

Shogren:2001:JAT

[Sho01] Carl Shogren. Jack-of-all-trades, master of none. ;login: the USENIX Association newsletter, 26(8):??, December 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-12/pdfs/shogren.pdf.

Stein:2001:UFS

[SHS01] Christopher A. Stein, John H. Howard, and Margo I. Seltzer. Unifying file system protection. In USE-NIX [USE01b], page ?? ISBN 1-880446-09-X. LCCN QA76.8.U65 U84 2001. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/usenix01/
                                        stein.html.

Sifry:2000:OSG

[Sif00] David Sifry. Open-source group calendaring: GCTP and OpenFlock. In USE-NIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/als2000/
                                        sifry.html.

Simonsen:2000:NAS

[Sim00]  Keld Simonsen.  News

about standardization. ;login: the USENIX Association newsletter, 25(3): ??, June 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        standards/37.simonsen.html.

Simonsen:2001:SRS

[Sim01] Keld Simonsen. Standard report: Some standardization news. ;login: the USE-NIX Association newsletter, 26(2):??, April 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
standards/simonsen.pdf.

Simmons:2003:SHL

[Sim03]  Steve Simmons.  Sym-

links and hard links don’t belong in /etc. ;login: the USENIX Association newsletter, 28(1):44– 46, February 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-02/pdfs/simmons.pdf.

Singer:2000:T

[Sin00a]  Daniel Singer.  Toolman.

;login: the USENIX Association newsletter, 25(5): ??, August 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-8/features/toolman.
html.

Sinteur:2000:FH

[Sin00b] John Sinteur. Firewalls at home. ;login: the USE-NIX Association newsletter, 25(4):??, July 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-7/features/firewalls.
html.

Singer:2001:NPP

[Sin01] Abe Singer. No plaintext passwords. ;login: the USE-NIX Association newsletter, 26(7):??, November 2001. CODEN LOGNEM. ISSN 1044-6397. URL http://

                                        www.usenix.org/publications/
                                        login/2001-11/pdfs/singer.
                                        pdf.

Singer:2002:RIW

[Sin02] Abe Singer. The regional information watch. ;login: the USENIX Association newsletter, 27(6):??, December 2002. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2002-12/pdfs/singer.pdf.

Singer:2003:LF

[Sin03] Abe Singer. Life without firewalls. ;login: the USENIX Association newsletter, 28(6): ??, December 2003. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2003-12/pdfs/singer.pdf.

Singer:2005:CPS

[Sin05a]  Abe Singer.  Conference

password sniffing: Legal and ethical issues. ;login: the USENIX Association newsletter, 30(4):??, August 2005. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2005-08/pdfs/singer.pdf.

Singer:2005:TF

[Sin05b] Abe Singer. Tempting fate. ;login: the USENIX Association newsletter, 30(1):??, February 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-02/pdfs/fate.pdf.

Sitaker:2002:PPW

[Sit02]  Kragen Sitaker.  Python

or Perl: Which is better? ;login: the USENIX Association newsletter, 27(3): ??, June 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-06/pdfs/sitaker.pdf.

Sitaker:2003:MWN

[Sit03a] Kragen Sitaker. Microsoft:

What’s next? ;login: the USENIX Association newsletter, 28(3): ??, June 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-06/pdfs/sitaker.pdf.

Sitaker:2003:NPF

[Sit03b]  Kragen Sitaker.  Non-

proprietary flash solutions. ;login: the USENIX Association newsletter, 28(3): ??, June 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-06/pdfs/sitaker2.pdf.

Sivonen:2001:SBC

[Siv01] Timo Sivonen. Setting up BIND8 in a change-rooted environment on Solaris. ;login: the USENIX Association newsletter, 26(4): ??, July 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-07/pdfs/sivonen.pdf.

Sivonen:2003:ICS

[Siv03] Timo Sivonen. IPv6 configuration on Solaris 9 and FreeBSD-4.x. ;login: the USENIX Association newsletter, 28(2): ??, April 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-04/pdfs/sivonen.pdf.

Sivonen:2006:MPF

[Siv06]  Timo Sivonen.  Measur-

ing performance of FreeBSD disk encryption. ;login: the USENIX Association newsletter, 31(5):??, October 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
october-2006-volume-31-
number-5/measuring-performance-
freebsd-disk-encryption.

Sarolahti:2002:CCL

[SK02] Pasi Sarolahti and Alexey Kuznetsov. Congestion control in Linux TCP. In USENIX [USE02d], page ?? ISBN 1-880446-01-4. LCCN QA76.8.U65 P765 2002.

URL http://www.usenix.

org/publications/library/
proceedings/usenix02/tech/
freenix/sarolahti.html.

Skoll:2000:PIL

[Sko00] David Skoll. A PPPoE implementation for Linux. In USENIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/als2000/
skoll.html.

Skvarcek:2003:RMS

[Skv03]  Jozef Skvarcek.  Remote

monitoring with SNMP: a practical example. ;login: the USENIX Association newsletter, 28(1):??, February 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-02/pdfs/skvarcek.pdf.

Lhee:2002:TAD

[sLC02] Kyung suk Lhee and Steve J.

Chapin. Type-assisted dynamic buffer overflow detection. In USENIX [USE02b], pages 81–89. ISBN 1-931971-00-5. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/sec02/
lhee.html.

Savarese:2002:RPA

[SLR02] Chris Savarese, Koen Langendoen, and Jan Rabaey. Robust positioning algorithms for distributed adhoc wireless sensor networks. In USENIX [USE02e], page ?? ISBN 1-880446-00-6. LCCN QA76.8.U65

U84 2002. URL http://www.
usenix.org/publications/
library/proceedings/usenix02/
savarese.html.

Sluyter:2005:GWY

[Slu05] Thomas Sluyter. Getting what you want: The fine art of proposal writing. ;login: the USENIX Association newsletter, 30(1):??, February 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-02/pdfs/sluyter.pdf.

Sluyter:2006:KTT

[Slu06]  Thomas Sluyter.  Keep-

ing track of time. ;login: the USENIX Association newsletter, 31(1):??, February 2006. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2006-02/pdfs/sluyter.pdf.

Sluyter:2007:AQW

[Slu07] Thomas Sluyter. All quiet on the Western Front: a discussion on the necessity of good reporting. ;login: the USE-NIX Association newsletter, 32(2):??, April 2007. CO-DEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/april-2007-volume-
32-number-2/all-quiet-western-

front-discussion-necessity-
good.

Sharp:2002:XXC

[SM02] Jamey Sharp and Bart Massey. XCL: An Xlib compatibility layer for XCB. In USENIX [USE02d], page ?? ISBN 1-880446-01-4. LCCN QA76.8.U65 P765 2002.

URL http://www.usenix.
org/publications/library/
proceedings/usenix02/tech/
freenix/sharp.html.

Schneider:2001:APM

[SMES01] Daniel Schneider, Bernd Mathiske, Matthias Ernst, and Matthew Seidl. Automatic persistent memory management for the Spotless Java^{T M} Virtual Machine on the Palm Connected Organizer. In USENIX [USE01g], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38

J42 2001. URL http://www.
usenix.org/publications/
library/proceedings/jvm01/
schneider.html.

Shah:2001:CNS

[SMF^{+}01] Hemal V. Shah, Dave B.

Minturn, Annie Foong, Gary L. McAlpine, Rajesh S. Madukkarumukumana, and Greg J. Regnier. CSP: a novel system architecture for scalable Internet and communication services. In USENIX [USE01c], page ?? ISBN 1-880446-14-6. LCCN TK5105.875.I57

U96 2001. URL http://www.

                                        usenix.org/publications/
                                        library/proceedings/usits01/
                                        shah.html.

Smith:2001:WEG

[Smi01] Richard M. Smith. Webenabled gadgets: Can we trust them?, 2001. Unpublished keynote address, Tenth USENIX Security Symposium, August 13–17, 2001, Washington, DC, USA.

Smart:2000:DTI

[SMJ00] Matthew Smart, G. Robert Malan, and Farnam Jahanian. Defeating TCP/IP stack fingerprinting. In USENIX [USE00i], page ?? ISBN 1-880446-18-9. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/sec2000/
                                        smart.html.

S:2002:SPI

[SMK02] Venugopal K. S., Geetha Manjunath, and Venkatesh Krishnan. sEc: a portable interpreter optimizing technique for embedded Java Virtual Machine. In USE-NIX [USE02f], page ?? ISBN 1-931971-01-3. LCCN QA76.73 .J38 J42 2002.

                                        URL http://www.usenix.
                                        org/publications/library/
                                        proceedings/javavm02/venugopal.
                                        html.

Sinderson:2004:MWS

[SMM04] Elias Sinderson, Vish Magapu, and Ronald Mak. Middleware and Web services for the collaborative information portal of NASA’s Mars Exploration Rovers Mission. In Jacobsen [Jac04], pages 1–?? CODEN LNCSD9. ISBN 3-540-23428-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.M54 I33 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.springerlink.
com/openurl.asp?genre=issue&
issn=0302-9743&volume=3231;
 http://www.springerlink.
com/openurl.asp?genre=volume&
id=doi:10.1007/b101561.

Small:2008:MST

[SMMM08] Sam Small, Joshua Mason, Ryan MacArthur, and Fabian Monrose. Masquerade: Simulating a thousand victims. ;login: the USENIX Association newsletter, 33(6):??, December 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
december-2008-volume-33-
number-6/masquerade-simulating-
thousand-victims.

Schermerhorn:2001:ULE

[SMRF01] Paul W. Schermerhorn, Robert J.

Minerick, Peter Rijks, and Vincent W. Freeh. Userlevel extensibility in the Mona file system. In USENIX [USE01f], page ?? ISBN 1-880446-10-3. LCCN QA76.8.U65 U84 2001. URL

http://www.usenix.org/publications/

                                        library/proceedings/usenix01/
                                        freenix01/schermerhorn.
                                        html.

Serra:2000:DAL

[SNC00] Albert Serra, Nacho Navarro, and Toni Cortes. DITools: Application-level support for dynamic extension and flexible composition. In USE-NIX [USE00a], pages 225–

238. ISBN 1-880446-22-7. LCCN ???? URL http:

                                        //www.usenix.org/events/
                                        usenix2000/general/serra.
                                        html.

Salim:2001:BS

[SNOK01] Jamal Hadi Salim, Znyx Networks, Robert Olsson, and Alexey Kuznetsov. Beyond Softnet. In USENIX [USE01d], page ?? ISBN ???? LCCN ???? URL http:

                                        //www.linuxshowcase.org/
                                        jamal.html.

Solomon:2008:UAI

[Sol08] Robert Solomon. A 36-user Asterisk installation. ;login: the USENIX Association newsletter, 33(5):??, October 2008. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        october-2008-volume-33-
                                        number-5/36-user-asterisk-
                                        installation.

Sood:2008:HPI

[Soo08a] Aditya K Sood. Hacking 802.11 protocol insecurities.

;login: the USENIX Association newsletter, 33(2): 38–44, April 2008. CO-DEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/april-2008-volume-
33-number-2/hacking-80211-
protocol-insecurities.

Sood:2008:IDX

[Soo08b] Aditya K Sood. Insecurities in designing XML signatures. ;login: the USENIX Association newsletter, 33(1): ??, February 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
february-2008-volume-33-
number-1/insecurities-designing-
xml-signatures.

Sorenson:2006:ERS

[Sor06] Eric Sorenson. Enterprise routing sinkholes using Linux and open source tools. ;login: the USENIX Association newsletter, 31(4):??, August 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2006-volume-31-number-
4/enterprise-routing-sinkholes-
using-linux-and-open.

Samuel:2008:CPM

[SPC08] Justin Samuel, Jeremy Plichta, and Justin Cappos. Centralized package management using Stork. ;login: the USENIX Association newsletter, 33(1):??, February 2008. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        february-2008-volume-33-
                                        number-1/centralized-package-
                                        management-using-stork.

Spinellis:2000:OUT

[Spi00] Diomidis D. Spinellis. Outwit: UNIX tool-based programming meets the Windows world. In USE-NIX [USE00a], pages 149–

158. ISBN 1-880446-22-7. LCCN ???? URL http:

                                        //www.usenix.org/events/
                                        usenix2000/general/spinellis.
                                        html.

Spitzner:2002:HHS

[Spi02]  Lance Spitzner.  HOSUS

(Honeypot Surveillance System). ;login: the USENIX Association newsletter, 27(6): ??, December 2002. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2002-12/pdfs/spitzner.pdf.

Spinellis:2003:OPF

[Spi03] Diomidis D. Spinellis. Organized pruning of file sets. ;login: the USENIX Association newsletter, 28(3): ??, June 2003. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2003-06/pdfs/spinellis.
                                        pdf.

Spinellis:2006:STM

[Spi06] Diomedis Spinellis. Some types of memory are more equal than others. ;login: the USENIX Association newsletter, 31(2):29– 36, April 2006. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2006-04/pdfs/spinellis.
pdf.

Subramanian:2002:GPI

[SPK02] Lakshminarayanan Subramanian, Venkata N. Padmanabhan, and Randy H. Katz. Geographic properties of Internet routing. In USENIX [USE02e], page ?? ISBN 1-880446-00-6. LCCN QA76.8.U65

U84 2002. URL http://www.
usenix.org/publications/
library/proceedings/usenix02/
subramanian.html.

Staniford:2002:HOI

[SPW02] Stuart Staniford, Vern Paxson, and Nicholas Weaver. How to own the Internet in your spare time. In USENIX [USE02b], pages 149–167. ISBN 1-931971-00-5. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/sec02/
staniford.html.

Sullivan:2000:IFR

[SS00] David G. Sullivan and Margo I. Seltzer. Isolation with flexibility: a resource management framework for central servers. In USE-NIX [USE00a], pages 337–

350. ISBN 1-880446-22-7. LCCN ???? URL http:

//www.usenix.org/events/
usenix2000/general/sullivan.
html.

Sato:2002:EOS

[SS02] Hiroki Sato and Keitaro Sekine. Experiences on an open source translation effort in Japan. In USE-NIX [USE02a], pages 19–

26. ISBN 1-880446-02-2. LCCN QA76.76.O63 B736

2002. URL http://www.
usenix.org/publications/
library/proceedings/bsdcon02/
sato.html.

Salus:2005:MAB

[SS05] Emily W. Salus and Peter H. Salus. Marketing after the bubble. ;login: the USENIX Association newsletter, 30(4): ??, August 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-08/pdfs/saluses.pdf.

Speight:2000:WLA

[SSB00] Evan Speight, Hazim Shafi, and John K. Bennett. WS-DLite: a lightweight alternative to Windows sockets direct path. In USE-NIX [USE00d], page ?? ISBN 1-880446-20-0. LCCN

????  URL http://db.
usenix.org/publications/

library/proceedings/usenix-
win2000/speight.html.

Salmon:2009:PSD

[SSCG09] Brandon Salmon, Steven W.

Schlosser, Lorrie Faith Cranor, and Gregory R. Ganger. Perspective: Semantic data management for the home. ;login: the USENIX Association newsletter, 34(5): ??, October 2009. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
october-2009-volume-34-
number-5/perspective-semantic-
data-management-home.

Stacey:2000:RWN

[Sta00] Craig Stacey. Remote Windows NT administration using Windows CE handheld devices. In USE-NIX [USE00b], page ?? ISBN 1-880446-19-7. LCCN

????  URL http://db.
usenix.org/publications/
library/proceedings/lisa-
nt2000/stacey.html.

Sterling:2001:BCC

[Ste01]  Thomas Sterling.  Be-

owulf cluster computing at the Third Plateau. ;login: the USENIX Association newsletter, 26(5):??, August 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-08/pdfs/sterling.pdf.

Shankar:2001:DFS

[STFW01] Umesh Shankar, Kunal Talwar, Jeffrey S. Foster, and David Wagner. Detecting format string vulnerabilities with type qualifiers. In USE-NIX [USE01h], page ?? ISBN 1-880446-18-9, 1-880446-07-3. LCCN QA76.9.A25 U565

                                        2001. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/sec01/
                                        shankar.html.

Stone:2000:CIO

[Sto00a] Robert Stone. CenterTrack: An IP overlay network for tracking DoS floods. In USENIX [USE00i], page ?? ISBN 1-880446-18-9. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/sec2000/
                                        stone.html.

Stoughton:2000:NE

[Sto00b] Nick Stoughton. A new editor. ;login: the USENIX Association newsletter, 25(1): ??, February 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        standards/31.neweditor.
                                        html.

Stoughton:2002:DPA

[Sto02] Nick Stoughton. Developing portable applications. In USENIX [USE02e], page 355. ISBN 1-880446-00-6. LCCN QA76.8.U65 U84

                                        2002. URL http://www.

usenix.org/publications/
library/proceedings/usenix02/
tech/techonefile.html. Un-

published guru session talk, 2002 USENIX Annual Technical Conference, June 10-15, 2002, Monterey Conference Center, Monterey, CA.

Stoughton:2005:SUS

[Sto05]  Nick Stoughton.  Stan-

dards: Update on standards. ;login: the USE-NIX Association newsletter, 30(2):??, April 2005. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2005-04/openpdfs/
standards2004.pdf.

Stoughton:2006:SUS

[Sto06a] Nicholas M. Stoughton. Standards: USENIX standards activities. ;login: the USE-NIX Association newsletter, 31(2):72–75, April 2006. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2006-04/openpdfs/
standards05.pdf.

Stoughton:2006:US

[Sto06b] Nicholas M. Stoughton. An update on standards. ;login: the USENIX Association newsletter, 31(6):??, December 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/

                                        december-2006-volume-31-
                                        number-6/update-standards.

Stoughton:2006:USD

[Sto06c] Nick Stoughton. An update on standards: Diary of a standard geek. ;login: the USENIX Association newsletter, 31(5):??, October 2006. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        october-2006-volume-31-
                                        number-5/update-standards-
                                        diary-standard-geek.

Stoughton:2007:WC

[Sto07a] Nicholas M. Stoughton. Whither C++? ;login: the USENIX Association newsletter, 32(5): ??, October 2007. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        october-2007-volume-32-
                                        number-5/whither-c.

Stoughton:2007:WS

[Sto07b] Nicholas M. Stoughton. Why standardize? ;login: the USENIX Association newsletter, 32(2):??, April 2007. CODEN LOGNEM. ISSN 1044-6397. URL

                                        https://www.usenix.org/
                                        publications/login/april-
                                        2007-volume-32-number-2/
                                        why-standardize.

Stoughton:2007:USR

[Sto07c] Nick Stoughton. An update on standards: Revision fever. ;login: the USENIX Association newsletter, 32(4): ??, August 2007. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2007-volume-32-number-
4/update-standards-revision-
fever.

Stoughton:2008:TA

[Sto08a] Nick Stoughton. Toward attributes. ;login: the USENIX Association newsletter, 33(1): ??, February 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
february-2008-volume-33-
number-1/toward-attributes.

Stoughton:2008:USD

[Sto08b] Nick Stoughton. Update on standards: Diary of a standards geek, part 2. ;login: the USENIX Association newsletter, 33(6):??, December 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
december-2008-volume-33-
number-6/update-standards-
diary-standards-geek-part-
2.

Stoughton:2008:USUb

[Sto08c]  Nick Stoughton.  Update

on standards: The USE-NIX standards project. ;login: the USENIX Association newsletter, 33(5):??, October 2008. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        october-2008-volume-33-
                                        number-5/update-standards-
                                        usenix-standards-project.

Stoughton:2008:USUa

[Sto08d]  Nick Stoughton.  Up-

date on standards: Undue influence? ;login: the USENIX Association newsletter, 33(3):??, June 2008. CODEN LOGNEM. ISSN 1044-6397. URL

                                        https://www.usenix.org/
                                        publications/login/june-
                                        2008-volume-33-number-3/
                                        update-standards-undue-
                                        influence.

Strauss:2002:LOS

[Str02] Daryll Strauss. Life in an open source startup, 2002. Unpublished invited talk, 2002 USENIX Annual Technical Conference, June 10– 15, 2002, Monterey, California, USA.

Sullivan:2000:IF

[Sul00] David G. Sullivan. Isolation with flexibility. ;login: the USENIX Association newsletter, 25(4): ??, July 2000. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2000-7/features/isolation.
                                        html.

Suri:2001:SCR

[Sur01] Niranjan Suri. State capture and resource control for Java: The design and implementation of the Aroma Virtual Machine. In USENIX [USE01g], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38

J42 2001. URL http://www.
usenix.org/publications/
library/proceedings/jvm01/
JVM_wips/S15.pdf.

Sanchez:2002:SMX

[SV02a] Wilfredo Sánchez and Kevin Van Vechten. System-Starter and the Mac OS X startup process. In USE-NIX [USE02a], pages 123–

129. ISBN 1-880446-02-2. LCCN QA76.76.O63 B736

2002. URL http://www.
usenix.org/publications/
library/proceedings/bsdcon02/
sanchez.html.

Shapiro:2002:AIC

[SV02b] Jonathan S. Shapiro and John Vanderburgh. Access and integrity control in a public-access, highassurance configuration management system. In USENIX [USE02b], pages 109–120. ISBN 1-931971-00-5. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/sec02/
shapiro.html.

Sanchez:2001:DGC

[SVF01] Alfonso Sánchez, Luís Veiga, and Paulo Ferreira. Distributed garbage collection for wide area replicated memory. In USE-NIX [USE01e], page ?? ISBN 1-880446-12-X. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/coots01/
                                        sanchez.html.

Sugerman:2001:VDV

[SVL01] Jeremy Sugerman, Ganesh Venkitachalam, and Beng-Hong Lim. Virtualizing I/O devices on VMware Workstation’s hosted virtual machine monitor. In USE-NIX [USE01b], page ?? ISBN 1-880446-09-X. LCCN QA76.8.U65 U84 2001. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/usenix01/
                                        sugerman.html.

Sluyter:2005:WDS

[SvM05] Thomas Sluyter and Roland van Maarschalkerweerd. When disaster strikes: Cailin and Roland discuss crisis management. ;login: the USE-NIX Association newsletter, 30(3):??, June 2005. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2005-06/pdfs/sluyter0506.
                                        pdf.

Smaragdakis:2000:PRM

[SW00] Yannis Smaragdakis and Paul Wilson. Performing replacement in modem pools. In USENIX [USE00a], pages 277–291. ISBN 1-880446-22-

7. LCCN ???? URL http:

//www.usenix.org/events/
usenix2000/general/smaragdakis.
html.

Siebert:2001:DEJ

[SW01] Fridtjof Siebert and Andy Walter. Deterministic execution of Java’s primitive bytecode operations. In USENIX [USE01g], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38 J42 2001. URL

http://www.usenix.org/publications/
library/proceedings/jvm01/
siebert.html.

Swartz:2001:NTR

[Swa01] Ray Swartz. A new twist on random number generators. ;login: the USENIX Association newsletter, 26(1): ??, February 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-02/pdfs/swartz.pdf.

Swartz:2002:PMR

[Swa02a]  Ray Swartz.  Planning

a model retirement. ;login: the USENIX Association newsletter, 27(2): ??, April 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-04/pdfs/swartz.pdf.

Swartz:2002:WDY

[Swa02b]  Ray Swartz.  When do

you plan to retire? ;login: the USENIX Association newsletter, 27(1):??, February 2002. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2002-02/openpdfs/swartz.
                                        pdf.

Swartz:2003:ILR

[Swa03a] Ray Swartz. Isn’t that a little risky? ;login: the USE-NIX Association newsletter, 28(2):??, April 2003. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2003-04/pdfs/swartz.pdf.

Swartz:2003:WCF

[Swa03b] Ray Swartz. Where can I find a good investment? ;login: the USENIX Association newsletter, 28(4): ??, August 2003. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2003-08/pdfs/swartz.pdf.

Song:2001:TAK

[SWT01] Dawn Xiaodong Song, David Wagner, and Xuqing Tian. Timing analysis of keystrokes and timing attacks on SSH. In USENIX [USE01h], page ?? ISBN 1-880446-18-9, 1-880446-07-3. LCCN QA76.9.A25 U565 2001. URL http://

                                        www.usenix.org/publications/
                                        library/proceedings/sec01/
                                        song.html.

Shen:2001:NSR

[SYC^{+}01] Kai Shen, Tao Yang, Lingkun Chu, JoAnne L. Holliday, Douglas A. Kuschner, and Huican Zhu. Neptune: Scalable replication management and programming support for cluster-based network services. In USENIX [USE01c], page ?? ISBN 1-880446-14-6. LCCN TK5105.875.I57

U96 2001. URL http://www.
usenix.org/publications/
library/proceedings/usits01/
shen.html.

Suganuma:2002:ESM

[SYN02] Toshio Suganuma, Toshiaki Yasue, and Toshio Nakatani. An empirical study of method in-lining for a Java just-in-time compiler. In USENIX [USE02f], page ?? ISBN 1-931971-01-3. LCCN QA76.73 .J38

J42 2002. URL http://www.
usenix.org/publications/
library/proceedings/javavm02/
suganuma.html.

Syverson:2001:DAT

[Syv01]  Paul Syverson.  Design-

ing against traffic analysis, 2001. Unpublished invited talk, Tenth USENIX Security Symposium, August 13–17, 2001, Washington, DC, USA.

Spadavecchia:2002:ENC

[SZ02] Joseph Spadavecchia and Erez Zadok. Enhancing NFS cross-administrative domain access. In USENIX [USE02d], page ?? ISBN 1-880446-01-

4. LCCN QA76.8.U65 P765

2002. URL http://www.
usenix.org/publications/
library/proceedings/usenix02/
tech/freenix/spadavecchia.
html.

Tarditi:2002:ITR

[Tar02]  David Tarditi.  Invited

talk: Research opportunities and research challenges. In USENIX [USE02f], page ?? ISBN 1-931971-01-3. LCCN QA76.73 .J38

J42 2002. URL http://www.
usenix.org/publications/
library/proceedings/javavm02/
talk/index.html.

Taylor:2001:XXC

[Tay01] Owen Taylor. XSETTINGS/ XEMBED cross-toolkit protocols. In USENIX [USE01d], page ?? ISBN ???? LCCN ???? URL http:

//www.linuxshowcase.org/
taylor.html.

USENIX-Board:2004:OLU

[The04] The USENIX Board of Directors. An open letter from the USENIX Association rebutting SCO’s position on open source software. ;login: the USENIX Association newsletter, 29(2): ??, April 2004. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2004-04/openpdfs/Usenix.
pdf.

Thomas:2000:BBL

[Tho00a] Guignon Thomas. BLASTH, a BLAS library for dual SMP computers. In USE-NIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/als2000/
thomas.html.

Thornton:2000:LS

[Tho00b]  Adam Thornton.  Linux

on the System/390. In USENIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/als2000/
thornton.html.

Threadgill:2001:EM

[Thr01a] Richard Threadgill. Easy management. ;login: the USENIX Association newsletter, 26(3):??, June 2001. CODEN LOGNEM. ISSN 1044-6397. URL http://

www.usenix.org/publications/
login/2001-06/pdfs/threadgill.
pdf.

Throwe:2001:RCF

[Thr01b] Tom Throwe. The RHIC computing facility Linux farms. ;login: the USENIX Association newsletter, 26(5): ??, August 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-08/pdfs/throwe.pdf.

Tierney:2001:TTG

[Tie01] Brian L. Tierney. TCP tuning guide for distributed application on wide area networks. ;login: the USENIX Association newsletter, 26(1): ??, February 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-02/pdfs/tierney.pdf.

Toledo:2001:FSD

[TK01] Sivan Toledo and Lars Knoll. Font subsetting and downloading in the PostScript printer driver of Qt/X11. In USENIX [USE01d], page ?? ISBN ???? LCCN ???? URL http:/

/www.linuxshowcase.org/
toledo.html.

Tai:2004:CCW

[TKM04] Stefan Tai, Rania Khalaf, and Thomas Mikalsen. Composition of coordinated Web services. In Jacobsen [Jac04], pages 294–?? CODEN LNCSD9. ISBN 3-540-23428-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.M54 I33 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.springerlink.
com/openurl.asp?genre=issue&
issn=0302-9743&volume=3231;
 http://www.springerlink.
com/openurl.asp?genre=volume&
id=doi:10.1007/b101561.

Teigland:2001:VML

[TM01] David Teigland and Heinz Mauelshagen. Volume managers in Linux. In USE-NIX [USE01f], page ?? ISBN 1-880446-10-3. LCCN QA76.8.U65 U84 2001. URL

http://www.usenix.org/publications/
library/proceedings/usenix01/
freenix01/teigland.html.

Tomlinson:2000:INE

[Tom00] Gary Tomlinson. ICAP at network edge caches. In USENIX [USE00j], page ?? ISBN ???? LCCN

????  URL http://db.
usenix.org/publications/
library/proceedings/ine2000/
tomlinson_slides.html;
http://db.usenix.org/publications/
library/proceedings/ine2000/
tomlinson_slides.ppt.

Toomey:2009:REU

[Too09] Warren Toomey. The restoration of early UNIX artifacts. In ????, editor, Proceedings of the 2009 USENIX Annual Technical Conference, June 14–19, San Diego, CA, USA, page ?? USENIX, Berkeley, CA, USA, 2009. ISBN ???? LCCN ???? URL http:

//www.usenix.org/events/
usenix09/tech/full_papers/
toomey/toomey.pdf; http:
//www.usenix.org/events/
usenix09/tech/slides/toomey.
pdf;   http://www.usenix.
org/multimedia/atc09toomey.

Traugott:2001:IA

[Tra01]  Steve Traugott.  Infras-

tructure architecture, 2001.

URL http://db.usenix.
org/publications/library/
proceedings/lisa2001/tech/

. Unpublished invited talk, LISA 2001: 15th Systems Administration Conference, December 2–7, 2001, Town and Country Resort Hotel, San Diego, CA.

Traugott:2002:ASA

[Tra02] Steve Traugott. Automated system administration. In USENIX [USE02e], page

355. ISBN 1-880446-00-

6. LCCN QA76.8.U65 U84

2002. URL http://www.
usenix.org/publications/
library/proceedings/usenix02/
tech/techonefile.html. Un-

published guru session talk, 2002 USENIX Annual Technical Conference, June 10-15, 2002, Monterey Conference Center, Monterey, CA.

Tilevich:2004:PED

[TS04] Eli Tilevich and Yannis Smaragdakis. Portable and efficient distributed threads for Java. In Jacobsen [Jac04], pages 478–492. CODEN LNCSD9. ISBN 3-540-23428-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.M54 I33 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.springerlink.

com/openurl.asp?genre=issue&
issn=0302-9743&volume=3231;
 http://www.springerlink.
com/openurl.asp?genre=volume&
id=doi:10.1007/b101561.

Tsafrir:2008:MIC

[TSW08] Dan Tsafrir, Dilma Da Silva, and David Wagner. The murky issue of changing process identity: Revising ‘setuid demystified’. ;login: the USE-NIX Association newsletter, 33(3):??, June 2008. CO-DEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/june-2008-volume-
33-number-3/murky-issue-
changing-process-identity-
revising.

Tso:2002:PEL

[TT02] Theodore Ts’o and Stephen Tweedie. Planned extensions to the Linux Ext2/ Ext3 filesystem. In USE-NIX [USE02d], page ?? ISBN 1-880446-01-4. LCCN QA76.8.U65 P765 2002.

URL http://www.usenix.
org/publications/library/
proceedings/usenix02/tech/
freenix/tso.html.

Turoff:2002:PPa

[Tur02a] Adam Turoff. Practical Perl.

;login: the USENIX Association newsletter, 27(3): ??, June 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-06/pdfs/turoff.pdf.

Turoff:2002:PPb

[Tur02b] Adam Turoff. Practical Perl. ;login: the USENIX Association newsletter, 27(4): ??, August 2002. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2002-08/pdfs/turoff.pdf.

Turoff:2002:PPc

[Tur02c] Adam Turoff. Practical Perl. ;login: the USENIX Association newsletter, 27(5): ??, October 2002. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2002-10/pdfs/turoff.pdf.

Turoff:2003:PPCb

[Tur03a] Adam Turoff. Practical Perl: Cleaning up with modules. ;login: the USENIX Association newsletter, 28(4): ??, August 2003. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2003-08/pdfs/turoff.pdf.

Turoff:2003:PPCa

[Tur03b]  Adam Turoff.  Practical

Perl: CPAN, modules, and the CPAN SHELL. ;login: the USENIX Association newsletter, 28(1):??, February 2003. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2003-02/pdfs/turoff.pdf.

Turoff:2003:PPF

[Tur03c]  Adam Turoff.  Practical

Perl: Fixing broken modules. ;login: the USENIX Association newsletter, 28(2): ??, April 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-04/pdfs/turoff.pdf.

Turoff:2003:PPK

[Tur03d]  Adam Turoff.  Practical

Perl: Keeping it simple. ;login: the USENIX Association newsletter, 28(3): ??, June 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-06/pdfs/turoff.pdf.

Turoff:2003:PPU

[Tur03e]  Adam Turoff.  Practical

Perl: Using object factories. ;login: the USENIX Association newsletter, 28(5): ??, October 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-10/pdfs/turoff.pdf.

Turoff:2004:PPIb

[Tur04a] Adam Turoff. Practical Perl:

An introduction to Web services. ;login: the USENIX Association newsletter, 29(2): ??, April 2004. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2004-04/pdfs/Turoff.pdf.

Turoff:2004:PPD

[Tur04b]  Adam Turoff.  Practi-

cal Perl: Database modeling with Class::DBI. ;login: the USENIX Association newsletter, 29(4):??, August 2004. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2004-08/pdfs/turoff.pdf.

Turoff:2004:PPIa

[Tur04c] Adam Turoff. Practical Perl: Integration with inline. ;login: the USENIX Association newsletter, 29(1):??, February 2004. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2004-02/pdfs/turoff.pdf.

Turoff:2004:PPW

[Tur04d]  Adam Turoff.  Practical

Perl: Web automation. ;login: the USENIX Association newsletter, 29(3): ??, June 2004. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2004-06/pdfs/turoff.pdf.

Turoff:2005:PPDa

[Tur05a] Adam Turoff. Practical Perl: Date and time formatting in Perl. ;login: the USE-NIX Association newsletter, 30(3):??, June 2005. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.

org/publications/login/
2005-06/pdfs/turoff0506.
pdf.

Turoff:2005:PPDb

[Tur05b]  Adam Turoff.  Practical

Perl: Defensive CGI programming with Taint mode and CGI::Untaint. ;login: the USENIX Association newsletter, 30(4):??, August 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-08/pdfs/turoff.pdf.

Turoff:2005:PPE

[Tur05c] Adam Turoff. Practical Perl:

Error handling patterns in Perl. ;login: the USENIX Association newsletter, 30(1): ??, February 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-02/pdfs/turoff.pdf.

Turoff:2005:PPP

[Tur05d] Adam Turoff. Practical Perl:

Programs to write programs. ;login: the USENIX Association newsletter, 30(2): ??, April 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-04/pdfs/turoff0504.
pdf.

Tylock:2001:ICB

[Tyl01] Steven M. Tylock. If computers had blood, we’d be called doctors, part 1. ;login: the USENIX Association newsletter, 26(8):??, December 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-12/pdfs/tylock.pdf.

Tylock:2002:GVL

[Tyl02a] Steven M. Tylock. GVSAGE visits the local technology show. ;login: the USENIX Association newsletter, 27(1): ??, February 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-02/pdfs/tylock.pdf.

Tylock:2002:ICB

[Tyl02b] Steven M. Tylock. If computers had blood, we’d be called doctors, part 2. ;login: the USENIX Association newsletter, 27(1):??, February 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-02/pdfs/tylock2.pdf.

Tylock:2002:ISW

[Tyl02c] Steven M. Tylock. If systems were cars, would yours be doubled parked? ;login: the USENIX Association newsletter, 27(4):??, August 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-08/pdfs/tylock.pdf.

Tylock:2002:SCL

[Tyl02d]  Steven M. Tylock.  The

SysAdmin’s Court: Liz Action vs. Jack Trades. ;login: the USENIX Association newsletter, 27(4):??, August 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-08/pdfs/tylock2.pdf.

Tylock:2003:CST

[Tyl03a] Steven M. Tylock. Confessions of a sysadmin turned salesman. ;login: the USE-NIX Association newsletter, 28(1):??, February 2003. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2003-02/pdfs/tylock.
pdf.

Tylock:2003:PHG

[Tyl03b]  Steven M. Tylock.  Pri-

vate, HMO, and group support. ;login: the USENIX Association newsletter, 28(3): ??, June 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-06/pdfs/tylock.pdf.

Uhley:2003:WBV

[Uhl03]  Peleus G. Uhley.  Web

Browser Vulnerabilities 101. ;login: the USENIX Association newsletter, 28(1):??, February 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-02/pdfs/uhley.pdf.

Uris:2006:DBC

[Uri06] Mark Uris. DataMonster: Building a cost-effective, high-speed, heterogeneous shared file system. ;login: the USENIX Association newsletter, 31(4):??, August 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2006-volume-31-number-
4/datamonster-building-
cost-effective-high-speed.

Uthayopas:2001:PCK

[US01] Putchong Uthayopas and Somsak Sriprayoonsakul. The PIRUN cluster at Kasetsart University: The most powerful supercomputer in Thailand. ;login: the USENIX Association newsletter, 26 (5):??, August 2001. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2001-08/pdfs/utha.
pdf.

USENIX:2000:UAT

[USE00a] USENIX, editor. 2000 USE-NIX Annual Technical Conference: San Diego, CA, USA, June 18–23, 2000. USENIX, Berkeley, CA, USA, 2000. ISBN 1-880446-22-7. LCCN ????

URL http://www.usenix.

org/publications/library/
proceedings/usenix2000.

USENIX:2000:PLI

[USE00b] USENIX, editor. Proceedings of the 3rd Large Installation System Administration of Windows NT/2000 Conference: August 1–2, 2000, Seattle, Washington, USA. USENIX, Berkeley, CA, USA, 2000. ISBN 1-880446-19-7. LCCN ????

URL http://db.usenix.
org/publications/library/
proceedings/lisa-nt2000/

.

USENIX:2000:PAL

[USE00c]  USENIX, editor.  Pro-

ceedings of the 4th Annual Linux Showcase and Conference, Atlanta, October 10–14, 2000, Atlanta, Georgia, USA. USENIX, Berkeley, CA, USA, 2000. ISBN 1-880446-17-0. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/als2000/

.

USENIX:2000:PUW

[USE00d] USENIX, editor. Proceedings of the 4th USENIX Windows Systems Symposium: August 3–4, 2000, Seattle, Washington, USA. USENIX, Berkeley, CA, USA, 2000. ISBN 1-880446-20-0. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/usenix-
win2000.

USENIX:2000:PUT

[USE00e] USENIX, editor. Proceedings of the 7th USENIX Tcl/Tk Conference (Tcl/2k): February 14–18, 2000, Austin, Texas, USA. USENIX, Berkeley, CA, USA, 2000. ISBN 1-880446-24-3. LCCN ????

                                        URL http://db.usenix.
                                        org/publications/library/
                                        proceedings/tcl2k/.

USENIX:2000:PFW

[USE00f] USENIX, editor. Proceedings of the First Workshop on Industrial Experiences with Systems Software (WIESS 2000), October 22, 2000, Paradise Point Resort, San Diego, California, USA. USENIX, Berkeley, CA, USA, 2000. ISBN 1-880446-15-4. LCCN QA76.76.S95 W67 2000.

                                        URL http://www.usenix.
                                        org/publications/library/
                                        proceedings/osdi2000/wiess2000/

.

USENIX:2000:PFSa

[USE00g] USENIX, editor. Proceedings of the Fourteenth Systems Administration Conference (LISA XIV), December 3–8, 2000, New Orleans, Louisiana, USA. USENIX, Berkeley, CA, USA, 2000. ISBN 1-880446-13-8. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/lisa2000/
                                        technical.html.

USENIX:2000:PFSb

[USE00h] USENIX, editor. Proceedings of the Fourth Symposium on Operating Systems Design and Implementation (OSDI 2000), October 23–25, 2000, San Diego, California, USA. USENIX, Berkeley, CA, USA, 2000. ISBN 1-880446-16-2. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/osdi2000/

.

USENIX:2000:PNU

[USE00i] USENIX, editor. Proceedings of the Ninth USE-NIX Security Symposium, August 14–17, 2000, Denver, Colorado. USENIX, Berkeley, CA, USA, 2000. ISBN 1-880446-18-9. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/sec2000.

USENIX:2000:PSW

[USE00j] USENIX, editor. Proceedings of the Special Workshop on Intelligence at the Network Edge: March 20, 2000, Holiday Inn — Fisherman’s Wharf, San Francisco, CA. USENIX, Berkeley, CA, USA, 2000. ISBN ???? LCCN

????  URL http://db.
usenix.org/publications/
library/proceedings/ine2000/

.

USENIX:2001:PLS

[USE01a] USENIX, editor. Proceedings of LISA 2001: 15th Systems Administration Conference, December 2–7, 2001, Town and Country Resort Hotel, San Diego, CA. USENIX, Berkeley, CA, USA, 2001. ISBN 1-880446-05-7. LCCN

????  URL http://db.
usenix.org/publications/
library/proceedings/lisa2001/
tech/.

USENIX:2001:PUA

[USE01b] USENIX, editor. Proceedings of the 2001 USENIX Annual Technical Conference: June 25–30, 2001, Marriott Copley Place Hotel, Boston, Massachusetts, USA. USE-NIX, Berkeley, CA, USA, 2001. ISBN 1-880446-09-X. LCCN QA76.8.U65 U84

2001. URL http://www.
usenix.org/publications/
library/proceedings/usenix01/
technical.html.

USENIX:2001:PUS

[USE01c] USENIX, editor. Proceedings of the 3rd USENIX Symposium on Internet Technologies and Systems (USITS ’01), March 26–28, 2001, San Francisco, California, USA. USENIX, Berkeley, CA, USA, 2001. ISBN 1-880446-14-6. LCCN TK5105.875.I57

U96 2001. URL http://www.
usenix.org/publications/
library/proceedings/usits01/
technical.html.

USENIX:2001:PAL

[USE01d] USENIX, editor. Proceedings of the 5th Annual Linux Showcase and Conference, November 5–10, 2001, Oakland, CA. USENIX, Berkeley, CA, USA, 2001. ISBN ???? LCCN ???? URL http:

//www.linuxshowcase.org/
tech.html.

USENIX:2001:PUC

[USE01e] USENIX, editor. Proceedings of the 6th USENIX Conference on Object-Oriented Technologies and Systems, January 29–February 2, 2001, San Antonio, Texas, USA. USENIX, Berkeley, CA, USA, 2001. ISBN 1-880446-12-X. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/coots01/

.

USENIX:2001:PFT

[USE01f] USENIX, editor. Proceedings of the FREENIX Track: 2001 USENIX Annual Technical Conference: June 25– 30, 2001, Marriott Copley Place Hotel, Boston, Massachusetts, USA. USE-NIX, Berkeley, CA, USA, 2001. ISBN 1-880446-10-

3. LCCN QA76.8.U65 U84

2001. URL http://www.
usenix.org/publications/
library/proceedings/usenix01/
freenix01/technical.html.

USENIX:2001:PJV

[USE01g] USENIX, editor. Proceedings of the Java Virtual Machine Research and Technology Sy[m]posium (JVM ’01): April 23–24, 2001, Monterey, California, USA. Berkeley, CA. USENIX, Berkeley, CA, USA, 2001. ISBN 1-880446-11-1. LCCN QA76.73.J38

                                        J42 2001. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/jvm01/

.

USENIX:2001:PTU

[USE01h] USENIX, editor. Proceedings of the Tenth USENIX Security Symposium, August 13–17, 2001, Washington, DC, USA. USENIX, Berkeley, CA, USA, 2001. ISBN 1-880446-18-9, 1-880446-07-

3. LCCN QA76.9.A25 U565

                                        2001. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/sec01/
                                        technical.html.

USENIX:2002:PBF

[USE02a] USENIX, editor. Proceedings of BSDCon 2002: February 11–14, 2002, Cathedral Hill Hotel, San Francisco, CA. USENIX, Berkeley, CA, USA, 2002. ISBN 1-880446-02-2. LCCN QA76.76.O63 B736 2002. URL http://

                                        www.usenix.org/publications/
                                        library/proceedings/bsdcon02/
                                        tech.html.

USENIX:2002:PUS

[USE02b]  USENIX, editor.  Pro-

ceedings of the 11th USE-NIX Security Symposium 2002, August 5–9, 2002, San Francisco, CA. USENIX, Berkeley, CA, USA, 2002. ISBN 1-931971-00-5. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/sec02/

.

USENIX:2002:PFC

[USE02c] USENIX, editor. Proceedings of the FAST ’02 Conference on File and Storage Technologies: January 28-30, 2002, Monterey, California, USA. USENIX, Berkeley, CA, USA, 2002. ISBN 1-880446-03-0. LCCN QA76.9.F5 C65

2002. URL http://www.
usenix.org/publications/
library/proceedings/fast02/

.

USENIX:2002:PFT

[USE02d]  USENIX, editor.  Pro-

ceedings of the FREENIX Track: 2002 USENIX Annual Technical Conference, June 10–15, 2002, Monterey, California, USA. USE-NIX, Berkeley, CA, USA, 2002. ISBN 1-880446-01-

4. LCCN QA76.8.U65 P765

2002. URL http://www.
usenix.org/publications/
library/proceedings/usenix02/

.

USENIX:2002:PGT

[USE02e]  USENIX, editor.  Pro-

ceedings of the General Track: 2002 USENIX Annual Technical Conference, June 10–15, 2002, Monterey, California, USA. USE-NIX, Berkeley, CA, USA, 2002. ISBN 1-880446-00-

6. LCCN QA76.8.U65 U84

                                        2002. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/usenix02/

.

USENIX:2002:PJV

[USE02f] USENIX, editor. Proceedings of the Java Virtual Machine Research and Technology Symposium (JVM ’02): August 1–2, 2002, San Francisco, California, US. USE-NIX, Berkeley, CA, USA, 2002. ISBN 1-931971-01-

3. LCCN QA76.73 .J38 J42

                                        2002. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/javavm02/

.

USENIX:2006:PUA

[USE06] USENIX, editor. Proceedings of the 2006 USENIX Annual Technical Conference: May 30–June 3, 2006, Boston, MA, USA. USENIX, Berkeley, CA, USA, 2006. ISBN 1-931971-44-7. LCCN ????

VanEpp:2001:PWB

[Van01] Peter Van Epp. Psst, wanna buy some network insurance? ;login: the USENIX Association newsletter, 26(7):??, November 2001. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2001-11/pdfs/epp.pdf.

VanDrunen:2009:BP

[Van09a] Rudi Van Drunen. The basics of Power. ;login: the USE-NIX Association newsletter, 34(2):??, April 2009. CO-DEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/april-2009-volume-
34-number-2/basics-power.

VanDrunen:2009:HBN

[Van09b] Rudi Van Drunen. A homebuilt NTP appliance. ;login: the USENIX Association newsletter, 34(4):??, August 2009. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2009-volume-34-number-
4/home-built-ntp-appliance.

VanDrunen:2009:S

[Van09c]  Rudi Van Drunen.  Sig-

nals. ;login: the USENIX Association newsletter, 34 (3):??, June 2009. CO-DEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/june-2009-volume-
34-number-3/signals.

Vckovski:2000:MTS

[VB00] Andrej Vckovski and Jason Brazile. A multithreaded server for shared hash table access. In USE-NIX [USE00e], page ?? ISBN 1-880446-24-3. LCCN

????  URL http://db.
usenix.org/publications/
library/proceedings/tcl2k/
vckovski.html.

vonBehren:2002:NFN

[vBBB^{+}02] J. Robert von Behren, Eric A. Brewer, Nikita Borisov, Michael Chen, Matt Welsh, Josh MacDonald, Jeremy Lau, Steve Gribble, and David Culler. Ninja: a framework for network services. In USENIX [USE02e], page ?? ISBN 1-880446-00-6. LCCN QA76.8.U65

U84 2002. URL http://www.
usenix.org/publications/
library/proceedings/usenix02/
behren.html.

vanDoorn:2000:SJV

[vD00] Leendert van Doorn. A secure Java^{T M} Virtual Machine. In USENIX [USE00i], page ?? ISBN 1-880446-18-9. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/sec2000/
vandoorn.html.

vanDrunen:2009:CSE

[vD09] Rudi van Drunen. Chips and static electricity. ;login: the USENIX Association newsletter, 34(5):??, October 2009. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
october-2009-volume-34-
number-5/chips-and-static-
electricity.

vanderLinden:2002:PNA

[vdL02] Frank van der Linden. Porting NetBSD to the AMD x86-64: a case study in OS portability. In USENIX [USE02a], pages 1–9. ISBN 1-880446-02-

2. LCCN QA76.76.O63 B736

2002. URL http://www.
usenix.org/publications/
library/proceedings/bsdcon02/
linden.html.

Veglia:2007:SUB

[Veg07] Toni Veglia. Summary of USENIX Board of Directors meetings and actions. ;login: the USENIX Association newsletter, 32(4): ??, August 2007. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2007-volume-32-number-
4/summary-usenix-board-
directors-meetings-and.

Verbowski:2007:SLC

[Ver07] Chad Verbowski. The secret lives of computers exposed: Flight data recorder for Windows. ;login: the USE-NIX Association newsletter, 32(2):??, April 2007. CO-DEN LOGNEM. ISSN 1044-

6397. URL https://www.

                                        usenix.org/publications/
                                        login/april-2007-volume-
                                        32-number-2/secret-lives-
                                        computers-exposed-flight-
                                        data-recorder.

VanMeter:2000:LMS

[VG00] Rodney Van Meter and Minxi Gao. Latency management in storage systems. In USENIX [USE00h], page ?? ISBN 1-880446-16-2. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/osdi2000/
                                        vanmeter.html.

Vicente:2000:QPN

[Vic00] John Vicente. QoS provisioning at the network edge. In USENIX [USE00j], page ?? ISBN ???? LCCN

                                        ????  URL http://db.
                                        usenix.org/publications/
                                        library/proceedings/ine2000/
                                        vicente_slides.html;   http:
                                        //db.usenix.org/publications/
                                        library/proceedings/ine2000/
                                        vicente_slides.ppt.

Vixie:2002:SRR

[Vix02] Paul A. Vixie. Spam — roles and responsibilities,

                                        2002. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/bsdcon02/

tech.html. Unpublished invited talk, BSDCON2002: Growing the BSD Community, February 11–14, 2002, Cathedral Hill Hotel, San Francisco, CA.

Vo:2002:EDC

[VK02] Phong Vo and David Korn.

Engineering a differencing and compression data format. In USENIX [USE02e], page ?? ISBN 1-880446-00-6. LCCN QA76.8.U65

U84 2002. URL http://www.
usenix.org/publications/
library/proceedings/usenix02/
korn.html.

Vijaykrishnan:2001:EBJ

[VKK^{+}01] N. Vijaykrishnan, M. Kandemir, S. Kim, S. Tomar, A. Sivasubramaniam, and M. J. Irwin. Energy behavior of Java applications from the memory perspective. In USENIX [USE01g], page ?? ISBN 1-880446-11-1. LCCN QA76.73.J38

J42 2001. URL http://www.
usenix.org/publications/
library/proceedings/jvm01/
vijaykrishnan.html.

Voleynik:2000:EPT

[Vol00]  Jarra Voleynik.  eEMU:

a practical tool and language for system monitoring and event management. In USENIX [USE00g], page ?? ISBN 1-880446-13-8. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/lisa2000/
voleynik.html.

vanRiel:2001:PRL

[vR01a] Rik van Riel. Page replacement in Linux 2.4 memory management. In USE-NIX [USE01f], page ?? ISBN 1-880446-10-3. LCCN QA76.8.U65 U84 2001. URL

http://www.usenix.org/publications/
library/proceedings/usenix01/
freenix01/riel.html.

vanRooij:2001:RST

[vR01b] Guido van Rooij. Real stateful TCP packet filtering in Ip-filter, 2001. URL http:

//www.usenix.org/events/
sec01/invitedtalks/rooij.

pdf. Unpublished invited talk, Tenth USENIX Security Symposium, August 13–17, 2001, Washington, DC, USA.

Vrable:2009:CFB

[VSV09] Michael Vrable, Stefan Savage, and Geoffrey M. Voelker. Cumulus: Filesystem backup to the Cloud. ;login: the USENIX Association newsletter, 34(4):??, August 2009. CODEN LOGNEM. ISSN 1044-6397. URL

https://www.usenix.org/
publications/login/august-
2009-volume-34-number-4/
cumulus-filesystem-backup-
cloud.

Voigt:2001:KMS

[VTFM01] Thiemo Voigt, Renu Tewari, Douglas Freimuth, and Ashish Mehra. Kernel mechanisms for service differentiation in overloaded Web servers. In USENIX [USE01b], page ?? ISBN 1-880446-09-X. LCCN QA76.8.U65 U84 2001. URL

http://www.usenix.org/publications/

library/proceedings/usenix01/
voigt.html.

Wallach:2000:GLN

[Wal00] Ryan S. Wallach. Gemini Lite: a non-intrusive debugger for Windows NT. In USENIX [USE00d], page ?? ISBN 1-880446-20-0. LCCN

????  URL http://db.
usenix.org/publications/
library/proceedings/usenix-
win2000/wallach.html.

Wallman:2006:SFE

[Wal06] Marc Wallman. Spam filtering for the enterprise. ;login: the USENIX Association newsletter, 31(1):??, February 2006. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2006-02/pdfs/wallman.pdf.

Wallach:2007:CEW

[Wal07]  Dan Wallach.  Creating

the EVT Workshop. ;login: the USENIX Association newsletter, 32(5):??, October 2007. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
october-2007-volume-32-
number-5/creating-evt-workshop.

Walker:2008:BAE

[Wal08] Edward Walker. Benchmarking Amazon EC2 for highperformance scientific computing. ;login: the USENIX Association newsletter, 33(5):

??, October 2008. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        october-2008-volume-33-
                                        number-5/benchmarking-amazon-
                                        ec2-high-performance.

Ward:2001:PSE

[War01]  Brian Ward.  A practi-

cal scripting environment for mobile devices. In USENIX [USE01f], page ?? ISBN 1-880446-10-3. LCCN QA76.8.U65 U84 2001. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/usenix01/
                                        freenix01/ward.html.

Watson:2001:TAT

[Wat01] Robert N. M. Watson. TrustedBSD: Adding trusted operating system features to FreeBSD. In USE-NIX [USE01f], page ?? ISBN 1-880446-10-3. LCCN QA76.8.U65 U84 2001. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/usenix01/
                                        freenix01/watson.html.

Wayman:2002:BAT

[Way02] James L. Wayman. Biometric authentication technologies: Hype meets the test results,

                                        2002. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/sec02/

tech.html. Unpublished.

Weerawarana:2001:BML

[WCD^{+}01] Sanjiva Weerawarana, Francisco Curbera, Matthew J.

Duftler, David A. Epstein, and Joseph Kesselman. Bean Markup Language: a composition language for JavaBeans components. In USENIX [USE01e], page ?? ISBN 1-880446-12-X. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/coots01/
weerawarana.html.

Royes:2005:WDS

[wCfJRW05] Christopher Jordan (with Contributions from Jason Royes and Jesse Whyte). Writing detection signatures. ;login: the USENIX Association newsletter, 30(6):??, December 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-12/pdfs/jordan.pdf.

Wright:2002:LSM

[WCS^{+}02] Chris Wright, Crispin Cowan, Stephen Smalley, James Morris, and Greg Kroah-Hartman. Linux security modules: General security support for the Linux kernel. In USENIX [USE02b], pages 17–31. ISBN 1-931971-00-5. LCCN ????

URL http://www.usenix.
org/publications/library/
proceedings/sec02/wright.
html.

Westerlund:2001:HWK

[WD01] Assar Westerlund and Johan Danielsson. Heimdal and Windows 2000 Kerberos — how to get them to play together. In USENIX [USE01f], page ?? ISBN 1-880446-10-3. LCCN QA76.8.U65

                                        U84 2001. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/usenix01/
                                        freenix01/westerlund.html.

Weaver:2004:RW

[WE04] Nicholas Weaver and Dan Ellis. Reflections on witty. ;login: the USENIX Association newsletter, 29(3): ??, June 2004. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2004-06/pdfs/weaver.pdf.

Weaver:2006:WWD

[WE06] Nicholas Weaver and Dan Ellis. White worms don’t work. ;login: the USENIX Association newsletter, 31(6):??, December 2006. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        december-2006-volume-31-
                                        number-6/white-worms-dont-
                                        work.

Welch:2000:TWS

[Wel00]  Brent Welch.  The Tcl

Web server. In USE-NIX [USE00e], page ?? ISBN 1-880446-24-3. LCCN

                                        ????  URL http://db.
                                        usenix.org/publications/
                                        library/proceedings/tcl2k/
                                        welch.html.

Wettstein:2000:GML

[WG00] Greg Wettstein and Johannes Grosen. Gaining the middleground: a Linuxbased open-source middleware initiative. In USE-NIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/als2000/
wettstein.html.

Wang:2002:WNR

[WH02] Jun Wang and Yiming Hu.

WOLF — a novel reordering write buffer to boost the performance of logstructured file system. In USENIX [USE02c], page ?? ISBN 1-880446-03-0. LCCN QA76.9.F5 C65 2002. URL

http://www.usenix.org/publications/
library/proceedings/fast02/
wang.html.

White:2000:AW

[Whi00] Jeremy White. The aging of Wine, 2000. URL http://

www.usenix.org/publications/
library/proceedings/als2000/

. Unpublished talk at Annual Linux Showcase 2000.

Whitehead:2001:KIS

[Whi01]  Ellis Whitehead.  KDE:

Interface, standards, and inter-process communication. In USENIX [USE01d], page ?? ISBN ???? LCCN ???? URL http:

//www.linuxshowcase.org/
whitehead.html.

Williams:2002:ISA

[Wil02] Nathan J. Williams. An implementation of scheduler activations on the NetBSD operating system. In USE-NIX [USE02d], page ?? ISBN 1-880446-01-4. LCCN QA76.8.U65 P765 2002.

URL http://www.usenix.
org/publications/library/
proceedings/usenix02/tech/
freenix/williams.html.

Wippler:2000:SD

[Wip00] Jean-Claude Wippler. Scripted documents. In USENIX [USE00e], page ?? ISBN 1-880446-24-3. LCCN ????

URL http://db.usenix.
org/publications/library/
proceedings/tcl2k/wippler.
html.

Wang:2004:OBP

[WJL04] Jinling Wang, Beihong Jin, and Jing Li. An ontologybased publish/subscribe system. In Jacobsen [Jac04], pages 232–?? CODEN LNCSD9. ISBN 3-540-23428-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.M54 I33 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.springerlink.
com/openurl.asp?genre=issue&
issn=0302-9743&volume=3231;
 http://www.springerlink.
com/openurl.asp?genre=volume&
id=doi:10.1007/b101561.

Woodard:2000:PP

[WM00] Ben Woodard and Nick Moffitt. Perspective on printing. In USENIX [USE00c], page ?? ISBN 1-880446-17-0. LCCN ????

URL http://www.usenix.
org/publications/library/
proceedings/als2000/woodard.
html.

Wolf:2002:P

[Wol02] Gunnar Wolf. ProtoWrap.

;login: the USENIX Association newsletter, 27(3): ??, June 2002. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2002-06/pdfs/wolf.pdf.

Wonnacott:2001:UAF

[Won01]  David Wonnacott.  Us-

ing accessory functions to generalize dynamic dispatch in single-dispatch objectoriented languages. In USENIX [USE01e], page ?? ISBN 1-880446-12-X. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/coots01/
wonnacott.html.

Wood:2001:BOS

[Woo01a] John Wood. Building an open-source Solaris-compatible threads library. In USE-NIX [USE01f], page ?? ISBN 1-880446-10-3. LCCN QA76.8.U65 U84 2001. URL

http://www.usenix.org/publications/

library/proceedings/usenix01/
freenix01/wood.html.

Wool:2001:ALF

[Woo01b] Avishai Wool. Architecting the Lumeta Firewall Analyzer. In USENIX [USE01h], page ?? ISBN 1-880446-18-9, 1-880446-07-

3. LCCN QA76.9.A25 U565

2001. URL http://www.
usenix.org/publications/
library/proceedings/sec01/
wool.html.

Wool:2002:CPP

[Woo02] Avishai Wool. Combating the perils of port 80 at the firewall. ;login: the USENIX Association newsletter, 27(4): 44–45, August 2002. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2002-08/pdfs/wool.
pdf.

Ward:2002:AAS

[WOSW02] Julie Ward, Michael O’Sullivan, Troy Shahoumian, and John Wilkes. Appia: Automatic storage area network fabric design. In USENIX [USE02c], page ?? ISBN 1-880446-03-

0. LCCN QA76.9.F5 C65

2002. URL http://www.
usenix.org/publications/
library/proceedings/fast02/
ward.html.

Wang:2001:DPM

[WPSO01] Nanbor Wang, Kirthika Parameswaran, Douglas Schmidt, and Ossama Othman. The design and performance of meta-programming mechanisms for object request broker middleware. In USE-NIX [USE01e], page ?? ISBN 1-880446-12-X. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/coots01/
wang.html.

Wang:2000:TBD

[WRA00] Yi-Min Wang, Wilf Russell, and Anish Arora. A toolkit for building dependable and extensible home networking applications. In USENIX [USE00d], page ?? ISBN 1-880446-20-0. LCCN

????  URL http://db.
usenix.org/publications/
library/proceedings/usenix-
win2000/wang.html.

Waldman:2000:PRT

[WRC00] Marc Waldman, Aviel D. Rubin, and Lorrie Faith Cranor. Publius: a robust, tamperevident, censorship-resistant Web publishing system. In USENIX [USE00i], page ?? ISBN 1-880446-18-9. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/sec2000/
waldman.html.

Farrow:2007:EOG

[wRF07] Gary McGraw with Rik Farrow. Exploiting online games: An interview. ;login: the USENIX Association newsletter, 32(6):??, December 2007. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        december-2007-volume-32-
                                        number-6/exploiting-online-
                                        games-interview.

Wang:2002:CBP

[WRPK02] An-I Wang, Peter Reiher, Gerald Popek, and Geoffrey H. Kuenning. Conquest: Better performance through a disk/persistent-RAM hybrid file system. In USENIX [USE02e], page ?? ISBN 1-880446-00-6. LCCN QA76.8.U65 U84 2002. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/usenix02/
                                        wang.html.

Welch:2001:KUB

[WS01] Ian Welch and Robert J. Stroud. Kava — using byte code rewriting to add behavioural reflection to Java. In USE-NIX [USE01e], page ?? ISBN 1-880446-12-X. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/coots01/
                                        welch.html.

Watanabe:2002:AIF

[WS02] Takanori Watanabe and Michael Smith. ACPI implementation on FreeBSD. In USENIX [USE02d], page ?? ISBN 1-880446-01-4. LCCN QA76.8.U65 P765 2002.

                                        URL http://www.usenix.

org/publications/library/
proceedings/usenix02/tech/
freenix/watanabe.html.

Welch:2000:TEA

[WT00] Brent Welch and Michael Thomas. The Tcl Extension Architecture. In USENIX [USE00e], page ?? ISBN 1-880446-24-3. LCCN

????  URL http://db.
usenix.org/publications/
library/proceedings/tcl2k/
welchextension.html.

Wong:2002:MCY

[WW02] Theodore Wong and John Wilkes. My cache or yours? Making storage more exclusive. In USENIX [USE02e], page ?? ISBN 1-880446-00-6. LCCN QA76.8.U65

U84 2002. URL http://www.
usenix.org/publications/
library/proceedings/usenix02/
wong.html.

Wynn:2000:GSP

[Wyn00] Bruce Alan Wynn. Growing SysAdmin as a profession; local groups. ;login: the USE-NIX Association newsletter, 25(4):??, July 2000. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2000-7/sagenews.html#growing.

Wysopal:2003:LSQ

[Wys03] Chris Wysopal. Learning security QA from the vulnerability researchers. ;login: the USENIX Association newsletter, 28(6):??, December 2003. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2003-12/pdfs/wysopal.pdf.

Wysopal:2004:PTS

[Wys04]  Chris Wysopal.  Putting

trust in software code. ;login: the USENIX Association newsletter, 29(6):??, December 2004. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2004-12/pdfs/code.pdf.

Yardley:2008:SIV

[Yar08] Tim Yardley. SCADA: Issues, vulnerabilities, and future directions. ;login: the USENIX Association newsletter, 33(6): ??, December 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
december-2008-volume-33-
number-6/scada-issues-vulnerabilities-
and-future.

Yocum:2001:PCH

[YC01] Kenneth Yocum and Jeffrey Chase. Payload caching: High-speed data forwarding for network intermediaries. In USENIX [USE01b], page ?? ISBN 1-880446-09-X. LCCN QA76.8.U65

U84 2001. URL http://www.
usenix.org/publications/
library/proceedings/usenix01/
yocum.html.

Yaghmour:2000:MCS

[YD00] Karim Yaghmour and Michel R.

Dagenais. Measuring and characterizing system behavior using kernel-level event logging. In USE-NIX [USE00a], pages 13–

26. ISBN 1-880446-22-7. LCCN ???? URL http:

//www.usenix.org/events/
usenix2000/general/yaghmour.
html.

Yener:2000:SA

[Yen00]  Bülent Yener.  Smart-

box architecture. In USE-NIX [USE00j], page ?? ISBN ???? LCCN

????  URL http://db.
usenix.org/publications/
library/proceedings/ine2000/
yener.html.

Yu:2000:TCP

[YGC^{+}00] Xiang Yu, Benjamin Gum, Yuqun Chen, Randolph Y. Wang, Kai Li, Arvind Krishnamurthy, and Thomas E. Anderson. Trading capacity for performance in a disk array. In USE-NIX [USE00h], page ?? ISBN 1-880446-16-2. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/osdi2000/
yugum.html.

Yamamura:2002:SKS

[YHS^{+}02] Shuji Yamamura, Akira Hirai, Mitsuru Sato, Masao Yamamoto, Akira Naruse, and Kouichi Kumon. Speeding up the kernel scheduler by reducing cache misses. In USENIX [USE02d], page ?? ISBN 1-880446-01-4. LCCN QA76.8.U65 P765 2002.

URL http://www.usenix.
org/publications/library/
proceedings/usenix02/tech/
freenix/yamamura.html.

Young:2005:UNS

[YM05] Ellie Young and Tara Mulligan. USENIX notes: Summary of USENIX Board of Directors meetings. ;login: the USENIX Association newsletter, 30(6):??, December 2005. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
2005-12/openpdfs/board.
pdf.

Yang:2002:DAA

[YMC02] Chuan-Kai Yang, Tulika Mitra, and Tzi-Cker Chiueh. A decoupled architecture for application-specific file prefetching. In USE-NIX [USE02d], page ?? ISBN 1-880446-01-4. LCCN QA76.8.U65 P765 2002.

URL http://www.usenix.
org/publications/library/
proceedings/usenix02/tech/
freenix/yang.html.

Young:2001:RP

[YN01] John Young and Deborah Natsios. Reversing the panopticon, 2001. Unpublished invited talk, Tenth USENIX Security Symposium, August 13–17, 2001, Washington, DC, USA.

Yang:2001:MTC

[YNN01] S. Jae Yang, Jason Nieh, and Naomi Novik. Measuring thin-client performance using slow-motion benchmarking. In USENIX [USE01b], page ?? ISBN 1-880446-09-X. LCCN QA76.8.U65

U84 2001. URL http://www.
usenix.org/publications/
library/proceedings/usenix01/
yang.html.

Yang:2002:PRD

[YNST02] S. Jae Yang, Jason Nieh, Matt Selsky, and Nikhil Tiwari. The performance of remote display mechanisms for thin-client computing. In USENIX [USE02e], page ?? ISBN 1-880446-00-6. LCCN QA76.8.U65

U84 2002. URL http://www.
usenix.org/publications/
library/proceedings/usenix02/
yang.html.

Young:2003:EBD

[You03]  Ellie Young.  2004 elec-

tion for Board of Directors. ;login: the USENIX Association newsletter, 28(6): ??, December 2003. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2003-12/openpdfs/
usenixnews.pdf.

Young:2004:TOV

[You04a] Ellie Young. Thanks to our volunteers. ;login: the USENIX Association newsletter, 29(6):??, December 2004. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2004-12/openpdfs/unotes.
                                        pdf.

Young:2004:UAF

[You04b] Ellie Young. The USENIX Association financial report for 2003. ;login: the USE-NIX Association newsletter, 29(4):??, August 2004. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2004-08/openpdfs/
                                        unotes.pdf.

Young:2005:UNT

[You05a] Ellie Young. USENIX notes: Thanks to our volunteers. ;login: the USENIX Association newsletter, 30(6):??, December 2005. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2005-12/openpdfs/thanks.
                                        pdf.

Young:2005:UNV

[You05b] Ellie Young. USENIX notes: Vale, Rob, atque Ave, Rik. ;login: the USENIX Association newsletter, 30(3): ??, June 2005. CODEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2005-06/openpdfs/
ave0506.pdf.

Young:2006:TOV

[You06a] Ellie Young. Thanks to our volunteers. ;login: the USENIX Association newsletter, 31(6):??, December 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
december-2006-volume-31-
number-6/thanks-our-volunteers.

Young:2006:UAF

[You06b] Ellie Young. The USENIX Association financial report for 2005. ;login: the USENIX Association newsletter, 31(4): ??, August 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2006-volume-31-number-
4/usenix-association-financial-
report-2005.

Young:2006:UBD

[You06c] Ellie Young. USENIX Board of Directors meetings and actions. ;login: the USENIX Association newsletter, 31(4): ??, August 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2006-volume-31-number-
4/usenix-board-directors-
meetings-and-actions.

Young:2007:EUB

[You07a] Ellie Young. 2008 election of the USENIX Board of Directors. ;login: the USENIX Association newsletter, 32(6): ??, December 2007. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        december-2007-volume-32-
                                        number-6/2008-election-
                                        usenix-board-directors.

Young:2007:UNC

[You07b] Ellie Young. 2008 USE-NIX Nominating Committee. ;login: the USENIX Association newsletter, 32(5): ??, October 2007. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        october-2007-volume-32-
                                        number-5/2008-usenix-nominating-
                                        committee.

Young:2007:TOV

[You07c] Ellie Young. Thanks to our volunteers. ;login: the USENIX Association newsletter, 32(6):??, December 2007. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        december-2007-volume-32-
                                        number-6/thanks-our-volunteers.

Young:2007:UAF

[You07d] Ellie Young. USENIX Association financial report for 2006. ;login: the USENIX Association newsletter, 32(4):

??, August 2007. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2007-volume-32-number-
4/usenix-association-financial-
report-2006.

Young:2007:UBD

[You07e] Ellie Young. USENIX Board of Directors meetings and actions. ;login: the USENIX Association newsletter, 32(1): ??, February 2007. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
february-2007-volume-32-
number-1/usenix-board-directors-
meetings-and-actions.

Young:2008:SUB

[You08a] Ellie Young. Summary of USENIX Board of Directors meetings and actions. ;login: the USENIX Association newsletter, 33(1):??, February 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
february-2008-volume-33-
number-1/summary-usenix-
board-directors-meetings-
and.

Young:2008:TOV

[You08b] Ellie Young. Thanks to our volunteers. ;login: the USENIX Association newsletter, 33(6):??, December 2008. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        december-2008-volume-33-
                                        number-6/thanks-our-volunteers.

Young:2009:EUB

[You09a] Ellie Young. 2010 election for the USENIX Board of Directors. ;login: the USENIX Association newsletter, 34(6): ??, December 2009. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        december-2009-volume-34-
                                        number-6/2010-election-
                                        usenix-board-directors.

Young:2009:SUB

[You09b] Ellie Young. Summary of USENIX Board of Directors actions. ;login: the USENIX Association newsletter, 34(5): ??, October 2009. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        october-2009-volume-34-
                                        number-5/summary-usenix-
                                        board-directors-actions.

Young:2009:TJL

[You09c] Ellie Young. Tribute to Jay Lepreau, 1952–2008. ;login: the USENIX Association newsletter, 34(2): 87, April 2009. CODEN LOGNEM. ISSN 1044-

                                        6397. URL https://www.
                                        usenix.org/publications/
                                        login/april-2009-volume-
                                        34-number-2/tribute-jay-
                                        lepreau-1952%E2%80%932008.

Young:2009:UAF

[You09d] Ellie Young. USENIX Association financial report for 2008. ;login: the USENIX Association newsletter, 34(4): ??, August 2009. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2009-volume-34-number-
4/usenix-association-financial-
report-2008.

Ye:2002:TPB

[YS02] Zishuang (Eileen) Ye and Sean Smith. Trusted paths for browsers. In USENIX [USE02b], pages 263–279. ISBN 1-931971-00-5. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/sec02/
ye.html.

Yu:2002:SBC

[YST02] Dachuan Yu, Zhong Shao, and Valery Trifonov. Supporting binary compatibility with static compilation. In USENIX [USE02f], page ?? ISBN 1-931971-01-3. LCCN QA76.73 .J38

J42 2002. URL http://www.
usenix.org/publications/
library/proceedings/javavm02/
yu.html.

Yu:2000:DEC

[YV00] Haifeng Yu and Amin Vahdat. Design and evaluation of a continuous consistency model for replicated services. In USE-NIX [USE00h], page ?? ISBN 1-880446-16-2. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/osdi2000/
yuvahdat.html.

Zwicky:2006:BRa

[Z^{+}06a] Elizabeth Zwicky et al. Book reviews. ;login: the USE-NIX Association newsletter, 31(1):??, February 2006. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://www.
usenix.org/publications/
login/2006-02/openpdfs/
reviews.pdf.

etal:2006:BRb

[Z^{+}06b] Elizabeth Zwicky et al. Book reviews. ;login: the USENIX Association newsletter, 31(3):??, June 2006. CODEN LOGNEM. ISSN 1044-6397. URL

https://www.usenix.org/
publications/login/june-
2006-volume-31-number-3/
book-reviews.

etal:2006:BRc

[Z^{+}06c] Elizabeth Zwicky et al. Book reviews. ;login: the USENIX Association newsletter, 31(4): ??, August 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2006-volume-31-number-
4/book-reviews.

etal:2006:BRd

[Z^{+}06d] Elizabeth Zwicky et al. Book reviews. ;login: the USENIX Association newsletter, 31(5): ??, October 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
october-2006-volume-31-
number-5/book-reviews.

etal:2006:BRe

[Z^{+}06e] Elizabeth Zwicky et al. Book reviews. ;login: the USENIX Association newsletter, 31(6): ??, December 2006. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
december-2006-volume-31-
number-6/book-reviews.

etal:2007:BRa

[Z^{+}07a] Elizabeth Zwicky et al. Book reviews. ;login: the USENIX Association newsletter, 32(1): ??, February 2007. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
february-2007-volume-32-
number-1/book-reviews.

etal:2007:BRb

[Z^{+}07b] Elizabeth Zwicky et al. Book reviews. ;login: the USE-NIX Association newsletter, 32(2):??, April 2007. CO-DEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/april-2007-volume-
32-number-2/book-reviews.

etal:2007:BRc

[Z^{+}07c] Elizabeth Zwicky et al. Book reviews. ;login: the USENIX Association newsletter, 32(3):??, June 2007. CODEN LOGNEM. ISSN 1044-6397. URL

                                        https://www.usenix.org/
                                        publications/login/june-
                                        2007-volume-32-number-3/
                                        book-reviews.

etal:2007:BRd

[Z^{+}07d] Elizabeth Zwicky et al. Book reviews. ;login: the USENIX Association newsletter, 32(4): ??, August 2007. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        august-2007-volume-32-number-
                                        4/book-reviews.

etal:2007:BRe

[Z^{+}07e] Elizabeth Zwicky et al. Book reviews. ;login: the USENIX Association newsletter, 32(5): ??, October 2007. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        october-2007-volume-32-
                                        number-5/book-reviews.

etal:2007:BRf

[Z^{+}07f] Elizabeth Zwicky et al. Book reviews. ;login: the USENIX Association newsletter, 32(6): ??, December 2007. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/

december-2007-volume-32-
number-6/book-reviews.

etal:2008:BRa

[Z^{+}08a] Elizabeth Zwicky et al. Book reviews. ;login: the USE-NIX Association newsletter, 33(2):??, April 2008. CO-DEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/april-2008-volume-
33-number-2/book-reviews.

etal:2008:BRb

[Z^{+}08b] Elizabeth Zwicky et al.

Book reviews. ;login: the USENIX Association newsletter, 33(3):??, June 2008. CODEN LOGNEM. ISSN 1044-6397. URL

https://www.usenix.org/
publications/login/june-
2008-volume-33-number-3/
book-reviews.

etal:2008:BRc

[Z^{+}08c] Elizabeth Zwicky et al. Book reviews. ;login: the USENIX Association newsletter, 33(4): ??, August 2008. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2008-volume-33-number-
4/book-reviews.

etal:2008:BRd

[Z^{+}08d] Elizabeth Zwicky et al. Book reviews. ;login: the USENIX Association newsletter, 33(5): ??, October 2008. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        october-2008-volume-33-
                                        number-5/book-reviews.

etal:2008:BRe

[Z^{+}08e] Elizabeth Zwicky et al. Book reviews. ;login: the USENIX Association newsletter, 33(6): ??, December 2008. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        december-2008-volume-33-
                                        number-6/book-reviews.

etal:2009:BRa

[Z^{+}09a] Elizabeth Zwicky et al. Book reviews. ;login: the USENIX Association newsletter, 34(1): ??, February 2009. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        february-2009-volume-34-
                                        number-1/book-reviews.

etal:2009:BRb

[Z^{+}09b] Elizabeth Zwicky et al. Book reviews. ;login: the USE-NIX Association newsletter, 34(2):??, April 2009. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL https://www.
                                        usenix.org/publications/
                                        login/april-2009-volume-
                                        34-number-2/book-reviews.

etal:2009:BRc

[Z^{+}09c] Elizabeth Zwicky et al. Book reviews. ;login: the USENIX Association newsletter, 34(3):??, June 2009. CODEN LOGNEM. ISSN 1044-6397. URL

https://www.usenix.org/
publications/login/june-
2009-volume-34-number-3/
book-reviews.

etal:2009:BRd

[Z^{+}09d] Elizabeth Zwicky et al. Book reviews. ;login: the USENIX Association newsletter, 34(4): ??, August 2009. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
august-2009-volume-34-number-
4/book-reviews.

etal:2009:BRe

[Z^{+}09e] Elizabeth Zwicky et al. Book reviews. ;login: the USENIX Association newsletter, 34(5): ??, October 2009. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
october-2009-volume-34-
number-5/book-reviews.

etal:2009:BRf

[Z^{+}09f] Elizabeth Zwicky et al. Book reviews. ;login: the USENIX Association newsletter, 34(6): ??, December 2009. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
december-2009-volume-34-
number-6/book-reviews.

Zadok:2001:FIS

[ZABN01] Erez Zadok, Johan M. Andersen, Ion Badulescu, and Jason Nieh. Fast indexing: Support for size-changing algorithms in stackable file systems. In USENIX [USE01b], page ?? ISBN 1-880446-09-X. LCCN QA76.8.U65

U84 2001. URL http://www.
usenix.org/publications/
library/proceedings/usenix01/
zadok.html.

Zadok:2002:OAC

[Zad02]  Erez Zadok.  Overhaul-

ing Amd for the ’00s: a case study of GNU Autotools. In USENIX [USE02d], page ?? ISBN 1-880446-01-

4. LCCN QA76.8.U65 P765

2002. URL http://www.
usenix.org/publications/
library/proceedings/usenix02/
tech/freenix/zadok.html.

Zendra:2002:STC

[ZD02] Olivier Zendra and Karel Driesen. Stress-testing control structures for dynamic dispatch in Java. In USE-NIX [USE02f], page ?? ISBN 1-931971-01-3. LCCN QA76.73 .J38 J42 2002.

URL http://www.usenix.
org/publications/library/
proceedings/javavm02/zendra.
html.

Zhang:2002:UCS

[ZEJ02] Xiaolan Zhang, Antony Edwards, and Trent Jaeger. Using CQUAL for static analysis of authorization hook placement. In USE-NIX [USE02b], pages 33–48. ISBN 1-931971-00-5. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/sec02/
zhang.html.

Zhao:2000:DIA

[ZM00] Q. Alex Zhao and D. Scott McCrickard. Designing information awareness applications in Tcl/Tk. In USENIX [USE00e], page ?? ISBN 1-880446-24-3. LCCN

????  URL http://db.
usenix.org/publications/
library/proceedings/tcl2k/
full_papers/zhao_poster/
zhao_layout.pdf;   http:
//db.usenix.org/publications/
library/proceedings/tcl2k/
full_papers/zhao_poster/
zhao_poster.pdf.

Zadok:2000:FLS

[ZN00] Erez Zadok and Jason Nieh.

FiST: a language for stackable file systems. In USE-NIX [USE00a], pages 55–

70. ISBN 1-880446-22-7. LCCN ???? URL http:

//www.usenix.org/events/
usenix2000/general/zadok.
html.

Zhang:2000:DB

[ZP00a] Yin Zhang and Vern Paxson. Detecting backdoors. In USENIX [USE00i], page ?? ISBN 1-880446-18-9. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/sec2000/
zhangbackdoor.html.

Zhang:2000:DSS

[ZP00b] Yin Zhang and Vern Paxson. Detecting stepping stones. In USENIX [USE00i], page ?? ISBN 1-880446-18-9. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/sec2000/
                                        zhangstepping.html.

Zhou:2001:MQR

[ZPL01] Yuanyuan Zhou, James Philbin, and Kai Li. The multi-queue replacement algorithm for second level buffer caches. In USE-NIX [USE01b], page ?? ISBN 1-880446-09-X. LCCN QA76.8.U65 U84 2001. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/usenix01/
                                        zhou.html.

Zhang:2008:HPB

[ZPU08] Jian Zhang, Phillip Porras, and Johannes Ullrich. Highly predictive blacklisting. ;login: the USENIX Association newsletter, 33(6):??, December 2008. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        december-2008-volume-33-
                                        number-6/highly-predictive-
                                        blacklisting.

Zhang:2000:MLI

[ZS00] Yongguang Zhang and Bikramjit Singh. A multi-layer IPsec protocol. In USENIX [USE00i], page ?? ISBN 1-880446-18-9. LCCN ????

URL http://www.usenix.
org/publications/library/
proceedings/sec2000/zhangipsec.
html.

Zhang:2001:HJA

[ZS01] Xiaolan Zhang and Margo Seltzer. HBench:JGC — an application-specific benchmark suite for evaluating JVM garbage collector performance. In USE-NIX [USE01e], page ?? ISBN 1-880446-12-X. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/coots01/
zhang.html.

Zhao:2004:SPH

[ZSB04] Yuanyuan Zhao, Daniel Sturman, and Sumeer Bhola. Subscription propagation in highly-available publish/ subscribe middleware. In Jacobsen [Jac04], pages 274– ?? CODEN LNCSD9. ISBN 3-540-23428-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.76.M54 I33 2004. URL ftp://

uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;
http://www.springerlink.
com/openurl.asp?genre=issue&
issn=0302-9743&volume=3231;
 http://www.springerlink.
com/openurl.asp?genre=volume&
id=doi:10.1007/b101561.

Zwicky:2006:BRb

[ZSF06] Elizabeth Zwicky, Sam Stover, and Rik Farrow. Book reviews. ;login: the USE-NIX Association newsletter, 31(2):76–79, April 2006. CO-DEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2006-04/openpdfs/
                                        reviews.pdf.

Zwicky:2005:BR

[ZT05] Elizabeth Zwicky and Adam Turoff. Book reviews. ;login: the USENIX Association newsletter, 30(5):??, October 2005. CODEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2005-10/openpdfs/
                                        reviews.pdf.

Zwieback:2001:HAF

[Zwi01] Dave Zwieback. High availability firewalls/VPN with VRRP. ;login: the USENIX Association newsletter, 26(8): ??, December 2001. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2001-12/pdfs/zwieback.pdf.

Zwicky:2005:BRO

[Zwi05] Elizabeth Zwicky. Book reviews: Online only! ;login: the USENIX Association newsletter, 30(6):??, December 2005. CODEN LOGNEM. ISSN 1044-

                                        6397. URL http://www.
                                        usenix.org/publications/
                                        login/2005-12/openpdfs/
                                        reviews_addendum.pdf.

Zhang:2002:CSE

[ZYKW02] Chi Zhang, Xiang Yu, Arvind Krishnamurthy, and Randolph Y. Wang. Configuring and scheduling an eagerwriting disk array for a transaction processing workload. In USENIX [USE02c], page ?? ISBN 1-880446-03-0. LCCN QA76.9.F5 C65

2002. URL http://www.
usenix.org/publications/
library/proceedings/fast02/
zhang.html.